### v3

# Zhang, Liu, and Zhao

[…]
Zhang, Liu, and Zhao
[…]
Data-driven PADR for SP with Covariate Information
[…]
Data-driven Piecewise Affine Decision Rules for Stochastic Programming with Covariate Information
[…]
Focusing on stochastic programming (SP) with covariate information, this paper proposes an empirical risk minimization (ERM) method embedded within a nonconvex piecewise affine decision rule (PADR), which aims to learn the direct mapping from features to optimal decisions. We establish the nonasymptotic consistency result of our PADR-based ERM model for unconstrained problems and asymptotic consistency result for constrained ones. To solve the nonconvex and nondifferentiable ERM problem, we develop an enhanced stochastic majorization-minimization algorithm and establish the asymptotic convergence to (composite strong) directional stationarity along with complexity analysis. We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational tractability. Our numerical study demonstrates the superior performance of PADR-based ERM methods compared to state-of-the-art approaches under various settings, with significantly lower costs, less computation time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
[…]
1 Introduction
Pioneered by George Dantzig in the mid-1950s, stochastic programming (SP) provides a class of optimization models for decision-making under uncertainty by treating the uncertainty as a random variable whose probability distribution is usually unknown in practice. With the growing real-world applications and availability of large datasets, the field of SP has flourished with many theoretical and algorithmic advances (Birge and Louveaux 2011, Shapiro et al. 2021), in which the empirical probability distribution based on historical scenarios plays a central role in approximating the true distribution. However, the empirical distribution may provide poor estimation when the random variable is related to covariate information (also referred to as contextual information or features) which can be observed before the decision is made. An unconstrained SP with the covariate information $X=x\in\mathbb{R}^{p}$ is formulated as follows,
$\inf_{z\in\mathbb{R}^{d}}\quad\mathbb{E}_{Y}\left[\varphi({z};{Y})\mid X=x\right],$ (1)
where $z\in\mathbb{R}^{d}$ is a decision vector, $\varphi:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}$ is a cost function, and $Y$ is a random variable supported by $\mathcal{Y}$ whose distribution is related to a feature vector $X$ supported by $\mathcal{X}\subseteq\mathbb{R}^{p}$. A simple instance is an inventory planning problem, in which $z$ represents order decisions for $d$ products, $\varphi$ represents the inventory cost, $Y$ represents uncertain demands, and $X$ represents covariate information such as seasonality, promotion, and weather. With historical scenarios of $(X,Y)$, a classical sample average approximation (SAA) model disregards the dependency of $Y$ on $X$ and thus may lead to inconsistent decisions; such deficiency is also observed in numerical results of Ban and Rudin (2018) and Bertsimas and Koduri (2021). Motivated by the latent dependency of the optimal solution to (1) on the feature vector, this paper aims to construct a data-driven decision rule that directly maps features to optimal decisions so as to approximately solve the SP problem (1) for any possible observation of the feature.
1.1 Problem Setting
Solving problem (1) for any realization of $X$ can be generalized as seeking an optimal decision rule (DR) to the following problem
$\inf_{f\in\mathcal{F}}\quad R(f)\coloneqq\mathbb{E}_{X,Y}\left[\varphi({f(X)};{Y})\right],$ (2)
where $\mathcal{F}$ contains all measurable functions mapping from the feature space $\mathcal{X}$ to the decision space $\mathbb{R}^{d}$. By the interchangeability principle (Shapiro et al. 2021, Theorem 7.80), if a mapping $f^{*}\in\mathcal{F}$ is optimal to problem (2), then the decision $f^{*}(x)$ is optimal to (1) for almost every $x\in\mathcal{X}$. In general, problem (2) is intractable due to the infinite-dimensional function space and the unknown joint distribution of $X$ and $Y$. A common approximation method is to construct an empirical risk minimization (ERM) problem based on a dataset $S\coloneqq\{(x^{s},y^{s})\}_{s=1}^{n}$ with the decision rule $f$ restricted to a parametric DR class $\mathcal{H}$ (also known as hypothesis class):
$\inf_{f\in\mathcal{H}}\quad\widehat{R}_{S}(f)\coloneqq\frac{1}{n}\sum_{s=1}^{n}\varphi({f(x^{s})};{y^{s}}).$ (3)
In this paper, we adopt the above DR-based ERM framework, recognizing that the efficacy of such a framework is influenced by the trade-off between approximation accuracy of $\mathcal{H}$ to the optimal DR $f^{*}$ and computational complexity of solving (3). For example, problem (3) with a convex cost function can be solved efficiently with $\mathcal{H}$ being the class of linear decision rules (LDR) (Beutel and Minner 2012, Ban and Rudin 2018), while the LDR has difficulty describing the potential nonlinear relation between decisions and features. On the other side, a DR class consisting of neural networks could have a strong approximation capability (Qi et al. 2022); nevertheless, the resulting nonconvex problem (3) with a multi-layer composite structure is difficult to solve with theoretical optimality guarantees, and the model may face overfitting issues when the dataset is not large enough.
Therefore, this paper aims to address a critical question: what could be an amenable hypothesis class for the DR-based ERM framework so as to maintain both high reliability and computational efficiency for solving SP with covariate information? To this end, we employ in the ERM problem (3) a piecewise affine decision rule (PADR) that is broader than LDR but simpler than general nonlinear DRs. According to Scholtes (2012), any continuous piecewise affine (PA) function $f:\mathcal{X}\rightarrow\mathbb{R}$ can be expressed as the difference of max-affine functions:
$f(x;\theta)=\max_{k\in[K]}\left\{(\alpha^{k})^{\top}x+a^{k}\right\}-\max_{k\in[K]}\left\{(\beta^{k})^{\top}x+b^{k}\right\},$ (4)
where $\theta\coloneqq\left(\alpha^{k},a^{k},\beta^{k},b^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)}$ and $[K]\coloneqq\{1,\dots,K\}$. Throughout the theoretical analysis in this paper, we assume that the numbers of pieces in the two max-affine components of (4) are equal to simplify the exposition, which could be easily extended to the case with different numbers of pieces. With $\mathcal{H}$ being the class of continuous PA functions with each max-affine component having $K$ pieces and parameters $\{\theta_{\iota}\}_{\iota\in[d]}$ satisfying $\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu$ for $\iota\in[d]$, the ERM problem (3) is formulated as the following parameter estimation problem
$\min_{\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\forall\iota\in[d]}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi\left(\Big{(}\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x^{s}+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x^{s}+b^{k}_{\iota}\right\}\Big{)}_{\iota=1}^{d};y^{s}\right).$ (5)
In response to the aforementioned question, we intend to tackle two critical issues of the PADR-based ERM model in this paper: the consistency issue of the PADR-based ERM problem (5) and the computational issue for solving this large-scale, nonconvex, and nondifferentiable composite optimization problem. We address the first issue by quantifying the excess risk bound of our model, a standard procedure for model selection (Bartlett et al. 2002) in machine learning, and the second issue by developing a sampling-based algorithm with convergence guarantees, through which we demonstrate the ability of our proposed PADR-based ERM approach to effectively solve a broad class of SP problems with covariate information.
1.2 Literature Review
We present a literature review of data-driven methods for SP with covariate information and algorithmic advances in solving nonconvex optimization problems.
1.2.1 SP with Covariate Information.
There are two main schemes of data-driven approaches in the literature of SP dealing with the latent dependency on covariate information: the sequential and integrated schemes. In the sequential scheme, also referred to as the predict-then-optimize (PO) scheme (Deng and Sen 2022, Kannan et al. 2022), one first predicts uncertain parameters and then solves the optimization problem composite with the prediction model. Several works such as Kao et al. (2009), Donti et al. (2017), and Elmachtoub and Grigas (2021) propose modifications of the loss function by accounting for the cost error of the downstream optimization problem. Since those modified loss functions involve optimal solutions to the decision-making problems that are parameterized by suspicious prediction models, such frameworks usually lead to bi-level nonconvex optimization problems and thus encounter substantial computational challenges in dealing with general constrained nonlinear programs.
Approaches in the integrated scheme yield decisions via one-step data-driven optimization. Bertsimas and Kallus (2019) and Kallus and Mao (2022) approximate the conditional expected cost in (1) by a reweighted empirical cost. However, as claimed by Bertsimas and Koduri (2021), this framework is highly affected by the data with features close to the current observation and thus suffers from the curse of dimensionality. The DR-based ERM framework presented in (3) is another one-step approach in the integrated scheme. Due to its functional approximation in nature, this approach is relieved from the aforementioned issue.
Besides the classical LDR, several generalized LDRs are proposed for standard SP problems which lift random vectors into a higher-dimensional probability space to describe nonlinear relations (Chen et al. 2007, Bampou and Kuhn 2011, Georghiou et al. 2015). However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, except for the posterior estimation provided in Kuhn et al. (2011), their approximation accuracy lacks theoretical guarantees. The PADR with the parametric formula (4) employed in Bertsimas and Georghiou (2015) and this paper requires no prior designs of breakpoints as in Georghiou et al. (2015); instead, the breakpoints are implicitly determined by parameters in (4). Unlike the mixed-integer formulation in Bertsimas and Georghiou (2015) for multi-stage adaptive optimization problems, we treat the PADR-based ERM problem (5) as a special class of nonconvex and nondifferentiable optimization problems and exploit algorithmic advances to accommodate the large sample size and high-dimensional feature space.
Several works on DR-based ERM methods are most relevant to this paper. Ban and Rudin (2018) employ LDR for the Newsvendor problem with nonasymptotic optimality guarantees under the assumption of the linear demand model. Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems. Compared to the above references, our PADR-based ERM model is suitable for a much broader class of SP problems that can be nonconvex, and by following the statistical learning theory (Vapnik 1999, Mohri et al. 2018), we establish the nonasymptotic consistency result of our model without any parametric assumptions on the optimal decision rule or the latent dependency of the random variable $Y$ on the feature $X$.
1.2.2 Difference-of-convex (DC) and Composite DC Programming.
Due to the finite-sum composite structure of the objective in (5) with potentially large sample size $n$ and feature dimension $p$, it is prohibitively challenging to obtain a global minimizer to such a nonconvex and nondifferentiable optimization problem. Thus, our goal is to compute a promising candidate for a local minimizer via a sampling-based algorithm.
When the cost function $\varphi(\cdot;y)$ is piecewise affine (e.g., the Newsvendor cost), the overall composite objective of (5) is also piecewise affine, which is a special DC function. Deterministic DC algorithms have been extensively studied for solving DC programs over the last several decades (Le Thi and Pham Dinh 2018). When the concave part of the DC decomposition is nondifferentiable, the basic DC algorithm converges to a critical point that satisfies the necessary condition for local optimality. To obtain a d(irectional)-stationary point that is the sharpest first-order stationary point for general nonconvex problems, Pang et al. (2016) study a class of nondifferentiable DC programs whose concave part is the maximum of a finite number of smooth convex functions and propose an enhanced DC algorithm which iteratively solves convex subproblems constructed by linearizing the $\varepsilon$-active piece of the concave part. Though sampling-based algorithms for stochastic DC programs have been recently studied in Nitanda and Suzuki (2017), Xu et al. (2019), and Le Thi et al. (2022) addressing smooth or partially smooth objectives, there still lacks a comprehensive study of the enhanced surrogation technique in stochastic DC algorithms to seek d-stationary solutions to the stochastic DC programs with nondifferentiable concave parts.
With a more general cost function that is not necessarily piecewise affine, the finite-sum composite structure of the objective lies at the cornerstone of algorithmic development for solving the PADR-based ERM problem (5). Cui et al. (2018) study the ERM problem with a convex composite DC structure and develop a majorization-minimization (MM) algorithm with an enhanced surrogation technique similar to Pang et al. (2016); however, such a deterministic algorithm requires iteratively solving finite-sum convex subproblems over all the data, raising concerns about its computational efficiency for solving large-scale ERM problems. The composite DC problem (5) can be regarded as a special class of compound SP studied in Liu et al. (2022) with finite support and solved by the stochastic majorization-minimization (SMM) algorithm proposed by the authors. Nonetheless, without the enhanced surrogation technique, SMM can only asymptotically compute a compound critical point when the inner concave function is nondifferentiable. Finally, we refer readers to Cui and Pang (2021) for a comprehensive study of modern nonconvex and nondifferentiable optimization, especially Chapters 6 and 7 for the study of stationary concepts and surrogation-based algorithms for nonconvex composite optimization problems.
1.3 Contributions
We summarize our major contributions as follows.
1. We propose a PADR-based ERM method for a general class of SP problems with covariate information, where both cost and constraint functions can be nonconvex. We quantify the approximation accuracy of the PA function class and establish the nonasymptotic consistency result of our framework for unconstrained problems and asymptotic consistency result for deterministic-constrained problems without any parametric assumptions on the optimal decision rule or the latent dependency.
2. For learning PADR from the ERM problem, we develop an enhanced SMM algorithm with asymptotic convergence to composite strong d-stationarity, which is a sharper stationary concept than d-stationarity specialized to the composite DC programs with PA inner functions. We further provide nonasymptotic convergence of the algorithm to d-stationarity, which to the best of our knowledge is the first complexity result for stochastic nonsmooth composite DC programs. The enhanced SMM algorithm is also extended with exact penalization for solving the PADR-based ERM problem with constraints.
3. We test our method through numerical experiments on the Newsvendor problem and its variations with comparisons to the PO methods and other prominent DR-based methods, including the (generalized) LDR and RKHS-DR. The results demonstrate the superior performance of our method with significantly lower costs and less computation time in most instances, particularly those with high-dimensional redundant features or highly nonlinear models.
The rest of this paper is organized as follows. In Section 2, we provide the theoretical consistency result of the PADR-based ERM model. We then develop the enhanced SMM algorithm for learning PADR with surrogation construction of composite PA objectives in Section 3 and present the convergence analysis in Section 4. In Section 5, we extend our framework to deterministic-constrained SP problems. In Section 6, we report numerical results. We end with conclusions in Section 7.
2 Theoretical Guarantees for PADR-based ERM
In this section, we provide theoretical guarantees of the PADR-based ERM model by first quantifying the approximation error of the PA function class to Lipschitz continuous functions and then presenting the nonasymptotic consistency result of the PADR-based ERM model via the excess risk bound. At the end of this section, we compare our theoretical results with prominent ERM models employing the LDR and RKHS-DR.
2.1 Approximation Accuracy of Piecewise Affine Function Class
PA functions are known to asymptotically approximate Lipschitz continuous functions. Balazs et al. (2015) and Siahkamari et al. (2020) quantify the approximation errors of the PA function class to convex and DC functions, respectively. We provide the approximation bound of the PA function class to Lipschitz continuous functions, which to the best of our knowledge is the first quantification result for such a general case. Specifically, with $L>0$ and $M>0$, let
[…]
where $\left\lVert f\right\rVert_{\infty}\coloneqq\sup_{x\in\mathcal{X}}\max_{\iota\in[d]}|f_{\iota}(x)|$. We define the PADR class ${\mathcal{H}}_{\mu}^{K}$ as
$\mathcal{H}^{K}_{\mu}\coloneqq\left\{\left(f_{\iota}(\,\cdot\,;\theta_{\iota}):\mathcal{X}\rightarrow\mathbb{R}\right)_{\iota=1}^{d}\ \middle|\ \begin{aligned} &f_{\iota}(x;\theta_{\iota})=\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x+b^{k}_{\iota}\right\},\\ &\theta_{\iota}\coloneqq\left(\alpha_{\iota}^{k},a_{\iota}^{k},\beta_{\iota}^{k},b_{\iota}^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)},\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\,\forall\,\iota\in[d]\end{aligned}\ \right\}.$ (6)
When $K=1$, $\mathcal{H}^{1}_{\mu}$ reduces to the LDR class. We provide the approximation bound of the PA function class to Lipschitz continuous functions below with the proof given in Appendix 8.1.
Proposition 2.1 (Approximation Bound of ${\mathcal{H}}_{\mu}^{K}$)
Suppose that $\mathcal{X}\subseteq\mathbb{R}^{p}$ is a compact set with $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}=\bar{X}$. Then given $L_{0}>0$ and $M_{0}>0$, for any $f_{0}\in\mathcal{F}_{L_{0},M_{0}}$ and ${\mathcal{H}}_{\mu}^{K}$ defined in (6) with $K\in\mathbb{N}$ and $\mu\geq\max\left\{\frac{1}{2}L_{0}K^{1/p},\frac{1}{4}pL_{0}\bar{X}K^{1/p}+\frac{1}{2}M_{0}\right\}$, we have
$\min_{f\in\mathcal{H}_{\mu}^{K}}\left\lVert f-f_{0}\right\rVert_{\infty}\leq 2(p^{1/2}+3)p^{1/2}L_{0}\bar{X}K^{-1/p}.$
The bound $\mathcal{O}(K^{-1/p})$ provided in Proposition 2.1 is weaker than $\mathcal{O}(K^{-2/p})$ given by Balazs et al. (2015) and Siahkamari et al. (2020) in approximating convex and DC functions. Regardless, our bound is applicable to general PA-based statistical learning problems in which the underlying true model may not be convex or DC; more importantly, it renders the succeeding nonasymptotic consistency result of the PADR-based ERM model without any parametric or convex assumptions on the optimal decision rule or latent dependency.
2.2 Nonasymptotic Consistency of PADR-based ERM
We impose the following assumptions for the consistency analysis:
  1. (A1)
    Samples in $S=\{(x^{s},y^{s})\}_{s=1}^{n}$ of $(X,Y)$ are independently and identically distributed.
  2. (A2)
    For some $\bar{X}>0$, the support set $\mathcal{X}\subseteq\mathbb{R}^{p}$ of $X$ is a compact set with $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}=\bar{X}$.
  3. (A3)
    For some $L_{\varphi}>0$, $C_{\varphi}>0$, $z_{0}\in\mathbb{R}^{d}$, and any $y\in\mathcal{Y}$, the cost function $\varphi({\cdot};{y})$ is Lipschitz continuous on $\mathbb{R}^{d}$ with modulus $L_{\varphi}$ and satisfies $\left\lvert\varphi(z_{0};y)\right\rvert\leq C_{\varphi}$.
  4. (A4)
    There exists a Lipschitz continuous function $f^{*}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ that is optimal to problem (2).
The boundedness of $\mathcal{X}$ in Assumption (A2) is common in practical applications. Under Assumption (A2), the optimal DR $f^{*}$ in Assumption (A4) is bounded on $\mathcal{X}$, and thus $f^{*}\in\mathcal{F}_{L_{z},M_{z}}$ for some $L_{z}>0$ and $M_{z}>0$.
With the optimal solution to (5) denoted by $f^{*}_{S,{\mathcal{H}}_{\mu}^{K}}$, the consistency of the PADR-based ERM model is evaluated by the excess risk $R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\underset{f\in\mathcal{F}}{\inf}\,R(f)$, which is the gap between the expected cost of the optimal empirical PADR and the minimal cost. With problem (2) restricted to ${\mathcal{H}}_{\mu}^{K}$, we decompose the excess risk into the approximation and estimation errors as follows,
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)}_{\text{estimation error}}+\underbrace{\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)-\inf_{f\in\mathcal{F}}R(f)}_{\text{approximation error}}.$ (7)
[…]
Proposition 2.2 (Uniform Generalization Bound of PADR-based ERM)
Under Assumptions (A1)-(A3), with $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert\leq\frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
The proof based on the Rademacher complexity theory is presented in Appendix 8.2. Combining the approximation and estimation error bounds, we obtain the excess risk bound of PADR-based ERM in Theorem 2.3 with the proof in Appendix 8.2.
Theorem 2.3 (Excess Risk Bound of PADR-based ERM)
Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})\,\leq\,$ $\inf_{f\in\mathcal{F}}R(f)+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}\bar{X}K^{-1/p}$
$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
With $\mu=\mathcal{O}(K^{1/p})$, the complexity of the excess risk bound with respect to the sample size $n$ and parameter $K$ is
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{\widetilde{\mathcal{O}}\left(K^{\frac{p+2}{2p}}n^{-\frac{1}{2}}\right)}_{\begin{subarray}{c}\text{estimation}\\ \text{error bound}\end{subarray}}+\underbrace{\mathcal{O}\left(K^{-\frac{1}{p}}\right)}_{\begin{subarray}{c}\text{approximation}\\ \text{error bound}\end{subarray}},$ (8)
where $\widetilde{\mathcal{O}}(\cdot)$ suppresses logarithmic factors. It can be seen that the sample size $n$ only affects the estimation error, whereas $K$, the number of max-affine pieces for PA functions in ${\mathcal{H}}_{\mu}^{K}$, plays the central role of balancing the two error terms. As $K$ grows, the PA function class becomes larger, so the approximation error decreases due to the stronger approximation ability of ${\mathcal{H}}_{\mu}^{K}$, and the estimation error increases due to the more pronounced overfitting issue. By setting $K=\mathcal{O}(n^{p/(p+4)})$, we derive the optimal excess risk bound $\widetilde{\mathcal{O}}(n^{-{1}/(4+p)})$.
2.3 Consistency Comparison with Existing DR-based Methods
[…]
In comparison, our PADR-based ERM model is nonasymptotically consistent for general SP problems, including nonconvex ones. Moreover, if there exists a convex or DC optimal DR, based on the approximation bounds $\mathcal{O}(K^{-2/p})$ of ${\mathcal{H}}_{\mu}^{K}$ from Balazs et al. (2015) and Siahkamari et al. (2020), our excess risk bound can be improved to $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$ with $K=\mathcal{O}(n^{p/(p+6)})$, which is close to the bound of NV-ERM1 (a special case of our model) when $p$ is large. This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$ for SP with a box-constrained set $[-M,M]^{d}$, by employing a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{\mu}$ in Theorem 2.3 could be independent of $\mu$.
Though obtaining the global optimum $f^{*}_{S,{\mathcal{H}}_{\mu}^{K}}$ of the nonconvex ERM problem (5) is prohibitively challenging, we show in Sections 3 and 4 how to asymptotically compute a stationary solution that could be globally optimal to problem (5) under particular algorithmic settings, upon which our PADR-based ERM framework is asymptotically optimal.
3 Enhanced SMM Algorithm for Learning PADR
In this section, we focus on learning PADR by solving problem (5). Following the majorization-minimization (MM) framework, we develop a stochastic algorithm that iteratively solves convex subproblems constructed with sampling-based enhanced surrogation functions, referred to as an enhanced SMM algorithm. The highlight of our algorithm that distinguishes itself from the deterministic MM-related algorithms in Pang et al. (2016) and Cui et al. (2018) and stochastic ones in Xu et al. (2019), Le Thi et al. (2022), and Liu et al. (2022) is that it combines the enhanced $\varepsilon$-active surrogation technique with sequential sampling so as to converge to d-stationary solutions with a significant improvement of computational efficiency.
We first make some remarks on notations to facilitate algorithmic development and convergence analysis. With a slight abuse of notation, to emphasize the role of PADR as a function of $\theta$, we represent the PA function as $f(\theta;\xi)=g(\theta;\xi)-h(\theta;\xi)$ with $\xi\in S=\{(x^{s},y^{s})\}_{s=1}^{n}$, where $g(\theta;\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta;\xi)$ and $h(\theta;\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta;\xi)$ are maximum of linear functions. With $\widetilde{\xi}_{n}$ representing a random variable uniformly distributed on $S$, the PADR-based ERM problem (5) can be generalized as the following stochastic program with finite support:
$\min_{\theta\in\Theta}\quad F(\theta)\coloneqq\mathbb{E}\,\left[\varphi(f(\theta;\widetilde{\xi}_{n});\widetilde{\xi}_{n})\right],$ (9)
where $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set. In particular, in problem (5) with $\Theta=[-\mu,\mu]^{q}$ and $q=2dK(p+1)$, the outer function $\varphi(\cdot;\xi)$ with $\xi=(x,y)$ depends solely on $y$, and the inner PA function $f(\cdot;\xi)$ depends solely on $x$.
3.1 Construction of Surrogation Functions
Similar to other MM-related algorithms, the cornerstone of the enhanced SMM algorithm is the construction of surrogation functions of $F(\cdot;\xi)\coloneqq\varphi(f(\cdot;\xi);\xi)$. For a directionally differentiable function $\phi:\Theta\to\mathbb{R}$, that is, the directional derivative $\phi^{\prime}(\theta;d)\coloneqq\lim_{\tau\to 0}\frac{\phi(\theta+\tau d)-\phi(\theta)}{\tau}$ exists for any $\theta\in\Theta$ and $d\in\mathbb{R}^{q}$, we are interested in constructing a surrogation function $\widehat{\phi}(\cdot;\theta^{\prime})$ at a reference point $\theta^{\prime}\in\Theta$ that may satisfy the following conditions:
  1. (P1)
    Touching condition: $\widehat{\phi}(\theta^{\prime};\theta^{\prime})=\phi(\theta^{\prime})$.
  2. (P2)
    Majorizing condition: $\widehat{\phi}({\theta};{\theta^{\prime}})\geq\phi(\theta)$ for any $\theta\in\Theta$.
  3. (P3)
    Convexity: $\widehat{\phi}({\cdot};{\theta^{\prime}})$ is convex on $\Theta$.
  4. (P4)
    Directional derivative consistency: $\widehat{\phi}(\cdot;\theta^{\prime})$ is directionally differentiable on $\Theta$ and satisfies $\widehat{\phi}(\cdot;\theta^{\prime})^{\prime}(\theta^{\prime};d)=\phi^{\prime}(\theta^{\prime};d)$ for any $d\in\mathbb{R}^{q}$.
  5. (P5)
    Quadratic surrogation gap: for some $\kappa>0$, $\phi(\theta)+\displaystyle{\frac{\kappa}{2}}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\geq\widehat{\phi}(\theta;\theta^{\prime})$ for any $\theta\in\Theta$.
A function satisfying the first two conditions is referred to as an upper surrogation function according to Cui and Pang (2021, Definition 7.1.1). Though the enhanced $\varepsilon$-active surrogation functions to be constructed generally do not satisfy (P1), we still refer to ones solely satisfying (P2) as upper surrogation functions and similarly for lower surrogation functions. Condition (P3) ensures that surrogation-based subproblems in the algorithm can be solved by convex-programming solvers. Condition (P4) connecting the directional derivatives of surrogation and original functions is used to establish stationarity conditions of the limit points produced by surrogation-based algorithms to the original problem. Condition (P5) originating from Liu et al. (2022) is a technical condition for the nonasymptotic convergence analysis of the enhanced SMM algorithm. Note that Condition (P5) implies (P4) provided that $\widehat{\phi}(\cdot;\theta^{\prime})$ is directionally differentiable and satisfies (P2).
We first present upper and lower surrogation functions of the inner PA function $f(\cdot;\xi)$, which provide the basis of surrogation construction of $F(\cdot;\xi)$ and also play an essential role in the randomized index-mapping selection in the enhanced SMM algorithm. For $d=1$ and $\varepsilon\geq 0$, let $\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ denote the set of $\varepsilon$-active index mappings for functions $\{g(\cdot;\xi)\}_{\xi\in S}$ at $\theta^{\prime}\in\Theta$, i.e.,
$\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})\coloneqq\Big{\{}I_{1}:S\rightarrow[K]\mid\,g_{I_{1}(\xi)}(\theta^{\prime};\xi)\geq g(\theta^{\prime};\xi)-\varepsilon,\,\forall\,\xi\in S\Big{\}},$
and the set $\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ is similarly defined for $\{h(\cdot;\xi)\}_{\xi\in S}$. Each mapping $I_{1}\in\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ assigns an $\varepsilon$-active index for each $\xi\in S$, and thus $\{I_{1}(\xi)\}_{\xi\in S}$ constitutes an $\varepsilon$-active index combination of $\{g(\cdot;\xi)\}_{\xi\in S}$ at $\theta^{\prime}$. It is clear to see that $\widehat{f}(\theta;\xi,I_{2})\coloneqq g(\theta;\xi)-h_{I_{2}(\xi)}(\theta;\xi)$ with $I_{2}\in\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ is an upper surrogation function of $f(\cdot;\xi)$ satisfying Conditions (P2) and (P3) with $\varepsilon\geq 0$, and additionally (P1) with $\varepsilon=0$. The concave lower surrogation function $\widecheck{f}(\theta;\xi,I_{1})$ with $I_{1}\in\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ can be similarly defined. When $d>1$, the index mapping sets and surrogates are constructed componentwisely.
Now, we can construct the upper surrogation function of $F(\cdot;\xi)$ at a reference point $\theta^{\prime}$ based on the surrogation of inner PA functions with the index mapping $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})\coloneqq\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})\times\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$, which we denote by $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$. To facilitate the convergence analysis of the enhanced SMM algorithm, we impose the following assumptions on the family of surrogates $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$:
  1. [label=(B0), series=B]
  2. 1.
    The family of surrogates $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ for any $\xi\in S$ satisfy the following properties:
    1. (a)
      $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies Conditions (P2) and (P3) for every $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ with $\varepsilon\geq 0$;
    2. (b)
      $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies Condition (P1) for every $I\in\mathcal{I}^{0}(\theta^{\prime})$;
    3. (c)
      $\widehat{F}(\cdot;\cdot,\xi,I)$ is continuous on $\Theta\times\Theta$ given a fixed index mapping $I$;
    4. (d)
      the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,I)$ satisfies Condition (P4) for any $\varepsilon\geq 0$.
  3. 2.
    The minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)$ satisfies Condition (P5) for any $\varepsilon\geq 0$ and $\xi\in S$.
To illustrate the specific construction of surrogation functions satisfying the above assumptions, we consider the following three cases of the outer function $\varphi$:
  1. Case 1.
    $\varphi(\cdot;\xi)$ is piecewise affine for any $\xi\in S$.
  2. Case 2.
    $\varphi(\cdot;\xi)$ can be decomposed into monotonic convex functions for any $\xi\in S$.
  3. Case 3.
    $\varphi(\cdot;\xi)$ is smooth with Lipschitz gradient for any $\xi\in S$.
In Appendix 9, we provide the surrogation construction of $F(\cdot;\xi)$ for these cases with verification of Assumptions 1 and 2. The three cases cover a broad class of cost functions. In particular, Case 2 is a multivariate extension of the one considered in Cui et al. (2018), which shows that any univariate convex function can be written as the sum of monotonic convex functions. The surrogation construction for Case 3 is new and suitable for nonconvex smooth cost functions.
3.2 Enhanced SMM Algorithm
We present the enhanced SMM algorithm in Algorithm 1.
Input: initial point $\theta^{0}$, iteration length $T$, parameters $\{\eta,\varepsilon\}$, and increasing sequence $\{N_{\nu}\}_{\nu}$;
1 for $\nu=0$ to $T-1$ do
2       for $s=1,\dots,N_{\nu}$, draw $\xi_{\nu}^{s}$ independently and uniformly from $S$;
3       select $I_{\nu}^{\varepsilon}$ uniformly from $\mathcal{I}^{\varepsilon}(\theta^{\nu})$;
4       compute $\theta^{\nu+\frac{1}{2}}\in\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}_{N_{\nu}}(\theta;\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\left\lVert\theta-\theta^{\nu}\right\rVert^{2}\right\}$;
5       update the iterate $\theta^{\nu+1}=\begin{cases}\theta^{\nu+\frac{1}{2}},&\text{if}\ \widehat{F}_{N_{\nu}}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{N_{\nu}}(\theta^{\nu}),\\ \theta^{\nu},&\text{otherwise.}\end{cases}$
6 end for
Output: $\widetilde{\theta}^{\,T}$ uniformly sampled from the iterates $\{\theta^{\nu}\}_{\nu=0}^{T-1}$.
Algorithm 1 Enhanced SMM Algorithm
At iteration $\nu$, with samples $\{\xi_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ and an $\varepsilon$-active index mapping $I^{\varepsilon}_{\nu}\in\mathcal{I}^{\varepsilon}(\theta^{\nu})$ randomly selected in Steps 2 and 3, we compute in Step 4 the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ with respect to the upper surrogation function $\widehat{F}_{N_{\nu}}(\cdot;\theta^{\nu},I_{\nu}^{\varepsilon})\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}\widehat{F}(\cdot;\theta^{\nu},\xi_{\nu}^{s},I^{\varepsilon}_{\nu})$ of the sample average approximation function $F_{N_{\nu}}(\cdot)\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}F(\cdot;\xi_{\nu}^{s})$. Since the surrogate function value $\widehat{F}_{N_{\nu}}(\theta^{\nu};\theta^{\nu},I_{\nu}^{\varepsilon})$ could be strictly larger than $F_{N_{\nu}}(\theta^{\nu})$, the candidate point $\theta^{\nu+\frac{1}{2}}$ may not lead to a valid update. Hence, it is accepted in Step 5 only when sufficient descent is satisfied. In the special case of $\varepsilon=0$, the touching condition (P1) is satisfied. Thus the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ naturally satisfies the sufficient descent inequality in Step 5, and we directly update $\theta^{\nu+1}=\theta^{\nu+\frac{1}{2}}$.
It is worth noticing that the randomized index-mapping selection in Step 3 is vital for the ERM problem, without which the number of subproblems to be solved at each iteration, identical with the cardinality of $\mathcal{I}^{\varepsilon}(\theta^{\nu})$, would grow exponentially with the sample size. Furthermore, by combining the enhanced $\varepsilon$-active surrogation with incremental sample averaging, the enhanced SMM can achieve satisfactory descent progress in early iterations with relatively small sample sizes, thus improving the computational efficiency; yet such a combination raises the major challenges of convergence analysis to be addressed in the next section.
4 Convergence Analysis of Enhanced SMM
In this section, we first introduce preliminaries on stationarity concepts and then analyze the convergence properties of the enhanced SMM algorithm for solving problem (9). Since much of the convergence analysis is quite technical, we present the main theoretical results with essential discussions while leaving analytical details in Appendix 10.
4.1 Preliminaries on Stationarity
Suppose that the outer function $\varphi$ is Bouligand-differentiable (B-differentiable), i.e., it is locally Lipschitz continuous and directionally differentiable on $\mathbb{R}^{d}$. Then the composite function $F$ is also B-differentiable according to Cui and Pang (2021, Proposition 4.1.2). We say that $\theta^{d}\in\Theta$ is a d-stationary point of problem (9) if $F^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for any $\theta\in\Theta$. For a general nonconvex problem on a convex set, d-stationarity is identified as the sharpest among the first-order stationarity concepts, which gives a necessary condition for local optimality. Specialized to the composite DC program (9), it is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2 (Cui et al. 2018, Proposition 2), and more generally we further define the following strengthened d-stationarity.
Definition 4.1
Given $\varepsilon\geq 0$ and $\rho>0$, we say that $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (9) if ${\{\theta^{d}\}=\mathop{\arg\min}_{\theta\in\Theta}\left\{\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}\widehat{F}(\theta;\theta^{d},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}}$, where $\widehat{F}(\theta;\theta^{d},I)\coloneqq\mathbb{E}\,[\widehat{F}(\theta;\theta^{d},\widetilde{\xi}_{n},I)\,]$, and $\{\widehat{F}(\cdot;\theta^{d},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}$ satisfy Assumption 1 for every $\xi\in S$.
This concept is generalized from the $(\alpha,\eta)$-d-stationarity defined in Lu et al. (2019) for structured DC programs. Due to the requirement of the unique proximal mapping point in the defining condition, our concept is slightly stronger for a reduced composite DC program with the identity outer function $\varphi$. Moreover, it is weaker than the composite $\varepsilon$-strong d-stationarity introduced in Qi et al. (2021) (see also Cui and Pang (2021, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition that is necessary for the follow-up convergence analysis.
Let $\mathcal{S}^{d}$, $\mathcal{S}^{d}_{\varepsilon,\rho}$, and $\mathcal{S}^{*}$ denote the sets of d-stationary points, composite $(\varepsilon,\rho)$-strongly d-stationary points, and global minimizers of (9), respectively. We show below that the composite strong d-stationarity is stronger than d-stationarity and necessary for global optimality.
Proposition 4.2
Under Assumption 1 for the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ for each $\xi\in S$, it holds for problem (9) that $\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$ for any $\varepsilon\geq 0$ and $\rho>0$. In particular, $\mathcal{S}^{d}_{0,\rho}=\mathcal{S}^{d}$ for any $\rho>0$.
The proof is provided in Appendix 10.1. Furthermore, we can show that $\mathcal{S}^{*}=\mathcal{S}^{d}_{\varepsilon,\rho}$ for sufficiently large $\varepsilon$ and suitable $\rho$ under Assumption 2 and a regularity condition that d-stationary values of problem (9) are isolated from the optimal value, which naturally hold when $\varphi$ is in Case 1 or Case 2; this is formally stated as Proposition 10.2 in Appendix 10.1.
4.2 Convergence Analysis of Enhanced SMM Algorithm
Now we establish the convergence analysis of Algorithm 1 in terms of (composite strong) d-stationarity for problem (9). Throughout this section, we make several blanket assumptions that the outer function $\varphi$ is B-differentiable, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ satisfies Assumption 1 for each $\xi\in S$.
Since the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{d})\coloneqq\underset{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}{\min}\widehat{F}(\theta;\theta^{d},I)$ is only of theoretical interest in defining composite strong d-stationarity, we need an auxiliary residual as a bridge to connect with the sampling-based surrogation function $\widehat{F}_{N_{\nu}}(\cdot;\theta^{\nu},I_{\nu}^{\varepsilon})$ utilized in Algorithm 1 for convergence analysis. With $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ and the corresponding surrogate $\widehat{F}(\cdot;\theta^{\prime},I)$ at $\theta^{\prime}\in\Theta$, we denote the proximal map $\mathrm{prox}_{\widehat{F}/\rho}(\theta^{\prime},I)\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}(\theta;\theta^{\prime},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}$ and further define
$P^{\rho}_{I}(\theta^{\prime})\coloneqq\begin{cases}\mathrm{prox}_{\widehat{F}/\rho}(\theta^{\prime},I)&\text{if}\ \min_{\theta\in\Theta}\,\left\{\widehat{F}(\theta;\theta^{\prime},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}\leq F(\theta^{\prime}),\\ \theta^{\prime}&\text{otherwise}.\end{cases}$ (10)
The map $P^{\rho}_{I}(\theta^{\prime})$ can be regarded as an iterative update of Algorithm 1 accounting for all $n$ scenarios, with which we define the averaged residual
$\bar{r}^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\frac{1}{\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert}\sum_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\left\lVert\theta^{\prime}-P^{\rho}_{I}(\theta^{\prime})\right\rVert.$
The following proposition gives convergence rates of Algorithm 1 in terms of $\bar{r}^{\varepsilon,\rho}$ in expectation with the proof provided in Appendix 10.5.
Proposition 4.3
Let $\widetilde{\theta}^{\,T}$ be the output of Algorithm 1 after $T$ iterations with $\varepsilon\geq 0$, $\eta\geq 0$, and $\{N_{\nu}\}_{\nu}$. Then there exists $C_{1}>0$ such that for any $\rho>\eta$,
$\big{(}\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]\,\big{)}^{2}\leq\frac{F(\theta^{0})-\min_{\theta\in\Theta}F(\theta)+C_{1}\cdot\sum_{\nu=0}^{T-1}\sqrt{\log N_{\nu}/N_{\nu}}}{\frac{\rho-\eta}{2}\cdot T}.$
In particular, with $\varepsilon=0$ and $\eta>0$, there exists $C_{2}>0$ such that for any $\rho>\eta$,
$\big{(}\mathbb{E}\,[\bar{r}^{0,\rho}(\widetilde{\theta}^{\,T})]\,\big{)}^{2}\leq\frac{F(\theta^{0})-\min_{\theta\in\Theta}F(\theta)+C_{2}\cdot\sum_{\nu=0}^{T-1}(1/N_{\nu})}{\frac{\rho-\eta}{2}\cdot T}.$
By relating $\bar{r}^{\varepsilon,\rho}$ to another residual that measures the proximal mapping gap with respect to $\widehat{F}_{\min}^{\varepsilon}(\cdot;\theta^{\prime})$, we can establish the convergence of Algorithm 1 to composite strong d-stationarity. The error-bound analysis for the nonasymptotic convergence to d-stationarity needs the following assumption generalized from Drusvyatskiy and Lewis (2018). We denote the closure operator of a set by $\mathrm{cl}(\cdot)$ and $\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)\coloneqq\mathop{\arg\min}_{\theta^{\prime}\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta^{\prime};\theta)+\frac{\rho}{2}\left\lVert\theta^{\prime}-\theta\right\rVert^{2}\right\}$.
  1. [label=(B0), resume*=B]
  2. 1.
    Generalized quadratic growth: given $\varepsilon>0$ and $\rho>0$, there exist positive constants $\delta_{q}$, $\lambda_{q}$, and $c_{q}$ such that for any $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$ and $\max_{\theta^{\prime}\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)}\left\lVert\theta^{\prime}-\theta\right\rVert\leq\lambda_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(t;\theta^{d},\theta)\coloneqq F(\theta^{d}+t(\theta-\theta^{d}))$ is convex on $[0,1]$ and
    $\bar{F}(t;\theta^{d},\theta)\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{q}}{2}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2},\quad\forall\,t\in[0,1].$
The condition holds, for example, when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope, which is verified in Appendix 10.4. It can be expected to hold in more general cases since it is a local property only requiring the convexity and quadratic growth of $F$ near $\mathcal{S}^{d}$.
Theorem 4.4 (Convergence of Enhanced SMM)
For problem (9), let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon>0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$. Then the following results hold:
  1. (a)
    Any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of problem (9) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$.
  2. (b)
    Under Assumptions 2 and 1, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$. In particular, if $\varphi$ is piecewise affine and $\Theta$ is a polytope, the result holds without the two assumptions.
Theorem 4.4 shows the asymptotic convergence of the enhanced SMM algorithm to composite strong d-stationarity and nonasymptotic convergence rate of the distance to the d-stationary set. We speculate that it is possible, albeit challenging, to achieve nonasymptotic results in terms of composite strong d-stationarity with carefully designed residuals and error-bound analysis, which we will leave for future study. When $\varphi$ is piecewise affine and $\Theta$ is a polytope, Assumption 2 holds naturally; furthermore, the nonasymptotic result holds without the need for Assumption 1 since a piecewise affine objective on a polytope has maximum descent rates bounded away from 0 at all points that are not d-stationary. Leveraging this special property, we can further establish the nonasymptotic convergence to d-stationarity (also local optimality in this case) of Algorithm 1 with $\varepsilon=0$, which reduces to the standard SMM with randomized index-mapping selection and achieves a faster rate with $\eta>0$. Details of the error-bound analysis and convergence proofs can be found in Appendix 10.3.
Theorem 4.5 (Convergence of SMM)
For problem (9) with $\varphi$ being piecewise affine and $\Theta$ being a polytope, under the same settings of Algorithm 1 specified in Theorem 4.4 except that $\varepsilon=0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$. Moreover, with $\eta>0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\mathcal{O}(T^{-\frac{1}{2}\min\left\{\alpha,1\right\}})$, which is up to a logarithmic factor when $\alpha=1$.
Finally, we make several remarks on the convergence results.
Remark 4.6
The enhanced $\varepsilon$-active surrogation with positive $\varepsilon$ is essential for general cases. As noted in Liu et al. (2022), due to the composite structure of objective functions, the SMM algorithm with $\varepsilon=0$ may yield convergence to compound criticality, which is weaker than d-stationarity. Pang et al. (2016, Example 4) also illustrates that an MM algorithm with $\varepsilon=0$ minimizing a piecewise quadratic objective could converge to a point isolated from the d-stationary set.
Remark 4.7
According to Theorems 4.4 and 4.5, the optimal iteration complexity of Algorithm 1 for $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]\leq\epsilon$ is $\mathcal{O}(\epsilon^{-2})$, which is the same as that of the stochastic DC algorithm in Le Thi et al. (2022) with incremental sampling for differentiable stochastic DC programs, provided that convex subproblems are solved to the global optimum. Notably, the optimal complexity of Algorithm 1 is achieved with the sampling rate $N_{\nu}=\mathcal{O}(\nu^{2})$ that is faster than $\mathcal{O}(\nu)$ in Le Thi et al. (2022) due to the nondifferentiable composite structure, and the convergence results of our algorithm hold for a much wider range of nonconvex and nondifferentiable composite problems. Moreover, with minor modifications of the convergence analysis, we may obtain the same convergence results for an inexact version of Algorithm 1, in which a $\delta_{\nu}$-suboptimal solution $\theta^{\nu+\frac{1}{2}}$ is computed in Step 4 and only accepted in Step 5 when it satisfies $\widehat{F}_{N_{\nu}}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{N_{\nu}}(\theta^{\nu})+\delta_{\nu}$ with $\delta_{\nu}=\mathcal{O}(\nu^{-1})$. In such a plausible extension, it is viable to design a two-loop algorithmic scheme addressing the computational complexity in solving convex subproblems. Together with our iteration complexity result, it may enable characterization of the actual work complexity of the entire algorithm similar to Pasupathy and Song (2021) and complexity comparisons with other first-order algorithms for nonconvex SP.
Remark 4.8
Since the output $\widetilde{\theta}^{\,t}$ of Algorithm 1 is uniformly sampled from the iterate sequence $\{\theta^{\nu}\}_{\nu=0}^{t-1}$, the almost sure convergence of $\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,t})$ to 0 in Proposition 4.3 implies that there exists a subsequence of $\{\theta^{\nu}\}_{\nu}$ with $\bar{r}^{\varepsilon,\rho}(\theta^{\nu})$ converging to 0 almost surely. With similar convergence analysis, we can obtain the almost sure existence of the limit point of $\{\theta^{\nu}\}_{\nu}$, which satisfies the same asymptotic convergence properties as the limit points of $\{\widetilde{\theta}^{\,t}\}_{t}$.
Remark 4.9
By Theorem 4.4 and the equivalence between $\mathcal{S}^{d}_{\varepsilon,\rho}$ and $\mathcal{S}^{*}$ stated in Proposition 10.2, with sufficiently large $\varepsilon$ and the proximal parameter $\eta$ close or equal to 0, any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a global minimizer of (9) with probability 1. This implies that the PADR-based ERM model together with the enhanced SMM algorithm under such settings is asymptotically optimal to the original problem (2). However, evidenced by the numerical implementations, the candidate solutions produced in Step 4 of Algorithm 1 with large $\varepsilon$ may be rejected in Step 5 very often, which results in inefficient computation and slow convergence. Such a computational issue is resolved by a shrinking-$\varepsilon$ strategy in Section 6.
5 Extension: Constrained SP with Covariate Information
[…]
6 Numerical Experiments
In this section, we evaluate the performance of the PADR-based ERM method through synthetic experiments on the Newsvendor problem and its variations with nonconvex costs or nonconvex constraints. We compare our method with PO methods and other two prominent DR-based methods. PO methods apply linear and PA regression models (referred to as PO-L and PO-PA respectively) and the squared error loss in the prediction stage; DR-based methods include NV-ERM1 method from Ban and Rudin (2018) that employs LDR and the optimizer prediction method from Bertsimas and Koduri (2021) that employs RKHS-DR. We omit comparisons with other data-driven approaches in the integrated scheme such as the objective prediction method from Bertsimas and Koduri (2021) and prescriptive methods from Bertsimas and Kallus (2019) because numerical results in Bertsimas and Koduri (2021) indicate that both two are outperformed by the optimizer prediction method in various cases.
The Newsvendor problem with covariate information $X=x$ is formulated as follows,
$\min_{z\geq 0}\quad\mathbb{E}\left[c_{b}(Y-z)_{+}+c_{h}(z-Y)_{+}\mid X=x\right],$ (11)
where $c_{b}$ and $c_{h}$ are the unit back-order and holding costs respectively, and $(\cdot)_{+}$ denotes $\max\{\cdot,0\}$. We set the true demand as $Y=\overline{Y}(X)+\epsilon$, where $\overline{Y}(\cdot)$ is the mean model, and $\epsilon$ is a random variable with the standard normal distribution independent of $X$. Throughout all experiments, we generate the feature data uniformly from $\mathcal{X}=[-1,1]^{p}$, and we set the mean model satisfying $\min_{x\in\mathcal{X}}\overline{Y}(x)\geq 5$ to keep all corresponding demand data positive. Unless otherwise noted, both training and test data sets contain 1000 samples.
We clarify the settings throughout this section. The PADR model with two max-affine components consisting of $K_{1}$ and $K_{2}$ pieces respectively is referred to as PADR$(K_{1},K_{2})$, and we implement multiple settings of $(K_{1},K_{2})$ from $\{0,1,2,3\}^{2}$. We set the uniform scale $\mu=50$ for parameters of PADR, PO, and LDR methods. In the implementation of the enhanced SMM algorithm, initial points are generated uniformly from $[-\mu,\mu]^{d(K_{1}+K_{2})(p+1)}$, and the sequence $\{N_{\nu}\}_{\nu}$ satisfies $N_{\nu}=\beta_{1}\nu+\beta_{2}$ with $\beta_{1},\beta_{2}>0$. Besides Algorithm 1 with constant $\varepsilon$, we also implement a version with a shrinking-$\varepsilon$ strategy, in which $\varepsilon$ is set as $\varepsilon_{0}$ at the first $T_{0}$ iterations and $\varepsilon_{1}(<\varepsilon_{0})$ at the remaining ones. We set the iterate point that obtains the minimum ERM cost in $T=10$ iterations as the output of the algorithm. For each instance, the algorithm is executed for $10$ rounds with random initialization, and we report the best performance unless otherwise noted. The hyperparameters are selected via a validation process with searching ranges listed in Table 6.
[…]
Hyperparameter Selection for PADR-based ERM and Enhanced SMM. Hyperparameters Description Searching Ranges $\gamma$ the parameter for the constraints in (5) $[0,1]$ $\lambda$ the penalty parameter for exact penalization $[0,10^{3}]$ $\varepsilon_{0},\varepsilon_{1},T_{0}$ parameters for the shrinking-$\varepsilon$ strategy $[0,10^{4}],[0,10^{4}],\{1,2,\dots,6\}$ $\beta_{1},\beta_{2}$ parameters for the sampling strategy $[5,50],[10,40]$ $\eta$ the proximal parameter $[0,1]$
We use the Gaussian kernel in RKHS-DR, and the diagonal of the kernel is added by values no larger than $10^{-3}$ so as to resolve numerical issues without undermining its performance. The kernel bandwidth, regularization parameter, and penalty parameter are selected via a validation procedure searching from $[10^{-1},10^{3}]$, $[10^{-8},1]$, and $[0,10^{3}]$ respectively. We approximate the optimal cost by the empirical cost of the theoretical (in Section 6.1) or simulated (in Sections 6.2 and 6.3) optimal decision, which is treated as the baseline and referred to as SIMOPT. For the training time comparisons, we report the total time of the enhanced SMM algorithm for 10 rounds with predetermined hyperparameters and the average time of RKHS-DR over all hyperparameter settings since the running time of RKHS-DR is highly related to the generated kernel and thus quite unstable. The training time of PO-L and LDR for each instance is less than 0.5 seconds and thus is omitted.
All methods are implemented in Python 3.9.7 on a MacBook Pro with a macOS Monterey system and an Apple M1 Pro chip. All programs involved are built and solved by the optimizer Gurobi, version 9.5.0.
6.1 Unconstrained Newsvendor Problem
For the Newsvendor problem (11), we set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $k=1$ unless otherwise specified:
$\overline{Y}(X)=k\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (12)
a. Basic Results. In Figure 1a, we report comparison results with the training sample size $n$ increasing from 50 to 1000. For a fair comparison, we set the PA function class with $(K_{1},K_{2})=(3,0)$ as the regression model in PO-PA, which contains the true model $\overline{Y}(\cdot)$; the parameters of the PA regression model are estimated by solving the corresponding ERM problem via the enhanced SMM algorithm with the same hyperparameter selection procedure as PADR methods. We also implement the quadratic DR and cubic DR as the generalized LDRs, denoted by GLDR-2 and GLDR-3 respectively, the running time of which is less than 0.5 seconds and thus is omitted.
(a) Basic Results
(b) Enhanced SMM with Different $\varepsilon$-Strategies
Unconstrained Newsvendor Problem: Basic Results We also test the performance of PO-L and LDR in this instance, whose costs (around 14 and 9, respectively) are excluded in (a) for readability. In (b), solid and dotted lines represent constant-$\varepsilon$ strategies; dashed lines represent shrinking-$\varepsilon$ strategies with $T_{0}=3$ and $\varepsilon_{1}=0$.
The costs of PADR(3,0) and PADR(3,3) converge to SIMOPT because the optimal DR in this case is a max-affine function with three pieces, and thus PADR(3,0) and PADR(3,3) have no approximation error in principle. The RKHS-DR method converges slower than PADR(3,0) and PADR(3,3); moreover, its performance with $n=1000$ is slightly worse than that of PADR(3,0) with $n=400$. For PO-PA, given the observation that the prediction model provided by the enhanced SMM algorithm is quite close to the true model $\overline{Y}(\cdot)$, we conjecture that its gap to SIMOPT is mainly due to inconsistency between the prediction loss and Newsvendor cost with imbalanced $c_{b}$ and $c_{h}$. This phenomenon is more evident from the comparison between LDR and PO-L, the two methods with misspecified models. The training time of RKHS-DR grows roughly exponentially with respect to the sample size, whereas that of PA-related methods is quite stable and significantly shorter when $n=1000$, which can be mainly credited to the sampling strategy used in the enhanced SMM algorithm. The numerical results with the parameter ratio $c_{b}:c_{h}$ varying from 1:9 to 9:1 exhibit similar patterns and thus are omitted.
Figure 1b displays sorted training cost curves of PADR(3,3) provided by the enhanced SMM with constant-$\varepsilon$ and shrinking-$\varepsilon$ strategies over 30 rounds. The shrinking-$\varepsilon$ strategy significantly improves the performance of the algorithm compared to the constant-$\varepsilon$ strategy, which typically does not reach satisfactory convergence within 10 iterations. This indicates that despite the theoretical convergence of the algorithm to global optimality with sufficiently large $\varepsilon$, it needs to engage with a huge number of index mappings when the sample size is large and thus can hardly make a valid update in each iteration. In contrast, the shrinking-$\varepsilon$ strategy with $\varepsilon_{0}=3000$ approximately attains the optimal value in 18 rounds. The significant efficiency improvement is because the shrinking-$\varepsilon$ strategy enables flexible adjustments of iterate points at early iterations as a warm start and afterward accelerates the convergence by ensuring valid updates with $\varepsilon=0$.
b. Varying Feature Dimensions. We study the effect of the feature dimension $p$ on the performance of PADR methods under sparse and dense demand models $\overline{Y}$ following the settings in Bertsimas and Koduri (2021) and report the results in Figure 1d.
(c) Sparse Demand Model
(d) Dense Demand Model
Unconstrained Newsvendor Problem with Varying Feature Dimensions We also test the performance of PO-L and LDR under dense demand models, whose costs (decreasing from 15 to 5 and 9 to 4, respectively) are excluded in (b) for readability.
In the sparse case, only the first two components of $X$ contribute to $\overline{Y}$ following the demand model (12); in the dense case, we replace $X_{1}$ in (12) with the average of the first half of the features $\frac{2}{p}\sum_{i=1}^{p/2}X_{i}$ and $X_{2}$ with the average of the other half. Figure 1d implies that PADR and PO-PA are generally more robust to redundant features than RKHS-DR in the sparse case, and that PADR performs better than RKHS-DR and PO-PA in the dense case. The performance of PADR(3,3) and PADR(3,0) deteriorates as $p$ increases due to the growing estimation error and optimization difficulty. The simpler PADR(2,0) performs more stable in both cases, despite the existence of an approximation error.
c. Nonlinearity in Demand Model. We study the effect of nonlinearity in the demand model (12) with varying $k$ and report results in Figure 1f. It indicates that the performance of PADR(3,0), PADR(3,3), and PO-PA is irrelevant to the nonlinearity level of the demand model, whereas that of PADR(2,0) and RKHS-DR deteriorates as the nonlinearity level increases.
(e) $p=2\phantom{xxxxxxx}$
(f) $p=20$
Unconstrained Newsvendor Problem with Varying Nonlinearity in Demand Model In (b), the demand model is of the dense type. We also test the performance of PO-L and LDR, whose costs are excluded for readability. When $p=2$, PO-L and LDR increase from 5 to 40 and 3 to 30, respectively; when $p=20$, PO-L and LDR increase from 4 to 16 and 3 to 13, respectively.
Besides the piecewise affine demand model (12), we also test the performance of the DR-based methods with another nonlinear demand model that includes a sine function to account for seasonality. The numerical results are provided in Appendix 12.
6.2 Constrained Newsvendor Problem
We next consider a two-product Newsvendor problem with a total inventory capacity $C_{0}$. Assume that the uncensored historical demand data is available. We denote the unit back-order cost, unit holding cost, and order decision by $c_{bi}$, $c_{hi}$, and $z_{i}$ respectively for product $i=1,2$. The capacity constraint is formulated as $z_{1}+z_{2}\leq C_{0}$. In this subsection, we set $(c_{b1},c_{h1},c_{b2},c_{h2})=(8,2,2,8)$, $C_{0}=60$, and the mean demand models for the two products as
$\overline{Y}_{1}(X)=15X_{1}-5X_{2}+30,\quad\overline{Y}_{2}(X)=15X_{1}+5X_{2}+30,$
under which the probability that the constraint is active at the optimal decision is around 0.5. We compare PADR(2,2) with PO-L and the penalized RKHS-DR method from Bertsimas and Koduri (2021, Section 6.2).
Since DR-based methods may yield infeasible decisions, in Figure 1i we report the feasible frequency of decisions provided by PADR(2,2) and RKHS-DR. By projecting the decisions to the convex feasible region when needed, we report the test cost of the feasible decisions in Figure 1i.
(g) Basic Results
(h) Sparse Demand Model
(i) Dense Demand Model
Newsvendor Problem with a Capacity Constraint
With the basic demand model, PADR(2,2) converges to SIMOPT and attains 100% feasibility when $n\geq 200$. In comparison, RKHS-DR and PO-L display notable gaps to SIMOPT, and the feasible frequency of RKHS-DR is only around 80%. Figures 1h and 1i display numerical results with varying feature dimensions. In the sparse case, PO-L is markedly better than the other two methods when $p=50$, indicating that regularization for sparsity is needed for learning PADR and RKHS-DR. In the dense case, PADR(2,2) performs the best except when $p=50$.
6.3 Nonconvex Newsvendor Problem
In this section, we consider two variations of the Newsvendor problem by adding a concave cost $C(z)$ in the objective or as a constraint. The cost $C(z)$ added in the objective can be interpreted as the capacity acquisition cost (Atamtürk and Hochbaum 2001); it also can be restricted in a constraint as a resource limitation. In the following experiments, we set the concave cost as
$C(z)=\min\{z,0.6z+0.8,0.4z+15.6\},$
and then both two variant problems of interest are nonconvex. We compare the PADR methods with RKHS-DR and use the deterministic MM algorithm to solve the nonconvex ERM problem for learning RKHS-DR, referred to as RKHS-DR(MM). The baseline SIMOPT is computed by solving simulated nonconvex problems via a solver of mix-integer linear programming.
a. Nonconvex Objective. We first consider a generalized single-product Newsvendor with $(c_{b},c_{h})=(5,5)$ and $C(z)$ added in the objective of (11). The numerical results in Figure 1l indicate that PADR(3,0) performs the best in almost all the settings, which is consistent with the convex case in Section 6.1; moreover, the gap between PADR(3,0) and RKHS-DR(MM) in the sparse case with $p=50$ is much more significant.
(j) Basic Results
(k) Sparse Demand Model
(l) Dense Demand Model
Newsvendor Problem with a Nonconvex Objective
b. Nonconvex Constraint. We next consider the 2-product Newsvendor problem given in Section 6.2 with parameters $(c_{b1},c_{h1},c_{b2},c_{h2})=(7,7,3,3)$ and the replacement of the convex constraint by $C(z_{1})+C(z_{2})\leq C_{0}$. We set $C_{0}=50$ to keep the same active probability for the constraint. Since the projection of infeasible decisions to the nonconvex feasible region may not yield meaningful solutions, we only report the average test cost of feasible decisions in Figure 1o, with a careful hyperparameter selection to ensure high feasible frequencies.
(m) Basic Results
(n) Sparse Demand Model
(o) Dense Demand Model
Newsvendor Problem with a Nonconvex Constraint In (b) and (c) with $p=50$, RKHS-DR(MM) takes more than 500 seconds on average, while PADR(2,2) takes less than 200 seconds.
For the problem with the basic demand model (12), PADR(2,2) performs comparably with RKHS-DR(MM), with a notable gap to SIMOPT. Figures 1n and 1o indicate the superiority of PADR(2,2) over the other two methods under both sparse and dense demand models with varying feature dimensions. It’s also worth noting that the simpler PADR(1,0) (equivalent to LDR), with poor performance in general, performs comparably well with the others in the high-dimensional case. It is more difficult to obtain feasible solutions for cases with sparse models than those with dense models; moreover, with sparse models, PADR(1,0) is more likely to produce feasible solutions than the others.
Finally, we also conduct experiments on constrained Newsvendor problems in Sections 6.2 and 6.3 with multiple $C_{0}$ values and Newsvendor cost parameters. The numerical results exhibit similar patterns and thus are omitted.
7 Conclusion
We propose the PADR-based ERM method together with the enhanced SMM algorithm to solve a broad class of SP problems with covariate information. For our PADR-based ERM framework, we provide the approximation bound of the PA function class and further establish the nonasymptotic consistency result for unconstrained problems and the asymptotic consistency result for deterministic-constrained ones. To solve the ERM problem with a composite DC structure, we develop the enhanced SMM algorithm by combining enhanced $\varepsilon$-active surrogation with sequential sampling. We provide the asymptotic convergence of the enhanced SMM algorithm to composite strong d-stationarity, and the nonasymptotic convergence to d-stationarity via the error-bound analysis. Numerical experiments indicate the superiority of our method over PO methods and prominent DR-based methods in the Newsvendor problem regarding the test cost, computational time, and robustness to feature dimensions and nonlinearity of the underlying dependency. In conclusion, the PADR is a favorable hypothesis class for DR-based ERM methods, supported by theoretical consistency, algorithmic convergence, and computational efficiency for solving SP with covariate information. Extensions of the PADR-based ERM framework for more complicated SP problems are deserved to be explored in the future, such as problems with chance constraints or endogenous random variables, which necessitate a further investigation of model construction, consistency analysis, and data-driven algorithmic development for problems with more complex composite structures.
Supplementary Material for
Data-driven Piecewise Affine Decision Rules for Stochastic Programming with Covariate Information
{APPENDICES}
8 Proofs of Theoretical Consistency Results in Section 2
8.1 Proof of Proposition 2.1
The idea of the proof is to construct an interpolation PA function to a set of points related to the target function $f_{0}$ and a cover of the set $\mathcal{X}$. The structure of the interpolation function is motivated by Siahkamari et al. (2020) but with a different parameter, which ensures the Lipschitz continuity of the function with modulus independent of the approximation accuracy.
Proof 8.1
[…]
$\left(\frac{\sqrt{p}\bar{X}}{\epsilon}\right)^{p}\leq\left\lvert\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\right\rvert\leq\left(\frac{\sqrt{p}\bar{X}}{\epsilon}+1\right)^{p}\leq\left(\frac{2\sqrt{p}\bar{X}}{\epsilon}\right)^{p}.$
Given $f_{0}\in\mathcal{F}_{L_{0},M_{0}}$, by McShane (1934, Theorem 1), we can extend $f_{0}$ to an $L_{0}$-Lipschitz function $\bar{f}_{0}:\bar{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying $\left\lVert\bar{f}_{0}\right\rVert_{\infty}\leq M_{0}$ and $\bar{f}_{0}(x)=f_{0}(x)$ for any $x\in\mathcal{X}$. Next, we construct a piecewise affine interpolation function to a set of points $\{(\widehat{x},\bar{f}_{0}(\widehat{x}))\mid\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\}$:
$f_{\widehat{\mathcal{C}}}(x)\coloneqq$ $\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}-\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\},$ (13)
where $C\coloneqq L_{0}/d_{\min}=L_{0}\sqrt{p}/(2\epsilon)$. The function $f_{\widehat{\mathcal{C}}}$ indeed interpolates $\{(\widehat{x},\bar{f}_{0}(\widehat{x})):\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\}$. Specifically, for any $\widehat{y}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$,
$\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(\widehat{y}-\widehat{x})+\frac{1}{2}C\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$ $=\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}C\left\lVert\widehat{x}-\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
$\leq\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}L_{0}\left\lVert\widehat{x}-\widehat{y}\right\rVert+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
$\leq\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}\left\lvert\bar{f}_{0}(\widehat{x})-\bar{f}_{0}(\widehat{y})\right\rvert+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
$\leq\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{y}),$
where the first inequality holds since $C=L_{0}/d_{\min}\geq L_{0}/\left\lVert\widehat{x}-\widehat{y}\right\rVert$. Moreover, this upper bound is attained by taking $\widehat{x}=\widehat{y}$ in the max operator. Similarly, the value of the second max-affine component of $f_{\widehat{\mathcal{C}}}$ at $\widehat{y}$ is $\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{y})$. Hence, we obtain $f_{\widehat{\mathcal{C}}}(\widehat{y})=\bar{f}_{0}(\widehat{y})$, verifying the interpolation. With $K\coloneqq\left\lvert\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\right\rvert$, for any $\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$, we obtain $\left\lVert C\widehat{x}\right\rVert_{\infty}\leq\frac{L_{0}\sqrt{p}}{2\epsilon}\bar{X}\leq\frac{L_{0}}{2}K^{1/p}$ and
$\left\lvert-C\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}C\left\lVert\widehat{x}\right\rVert^{2}\pm\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\rvert\leq\frac{1}{2}Cp\left\lVert\widehat{x}\right\rVert_{\infty}^{2}+\frac{1}{2}M_{0}\leq\frac{1}{4}pL_{0}\bar{X}K^{1/p}+\frac{1}{2}M_{0}.$
Therefore, with $\mu\geq\max\left\{\frac{1}{2}L_{0}K^{1/p},\frac{1}{4}pL_{0}\bar{X}K^{1/p}+\frac{1}{2}M_{0}\right\}$, we have $f_{\widehat{\mathcal{C}}}\in{\mathcal{H}}_{\mu}^{K}$.
Note that for any $w\in\mathcal{X}$, there exists $\widehat{x}(w)\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$ such that $\left\lVert w-\widehat{x}(w)\right\rVert\leq\epsilon$ and $f_{\widehat{\mathcal{C}}}(\widehat{x}(w))=\bar{f}_{0}(\widehat{x}(w))$. We next show that $f_{\widehat{\mathcal{C}}}$ is Lipschitz continuous with modulus $(\sqrt{p}+2)L_{0}$, with which we obtain the desired inequality,
$\begin{array}[]{ll}\displaystyle{\min_{f\in{\mathcal{H}}_{\mu}^{K}}}\max_{w\in\mathcal{X}}\left\lvert f(w)-f_{0}(w)\right\rvert&\leq\displaystyle{\max_{w\in\mathcal{X}}}\left\lvert f_{\widehat{\mathcal{C}}}(w)-f_{0}(w)\right\rvert\leq\displaystyle{\max_{w\in\mathcal{X}}}\left\{\left\lvert f_{\widehat{\mathcal{C}}}(w)-f_{\widehat{\mathcal{C}}}(\widehat{x}(w))\right\rvert+\left\lvert\bar{f}_{0}(\widehat{x}(w))-\bar{f}_{0}(w)\right\rvert\right\}\\ &\leq 2(p^{1/2}+3)p^{1/2}L_{0}\bar{X}K^{-1/p},\end{array}$
where the last inequality follows from the Lipschitz continuity of $f_{\widehat{\mathcal{C}}}$ and $\bar{f}_{0}$ and the relation $K\leq(2\sqrt{p}\bar{X}/\epsilon)^{p}$. For any $x\in\bar{\mathcal{X}}$, let
$\begin{array}[]{ll}&\mathcal{A}_{1}(x)\coloneqq\displaystyle{\mathop{\arg\max}_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\},\\ &\mathcal{A}_{2}(x)\coloneqq\displaystyle{\mathop{\arg\max}_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}.\end{array}$
It suffices to show that $\left\lVert C(\widehat{x}_{1}^{*}-\widehat{x}_{2}^{*})\right\rVert\leq(\sqrt{p}+2)L_{0}$ for any $\widehat{x}_{1}^{*}\in\mathcal{A}_{1}(x)$ and $\widehat{x}_{2}^{*}\in\mathcal{A}_{2}(x)$. Let $\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$ satisfy $\left\lVert x-\widehat{x}\right\rVert\leq\epsilon$. For any $\widehat{y}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$ satisfying $\left\lVert x-\widehat{y}\right\rVert>\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$, we have that
$C\widehat{y}^{\top}(x-\widehat{y})+\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{y})$ $=\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}-C\left\lVert x-\widehat{y}\right\rVert^{2}+\bar{f}_{0}(\widehat{y})\right]$
$=\frac{1}{2}\left[\bar{f}_{0}(\widehat{y})-\bar{f}_{0}(x)-C\left\lVert x-\widehat{y}\right\rVert^{2}\right]+\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}+\bar{f}_{0}(x)\right]$
$\leq\frac{1}{2}\left[L_{0}\left\lVert x-\widehat{y}\right\rVert-\frac{\sqrt{p}L_{0}}{2\epsilon}\left\lVert x-\widehat{y}\right\rVert^{2}\right]+\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}+\bar{f}_{0}(x)\right]$
$<\frac{1}{2}\left[L_{0}\cdot\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon-\frac{\sqrt{p}L_{0}}{2\epsilon}\left(\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon\right)^{2}\right]+\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}+\bar{f}_{0}(x)\right]$
$=\frac{1}{2}\left[-L_{0}\cdot\epsilon-\frac{\sqrt{p}L_{0}}{2\epsilon}\epsilon^{2}\right]+\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}+\bar{f}_{0}(x)\right]$
$\leq\frac{1}{2}\left[\bar{f}_{0}(\widehat{x})-\bar{f}_{0}(x)-C\left\lVert x-\widehat{x}\right\rVert^{2}\right]+\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}+\bar{f}_{0}(x)\right]$
$=C\widehat{x}^{\top}(x-\widehat{x})+\frac{1}{2}C\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x}),$
where the second inequality is derived by the strict monotonicity of the quadratic function $t\mapsto L_{0}t-\frac{\sqrt{p}L_{0}}{2\epsilon}t^{2}$ on $[\epsilon/\sqrt{p},+\infty)$. The above result shows $\left\lVert\widehat{x}_{1}^{*}-x\right\rVert\leq\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$ for all $\widehat{x}_{1}^{*}\in\mathcal{A}_{1}(x)$. Similarly, we have $\left\lVert\widehat{x}_{2}^{*}-x\right\rVert\leq\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$ for all $\widehat{x}_{2}^{*}\in\mathcal{A}_{2}(x)$. It follows that
$\left\lVert C(\widehat{x}_{1}^{*}-\widehat{x}_{2}^{*})\right\rVert\leq C\left\lVert\widehat{x}_{1}^{*}-x\right\rVert+C\left\lVert\widehat{x}_{2}^{*}-x\right\rVert\leq 2\frac{\sqrt{p}L_{0}}{2\epsilon}\cdot\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon\leq(\sqrt{p}+2)L_{0},$
which completes the proof.∎
8.2 Proofs of Proposition 2.2 and Theorem 2.3
8.2.1 Preliminaries on Rademacher Complexity.
We start with a brief introduction to the Rademacher complexity, which is utilized to quantify the generalization error. The empirical Rademacher complexity of a function class is defined as follows.
Definition 8.2
Let $\Phi$ be a class of functions mapping from $\mathcal{U}$ to $\mathbb{R}$ and $S_{N}=\{u^{s}\}_{s=1}^{N}\subseteq\mathcal{U}$ be a sample set of size $N$. Then the empirical Rademacher complexity of $\Phi$ with respect to $S_{N}$ is defined as
^$\widehat{\mathcal{R}}_{S_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
where ${\boldsymbol{\sigma}}=(\sigma_{1},\dots,\sigma_{N})^{\top}$ is a uniform random vector taking values in $\{-1,+1\}^{N}$.
The Rademacher complexity of $\Phi$ is defined as the expectation of the empirical Rademacher complexity over all samples of size $N$ drawn from the underlying distribution:
$\mathcal{R}_{N}(\Phi)\coloneqq\mathbb{E}_{S_{N}}[\widehat{\mathcal{R}}_{S_{N}}(\Phi)].$
It captures the richness of the class $\Phi$ by measuring the degree to which a hypothesis set can fit random noise. The following lemma quantifies the Rademacher complexity of a class of parametric functions, which is generalized from Ermoliev and Norkin (2013, Lemma B.2) with minor modifications made to accommodate the developments in this paper.
Lemma 8.3
Let $\widehat{\Phi}\coloneqq\{\phi(\cdot;\theta):\mathcal{U}\rightarrow\mathbb{R}\mid\theta\in\Theta\}$ be a class of functions parameterized by vectors in a compact set $\Theta\subseteq\mathbb{R}^{q}$ with $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Suppose that there exist $M_{\widehat{\Phi}}>0$ and $L_{\widehat{\Phi}}>0$ such that $\max_{\theta\in\Theta}\left\lVert\phi(\cdot;\theta)\right\rVert_{\infty}\leq M_{\widehat{\Phi}}$ and $\left\lVert\phi(\cdot;\theta_{1})-\phi(\cdot;\theta_{2})\right\rVert_{\infty}\leq L_{\widehat{\Phi}}\left\lVert\theta_{1}-\theta_{2}\right\rVert$ for any $\theta_{1},\theta_{2}\in\Theta$. Then
$\mathcal{R}_{N}(\widehat{\Phi})\leq\frac{2\mu L_{\widehat{\Phi}}\sqrt{q}}{N}+M_{\widehat{\Phi}}\sqrt{\frac{q\log N}{N}}.$
Proof 8.4
Proof of Lemma 8.3 Given a sample set $S_{N}=\{u^{s}\}_{s=1}^{N}$ of size $N$, following the proof of Ermoliev and Norkin (2013, Lemma B.2), we derive that for any $\epsilon\in(0,2\mu]$,
^$\widehat{\mathcal{R}}_{S_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
By setting $\epsilon=2\mu/N$ and taking expectation over $S_{N}$, we obtain the desired upper bound.∎
The following lemma gives the generalization bound based on the Rademacher complexity.
Lemma 8.5 (Mohri et al. (2018), Theorem 3.3)
Let $\Phi$ be a function class mapping from $\mathcal{U}$ to $[-M,M]$. Then for any $\delta>0$, with probability at least $1-\delta$ over the draw of an independently and identically distributed (i.i.d.) sample $S_{N}$ of size $N$, the following holds true for all $\phi\in\Phi$:
$\mathbb{E}[\phi(\tilde{u})]\leq\frac{1}{N}\sum_{s=1}^{N}\phi(u^{s})+2{\mathcal{R}}_{N}(\Phi)+\sqrt{2}M\sqrt{\frac{\log(1/\delta)}{N}}.$
8.2.2 Proofs of Proposition 2.2 and Theorem 2.3.
Proof 8.6
Proof of Proposition 2.2. We first quantify the Rademacher complexity of the composite function class $\Phi_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\left\{\phi:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}\mid\phi(x,y)=\varphi({f(x)};{y}),\,f\in{\mathcal{H}}_{\mu}^{K}\right\}$. Notice that any function in $\Phi_{{\mathcal{H}}_{\mu}^{K}}$ is parameterized by some $\theta\in[-\mu,\mu]^{2dK(p+1)}$. For any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\theta_{1},\theta_{2}\in[-\mu,\mu]^{2dK(p+1)}$,
$\left\lvert\varphi(f(x;\theta_{1});y)-\varphi(f(x;\theta_{2});y)\right\rvert\leq L_{\varphi}\left\lVert f(x;\theta_{1})-f(x;\theta_{2})\right\rVert\leq 2L_{\varphi}\sqrt{p\bar{X}^{2}+1}\left\lVert\theta_{1}-\theta_{2}\right\rVert,$
where the second inequality follows from Assumption (A2). Moreover, for any $f\in{\mathcal{H}}_{\mu}^{K}$,
$\left\lvert\varphi(f(x);y)\right\rvert\leq\left\lvert\varphi(z_{0};y)\right\rvert+\left\lvert\varphi(f(x);y)-\varphi(z_{0};y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{\mu}.$ (14)
Therefore, by Lemma 8.3, we deduce
$\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{4\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}.$ (15)
Now we prove the uniform generalization bound. By Lemma 8.5 and the inequality (14), with probability at least $21-\delta/$ 2,
$\max_{\phi\in\Phi_{{\mathcal{H}}_{\mu}^{K}}}\left(\mathbb{E}_{X,Y}\phi(X,Y)-\frac{1}{n}\sum_{s=1}^{n}\phi(x^{s},y^{s})\right)\leq 2{\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
For $\Phi_{{\mathcal{H}}_{\mu}^{K}}^{-}\coloneqq\{-\phi:\phi\in\Phi_{{\mathcal{H}}_{\mu}^{K}}\}$, we have ${\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}}^{-})={\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})$. Therefore, by applying Lemma 8.5 again to $\Phi_{{\mathcal{H}}_{\mu}^{K}}^{-}$ and using the union bound, we derive that with probability at least $1-\delta$,
$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$ (16)
Finally, by combining (15) and (16), we complete the proof.∎
Proof 8.7
Proof of Theorem 2.3. Let $f^{*}_{{\mathcal{H}}_{\mu}^{K}}$ denote the optimal solution to problem (2) restricted to ${\mathcal{H}}_{\mu}^{K}$. Following the excess risk decomposition (7), we derive the excess risk bound by bounding the approximation and estimation errors respectively. With $f_{app}\in\mathop{\arg\min}_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lVert f-f^{*}\right\rVert_{\infty}$, the approximation error can be bounded as follows:
$R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})-\min_{f\in\mathcal{F}}R(f)$ $\leq R(f_{app})-R(f^{*})$
$=\mathbb{E}_{X,Y}\left[\varphi({f_{app}(X)};{Y})-\varphi({f^{*}(X)};{Y})\right]$
$\leq L_{\varphi}\sup_{x\in\mathcal{X}}\left\lVert f_{app}(x)-f^{*}(x)\right\rVert_{2}$
$\leq 2L_{\varphi}\sqrt{d}(p^{1/2}+3)p^{1/2}L_{z}\bar{X}K^{-1/p},$
where the last inequality follows from Proposition 2.1. For the estimation error, we have
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$ $=R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
$\leq R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{S}(f^{*}_{{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
$\leq 2\left(\max_{f\in{\mathcal{H}}_{\mu}^{K}}\,|R(f)-\widehat{R}_{S}(f)|\right).$
By Proposition 2.2, with probability at least $1-\delta$, we have
$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
By combining these two error bounds, we obtain the desired inequality. ∎
9 Surrogation Construction of $F(\cdot;\xi)$ for Case 1 to Case 3
In this section, we first illustrate the construction of upper surrogation functions of $F(\cdot;\xi)\coloneqq\varphi(f(\cdot;\xi);\xi)$ for Case 1 to Case 3, which are directionally differentiable and satisfy the first three properties in Assumption 1. Then we show in Claim 1 that $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,I)$ with $\varepsilon\geq 0$ satisfies Condition (P5) for the three cases, which also validates Condition (P4) since $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ by construction is directionally differentiable and satisfies (P2).
Case 1:
since $f(\cdot;\xi)$ and $\varphi(\cdot;\xi)$ are both piecewise affine, the composite function $F(\cdot;\xi)=\varphi(f(\cdot;\xi);\xi)$ is also piecewise affine. Thus, following the surrogation construction of $f(\cdot;\xi)$ with the difference-of-max-affine representation in Section 3.1, we can construct the upper surrogation function of $F(\cdot;\xi)$, which satisfies the first three properties in Assumption 1. Specifically, the first two properties clearly hold, and the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$ is justified since given a fixed index mapping $I$, the surrogate of a PA function is independent of the reference point $\theta^{\prime}$.
Case 2:
by assumption, $\varphi(\cdot;\xi)$ can be written as $\varphi(\cdot;\xi)=\varphi^{\uparrow}(\cdot;\xi)+\varphi^{\downarrow}(\cdot;\xi)$, where $\varphi^{\uparrow}(\cdot;\xi)$ and $\varphi^{\downarrow}(\cdot;\xi)$ are nondecreasing and nonincreasing convex functions respectively, that is, for any $\theta_{1},\theta_{2}\in\Theta$ satisfying $\theta_{1}\leq\theta_{2}$ componentwisely, we have $\varphi^{\uparrow}(\theta_{2};\xi)\geq\varphi^{\uparrow}(\theta_{1};\xi)$ and $\varphi^{\downarrow}(\theta_{2};\xi)\leq\varphi^{\downarrow}(\theta_{1};\xi)$. With $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$, the surrogation function of $F(\cdot;\xi)$ at $\theta^{\prime}$ is constructed as
$\widehat{F}(\theta;\theta^{\prime},\xi,I)=\varphi^{\uparrow}(\widehat{f}(\theta;\xi,I_{2});\xi)+\varphi^{\downarrow}(\widecheck{f}(\theta;\xi,I_{1});\xi),$
where $\widehat{f}(\theta;\xi,I_{2})$ and $\widecheck{f}(\theta;\xi,I_{1})$ are the upper and lower surrogation functions of $f(\theta,\xi)$ respectively. By monotonicity of $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$, $\widehat{F}(\theta;\theta^{\prime},\xi,I)\geq\varphi^{\uparrow}(f(\theta;\xi);\xi)+\varphi^{\downarrow}(f(\theta;\xi);\xi)=F(\theta;\xi)$ for any $\theta\in\Theta$, and thus the majorizing condition (P2) holds. Moreover, $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies the convex condition (P3) since a nondecreasing convex function composite with a convex function and a nonincreasing convex function composite with a concave function are both convex. The touching condition (P1) clearly holds when $\varepsilon=0$. With a fixed index mapping $I$, $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ in this case is also independent of $\theta^{\prime}$, validating the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$.
Case 3:
since $S$ is finite, for any $\xi\in S$, $\varphi(\cdot;\xi)$ is Lipschitz smooth with uniform modulus denoted by $L_{\nabla\varphi}$, and $f(\cdot;\xi)$ is Lipschitz continuous with uniform modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$. Then the surrogation function of $F(\cdot;\xi)$ at $\theta^{\prime}$ with $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ is constructed by surrogating the inner PA function according to componentwise signs of the gradient $\nabla\varphi(f(\theta^{\prime};\xi);\xi)$ as follows,
$\widehat{F}(\theta;\theta^{\prime},\xi,I)$ $=\max\{\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widehat{f}(\theta;\xi,I_{2})-\max\{-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widecheck{f}(\theta;\xi,I_{1})$
$\phantom{=,}-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi)^{\top}f(\theta^{\prime};\xi)+\varphi(f(\theta^{\prime};\xi);\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}.$
The convex condition (P3) holds evidently; the touching condition (P1) also holds when $\varepsilon=0$. The surrogate $\widehat{F}(\cdot;\cdot,\xi,I)$ is continuous on $\Theta\times\Theta$ since $\nabla\varphi(\cdot;\xi)$ is continuous. Based on the quadratic upper bound of a Lipschitz smooth function, the majorizing condition (P2) is verified as follows. Since $\varphi(\cdot;\xi)$ is $L_{\nabla\varphi}$-smooth, we have for any $z_{1},z_{2}\in\mathbb{R}^{d}$ that
$\varphi(z_{1};\xi)\leq\varphi(z_{2};\xi)+\nabla\varphi(z_{2};\xi)^{\top}(z_{1}-z_{2})+\frac{L_{\nabla\varphi}}{2}\left\lVert z_{1}-z_{2}\right\rVert^{2}.$
By setting $z_{1}=f(\theta;\xi)$ and $z_{2}=f(\theta^{\prime};\xi)$, we obtain that for any $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$,
$F(\theta;\xi)$ $\leq F(\theta^{\prime};\xi)+\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}(f(\theta;\xi)-f(\theta^{\prime};\xi))+\frac{L_{\nabla\varphi}}{2}\left\lVert f(\theta;\xi)-f(\theta^{\prime};\xi)\right\rVert^{2}$
$\leq F(\theta^{\prime};\xi)+\max\{\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widehat{f}(\theta;\xi,I_{2})$
$\phantom{\leq,}-\max\{-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widecheck{f}(\theta;\xi,I_{1})-\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}f(\theta^{\prime};\xi)+\frac{L_{\nabla\varphi}L_{f}^{2}}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}$
$=\widehat{F}(\theta;\theta^{\prime},\xi,I).$
Claim 1
Let $\Theta$ be a compact set. Given $\varepsilon\geq 0$ and $\theta^{\prime}\in\Theta$, with the surrogation construction for Case 1 to Case 3, $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ satisfies Condition (P5) for any $\xi\in S$.
Proof 9.1
Proof of Claim 1. Since $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ also satisfies the majorizing condition (P2), by the compactness of $\Theta$, it suffices to show that
$\widehat{\kappa}\coloneqq\limsup_{\theta\rightarrow\theta^{\prime},\theta\neq\theta^{\prime}}\frac{2(\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)-F(\theta;\xi))}{\left\lVert\theta-\theta^{\prime}\right\rVert^{2}}<+\infty.$
For the piecewise affine function $f(\theta;\xi)\coloneqq g(\theta;\xi)-h(\theta;\xi)$, we have $\mathcal{I}_{h}^{0}(\theta)\subseteq\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ when $\theta\in\Theta$ is sufficiently close to $\theta^{\prime}$. It follows from (P2) that for any $\theta\in\Theta$ sufficiently close to $\theta^{\prime}$,
$f(\theta;\xi)\leq\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2})\leq\min_{I_{2}\in\mathcal{I}_{h}^{0}(\theta)}\widehat{f}(\theta;\xi,I_{2})=f(\theta;\xi).$
Therefore, by similar arguments for lower surrogation functions of $f(\cdot;\xi)$, we obtain that for any $\theta\in\Theta$ sufficiently close to $\theta^{\prime}$,
$\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2})=f(\theta;\xi)=\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\widecheck{f}(\theta;\xi,I_{1}).$
For Case 1, since the objective is piecewise affine, we directly obtain $\widehat{\kappa}=0$. For Case 2, when $\theta$ is sufficiently close to $\theta^{\prime}$, we have
$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)$ $\coloneqq\min_{(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,(I_{1},I_{2}))$
$=\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\varphi^{\uparrow}(\widehat{f}(\theta;\xi,I_{2});\xi)+\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\varphi^{\downarrow}(\widecheck{f}(\theta;\xi,I_{1});\xi)$
$=\varphi^{\uparrow}\left(\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2});\xi\right)+\varphi^{\downarrow}\left(\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\widecheck{f}(\theta;\xi,I_{1});\xi\right)$
$=F(\theta;\xi),$
also implying $\widehat{\kappa}=0$. For Case 3, when $\theta$ is sufficiently close to $\theta^{\prime}$, we obtain
$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)=\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}(f(\theta;\xi)-f(\theta^{\prime};\xi))+\varphi(f(\theta^{\prime};\xi);\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}.$
Thus, by the second-order Taylor expansion of $\varphi(\cdot;\xi)$ at $f(\theta^{\prime};\xi)$, we obtain
$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)-F(\theta;\xi)=\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}-\frac{1}{2}v^{\top}\nabla^{2}\varphi(f(\theta^{\prime};\xi);\xi)v+o(\left\lVert v\right\rVert^{2}),$
[…]
10 Details of Convergence Analysis in Section 4
10.1 Proof of Proposition 4.2 and Equivalence Between $\mathcal{S}^{d}_{\varepsilon,\rho}$ and $\mathcal{S}^{*}$
Proof 10.1
Proof of Proposition 4.2. Given $\varepsilon\geq 0$ and $\rho>0$, for any $\theta^{*}\in\mathcal{S}^{*}$ it holds that $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{*})+\frac{\rho}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}>F(\theta)\geq F(\theta^{*})=\widehat{F}^{\varepsilon}_{\min}(\theta^{*};\theta^{*})$ for any $\theta\neq\theta^{*}$, showing $\theta^{*}\in\mathcal{S}^{d}_{\varepsilon,\rho}$ and thus $\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}$. For any $\theta^{d}\in\mathcal{S}^{d}_{\varepsilon,\rho}$, from Condition (P4) and the first-order optimality condition, it follows that $F^{\prime}(\theta^{d};\theta-\theta^{d})=\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{d})^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for all $\theta\in\Theta$, yielding $\theta^{d}\in\mathcal{S}^{d}$ and thus $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$.
It remains to show $\mathcal{S}^{d}\subseteq\mathcal{S}^{d}_{0,\rho}$ for any $\rho>0$. Assume by contradiction that there exists $\theta^{d}\in\mathcal{S}^{d}\backslash\mathcal{S}^{d}_{0,\widehat{\rho}}$ for some $\widehat{\rho}>0$. Then there exist $\widehat{I}\in\mathcal{I}^{0}(\theta^{d})$ and $\widehat{\theta}(\neq\theta^{d})\in\Theta$ such that $\widehat{\theta}\in\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}(\theta;\theta^{d},\widehat{I}\,)+\frac{\widehat{\rho}}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}$. By the strong convexity of the function $\theta\mapsto\widehat{F}(\theta;\theta^{d},\widehat{I}\,)+\frac{\widehat{\rho}}{2}\,\left\lVert\theta-\theta^{d}\right\rVert^{2}$, we have $\widehat{F}(\cdot;\theta^{d},\widehat{I})^{\prime}(\theta^{d};\widehat{\theta}-\theta^{d})<0$. Thus, there exists $t_{0}>0$ such that
$F(\theta^{d}+t(\widehat{\theta}-\theta^{d}))\leq\widehat{F}(\theta^{d}+t(\widehat{\theta}-\theta^{d});\theta^{d},\widehat{I})<\widehat{F}(\theta^{d};\theta^{d},\widehat{I})=F(\theta^{d}),\quad\forall\,t\in(0,t_{0}],$
where the equality holds by the touching condition (P1). This contradicts that $\theta^{d}$ is a d-stationary point. Therefore, we have $\mathcal{S}^{d}\subseteq\mathcal{S}^{d}_{0,\rho}$, which completes the proof.∎
As an extension, we show the equivalence between $\mathcal{S}^{d}_{\varepsilon,\rho}$ and $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ and suitable $\rho$ in Proposition 10.2. The following assumption is needed:
  1. [label=(B0), resume*=B, labelwidth=!, labelindent=0pt]
  2. 1.
    There exists $M_{\Delta}>0$ such that $F(\theta^{d})\geq\min_{\theta\in\Theta}F(\theta)+M_{\Delta}$ for any $\theta^{d}\in\mathcal{S}^{d}\backslash\mathcal{S}^{*}$.
The assumption is naturally satisfied when the outer function of $F$ is in Case 1 or Case 2. Specifically, from the proof of Claim 2 afterward, we see that in these two cases, $\Theta$ can be decomposed into finitely many convex regions, on each of which any d-stationary point $\theta^{d}\in\mathcal{S}^{d}$ is a minimizer of $F$ over that region. Since the number of such regions is finite, the function $F$ takes finitely many values on $\mathcal{S}^{d}$, verifying Assumption 1. Discussions on the finite number of d-stationary values for more general nonconvex problems with piecewise and composite structures can be seen in Cui and Pang (2018).
Proposition 10.2
For problem (9) with $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$ and Assumption 1 holding for the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ for each $\xi\in S$, the following statements hold:
  1. (a)
    With $\varphi$ in Case 1 or Case 2, $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$, where $\varepsilon_{0}\coloneqq 2\mu\sqrt{q}\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, and $M_{\Delta}$ is the constant defined in 1.
  2. (b)
    Suppose that Assumptions 2 and 1 additionally hold with $M_{\Delta}>2\kappa q\mu^{2}$. It holds that $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}-\kappa]$.
Proof 10.3
Proof of Proposition 10.2. By Proposition 4.2, it suffices to prove the inclusion $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{*}$. Since the two max-affine components of $f(\cdot;\xi)$ for any $\xi\in S$ are Lipschitz continuous with modulus $L_{g,h}\coloneqq\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, we see that with $\varepsilon\geq\varepsilon_{0}=2L_{g,h}\sqrt{q}\mu$, the set $\mathcal{I}^{\varepsilon}(\theta^{\prime})$ contains all index mappings, that is, $\mathcal{I}^{\varepsilon}(\theta^{\prime})=\{(I_{1},I_{2})\mid I_{1}:S\rightarrow[K],I_{2}:S\rightarrow[K]\}$ for any $\theta^{\prime}\in\Theta$.
For the first statement, assume by contradiction that there exists $\widehat{\theta}^{d}\in\mathcal{S}^{d}_{\widehat{\varepsilon},\widehat{\rho}}\backslash\mathcal{S}^{*}$ for some $\widehat{\varepsilon}\in[\varepsilon_{0},+\infty)$ and $\widehat{\rho}\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$. Since the set $\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ contains all index mappings for the inner PA functions, by the surrogation construction for Case 1 and Case 2 in Appendix 9, for any $\theta^{*}\in\mathcal{S}^{*}$ there exists $I^{\widehat{\varepsilon}}\in\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ such that $\widehat{F}^{\widehat{\varepsilon}}_{\min}(\theta^{*};\widehat{\theta}^{d})=\widehat{F}(\theta^{*};\widehat{\theta}^{d},I^{\widehat{\varepsilon}})=\min_{\theta\in\Theta}F(\theta)$. It follows that
$\widehat{F}^{\widehat{\varepsilon}}_{\min}(\theta^{*};\widehat{\theta}^{d})+\frac{\widehat{\rho}}{2}\,\big{\lVert}\,{\theta^{*}-\widehat{\theta}^{d}}\,\big{\rVert}^{2}\leq\min_{\theta\in\Theta}F(\theta)+M_{\Delta}\leq F(\widehat{\theta}^{d})=\widehat{F}^{\widehat{\varepsilon}}_{\min}(\widehat{\theta}^{d};\widehat{\theta}^{d}),$
which contradicts $\widehat{\theta}^{d}\in\mathcal{S}^{d}_{\widehat{\varepsilon},\widehat{\rho}}$. For the second statement, in a similar way, assume that there exists $\widehat{\theta}^{d}\in\mathcal{S}^{d}_{\widehat{\varepsilon},\widehat{\rho}}\backslash\mathcal{S}^{*}$ for some $\widehat{\varepsilon}\in[\varepsilon_{0},+\infty)$ and $\widehat{\rho}\in(0,\frac{M_{\Delta}}{2q\mu^{2}}-\kappa]$. By Condition (P5), we have that for $\theta^{*}\in\mathcal{S}^{*}$,
$\widehat{F}^{\widehat{\varepsilon}}_{\min}(\theta^{*};\widehat{\theta}^{d})+\frac{\widehat{\rho}}{2}\,\big{\lVert}\,{\theta^{*}-\widehat{\theta}^{d}}\,\big{\rVert}^{2}\leq F(\theta^{*})+\frac{\widehat{\rho}+\kappa}{2}\,\big{\lVert}\,{\theta^{*}-\widehat{\theta}^{d}}\,\big{\rVert}^{2}\leq\min_{\theta\in\Theta}F(\theta)+M_{\Delta}\leq F(\widehat{\theta}^{d}).$
Again, it leads to a contradiction, which completes the proof.∎
10.2 Proof of Proposition 4.3
The following lemma from Deng and Gao (2021) is rephrased in our context for the proof of Proposition 4.3.
Lemma 10.4 (Deng and Gao (2021), Theorem 3.2)
Suppose that the stochastic function $\widehat{F}(\cdot;\theta^{\prime},\xi)$ is convex and Lipschitz continuous on $\Theta$ with modulus $L_{\widehat{F}}$ for any $\xi\in S$ and $\theta^{\prime}\in\Theta$, and that samples $\Xi_{N}\coloneqq\{\xi^{s}\}_{s=1}^{N}$ are i.i.d. generated from $S$. For $\eta>0$, denote $\theta^{\prime}_{+}\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\frac{1}{N}\sum_{s=1}^{N}\widehat{F}(\theta;\theta^{\prime},\xi^{s})+\frac{\eta}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}$. Then it holds that
$\left\lvert\mathbb{E}_{\Xi_{N}}\left[\frac{1}{N}\sum_{s=1}^{N}\widehat{F}(\theta^{\prime}_{+};\theta^{\prime},\xi^{s})-\mathbb{E}_{\widetilde{\xi}_{n}}\left[\widehat{F}(\theta^{\prime}_{+};\theta^{\prime},\widetilde{\xi}_{n})\right]\right]\right\rvert\leq\frac{2L_{\widehat{F}}^{2}}{\eta N}.$
Proof 10.5
Proof of Proposition 4.3. Let $\widetilde{I}^{\varepsilon}_{\nu}$ be a random mapping uniformly drawn from $\mathcal{I}^{\varepsilon}(\theta^{\nu})$, $\{\widetilde{\xi}^{s}_{\nu}\}_{s=1}^{N_{\nu}}$ be random vectors independently and uniformly drawn from $S$, $\mathcal{G}_{\nu-1}$ denote the $\sigma$-algebra generated by all past samples $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and mappings $\{\widetilde{I}^{\varepsilon}_{t}\}_{t=0}^{\nu-1}$, and $\mathcal{G}_{\nu-1}^{+}$ denote the $\sigma$-algebra generated by $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and $\{\widetilde{I}^{\varepsilon}_{t}\}_{t=0}^{\nu}$. Given the current iterate $\theta^{\nu}$ and the index mapping $I^{\varepsilon}_{\nu}\in\mathcal{I}^{\varepsilon}(\theta^{\nu})$, by the definition of $P_{I^{\varepsilon}_{\nu}}^{\rho}(\cdot)$ in (10), we have
$\min\left\{\widehat{F}(P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},I^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}+\frac{\rho}{2}\,\big{\lVert}\,{P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{\nu}).$ (17)
With the notation $\mathbb{E}_{N_{\nu}}$ representing the expectation with respect to $\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$, by rearranging terms of (17) and taking the conditional expectation on both sides, we obtain that
$\frac{\rho}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$ $\leq F(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=\underbrace{F(\theta^{\nu})-W(\theta^{\nu})}_{\mathcal{E}_{1}^{\nu}}+\underbrace{W(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]}_{\mathcal{E}_{2}^{\nu}},$ (18)
where $W(\theta^{\nu})\coloneqq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$.
From the update rule in Step 5 of Algorithm 1, we have $\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\geq F_{N_{\nu}}(\theta^{\nu+1})$ and $F_{N_{\nu}}(\theta^{\nu})\geq F_{N_{\nu}}(\theta^{\nu+1})$. It follows that
$W(\theta^{\nu})$ $\geq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\mid\mathcal{G}_{\nu-1}\right]=\mathbb{E}\left[F_{N\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right].$
With $\Theta$ being a compact set, let $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Define
$\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\{\phi(\cdot;\theta):S\rightarrow\mathbb{R}\mid\phi(\xi;\theta)=\varphi(f(\theta;\xi);\xi),\,\xi\in S,f(\theta;\cdot)\in{\mathcal{H}}_{\mu}^{K},\theta\in\Theta\}.$
From finiteness of $S$ and continuity of $\varphi(\cdot;\xi)$, it follows that for any $\phi\in\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}$, $\phi(\xi;\cdot)$ is Lipschitz continuous on the compact set $\Theta$ with uniform modulus, and that $\phi(\cdot;\theta)$ is uniformly bounded. Therefore, for the term $\mathcal{E}_{1}^{\nu}$, there exists a constant $C_{1}>0$ such that
$\mathcal{E}_{1}^{\nu}$ $\leq F(\theta^{\nu})-\mathbb{E}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
$=F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\sup_{\theta\in\Theta}\left\lvert F(\theta)-F_{N_{\nu}}(\theta)\right\rvert\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+2\mathcal{R}_{N_{\nu}}(\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}})$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+C_{1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}},$ (19)
where the third inequality follows from inequalities (3.8)-(3.13) in Mohri et al. (2018) that are part of the proof of Lemma 8.5, and the last inequality is derived by Lemma 8.3.
Next, we bound the term $\mathcal{E}_{2}^{\nu}$. Given $I^{\varepsilon}_{\nu}$ and $\{\xi_{\nu}^{s}\}_{s=1}^{N_{\nu}}$, by the optimality of ${\theta}^{\nu+\frac{1}{2}}$, we have
$\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\leq\widehat{F}_{N_{\nu}}(P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}.$
It follows that
$W(\theta^{\nu})$ $\leq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]+\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right],$
where the equality holds since both $\theta^{\nu}$ and $P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})$ are independent of $\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$. Then we obtain
$\mathcal{E}_{2}^{\nu}$ $\leq\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$ (20)
Combining bounds (10.5) and (20) of $\mathcal{E}_{1}^{\nu}$ and $\mathcal{E}_{2}^{\nu}$ and rearranging terms, we derive that
$\frac{\rho-\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+C_{1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
By applying Jensen’s inequality, taking full expectation on both sides, and summing inequalities up over $\nu=0,\dots,T-1$, we have
$\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\big{(}\,\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\theta^{\nu})]\,\big{)}^{2}\leq\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\mathbb{E}\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{0})-\mathbb{E}\left[F(\theta^{T})\right]+C_{1}\cdot\sum_{\nu=0}^{T-1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
Dividing both sides by $T$, applying Jensen’s inequality again, and using the fact that $\mathbb{E}\left[F(\theta^{T})\right]\geq\min_{\theta\in\Theta}F(\theta)$, we finally obtain that
$\left(\mathbb{E}\,\big{[}\,\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{T})\,\big{]}\right)^{2}\leq\frac{F(\theta^{0})-\min_{\theta\in\Theta}F(\theta)+C_{1}\cdot\sum_{\nu=0}^{T-1}\sqrt{\log N_{\nu}/N_{\nu}}}{\frac{\rho-\eta}{2}\cdot T}.$
Now we consider the special case with $\varepsilon=0$ and $\eta>0$. In such a case, ${\theta}^{\nu+\frac{1}{2}}=\theta^{\nu+1}$, and the intermediate term $W(\theta^{\nu})$ degenerates into
$W(\theta^{\nu})=\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$
The term $\mathcal{E}_{2}^{\nu}$ is still bounded by (20). For $\mathcal{E}_{1}^{\nu}$, we can derive a tighter bound in the complexity of the sampling size $N_{\nu}$. Specifically, there exists $C_{2}>0$ such that
$\mathcal{E}_{1}^{\nu}$ $=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})-\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\frac{C_{2}}{N_{\nu}},$
where the last inequality holds by Lemma 10.4, and the Lipschitz continuity of surrogates follows from the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$, compactness of $\Theta$, and finiteness of $S$. Then the second inequality in the proposition can be proved with arguments similar to those above.∎
10.3 Proofs of Theorems 4.4 and 4.5
In this section, we prove the convergence results in Theorems 4.4 and 4.5. As in Section 4.2, we suppose throughout this section that $\varphi$ is B-differentiable, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ satisfies Assumption 1 for each $\xi\in S$.
The convergence analysis requires the following two residual functions. The first residual of a reference point $\theta^{\prime}\in\Theta$ with $\varepsilon\geq 0$ and $\rho>0$ measures the proximal mapping gap with respect to the surrogation function $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime})\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},I)$, which is defined as
$r^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\max_{\theta\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta^{\prime})}\left\lVert\theta-\theta^{\prime}\right\rVert,\text{\ where\ }\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime})+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}.$ (21)
Clearly, $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (9) if and only if $r^{\varepsilon,\rho}(\theta^{d})=0$. The residual $r^{\varepsilon,\rho}$ works as a criterion for the subsequential convergence to composite strong d-stationarity and is also exploited to establish the error-bound analysis. The other useful residual is defined as $r_{d}(\theta)\coloneqq\underset{d\in\mathcal{T}(\theta;\Theta),\left\lVert d\right\rVert\leq 1}{\max}\left\{-F^{\prime}(\theta;d)\right\}$, where $\mathcal{T}(\theta;\Theta)$ is the tangent cone of $\Theta$ at $\theta\in\Theta$. This residual is utilized to derive the global error bound in terms of $r^{\varepsilon,\rho}$ in the second statement of Proposition 10.6. In Figure 10.3, we summarize a roadmap for the convergence analysis of the enhanced SMM algorithm.
A Roadmap for Convergence Analysis of Enhanced SMM.
We first provide in the following proposition the limiting property of the residual $r^{\varepsilon,\rho}$ and the global error bound of distance to $\mathcal{S}^{d}$ in terms of $r^{\varepsilon,\rho}$.
Proposition 10.6
For problem (9), the following statements hold:
  1. (a)
    For ${\varepsilon}>0$ and ${\rho}>0$, let $\theta_{\infty}$ be a limit point of a sequence $\{\theta_{t}\}_{t}$ in $\Theta$ satisfying $\lim_{t\rightarrow\infty}r^{{\varepsilon},{\rho}}(\theta_{t})=0$. Then $\theta_{\infty}$ is a composite $(\varepsilon^{\prime},\rho^{\prime})$-strongly d-stationary point for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho^{\prime}\in(\rho,+\infty)$.
  2. (b)
    Under Assumptions 2 and 1, for $\varepsilon>0$ and $\rho>0$, there exists a constant $\eta_{r}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}r^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$. In particular, if $\varphi$ is piecewise affine and $\Theta$ is a polytope, the error bound holds for any $\varepsilon\geq 0$ and $\rho>0$ without the two assumptions.
Before proving this proposition, we show that Theorem 4.4 could be derived straightforwardly by combining Propositions 4.3 and 10.6, similarly for Theorem 4.5.
Proof 10.7
Proof of Theorem 4.4. According to the definition in (21), for any $\theta\in\mathcal{P}^{\varepsilon,\rho}_{\min}(\theta^{\prime})$, there exists $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ such that $\theta=P^{\rho}_{I}(\theta^{\prime})$. Due to the finite-sum structure of the objective $F$, there exists a uniform constant $C_{\mathcal{I}}>0$ such that the cardinality $\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert\leq C_{\mathcal{I}}$ for any $\theta^{\prime}\in\Theta$, and thus ${r}^{\varepsilon,\rho}(\theta^{\prime})\leq\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert\bar{r}^{\varepsilon,\rho}(\theta^{\prime})\leq C_{\mathcal{I}}\bar{r}^{\varepsilon,\rho}(\theta^{\prime})$, which implies that $\mathbb{E}\,[{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ converges to 0 at least at the same rate as $\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ in terms of $\{N_{\nu}\}_{\nu}$ and $T$. Therefore, the statements of the theorem follow from Propositions 4.3 and 10.6.∎
We next present the proof of Proposition 10.6. The overall scheme of the error-bound analysis follows the error-bound theory developed in Liu et al. (2022) for the post-convergence analysis of the SMM algorithm. To justify the global error bound in the second statement of Proposition 10.6, we first give the error bounds of distance to $\mathcal{S}^{d}$ in terms of the residual $r_{d}$ in Proposition 10.8 and then relate $r_{d}$ with $r^{\varepsilon,\rho}$ in Proposition 10.10.
Proposition 10.8
For problem (9), the following statements hold:
  1. (a)
    If $\varphi$ is piecewise affine and $\Theta$ is a polyhedral set, then there exists $\eta_{d}>0$ such that
    $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{d}r_{d}(\theta),\quad\forall\,\theta\in\Theta.$
  2. (b)
    Under Assumption 1 with $\varepsilon>0$, $\rho>0$, and the corresponding constants $\delta_{q}$, $\lambda_{q}$, and $c_{q}$,
    $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\frac{2}{c_{q}}r_{d}(\theta),\quad\text{$\forall\,\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$ and $r^{\varepsilon,\rho}(\theta)\leq\lambda_{q}$.}$
Proof 10.9
Proof of Proposition 10.8. For the first statement, since the objective $F$ is a piecewise affine function on a polyhedral set $\Theta$, $\Theta$ can be decomposed into finitely many polyhedral sets such that $F$ is affine on each of them, showing that $r_{d}(\cdot)$ takes finitely many values on $\Theta$. Notice that $\theta\in\mathcal{S}^{d}$ if and only if $r_{d}(\theta)=\max_{d\in\mathcal{T}(\theta;\Theta),\left\lVert d\right\rVert\leq 1}\{-F^{\prime}(\theta;d)\}=0$. Hence, there exists $l_{0}>0$ such that $r_{d}(\theta)\geq l_{0}$ for any $\theta\in\Theta\backslash\mathcal{S}^{d}$. By setting $\eta_{d}\coloneqq\frac{2}{l_{0}}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, we deduce $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq 2\max_{\theta\in\Theta}\left\lVert\theta\right\rVert\leq\eta_{d}r_{d}(\theta)$.
For the second statement, the inequality holds trivially if $\theta\in\mathrm{cl}(\mathcal{S}^{d})$. By Assumption 1, for any $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$ satisfying $r^{\varepsilon,\rho}(\theta)\leq\lambda_{q}$ and $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(\cdot;\theta^{d},\theta)$ is convex and $\bar{F}(t;\theta^{d},\theta)\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{q}}{2}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}$ for all $t\in[0,1].$ According to Drusvyatskiy and Lewis (2018, Theorem 3.3), it follows that
$\left\lVert\theta-\theta^{d}\right\rVert\leq\frac{2}{c_{q}}\mathrm{dist}(0,\partial\bar{F}(1;\theta^{d},\theta))\leq\frac{2}{c_{q}}\left[-\bar{F}(\cdot;\theta^{d},\theta)^{\prime}(1;-1)\right]=\frac{2}{c_{q}}\left[-F^{\prime}\left(\theta;\frac{\theta^{d}-\theta}{\left\lVert\theta^{d}-\theta\right\rVert}\right)\right]\leq\frac{2}{c_{q}}r_{d}(\theta).$
Thus, we obtain $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\left\lVert\theta-\theta^{d}\right\rVert\leq\frac{2}{c_{q}}r_{d}(\theta)$, which completes the proof.∎
Proposition 10.10
Under Assumption 2, for $\varepsilon\geq 0$ and $\rho>0$, it holds that for any $\theta\in\Theta$, there exists $\bar{\theta}\in\Theta$ such that $\left\lVert\bar{\theta}-\theta\right\rVert\leq 2\,r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\bar{\theta})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}(\theta)$.
A similar result to Proposition 10.10 is provided in Liu et al. (2022, Lemma 4) under the assumption that the relevant surrogation function satisfies Condition (P5). With the surrogation function $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime})$ associated with $r^{\varepsilon,\rho}(\theta^{\prime})$ satisfying (P5), we can prove Proposition 10.10 following the lines of Liu et al. (2022, Lemma 4), and thus the details are omitted here.
Proof 10.11
Proof of Proposition 10.6. For the first statement, let $\{\theta_{t}\}_{t\in\mathcal{K}}$ be a convergent subsequence with the limit point $\theta_{\infty}$. Given $\varepsilon^{\prime}\in[0,\varepsilon)$, it follows from the Lipschitz continuity of $f$ that $\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\subseteq\mathcal{I}^{\varepsilon}(\theta_{t})$ for sufficiently large $t\in\mathcal{K}$. With $\widehat{P}^{\rho}_{I}(\theta_{t})\coloneqq\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta_{t},I)+\frac{\rho}{2}\left\lVert\theta-\theta_{t}\right\rVert^{2}\right\}$, by definition of the residual $r^{\varepsilon,\rho}$ there exists $I^{*}_{t}\in\underset{I\in\mathcal{I}^{\varepsilon}(\theta_{t})}{\arg\min}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon,\rho}(\theta_{t})=\,\big{\lVert}\,{\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}}\,\big{\rVert}$. Then for sufficiently large $t\in\mathcal{K}$, we have that
$\widehat{F}^{*}(\theta_{t})$ $\coloneqq\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
$\geq\min_{I\in\mathcal{I}^{\varepsilon}(\theta_{t})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
$=\widehat{F}(\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t});\theta_{t},I^{*}_{t})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}$
$\geq F(\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t}))+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}.$ (22)
Since $\lim_{t(\in\mathcal{K})\rightarrow\infty}r^{{\varepsilon},{\rho}}(\theta_{t})=0$, we deduce $\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})=\theta_{\infty}$. Additionally, for a fixed mapping $I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})$, the surrogation function $\widehat{F}(\cdot;\cdot,I)$ is continuous on $\Theta\times\Theta$ by assumption, and thus by Shapiro et al. (2021, Theorem 7.23), $\widehat{P}^{\rho}_{I}(\cdot)$ is continuous at $\theta_{\infty}$. It follows from (10.11) that
$F(\theta_{\infty})\leq\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{F}^{*}(\theta_{t})=\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{\infty});\theta_{\infty},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\},$
which leads to $F(\theta_{\infty})<\displaystyle{\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\backslash\mathcal{I}^{0}(\theta_{\infty})}}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{I}(\theta_{\infty});\theta_{\infty},I)+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,\widehat{P}^{\rho^{\prime}}_{I}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\}$ for $\rho^{\prime}>\rho$, yielding $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$.
We next prove the second statement. For the special case of $\varphi$ being piecewise affine, Assumption 2 holds naturally. Given $\varepsilon\geq 0$ and $\rho>0$, by Proposition 10.8 there exists $\eta_{d}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{d}r_{d}(\theta)$ for any $\theta\in\Theta$; by Proposition 10.10, for any $\theta\in\Theta$, there exists $\bar{\theta}\in\Theta$ such that $\left\lVert\bar{\theta}-\theta\right\rVert\leq 2r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\bar{\theta})\leq(2\rho+\frac{5\kappa}{2})r^{\varepsilon,\rho}(\theta)$. Setting $\eta_{r}=2+\eta_{d}\left(2\rho+\frac{5\kappa}{2}\right)$, we have that
$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\eta_{d}r_{d}(\bar{\theta})+2r^{\varepsilon,\rho}(\theta)\leq\eta_{r}r^{\varepsilon,\rho}(\theta).$
To prove the desired statement for general B-differentiable $\varphi$ under Assumptions 2 and 1, we first show that given $\varepsilon>\varepsilon^{\prime}\geq 0$ and $\rho^{\prime}>\rho>0$, for any $\lambda^{\prime}_{d}>0$ there exists $\lambda_{d}>0$ such that
$\Big{\{}\theta^{\prime}\in\Theta:r^{\varepsilon,\rho}(\theta)\leq\lambda_{d},\ \lVert\theta^{\prime}-\theta\rVert\leq 2\lambda_{d},\ \theta\in\Theta\Big{\}}\subseteq\left\{\theta\in\Theta:r^{\varepsilon^{\prime},\rho^{\prime}}(\theta)\leq\lambda_{d}^{\prime}\right\}.$ (23)
By contradiction, suppose that there exists $\bar{\lambda}^{\prime}_{d}>0$ such that for any $\lambda>0$, there exist points $\theta,\theta^{\prime}\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq\lambda$, $\left\lVert\theta-\theta^{\prime}\right\rVert\leq 2\lambda$, and $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime})>\bar{\lambda}^{\prime}_{d}$. Then there exist two sequences $\{\theta_{k}\}_{k}$ and $\{\theta^{\prime}_{k}\}_{k}$ in $\Theta$ with $\inf_{k\in\mathbb{N}}r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime}_{k})\geq\bar{\lambda}^{\prime}_{d}$ such that both $r^{\varepsilon,\rho}(\theta_{k})$ and $\left\lVert\theta_{k}-\theta^{\prime}_{k}\right\rVert$ converge to 0. Since $\Theta$ is compact, by passing to a subsequence if necessary, we can assume that both $\{\theta_{k}\}_{k}$ and $\{\theta^{\prime}_{k}\}_{k}$ converge to $\theta_{\infty}\in\Theta$. By the first statement of the proposition, we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$. Let $I_{k}\in{\arg\min}_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta^{\prime}_{k})}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{I}(\theta^{\prime}_{k});\theta^{\prime}_{k},I)+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{I}(\theta^{\prime}_{k})-\theta^{\prime}_{k}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime}_{k})=\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{I_{k}}(\theta^{\prime}_{k})-\theta^{\prime}_{k}}\,\big{\rVert}>\bar{\lambda}^{\prime}_{d}$. Since the total number of index mappings is finite, by passing to a subsequence if necessary, we can assume that $I_{k}\equiv\widehat{I}$ for any $k\in\mathbb{N}$. By continuity of $\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\cdot)$ at $\theta_{\infty}$, we derive $\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}\geq\bar{\lambda}^{\prime}_{d}$ and $\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty});\theta_{\infty},\widehat{I})+\frac{\rho^{\prime}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}^{2}\leq F(\theta_{\infty})$, implying $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})>0$, a contradiction.
Given $\varepsilon^{\prime}\in(0,\varepsilon)$ and $\rho^{\prime}\in(\rho,+\infty)$, let $\delta_{q}^{\prime},\lambda_{q}^{\prime}$, and $c_{q}^{\prime}$ be the corresponding constants in Assumption 1. There exists $\lambda_{q}>0$ such that (23) holds for $\lambda_{q}^{\prime}$. For $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\frac{\delta_{q}^{\prime}}{3}$ and $r^{\varepsilon,\rho}(\theta)\leq\lambda_{0}\coloneqq\min\left\{\lambda_{q},\frac{\delta_{q}^{\prime}}{3}\right\}$, by Proposition 10.10 let $\bar{\theta}$ be the point satisfying $\left\lVert\bar{\theta}-{\theta}\right\rVert\leq 2\,r^{\varepsilon,\rho}({\theta})\leq 2\lambda_{q}$ and $r_{d}(\bar{\theta})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}({\theta})$. Then we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta})\leq\lambda_{q}^{\prime}$ from (23) and
$\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})\leq\mathrm{dist}(\theta,\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\frac{\delta_{q}^{\prime}}{3}+2r^{\varepsilon,\rho}(\theta)\leq\delta_{q}^{\prime}.$
It follows from Proposition 10.8 that $\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})\leq\frac{2}{c_{q}^{\prime}}r_{d}(\bar{\theta})$. We further obtain that
$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\frac{2}{c_{q}^{\prime}}r_{d}(\bar{\theta})+2r^{\varepsilon,\rho}(\theta)\leq\left(\frac{4\rho+5\kappa}{c_{q}^{\prime}}+2\right)r^{\varepsilon,\rho}(\theta).$
Next, for $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})>\frac{\delta_{q}^{\prime}}{3}$, by the first statement of the proposition, there exists $\lambda_{1}>0$ such that $r^{\varepsilon,\rho}(\theta)\geq\lambda_{1}$. With $\Theta$ being a compact set, let $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Then we obtain $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq 2\sqrt{q}\mu\leq\frac{2\sqrt{q}\mu}{\lambda_{1}}r^{\varepsilon,\rho}(\theta)$. Finally, for $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)>\lambda_{0}$, we simply have $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq 2\sqrt{q}\mu<\frac{2\sqrt{q}\mu}{\lambda_{0}}r^{\varepsilon,\rho}(\theta)$. Therefore, we obtain $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}{r}^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$ with $\eta_{r}=\max\left\{\frac{4\rho+5\kappa}{c_{q}^{\prime}}+2,\frac{2\sqrt{q}\mu}{\lambda_{1}},\frac{2\sqrt{q}\mu}{\min\{\lambda_{q},\delta_{q}^{\prime}/3\}}\right\}$, which completes the proof.∎
10.4 Verification of Assumption 1
We verify that Assumption 1 holds when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope by proving a stronger result in Claim 2 with the following lemma.
Lemma 10.12 (Cui et al. (2018), Proposition 2 (ii))
Let $\Theta\subseteq\mathbb{R}^{q}$ be a closed convex set. Let $f:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ be piecewise affine and $\varphi:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be convex. It holds that every d-stationary point of the composite function $\varphi\circ f$ on $\Theta$ is a local minimizer.
Recall the notations $\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)\coloneqq\mathop{\arg\min}_{\theta^{\prime}\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta^{\prime};\theta)+\frac{\rho}{2}\left\lVert\theta^{\prime}-\theta\right\rVert^{2}\right\}$ and $\bar{F}(t;\theta_{1},\theta_{2})\coloneqq F(\theta_{1}+t(\theta_{2}-\theta_{1}))$ for $\theta_{1},\theta_{2}\in\Theta$. In the analysis below, let $\partial(S)$ denote the boundary of $S$, $\mathrm{int}(S)$ denote the interior of $S$, and $\mathrm{rint}(S)$ denote the relative interior of $S$.
Claim 2
Let $\Theta\subseteq\mathbb{R}^{q}$ be a polytope and $\varphi(\cdot;\xi)$ be a strongly convex function with modulus $c_{\varphi}$ for any $\xi\in S$. Given $\varepsilon>0$ and $\rho>0$, there exist $\lambda_{q}>0$ and $c_{q}>0$ such that for any $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\coloneqq\underset{\theta^{\prime}\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)}{\max}\,\left\lVert\theta-\theta^{\prime}\right\rVert\leq\lambda_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(\cdot;\theta^{d},\theta)$ is convex on $[0,1]$ and
$\bar{F}(t;\theta^{d},\theta)\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{q}}{2}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2},\quad\forall\,t\in[0,1].$
Proof 10.13
Proof of Claim 2.
For $\theta\in\mathrm{cl}(\mathcal{S}^{d})$, the claim holds clearly with $\theta^{d}=\theta$. For $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$, it suffices to consider the case with $n=1$, for which we denote $F(\theta)=\varphi(f(\theta))$, where $f:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ is piecewise linear and $\varphi$ is strongly convex. The polytope $\Theta$ can be partitioned into a finite number of $q$-polytopes, denoted by $\mathfrak{C}\coloneqq\{\mathcal{C}_{i}\}_{i=1}^{N}$, such that $f$ is linear on each of them and the intersection of any two polytopes in $\mathfrak{C}$ is their common face. Given $\mathcal{C}\in\mathfrak{C}$, $f$ restricted to $\mathcal{C}$ can be represented as $f(\theta)=A\theta$ for some $A\in\mathbb{R}^{d\times q}$. Since the strongly convex function $\varphi$ has a unique minimizer on $\{A\theta\in\mathbb{R}^{d}\mid\theta\in\mathcal{C}\}$, the set of minimizers of $F$ on $\mathcal{C}$ denoted by $\mathcal{S}^{*}_{\mathcal{C}}$ can be represented as $\mathcal{S}^{*}_{\mathcal{C}}=[\{\bar{\theta}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}$ for any $\bar{\theta}\in\mathcal{S}^{*}_{\mathcal{C}}$, where $\mathcal{N}_{A}$ denotes the null space of $A$.
We first show that $\mathcal{S}^{*}_{\mathcal{C}}=\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$ if $\mathcal{S}^{d}\bigcap\mathcal{C}\neq\emptyset$. By Lemma 10.12, $\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}\subseteq\mathcal{S}^{*}_{\mathcal{C}}$. We show the reverse inclusion by considering the two cases $\mathcal{S}^{d}\bigcap\mathrm{int}(\mathcal{C})\neq\emptyset$ and $\mathcal{S}^{d}\bigcap\mathrm{int}(\mathcal{C})=\emptyset$ respectively. If there exists $\widehat{\theta}_{1}^{d}\in\mathcal{S}^{d}\bigcap\mathrm{int}(\mathcal{C})$, then for any ${\theta}^{\,\prime}\in[\{\widehat{\theta}_{1}^{d}\}+\mathcal{N}_{A}]\bigcap\mathrm{int}(\mathcal{C})$,
$F^{\prime}(\theta^{\prime};\theta-{\theta}^{\prime})=\varphi^{\prime}(A\widehat{\theta}_{1}^{d}+A({\theta^{\prime}}-\widehat{\theta}_{1}^{d});A(\theta-\widehat{\theta}_{1}^{d})+A(\widehat{\theta}_{1}^{d}-\theta^{\prime}))=F^{\prime}(\widehat{\theta}_{1}^{d};\theta-\widehat{\theta}_{1}^{d})\geq 0,\quad\forall\,\theta\in\Theta,$
which indicates $\mathcal{S}^{*}_{\mathcal{C}}=[\{\widehat{\theta}_{1}^{d}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}\subseteq\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$. If $\mathcal{S}^{d}\bigcap\mathrm{int}(\mathcal{C})=\emptyset$, we must have $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\partial(\mathcal{C})$. Let $\widehat{\theta}_{2}^{d}\in\mathcal{S}^{d}\bigcap{\mathcal{C}}\subseteq\mathcal{S}^{*}_{\mathcal{C}}$. Since $\mathcal{S}^{*}_{\mathcal{C}}=[\{\widehat{\theta}_{2}^{d}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}$ is a polyhedral set, it is contained in a face of $\mathcal{C}$. Therefore, we have either $\mathrm{rint}(\mathcal{S}^{*}_{\mathcal{C}})\subseteq\mathcal{C}^{\prime}$ or $\mathrm{rint}(\mathcal{S}^{*}_{\mathcal{C}})\bigcap\mathcal{C}^{\prime}=\emptyset$ for any $\mathcal{C}^{\prime}\in\mathfrak{C}$. Then for any $\theta^{\prime}\in\mathrm{rint}(\mathcal{S}^{*}_{\mathcal{C}})$, by restricting to the common polytope and noticing $\widehat{\theta}_{2}^{d}-\theta^{\prime}\in\mathcal{N}_{A}$, we obtain $F^{\prime}({\theta}^{\prime};\theta-{\theta}^{\prime})=F^{\prime}(\widehat{\theta}_{2}^{d};\theta-\widehat{\theta}_{2}^{d})\geq 0$ for any $\theta$ sufficiently close to $\theta^{\prime}$. This implies $\mathrm{rint}(\mathcal{S}^{*}_{\mathcal{C}})\subseteq\mathcal{S}^{d}\bigcap\mathcal{C}$ and further $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$. Combining the above two cases, we obtain that $\mathcal{S}^{*}_{\mathcal{C}}=\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$ if $\mathcal{S}^{d}\bigcap\mathcal{C}\neq\emptyset$.
We next show that for any $\mathcal{C}\in\mathfrak{C}$ with $\mathcal{S}^{d}\bigcap\mathcal{C}=\emptyset$, given $\varepsilon>0$ and $\rho>0$, there exists $\widehat{\lambda}_{q}>0$ such that $r^{\varepsilon,\rho}(\theta)>\widehat{\lambda}_{q}$ for any $\theta\in\mathcal{C}$. For such $\mathcal{C}\in\mathfrak{C}$, we have $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\partial{(\mathcal{C})}$ and $\mathcal{S}^{*}_{\mathcal{C}}\bigcap\mathcal{S}^{d}=\emptyset$. Since $\mathcal{S}^{d}\bigcap\mathrm{int}(\mathcal{C})=\emptyset$, by arguments above there exists a polytope $\mathcal{C}^{\prime}\in\mathfrak{C}$ such that $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\mathcal{C}^{\prime}$ and $\min_{\theta\in\mathcal{S}^{*}_{\mathcal{C}}}\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}^{\prime}})>0$. Let $I_{\mathcal{C}^{\prime}}$ denote the index mapping of $f$ corresponding to $\mathcal{C}^{\prime}$. Due to the Lipschitz continuity of $F$ on $\Theta$, for any $\varepsilon>0$ there exist $\delta>0$ and $\Delta>0$ such that any $\theta\in\mathcal{C}$ with $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})\leq\delta$ satisfies $I_{\mathcal{C}^{\prime}}\in\mathcal{I}^{\varepsilon}(\theta)$ and $F(\theta)>\min_{\theta^{\prime}\in\mathcal{C}^{\prime}}F(\theta^{\prime})+\Delta$; consequently, for any given $\rho>0$, there exists $\lambda_{1}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{1}$. On the other hand, for any point $\theta\in\mathcal{C}$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})>\delta$, by the first statement of Proposition 10.6, there exists $\lambda_{2}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{2}$. Combining the above two cases, we obtain the desired statement. It yields that there exists $\lambda_{q}>0$ such that for any $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq{\lambda}_{q}$, we have $\mathcal{S}^{d}\bigcap\mathcal{C}\neq\emptyset$ and $\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}=\mathcal{S}^{*}_{\mathcal{C}}$ for any $\mathcal{C}\in\mathfrak{C}_{\theta}\coloneqq\{\mathcal{C}^{\prime}\in\mathfrak{C}\mid\theta\in\mathcal{C}^{\prime}\}$.
Now, consider a point $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$ satisfying $r^{\varepsilon,\rho}(\theta)\leq{\lambda}_{q}$. Given a polytope $\mathcal{C}\in\mathfrak{C}_{\theta}$ denoted by $\mathcal{C}\coloneqq\{\theta\in\mathbb{R}^{q}\mid B\theta\leq b\}$, since $\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}=\mathcal{S}^{*}_{\mathcal{C}}$, we also have $\theta\in\mathcal{C}\backslash\mathcal{S}^{*}_{\mathcal{C}}$. With $f(\theta)\coloneqq A\theta$ over $\mathcal{C}$ for some matrix $A$ and $a^{*}\coloneqq A\theta^{*}$ for $\theta^{*}\in\mathcal{S}^{*}_{\mathcal{C}}$, we obtain $\mathcal{S}^{*}_{\mathcal{C}}=\{\theta\in\mathbb{R}^{q}\mid B\theta\leq b,A\theta\leq a^{*},-A\theta\leq-a^{*}\}$. By compactness of $\mathcal{C}$ and Hoffman’s lemma (Hoffman 1952), there exists $c>0$ depending on $\mathcal{C}$ such that for any $\theta\in\mathcal{C}$, there exists $\theta^{d}\in\mathcal{S}^{*}_{\mathcal{C}}=\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$ satisfying
$\left\lVert\theta-\theta^{d}\right\rVert=\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})\leq c\,\big{(}\left\lVert(B\theta-b)_{+}\right\rVert+\left\lVert A\theta-a^{*}\right\rVert\big{)}=c\left\lVert A(\theta-\theta^{d})\right\rVert.$
Clearly, $\bar{F}(\cdot;\theta^{d},\theta)$ is convex on $[0,1]$. Moreover, since $\varphi$ is strongly convex, for any $t\in[0,1]$,
$\bar{F}(t;\theta^{d},\theta)=\varphi(f(\theta^{d}+t(\theta-\theta^{d})))$ $\geq\varphi(f(\theta^{d}))+\frac{c_{\varphi}}{2}\left\lVert f(\theta^{d}+t(\theta-\theta^{d}))-f(\theta^{d})\right\rVert^{2}$
$=\varphi(A\theta^{d})+\frac{c_{\varphi}}{2}t^{2}\left\lVert A(\theta-\theta^{d})\right\rVert^{2}$
$\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{\varphi}}{2c^{2}}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}.$
By the finiteness of $\mathfrak{C}$, we thus prove the desired claim.∎
11 Analysis of PADR-based ERM for Constrained SP
11.1 Asymptotic Consistency of Constrained PADR-based ERM
To establish the consistency result of the PADR-based ERM model for the constrained SP problem (5), we impose the following assumption on functions $\{\psi_{j}\}_{j\in[J]}$:
  1. (A5)
    For every $j\in[J]$, $\psi_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L_{\psi}$-Lipschitz continuous.
[…]
  1. (${}^{\prime}$)
    There exists a continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (11.1).
By Assumptions (A2) and (${}^{\prime}$), $f^{*}_{\gamma_{0}}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for some $L_{z}^{\prime}>0$ and $M_{z}^{\prime}>0$. We denote by ~${\widetilde{\mathcal{H}}}^{K}_{\mu,\gamma}\coloneqq{\mathcal{H}}_{\mu}^{K}\bigcap\left\{f\in\mathcal{F}\mid\psi_{j}(f(x^{s}))+\gamma\leq 0,\,s\in[n],\,j\in[J]\right\}$ and $f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ the feasible PADR class and solution to the constrained ERM problem (5) respectively. Then asymptotic consistency of the constrained PADR-based ERM model is presented below.
Theorem 11.1 (Asymptotic Consistency of Constrained PADR-based ERM)
Suppose that Assumptions (A1)-(A3) and (A5) hold, and that Assumption (${}^{\prime}$) holds for some $\gamma_{0}>\Gamma\coloneqq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}$. With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
$R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})\leq$ $R(f^{*}_{\gamma_{0}})+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}^{\prime}\bar{X}K^{-1/p}$
$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
Furthermore, with $\mu=\mathcal{O}(K^{1/p})$, $\gamma=\mathcal{O}(K^{-1/p})$, $\gamma_{0}=\mathcal{O}(K^{-1/p})$, and $K=\mathcal{O}(n^{p/(p+4)})$, suppose that there exists $\bar{\gamma}_{0}>0$ such that for any $\gamma\in[0,\bar{\gamma}_{0}]$, Assumption (${}^{\prime}$) holds and $f^{*}_{\gamma}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$. Then $R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})$ converges to the optimal cost of (5) in probability, as $n$ goes to infinity.
Proof 11.2
[…]
$R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$ $=R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
$\leq R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{S}(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
$\leq\underbrace{2\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert}_{\mathcal{E}_{1}}+\underbrace{R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})}_{\mathcal{E}_{2}},$
where the last inequality holds since ~$\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}\subseteq{\mathcal{H}}_{\mu}^{K}$. The first term $\mathcal{E}_{1}$ can be directly bounded by Proposition 2.2, that is, given $\delta>0$, with probability at least $1-\delta$ over the draw of $S$,
$\mathcal{E}_{1}\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
For the second term $\mathcal{E}_{2}$, we further decompose it by
$\mathcal{E}_{2}=\underbrace{R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f_{app,\gamma_{0}})}_{\mathcal{E}_{21}}+\underbrace{R(f_{app,\gamma_{0}})-R(f^{*}_{\gamma_{0}})}_{\mathcal{E}_{22}},$
where $f_{app,\gamma_{0}}\in\mathop{\arg\min}_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lVert f-f^{*}_{\gamma_{0}}\right\rVert_{\infty}$. Then for almost every $x\in\mathcal{X}$,
${\psi_{j}}(f_{app,\gamma_{0}}(x))+\gamma$ $={\psi_{j}}(f^{*}_{\gamma_{0}}(x))+\left({\psi_{j}}(f_{app,\gamma_{0}}(x))-{\psi_{j}}(f^{*}_{\gamma_{0}}(x))\right)+\gamma$
$\leq{\psi_{j}}(f^{*}_{\gamma_{0}}(x))+L_{\psi}\sqrt{d}\left\lVert f_{app,\gamma_{0}}-f^{*}_{\gamma_{0}}\right\rVert_{\infty}+\gamma$
$\leq{\psi_{j}}(f^{*}_{\gamma_{0}}(x))+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}+\gamma$
$<{\psi_{j}}(f^{*}_{\gamma_{0}}(x))+\gamma_{0},$
where the second inequality holds by Proposition 2.1. This implies that $\psi_{j}(f_{app,\gamma_{0}}(x^{s}))+\gamma\leq 0$ for any $s\in[n]$ and $j\in[J]$, i.e., $f_{app,\gamma_{0}}\in\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}$, with probability 1. It follows that $\mathcal{E}_{21}=\min_{f\in\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}R(f)-R(f_{app,\gamma_{0}})\leq 0$ with probability 1. Moreover, $\mathcal{E}_{22}$ can be bounded as
$\mathcal{E}_{22}\leq L_{\varphi}\sqrt{d}\left\lVert f_{app,\gamma_{0}}-f^{*}_{\gamma_{0}}\right\rVert_{\infty}\leq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}^{\prime}\bar{X}K^{-1/p}.$
By combining bounds of $\mathcal{E}_{1}$, $\mathcal{E}_{21}$, and $\mathcal{E}_{22}$, we derive the desired inequality in the theorem, and asymptotic consistency follows instantly.∎
11.2 Convergence of Enhanced SMM to B-stationarity
According to (Cui and Pang 2021), for a constrained problem with a B-differentiable objective, a B-stationary point is defined as follows.
Definition 11.3
We say that $\theta^{b}$ is a B-stationary point of the objective $F(\cdot)$ on a closed set $\bar{\Theta}\subseteq\Theta$ if $\theta^{b}\in\bar{\Theta}$ and $F^{\prime}(\theta^{b};v)\geq 0$ for any $v\in\mathcal{T}(\theta^{b};\bar{\Theta})$.
The B-stationarity coincides with d-stationarity if $\bar{\Theta}$ is convex. Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (5) are B-differentiable. With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (5) by $\mathcal{S}^{b}$.
With Assumptions 1-1 imposed in the context of the penalized problem (5), we can apply the enhanced SMM algorithm to solve (5) and obtain the same convergence results in terms of d-stationarity as in Theorems 4.4 and 4.5. To further obtain the convergence to B-stationarity, we intend to connect d-stationarity of (5) with B-stationarity of (5) by applying the following exact penalty result from Cui and Pang (2021). Note that by the finiteness of $S$ and compactness of $\Theta$, for any $\xi\in S$, $f(\cdot;\xi)$ is Lipschitz continuous with modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$ and uniformly bounded with $\left\lVert f(\cdot;\xi)\right\rVert_{\infty}\leq M_{f}\coloneqq L_{f}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, and $\varphi(\cdot;\xi)$ is Lipschitz continuous on $[-M_{f},M_{f}]^{d}$ with uniform modulus denoted by $L^{\prime}_{\varphi}$.
Lemma 11.4 (Cui and Pang (2021), Proposition 9.2.2)
Let $\Theta$ be a compact convex set. For $\gamma>0$, if there exists $l_{c}>0$ such that
$\sup_{\theta\in\Theta\setminus\bar{\Theta}}\left[\min_{v\in\mathcal{T}(\theta;\Theta);\left\lVert v\right\rVert\leq 1}G_{\gamma}^{\prime}(\theta;v)\right]\leq-l_{c},$ (24)
then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (5) is feasible to (5) and thus is a B-stationary point of (5).
Condition (24) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (5) are bounded away from 0. Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (5) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (5). Nonetheless, it is not easy to verify the technical condition (24) when $G_{\gamma}$ is in a composite structure. The following assumption on $\{\psi_{j}\}_{j\in[J]}$ provides a sufficient condition for (24):
  1. (B1)
    Given $\gamma>0$, there exist positive constants $c_{\psi}$ and $l_{\psi}$ such that for $z\in\mathbb{R}^{d}$, if $\psi_{j_{0}}(z)+\gamma\geq 0$ for some $j_{0}\in[J]$, then $\left\lVert z\right\rVert\geq c_{\psi}$ and $\psi_{j_{0}}^{\prime}\left(z;-\frac{z}{\left\lVert z\right\rVert}\right)\leq-l_{\psi}$.
Assumption (B1) is more intuitive than (24). For example, consider a decision-making problem with a constraint function $\psi(z)=\left\lVert z\right\rVert_{1}-C_{0}$, where $C_{0}>0$ represents the total (absolute) capacity. Then the constraint $\psi(z)+\gamma=\sum_{\iota=1}^{d}\left\lvert z_{\iota}\right\rvert-(C_{0}-\gamma)\leq 0$ with $\gamma<C_{0}$ satisfies (B1) with $c_{\psi}=(C_{0}-\gamma)/\sqrt{d}$ and $l_{\psi}=1$. Based on Lemma 11.4, we derive the following exact penalty result under Assumption (B1) by utilizing the piecewise linearity of the inner function.
Proposition 11.5
Suppose that Assumption (B1) holds, and that $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set with $0\in\Theta$ and $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Then condition (24) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$. Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (5) is a B-stationary point of problem (5).
Proof 11.6
Proof of Proposition 11.5. By Lemma 11.4, it suffices to show that condition (24) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$. For any $\xi\in S$ and $j\in[J]$, denote $G_{\gamma}(\theta;\xi,j)\coloneqq\max\left\{\psi_{j}(f(\theta;\xi))+\gamma,0\right\}$. For a point $\bar{\theta}\in\Theta$ infeasible to the constraint $G_{\gamma}(\theta)\coloneqq\mathbb{E}\left[\sum_{j=1}^{J}G_{\gamma}(\theta;\widetilde{\xi}_{n},j)\right]\leq 0$, there exist $\bar{j}\in[J]$ and $\bar{\xi}\in S$ such that $\psi_{\bar{j}}(f(\bar{\theta};\bar{\xi}))+\gamma>0$.
For any $\xi\in S$, we denote $\bar{z}=f(\bar{\theta};\xi)$ and $\bar{v}=-\frac{\bar{\theta}}{\left\lVert\bar{\theta}\right\rVert}$ for simplicity. Since $f(\cdot;\xi)$ is piecewise linear, there exists $A\in\mathbb{R}^{d\times q}$ dependent on $\xi$ and $\bar{\theta}$ such that $f(\bar{\theta};\xi)=A\bar{\theta}$ and $f(\cdot;\xi)^{\prime}(\bar{\theta};\bar{v})=A\bar{v}$. Let
$\mathcal{A}_{<}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma<0\},\ \mathcal{A}_{>}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma>0\},\text{ and }\mathcal{A}_{=}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma=0\}.$
For $j\in\mathcal{A}_{<}$, it is clear that $G_{\gamma}(\cdot;\xi,j)^{\prime}(\bar{\theta};\bar{v})=0$. By Assumption (B1), we have that for $j\in\mathcal{A}_{>}$,
$G_{\gamma}(\cdot;\xi,j)^{\prime}(\bar{\theta};\bar{v})=\psi_{j}^{\prime}(f(\bar{\theta};\xi);f(\cdot;\xi)^{\prime}(\bar{\theta};\bar{v}))=\psi_{j}^{\prime}\left(\bar{z};-\frac{\bar{z}}{\left\lVert\bar{\theta}\right\rVert}\right)\leq-\frac{c_{\psi}l_{\psi}}{\left\lVert\bar{\theta}\right\rVert}\leq-\frac{c_{\psi}l_{\psi}}{\sqrt{q}\mu},$
and that for $j\in\mathcal{A}_{=}$,
[…]
Since $\bar{j}\in A_{>}$, we obtain that
$\min_{v\in\mathcal{T}(\bar{\theta};\Theta);\left\lVert v\right\rVert\leq 1}G_{\gamma}^{\prime}(\bar{\theta};v)\leq G_{\gamma}^{\prime}(\bar{\theta};\bar{v})\leq\frac{1}{n}G_{\gamma}(\cdot;\bar{\xi},\bar{j})^{\prime}(\bar{\theta};\bar{v})\leq-\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu},$
which completes the proof.∎
Finally, from Proposition 11.5 and convergence results in Section 4, we summarize the convergence properties of the enhanced SMM with exact penalization in the theorem below, where Assumptions 1 and 2 are adjusted to surrogate families of $F(\cdot;\xi)$ and $\psi_{j}(f(\cdot;\xi))$ for any $\xi\in S$ and $j\in[J]$, and Assumption 1 is adjusted to $V(\cdot;\lambda)$. The proof is similar to those of Theorems 4.4 and 4.5 and thus is omitted here.
Theorem 11.7 (Convergence of Enhanced SMM with exact penalization)
Under Assumption 1 and the same settings of Proposition 11.5, for problem (5) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$. The following statements hold:
  1. (a)
    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (5) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (5).
  2. (b)
    With $\varepsilon>0$, under Assumptions 2 and 1, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\widetilde{\mathcal{O}}(T^{-\frac{1}{2}\min\{\frac{\alpha}{2},1\}})$. In particular, if $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ are piecewise affine and $\Theta$ is a polytope, the same convergence rate holds without the two assumptions.
  3. (c)
    With $\varepsilon=0$, functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ being piecewise affine, and $\Theta$ being a polytope, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\widetilde{\mathcal{O}}(T^{-\frac{1}{2}\min\{\frac{\alpha}{2},1\}})$; moreover, with $\eta>0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\mathcal{O}(T^{-\frac{1}{2}\min\{\alpha,1\}})$, which is up to a logarithmic factor when $\alpha=1$.
12 Additional Experimental Results
For the unconstrained Newsvendor problem in Section 6.1, we set $(c_{b},c_{h})=(5,5)$ and consider another demand model
$\overline{Y}^{\prime}(X)=4\sin(\pi X_{1})+\max\{16X_{2},-20X_{2}\}+10,$
where the sine function can be interpreted as the seasonal dependence of the demand on $X_{1}$, and the max-affine function represents another nonlinear trend with respect to $X_{2}$. In practice, we may distinguish seasonal features from the others and build the improved DR represented as the sum of multiple DRs that are constructed for separate groups of features. Thus, we employ the sum of two PADR(2,2) models and the sum of two RKHS-DRs as the improved DRs in this instance and denote them by PADR(2,2)+ and RKHS-DR+, respectively.
a. Basic Results. In Figure 1p, we report results with the training sample size $n$ increasing from 50 to 1000. All PADR and RKHS-DR methods except PADR(2,2) converge close to SIMOPT. Figure 1q shows the results of running PADR(4,4) and PADR(2,2)+ for 30 rounds each, where we report the mean cost of the top 15 with standard deviation for each trial. It can be seen that although PADR(4,4) has a stronger approximation ability, PADR(2,2)+, designed with prior information, obtains lower and stabler costs.
(p) Comparison with Benchmarks
(q) PADR(4,4) vs. PADR(2,2)+
Unconstrained Newsvendor Problem with $\overline{Y}^{\prime}$: Basic Results
b. Varying Feature Dimensions. We set sparse and dense models in the same way as in Section 6.1 and report the comparison results with RKHS-DR and RKHS-DR+ in Figure 1s. PADR(2,2)+ outperforms all the others in both cases. It is worth mentioning that PADR(2,2) is close to PADR(2,2)+ in high-dimensional cases, showing that the training difficulty gradually dominates compared to the value of prior information as the dimension grows. But RKHS-DR+ consistently performs better than RKHS-DR, and it performs better than PADR(2,2) when $p$ is small. As for the running time, PADR(4,4) and PADR(2,2)+ grow linearly with $p$ and are longer than RKHS-DR and RKHS-DR+ only when $p\geq 30$.
(r) Sparse Demand Model
(s) Dense Demand Model
Unconstrained Newsvendor Problem with $\overline{Y}^{\prime}$ and Varying Feature Dimensions

### v4

@@
 Zhang, Liu, and Zhao
+[…]
+[…]
 […]
@@
 Focusing on stochastic programming (SP) with covariate information, this paper proposes an empirical risk minimization (ERM) method embedded within a nonconvex piecewise affine decision rule (PADR), which aims to learn the direct mapping from features to optimal decisions.
-We establish the nonasymptotic consistency result of our PADR-based ERM model for unconstrained problems and asymptotic consistency result for constrained ones.
-To solve the nonconvex and nondifferentiable ERM problem, we develop an enhanced stochastic majorization-minimization algorithm and establish the asymptotic convergence to (composite strong) directional stationarity along with complexity analysis.
-We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational tractability.
-Our numerical study demonstrates the superior performance of PADR-based ERM methods compared to state-of-the-art approaches under various settings, with significantly lower costs, less computation time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
+We establish the nonasymptotic consistency result of our PADR-based ERM model for unconstrained problems, which illustrates the role of piece number in balancing the trade-off between the approximation and estimation errors.
+To solve the nonconvex and nondifferentiable ERM problem, we develop an enhanced stochastic majorization-minimization (ESMM) algorithm and establish the convergence to (composite strong) directional stationarity along with convergence rate analysis.
+We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational traceability.
+Our numerical study demonstrates the effectiveness of PADR-based ERM methods with comparisons to state-of-the-art approaches under various settings, with significantly lower costs, less computation time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
 […]
@@
 1 Introduction
-Pioneered by George Dantzig in the mid-1950s, stochastic programming (SP) provides a class of optimization models for decision-making under uncertainty by treating the uncertainty as a random variable whose probability distribution is usually unknown in practice.
-With the growing real-world applications and availability of large datasets, the field of SP has flourished with many theoretical and algorithmic advances (Birge and Louveaux 2011, Shapiro et al.
+With the fast-growing real-world applications and availability of large datasets, the field of stochastic programming (SP) has flourished with many theoretical and algorithmic advances (Birge and Louveaux 2011, Shapiro et al.
 2021), in which the empirical probability distribution based on historical scenarios plays a central role in approximating the true distribution.
-However, the empirical distribution may provide poor estimation when the random variable is related to covariate information (also referred to as contextual information or features) which can be observed before the decision is made.
-An unconstrained SP with the covariate information $X=x\in\mathbb{R}^{p}$ is formulated as follows,
-$\inf_{z\in\mathbb{R}^{d}}\quad\mathbb{E}_{Y}\left[\varphi({z};{Y})\mid X=x\right],$ (1)
+However, it is common in practice that some covariate information (also referred to as contextual information or features) that may affect the random variable could be observed before the decision is made.
+In such situations, the empirical probability distribution and classical SP methodologies ignoring the covariate information leads to the poor performance.
+Given the covariate information $X=x\in\mathbb{R}^{p}$, an unconstrained SP is formulated as follows,
+$\inf_{z\in\mathbb{R}^{d}}\quad\mathbb{E}_{Y}\left[\varphi({z},{Y})\mid X=x\right],$ (1)
 where $z\in\mathbb{R}^{d}$ is a decision vector, $\varphi:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}$ is a cost function, and $Y$ is a random variable supported by $\mathcal{Y}$ whose distribution is related to a feature vector $X$ supported by $\mathcal{X}\subseteq\mathbb{R}^{p}$.
-A simple instance is an inventory planning problem, in which $z$ represents order decisions for $d$ products, $\varphi$ represents the inventory cost, $Y$ represents uncertain demands, and $X$ represents covariate information such as seasonality, promotion, and weather.
+A simple instance is an inventory planning problem, in which $z$ represents order decisions for $d$ products, $\varphi$ represents the inventory cost, $Y$ represents uncertain demands, and $X$ represents covariate information such as seasonality, promotion, weather, past orders, etc.
 With historical scenarios of $(X,Y)$, a classical sample average approximation (SAA) model disregards the dependency of $Y$ on $X$ and thus may lead to inconsistent decisions; such deficiency is also observed in numerical results of Ban and Rudin (2018) and Bertsimas and Koduri (2021).
@@
 Solving problem (1) for any realization of $X$ can be generalized as seeking an optimal decision rule (DR) to the following problem
-$\inf_{f\in\mathcal{F}}\quad R(f)\coloneqq\mathbb{E}_{X,Y}\left[\varphi({f(X)};{Y})\right],$ (2)
+$\inf_{f\in\mathcal{F}}\quad R(f)\coloneqq\mathbb{E}_{X,Y}\left[\varphi({f(X)},{Y})\right],$ (2)
 where $\mathcal{F}$ contains all measurable functions mapping from the feature space $\mathcal{X}$ to the decision space $\mathbb{R}^{d}$.
@@
 In general, problem (2) is intractable due to the infinite-dimensional function space and the unknown joint distribution of $X$ and $Y$.
-A common approximation method is to construct an empirical risk minimization (ERM) problem based on a dataset $S\coloneqq\{(x^{s},y^{s})\}_{s=1}^{n}$ with the decision rule $f$ restricted to a parametric DR class $\mathcal{H}$ (also known as hypothesis class):
-$\inf_{f\in\mathcal{H}}\quad\widehat{R}_{S}(f)\coloneqq\frac{1}{n}\sum_{s=1}^{n}\varphi({f(x^{s})};{y^{s}}).$ (3)
-In this paper, we adopt the above DR-based ERM framework, recognizing that the efficacy of such a framework is influenced by the trade-off between approximation accuracy of $\mathcal{H}$ to the optimal DR $f^{*}$ and computational complexity of solving (3).
-For example, problem (3) with a convex cost function can be solved efficiently with $\mathcal{H}$ being the class of linear decision rules (LDR) (Beutel and Minner 2012, Ban and Rudin 2018), while the LDR has difficulty describing the potential nonlinear relation between decisions and features.
+It is common to consider an approximation method which constructs an empirical risk minimization (ERM) problem based on a dataset $\Xi\coloneqq\{(x^{s},y^{s})\}_{s=1}^{n}$ with the decision rule $f$ restricted to a parametric DR class $\mathcal{H}$ (also known as hypothesis class):
+$\inf_{f\in\mathcal{H}}\quad\widehat{R}_{\Xi}(f)\coloneqq\frac{1}{n}\sum_{s=1}^{n}\varphi({f(x^{s})},{y^{s}}).$ (3)
+In this paper, we adopt the above DR-based ERM framework, recognizing that the efficacy of such a framework is influenced by the trade-off between approximation accuracy of $\mathcal{H}$ to the optimal DR and computational complexity of solving (3).
+For example, problem (3) with a convex cost function and linear decision rules (LDR) can be solved efficiently (Beutel and Minner 2012, Ban and Rudin 2018), while the LDR has difficulty describing the potential nonlinear relation between decisions and features.
 On the other side, a DR class consisting of neural networks could have a strong approximation capability (Qi et al.
@@
 Therefore, this paper aims to address a critical question: what could be an amenable hypothesis class for the DR-based ERM framework so as to maintain both high reliability and computational efficiency for solving SP with covariate information?
-To this end, we employ in the ERM problem (3) a piecewise affine decision rule (PADR) that is broader than LDR but simpler than general nonlinear DRs.
+To this end, we employ a piecewise affine decision rule (PADR) in the ERM problem (3) that is broader than LDR but simpler than many general nonlinear DRs.
 According to Scholtes (2012), any continuous piecewise affine (PA) function $f:\mathcal{X}\rightarrow\mathbb{R}$ can be expressed as the difference of max-affine functions:
-$f(x;\theta)=\max_{k\in[K]}\left\{(\alpha^{k})^{\top}x+a^{k}\right\}-\max_{k\in[K]}\left\{(\beta^{k})^{\top}x+b^{k}\right\},$ (4)
+$f(x,\theta)=\max_{k\in[K]}\left\{(\alpha^{k})^{\top}x+a^{k}\right\}-\max_{k\in[K]}\left\{(\beta^{k})^{\top}x+b^{k}\right\},$ (4)
 where $\theta\coloneqq\left(\alpha^{k},a^{k},\beta^{k},b^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)}$ and $[K]\coloneqq\{1,\dots,K\}$.
+The PADR of such a form is beneficial because it requires minimal parameter tuning in multi-dimensional cases without the need to specify breakpoints for piecewise linear parts in each dimension, and it can asymptotically approximate any continuous functions according to Royset (2020).
 Throughout the theoretical analysis in this paper, we assume that the numbers of pieces in the two max-affine components of (4) are equal to simplify the exposition, which could be easily extended to the case with different numbers of pieces.
-With $\mathcal{H}$ being the class of continuous PA functions with each max-affine component having $K$ pieces and parameters $\{\theta_{\iota}\}_{\iota\in[d]}$ satisfying $\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu$ for $\iota\in[d]$, the ERM problem (3) is formulated as the following parameter estimation problem
-$\min_{\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\forall\iota\in[d]}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi\left(\Big{(}\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x^{s}+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x^{s}+b^{k}_{\iota}\right\}\Big{)}_{\iota=1}^{d};y^{s}\right).$ (5)
-In response to the aforementioned question, we intend to tackle two critical issues of the PADR-based ERM model in this paper: the consistency issue of the PADR-based ERM problem (5) and the computational issue for solving this large-scale, nonconvex, and nondifferentiable composite optimization problem.
-We address the first issue by quantifying the excess risk bound of our model, a standard procedure for model selection (Bartlett et al.
-2002) in machine learning, and the second issue by developing a sampling-based algorithm with convergence guarantees, through which we demonstrate the ability of our proposed PADR-based ERM approach to effectively solve a broad class of SP problems with covariate information.
+To address the aforementioned question, we intend to tackle two critical issues of the PADR-based ERM model for solving the unconstrained decision-making problem (1) in this paper: the consistency guarantees of the PADR-based ERM problem and the computational method for efficiently solving a large-scale, nonconvex, and nondifferentiable composite optimization problem.
+It is more challenging to extend the data-driven decision rule methods for constrained decision-making problems.
+Motivated by the scenario approximation method in Calafiore and Campi (2005), a straightforward approach is to impose the strengthened feasibility requirement over the observed scenarios in learning the data-driven decision rule.
+In Appendix 9, we illustrate that the obtained decision rule at a new scenario is feasible with high probability and is asymptotically consistent with the original SP problem, and we also discuss several special cases where the feasibility of the decision rule can be guaranteed.
+In the rest of this section, we give a brief literature review and present the distinguished features of the proposed PADR method over other relative work.
+We will also summarize major contributions of the present paper.
 1.2 Literature review
-We present a literature review of data-driven methods for SP with covariate information and algorithmic advances in solving nonconvex optimization problems.
-1.2.1 SP with Covariate Information.
 There are two main schemes of data-driven approaches in the literature of SP dealing with the latent dependency on covariate information: the sequential and integrated schemes.
@@
 (2009), Donti et al.
-(2017), and Elmachtoub and Grigas (2021) propose modifications of the loss function by accounting for the cost error of the downstream optimization problem.
-Since those modified loss functions involve optimal solutions to the decision-making problems that are parameterized by suspicious prediction models, such frameworks usually lead to bi-level nonconvex optimization problems and thus encounter substantial computational challenges in dealing with general constrained nonlinear programs.
+(2017), Elmachtoub and Grigas (2021), and Qi et al.
+(2021a) propose modifications of the loss function by considering the cost error of the downstream optimization problem.
+Since those modified loss functions involve optimal solutions to the decision-making problems that are parameterized by prediction models, such frameworks usually lead to bi-level nonconvex optimization problems and thus encounter substantial computational challenges in dealing with nonlinear constrained programs.
 Approaches in the integrated scheme yield decisions via one-step data-driven optimization.
-Bertsimas and Kallus (2019) and Kallus and Mao (2022) approximate the conditional expected cost in (1) by a reweighted empirical cost.
-However, as claimed by Bertsimas and Koduri (2021), this framework is highly affected by the data with features close to the current observation and thus suffers from the curse of dimensionality.
-The DR-based ERM framework presented in (3) is another one-step approach in the integrated scheme.
-Due to its functional approximation in nature, this approach is relieved from the aforementioned issue.
+Prescriptive methods by Bertsimas and Kallus (2019) and Kallus and Mao (2023) approximate the conditional expected cost in (1) through reweighing empirical costs.
+Elmachtoub et al.
+(2023) point out that, when the model class is well-specified and there is sufficient data, methods that directly estimate the conditional distribution outperform those that consider the downstream optimization problem; in a misspecified scenario, the reverse holds.
+However, as claimed by Bertsimas and Koduri (2021), the prescriptive methods are highly affected by the data with features close to the current observation and thus suffers from the curse of dimensionality.
+The DR-based ERM framework presented in (3), which is another one-step approach in the integrated scheme, is relieved from the aforementioned issue due to its functional approximation in nature.
+For constrained SP with general nonlinear constraints, the current study of decision rule methods could only guarantee feasibility with high probability, and the predict-then-optimize framework may be more suitable for solving the contextual problems when the feasibility is essential to decision-makers.
+Several works on DR-based ERM methods are most relevant to this paper.
+Ban and Rudin (2018) employ LDR for the Newsvendor problem with nonasymptotic optimality guarantees under the assumption of the linear demand model.
+Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems.
+Compared to the above existing DR methods, given the asymptotic approximation of PA functions to any continuous functions, the PADR method is applicable to more general decision-making problems with nonconvex cost function or nonlinear dependency with asymptotic optimality.
 Besides the classical LDR, several generalized LDRs are proposed for standard SP problems which lift random vectors into a higher-dimensional probability space to describe nonlinear relations (Chen et al.
@@
 2015).
-However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, except for the posterior estimation provided in Kuhn et al.
-(2011), their approximation accuracy lacks theoretical guarantees.
-The PADR with the parametric formula (4) employed in Bertsimas and Georghiou (2015) and this paper requires no prior designs of breakpoints as in Georghiou et al.
-(2015); instead, the breakpoints are implicitly determined by parameters in (4).
-Unlike the mixed-integer formulation in Bertsimas and Georghiou (2015) for multi-stage adaptive optimization problems, we treat the PADR-based ERM problem (5) as a special class of nonconvex and nondifferentiable optimization problems and exploit algorithmic advances to accommodate the large sample size and high-dimensional feature space.
-Several works on DR-based ERM methods are most relevant to this paper.
-Ban and Rudin (2018) employ LDR for the Newsvendor problem with nonasymptotic optimality guarantees under the assumption of the linear demand model.
-Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems.
-Compared to the above references, our PADR-based ERM model is suitable for a much broader class of SP problems that can be nonconvex, and by following the statistical learning theory (Vapnik 1999, Mohri et al.
-2018), we establish the nonasymptotic consistency result of our model without any parametric assumptions on the optimal decision rule or the latent dependency of the random variable $Y$ on the feature $X$.
-1.2.2 Difference-of-convex (DC) and Composite DC Programming.
-Due to the finite-sum composite structure of the objective in (5) with potentially large sample size $n$ and feature dimension $p$, it is prohibitively challenging to obtain a global minimizer to such a nonconvex and nondifferentiable optimization problem.
-Thus, our goal is to compute a promising candidate for a local minimizer via a sampling-based algorithm.
-When the cost function $\varphi(\cdot;y)$ is piecewise affine (e.g., the Newsvendor cost), the overall composite objective of (5) is also piecewise affine, which is a special DC function.
-Deterministic DC algorithms have been extensively studied for solving DC programs over the last several decades (Le Thi and Pham Dinh 2018).
-When the concave part of the DC decomposition is nondifferentiable, the basic DC algorithm converges to a critical point that satisfies the necessary condition for local optimality.
-To obtain a d(irectional)-stationary point that is the sharpest first-order stationary point for general nonconvex problems, Pang et al.
-(2016) study a class of nondifferentiable DC programs whose concave part is the maximum of a finite number of smooth convex functions and propose an enhanced DC algorithm which iteratively solves convex subproblems constructed by linearizing the $\varepsilon$-active piece of the concave part.
-Though sampling-based algorithms for stochastic DC programs have been recently studied in Nitanda and Suzuki (2017), Xu et al.
-(2019), and Le Thi et al.
-(2022) addressing smooth or partially smooth objectives, there still lacks a comprehensive study of the enhanced surrogation technique in stochastic DC algorithms to seek d-stationary solutions to the stochastic DC programs with nondifferentiable concave parts.
-With a more general cost function that is not necessarily piecewise affine, the finite-sum composite structure of the objective lies at the cornerstone of algorithmic development for solving the PADR-based ERM problem (5).
-Cui et al.
-(2018) study the ERM problem with a convex composite DC structure and develop a majorization-minimization (MM) algorithm with an enhanced surrogation technique similar to Pang et al.
-(2016); however, such a deterministic algorithm requires iteratively solving finite-sum convex subproblems over all the data, raising concerns about its computational efficiency for solving large-scale ERM problems.
-The composite DC problem (5) can be regarded as a special class of compound SP studied in Liu et al.
-(2022) with finite support and solved by the stochastic majorization-minimization (SMM) algorithm proposed by the authors.
-Nonetheless, without the enhanced surrogation technique, SMM can only asymptotically compute a compound critical point when the inner concave function is nondifferentiable.
-Finally, we refer readers to Cui and Pang (2021) for a comprehensive study of modern nonconvex and nondifferentiable optimization, especially Chapters 6 and 7 for the study of stationary concepts and surrogation-based algorithms for nonconvex composite optimization problems.
+However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, they usually lack theoretical guarantees on approximation accuracy except for the posterior estimation provided in Kuhn et al.
+(2011).
+Regarding solving the PADR-based ERM problem with potentially large sample size $n$, it is prohibitively challenging to obtain a global minimizer to such a coupled nonconvex and nondifferentiable optimization problem.
+As provided by Bertsimas and Georghiou (2015, Section 2.4), the PADR-based ERM problem can be formulated as a mixed-integer nonlinear program using Big-M method with additional $2ndK$ binary variables to describe the PADR of interest.
+For a PADR-based ERM problem of the modest size with data size $n=1000$, the obtained mixed-integer nonlinear program could have $10^{4}\sim 10^{5}$ binary variables and the state-of-the-art solvers are hard to solve within a few hours to achieve an approximately optimal solution.
+Instead of seeking global optimality, majorization-minimization (MM) methods provide a unified framework that is practically efficient and theoretically convergent for solving the PADR-based ERM problem with the finite-sum composite objective.
+MM methods iteratively construct convex subproblems with an upper surrogate objective function for obtaining “stationary” solutions.
+According to variational analysis by Rockafellar and Wets (2009), there are many kinds of stationary concepts which differ in the ability to characterize necessary optimality conditions for local/global optimality.
+As pointed in Cui and Pang (2021, Chapter 7), MM algorithms may converge to different types of stationary solutions depending on the properties of surrogate and objective functions.
+Algorithms with convergence to a stronger stationarity is favorable not only theoretically but also computationally because the obtained solution from numerical implementation with random initialization has the greater potential to be a global optimum.
+Difference-of-convex (DC) algorithms, which is a special MM method, have been extensively studied over the last several decades (Le Thi and Pham Dinh 2018) and can compute a d(irectional)-stationary point for DC programs with a differentiable concave component.
+When the cost function $\varphi(\cdot,y)$ is convex or piecewise affine, the PADR-based ERM problem can be formulated as a nondifferentiable DC program for which DC algorithms including Tao and An (1997) and Lipp and Boyd (2016) are applicable.
+However, due to the nondifferentiable concave component, DC algorithms for the ERM problem converge to a critical point whose defining condition is dependent on the dc decomposition form and is weaker than d(irectional)-stationarity.
+For nondifferentiable DC or composite DC problems, Pang et al.
+(2016) and Cui et al.
+(2018) propose an enhanced surrogation technique that iteratively solves convex subproblems that are constructed based on $\varepsilon$-active pieces of the concave component.
+However, the enhanced MM algorithm for large-scale ERM problems requires iteratively solving finite-sum convex subproblems over all the data, which raises the fundamental issue of computational efficiency.
+Sampling-based algorithms for stochastic DC programs have been recently studied in Nitanda and Suzuki (2017), Xu et al.
+(2019), Le Thi et al.
+(2022) addressing smooth or partially smooth objectives, and in Liu et al.
+(2022) addressing nonconvex and nonsmooth compound SP to compute critical solutions.
+There still lacks a comprehensive study on whether the $\varepsilon$-active set technique is amenable for sampling-based methods to enhance the convergence properties for nondifferentiable stochastic DC programs and PADR-based ERM problems.
 1.3 Contributions
@@
 1.
-We propose a PADR-based ERM method for a general class of SP problems with covariate information, where both cost and constraint functions can be nonconvex.
-We quantify the approximation accuracy of the PA function class and establish the nonasymptotic consistency result of our framework for unconstrained problems and asymptotic consistency result for deterministic-constrained problems without any parametric assumptions on the optimal decision rule or the latent dependency.
+We propose a PADR-based ERM method for a general class of SP problems with covariate information, where the cost function could be nonconvex and nondifferentiable.
+We quantify the approximation accuracy of the PA function class to any Lipschitz continuous function, which to the best of our knowledge is the first quantification result for such a general case.
+Without any parametric assumption on the optimal decision rule or the latent dependency, we establish the nonasymptotic consistency guarantee with the excess risk bound of our proposed framework, which illustrates the role of piece number in balancing the approximation and estimation errors.
 2.
-For learning PADR from the ERM problem, we develop an enhanced SMM algorithm with asymptotic convergence to composite strong d-stationarity, which is a sharper stationary concept than d-stationarity specialized to the composite DC programs with PA inner functions.
-We further provide nonasymptotic convergence of the algorithm to d-stationarity, which to the best of our knowledge is the first complexity result for stochastic nonsmooth composite DC programs.
-The enhanced SMM algorithm is also extended with exact penalization for solving the PADR-based ERM problem with constraints.
-3.
-We test our method through numerical experiments on the Newsvendor problem and its variations with comparisons to the PO methods and other prominent DR-based methods, including the (generalized) LDR and RKHS-DR.
-The results demonstrate the superior performance of our method with significantly lower costs and less computation time in most instances, particularly those with high-dimensional redundant features or highly nonlinear models.
+For efficiently solving the PADR-based ERM problem, we develop an enhanced sampling-based MM (ESMM) algorithm with the natural combination of sequential sampling and the enhanced technique over the $\varepsilon$-active set.
+With sequential sampling, the ESMM algorithm effectively reduces computational costs by solving subproblems of small sizes in the early iterations.
+We show that under the proper control of increasing sample rate, ESMM achieves the asymptotic convergence in terms of composite strong d-stationarity, which is strictly stronger than d-stationarity and thus has greater potential to be a global minimizer.
+Numerical results show that the ESMM algorithm could speed up to 4 times faster than the basic MM method.
+We also would like to highlight several distinguishing features of the proposed PADR method for contextual SP: (a) from the modeling aspect, the PADR method is suitable for contextual SP problems with nonlinear or nonconvex cost functions.
+All other relevant works on contextual SP with asymptotic optimality are restricted to linear programs or convex programs.
+For instance, the RKHS-based DR method by Bertsimas and Koduri (2021) requires convexity of the cost function for asymptotic optimality.
+The PADR method opens up the application to nonconvex contextual SP problems; (b) from the theoretical aspect, PADR provides a compact structure to asymptotically approximate non-parametric decision rule with minimal hyperparameter tuning.
+Though PA hypothesis class is not new in classical SP, we provide the first quantification of the approximation error and the excess risk bound of the data-driven PADR model for general nonconvex SP problems.
+Except for the LDR method, all other methods for contextual SP do not provide the excess risk bound, including the RKHS-based DR in Bertsimas and Koduri (2021) and Notz and Pibernik (2021); (c) regarding to the computational aspect, the proposed ESMM algorithm is developed for efficiently solving the PADR-based ERM problem of the large scale.
+Under the suitable sampling scheme, we show that the ESMM method significantly improves computational efficiency and enhances the convergence property compared to the existing MM methods in the literature; (d) moreover, the PADR method possesses several prominent numerical advantages.
+Numerous experiments indicate that it requires significantly fewer data points for satisfactory performance, thus is particularly suitable for cases with insufficient data sets.
+Additionally, it yields superior performance in terms of the test costs with both synthetic and real data sets when the feature dimension is less than 20.
+The nonlinearity level of the demand model has almost no effect on the performance of the PADR method, while other methods suffer significantly from the high level of nonlinearity.
 The rest of this paper is organized as follows.
 In Section 2, we provide the theoretical consistency result of the PADR-based ERM model.
-We then develop the enhanced SMM algorithm for learning PADR with surrogation construction of composite PA objectives in Section 3 and present the convergence analysis in Section 4.
-In Section 5, we extend our framework to deterministic-constrained SP problems.
-In Section 6, we report numerical results.
+We develop the ESMM algorithm for learning PADR and present the convergence analysis in Section 3.
+In Section 4 we report numerical results.
 We end with conclusions in Section 5.
-2 Theoretical Guarantees for PADR-based ERM
-In this section, we provide theoretical guarantees of the PADR-based ERM model by first quantifying the approximation error of the PA function class to Lipschitz continuous functions and then presenting the nonasymptotic consistency result of the PADR-based ERM model via the excess risk bound.
-At the end of this section, we compare our theoretical results with prominent ERM models employing the LDR and RKHS-DR.
-2.1 Approximation Accuracy of Piecewise Affine Function Class
-PA functions are known to asymptotically approximate Lipschitz continuous functions.
-Balazs et al.
+2 Consistency Guarantees for PADR-based ERM
+In this section, we aim to derive the consistency guarantees for the PADR-based ERM problem.
+Considering problem (3) equipped with a class of continuous PA functions with each max-affine component having $K$ pieces and parameters $\{\theta_{\iota}\}_{\iota\in[d]}$ satisfying $\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu$ for $\iota\in[d]$, we obtain the following ERM problem:
+$\min_{\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\forall\iota\in[d]}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi\left(\Big{(}\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x^{s}+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x^{s}+b^{k}_{\iota}\right\}\Big{)}_{\iota=1}^{d},y^{s}\right).$ (5)
+Excess risk is commonly used to quantify the performance of the ERM model in machine learning, which can be decomposed as the approximation error for the hypothesis class and the estimation error for the empirical risk.
+Intuitively, when the number of pieces in the PADR class increases, the decision rule class becomes more complex, which is expected to reduce the approximation error while increasing the estimation error from the empirical risk.
+In this section, assuming the existence of an optimal decision rule that is Lipschitz continuous, we provide consistency guarantees of the PADR-based ERM model (5) by quantifying the approximation and estimation errors, which sheds light upon the role of the number of pieces and sample size in balancing the two errors in the excess risk.
+To the best of our knowledge, the approximation error bound of the PA function class established in the succeeding section is the first result for Lipschitz continuous functions in the literature.
+At the end of this section, we discuss the excess risk in practical implementation and compare our theoretical results with other prominent ERM models employing the LDR and RKHS-DR.
+2.1 Approximation error of piecewise affine function class
+The approximation errors of the PA function class to convex and DC functions are quantified in Balazs et al.
 (2015) and Siahkamari et al.
-(2020) quantify the approximation errors of the PA function class to convex and DC functions, respectively.
-We provide the approximation bound of the PA function class to Lipschitz continuous functions, which to the best of our knowledge is the first quantification result for such a general case.
-Specifically, with $L>0$ and $M>0$, let
+(2020) respectively.
+PA functions which are of “nonparametric” nature are known to asymptotically approximate Lipschitz continuous functions (Royset 2020); nevertheless, the approximation error of PA functions with finite pieces remains unsolved.
+For the class of Lipschitz continuous functions that allow a more general setting for SP problems of interest, we provide the approximation error bound of the PA function class, which to the best of our knowledge is the first quantification result for such a general case.
+With $L>0$ and $M>0$, let
 […]
 where $\left\lVert f\right\rVert_{\infty}\coloneqq\sup_{x\in\mathcal{X}}\max_{\iota\in[d]}|f_{\iota}(x)|$. We define the PADR class ${\mathcal{H}}_{\mu}^{K}$ as
-$\mathcal{H}^{K}_{\mu}\coloneqq\left\{\left(f_{\iota}(\,\cdot\,;\theta_{\iota}):\mathcal{X}\rightarrow\mathbb{R}\right)_{\iota=1}^{d}\ \middle|\ \begin{aligned} &f_{\iota}(x;\theta_{\iota})=\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x+b^{k}_{\iota}\right\},\\ &\theta_{\iota}\coloneqq\left(\alpha_{\iota}^{k},a_{\iota}^{k},\beta_{\iota}^{k},b_{\iota}^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)},\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\,\forall\,\iota\in[d]\end{aligned}\ \right\}.$ (6)
-When $K=1$, $\mathcal{H}^{1}_{\mu}$ reduces to the LDR class. We provide the approximation bound of the PA function class to Lipschitz continuous functions below with the proof given in Appendix 8.1.
-Proposition 2.1 (Approximation Bound of ${\mathcal{H}}_{\mu}^{K}$)
+$\mathcal{H}^{K}_{\mu}\coloneqq\left\{\left(f_{\iota}(\,\cdot\,,\theta_{\iota}):\mathcal{X}\rightarrow\mathbb{R}\right)_{\iota=1}^{d}\ \middle|\ \begin{aligned} &f_{\iota}(x,\theta_{\iota})=\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x+b^{k}_{\iota}\right\},\\ &\theta_{\iota}\coloneqq\left(\alpha_{\iota}^{k},a_{\iota}^{k},\beta_{\iota}^{k},b_{\iota}^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)},\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\,\forall\,\iota\in[d]\end{aligned}\ \right\}.$ (6)
+When $K=1$, $\mathcal{H}^{1}_{\mu}$ reduces to the LDR class. The following proposition provides the bound for the approximation error of the PA function class to Lipschitz continuous functions.
+Proposition 2.1 (Approximation error bound of ${\mathcal{H}}_{\mu}^{K}$)
 Suppose that $\mathcal{X}\subseteq\mathbb{R}^{p}$ is a compact set with $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}=\bar{X}$.
@@
 $\min_{f\in\mathcal{H}_{\mu}^{K}}\left\lVert f-f_{0}\right\rVert_{\infty}\leq 2(p^{1/2}+3)p^{1/2}L_{0}\bar{X}K^{-1/p}.$
-The bound $\mathcal{O}(K^{-1/p})$ provided in Proposition 2.1 is weaker than $\mathcal{O}(K^{-2/p})$ given by Balazs et al.
+The idea is that for any target function $f_{0}\in\mathcal{F}_{L_{0},M_{0}}$ we construct an interpolation PA function motivated by Siahkamari et al.
+(2020).
+The difference is that with the absence of convexity, the parameters that define the interpolation functions are carefully chosen to ensure that the Lipschitz continuity modulus of the interpolation function is independent of the approximation accuracy.
+The constructive proof in detail is given in Online Appendix 6.1.
+From Proposition 2.1, the approximation error for Lipschitz continuous functions is bounded by $\mathcal{O}(K^{-1/p})$, which is weaker than the approximation error $\mathcal{O}(K^{-2/p})$ in approximating convex and DC functions by Balazs et al.
 (2015) and Siahkamari et al.
-(2020) in approximating convex and DC functions.
-Regardless, our bound is applicable to general PA-based statistical learning problems in which the underlying true model may not be convex or DC; more importantly, it renders the succeeding nonasymptotic consistency result of the PADR-based ERM model without any parametric or convex assumptions on the optimal decision rule or latent dependency.
+(2020).
+However, it applies to general PA-based statistical learning problems in which the underlying true model is not necessarily convex or DC.
+Moreover, it renders the succeeding nonasymptotic consistency result of the PADR-based ERM model without any parametric or convex assumptions on the optimal decision rule or latent dependency.
 2.2 Nonasymptotic consistency of PADR-based ERM
-We impose the following assumptions for the consistency analysis:
+We impose the following assumptions for the nonasymptotic consistency analysis:
   1. (A1)
-    Samples in $S=\{(x^{s},y^{s})\}_{s=1}^{n}$ of $(X,Y)$ are independently and identically distributed.
+    Samples in $\Xi=\{(x^{s},y^{s})\}_{s=1}^{n}$ of $(X,Y)$ are independently and identically distributed.
   2. (A2)
-    For some $\bar{X}>0$, the support set $\mathcal{X}\subseteq\mathbb{R}^{p}$ of $X$ is a compact set with $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}=\bar{X}$.
+    The support set $\mathcal{X}\subseteq\mathbb{R}^{p}$ of $X$ is a compact set with $\bar{X}\coloneqq\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}$.
   3. (A3)
-    For some $L_{\varphi}>0$, $C_{\varphi}>0$, $z_{0}\in\mathbb{R}^{d}$, and any $y\in\mathcal{Y}$, the cost function $\varphi({\cdot};{y})$ is Lipschitz continuous on $\mathbb{R}^{d}$ with modulus $L_{\varphi}$ and satisfies $\left\lvert\varphi(z_{0};y)\right\rvert\leq C_{\varphi}$.
+    For some $L_{\varphi}>0$, $C_{\varphi}>0$, $z_{0}\in\mathbb{R}^{d}$, and any $y\in\mathcal{Y}$, the cost function $\varphi({\cdot},{y})$ is Lipschitz continuous on $\mathbb{R}^{d}$ with modulus $L_{\varphi}$ and satisfies $\left\lvert\varphi(z_{0},y)\right\rvert\leq C_{\varphi}$.
   4. (A4)
@@
 Under Assumption (A2), the optimal DR $f^{*}$ in Assumption (A4) is bounded on $\mathcal{X}$, and thus $f^{*}\in\mathcal{F}_{L_{z},M_{z}}$ for some $L_{z}>0$ and $M_{z}>0$.
-With the optimal solution to (5) denoted by $f^{*}_{S,{\mathcal{H}}_{\mu}^{K}}$, the consistency of the PADR-based ERM model is evaluated by the excess risk $R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\underset{f\in\mathcal{F}}{\inf}\,R(f)$, which is the gap between the expected cost of the optimal empirical PADR and the minimal cost.
-With problem (2) restricted to ${\mathcal{H}}_{\mu}^{K}$, we decompose the excess risk into the approximation and estimation errors as follows,
-$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)}_{\text{estimation error}}+\underbrace{\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)-\inf_{f\in\mathcal{F}}R(f)}_{\text{approximation error}}.$ (7)
+Let $f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}}$ denote the optimal solution to the PADR-based ERM model (5).
+The consistency of the PADR-based ERM model is evaluated by the excess risk $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\underset{f\in\mathcal{F}}{\inf}\,R(f)$, which is the gap between the expected cost of the optimal empirical PADR and the minimal cost.
+With $\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)$ as an intermediate problem, the excess risk can be decomposed into the approximation and estimation errors as follows,
+$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)}_{\text{estimation error}}+\underbrace{\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)-\inf_{f\in\mathcal{F}}R(f)}_{\text{approximation error}}.$ (7)
 […]
 Proposition 2.2 (Uniform generalization bound of PADR-based ERM)
-Under Assumptions (A1)-(A3), with $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
-$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert\leq\frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
-The proof based on the Rademacher complexity theory is presented in Appendix 8.2.
-Combining the approximation and estimation error bounds, we obtain the excess risk bound of PADR-based ERM in Theorem 2.3 with the proof in Appendix 8.2.
+Under Assumptions (A1)-(A3), with $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
+$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq\frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
 Theorem 2.3 (Excess risk bound of PADR-based ERM)
-Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
-$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})\,\leq\,$ $\inf_{f\in\mathcal{F}}R(f)+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}\bar{X}K^{-1/p}$
+Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, with probability at least $1-\delta$ over the draw of $\Xi$, we have
+$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})\,\leq\,$ $\inf_{f\in\mathcal{F}}R(f)+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}\bar{X}K^{-1/p}$
 $+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
-With $\mu=\mathcal{O}(K^{1/p})$, the complexity of the excess risk bound with respect to the sample size $n$ and parameter $K$ is
-$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{\widetilde{\mathcal{O}}\left(K^{\frac{p+2}{2p}}n^{-\frac{1}{2}}\right)}_{\begin{subarray}{c}\text{estimation}\\ \text{error bound}\end{subarray}}+\underbrace{\mathcal{O}\left(K^{-\frac{1}{p}}\right)}_{\begin{subarray}{c}\text{approximation}\\ \text{error bound}\end{subarray}},$ (8)
+The above theorem implies that with $\mu=\mathcal{O}(K^{1/p})$, the complexity of the excess risk bound with respect to the sample size $n$ and parameter $K$ is
+$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{\widetilde{\mathcal{O}}\left(K^{\frac{p+2}{2p}}n^{-\frac{1}{2}}\right)}_{\begin{subarray}{c}\text{estimation}\\ \text{error bound}\end{subarray}}+\underbrace{\mathcal{O}\left(K^{-\frac{1}{p}}\right)}_{\begin{subarray}{c}\text{approximation}\\ \text{error bound}\end{subarray}},$ (8)
 where $\widetilde{\mathcal{O}}(\cdot)$ suppresses logarithmic factors.
@@
 As $K$ grows, the PA function class becomes larger, so the approximation error decreases due to the stronger approximation ability of ${\mathcal{H}}_{\mu}^{K}$, and the estimation error increases due to the more pronounced overfitting issue.
-By setting $K=\mathcal{O}(n^{p/(p+4)})$, we derive the optimal excess risk bound $\widetilde{\mathcal{O}}(n^{-{1}/(4+p)})$.
+With an appropriate choice of $K$ that balances these two errors, we can derive the optimal excess risk bound.
+Furthermore, with the existence of a convex or DC optimal DR, this bound can be improved based on the approximation bounds $\mathcal{O}(K^{-2/p})$ of ${\mathcal{H}}_{\mu}^{K}$ from Balazs et al.
+(2015) and Siahkamari et al.
+(2020).
+The extensions are summarized in the following corollary.
+Corollary 2.4
+Under the settings in Theorem 2.3, with $K=\mathcal{O}(n^{p/(p+4)})$, the optimal excess risk bound is $\widetilde{\mathcal{O}}(n^{-{1}/(4+p)})$.
+If there exists a convex or DC optimal decision rule to problem (2), the optimal excess risk bound is $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$ with $K=\mathcal{O}(n^{p/(p+6)})$.
+We make a remark that Proposition 2.1 provides a bound for the worst-case approximation error and thus the excess risk bound (8) could be conservative in practice.
+In the practical implementation of PADR in Section 4, we observe that a much more moderate number of $K$ is able to produce satisfactory results including the two-stage product placement problem and Newsvendor problems with real data.
+This is because the (near) optimal decision rule may possess some underlying structural property besides Lipschitz continuity so that it could be approximated well by the PADR with a moderate number of pieces, for which a smaller estimation error and excess risk could be achieved.
+The excess risk bound (8) illustrates the role of $K$ in balancing the two errors, which is further evidenced by numerical experiments summarized in the table below.
+Table 2.2 reports the validation and training costs provided by the PADR-based ERM method in solving a Newsvendor problem with the sine plus max-affine demand model (19) (details are given in Section 4.1), under multiple choices of $K_{1}$ and $K_{2}$ which represent the number of pieces for the two max-affine components in the PADR respectively.
+The validation cost, which is the empirical cost of the obtained PADR solution over the validation data set of size 1000, can be regarded as the sample average approximation of $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})$.
+Noticing that $\inf_{f\in\mathcal{F}}R(f)$ is irrelevant to the PA hypothesis class, the performance of excess risk could be revealed through the pattern of validation costs in the table below.
+For $K_{1}$ and $K_{2}$ being small, the validation cost decreases along with increasing values of $K_{1}$ and $K_{2}$ which may be credited to the decreasing approximation error.
+For $K_{1}$ and $K_{2}$ being larger than 4, the validation cost slightly increases, suggesting that as the PADR model becomes more complex, the increasing estimation error gradually dominates the decreasing approximation error.
+Hence, the best validation costs shall be achieved with moderate values of $K_{1}$ and $K_{2}$, which are usually considerably smaller than the ones implied by the theoretical results.
+[…]
+Validation and training costs for basic Newsvendor with demand model (19) $K_{1}\backslash K_{2}$ 1 2 3 4 5 6 1 24.06 (24.09) 23.62 (23.57) 23.60 (23.63) 23.64 (23.55) 23.72 (23.66) 23.64 (23.62) 2 8.68 (8.73) 7.01 (7.09) 7.01 (7.09) 7.02 (7.08) 7.00 (7.10) 7.02 (7.09) 3 7.82 (7.95) 5.41 (5.53) 5.45 (5.51) 5.44 (5.51) 5.43 (5.49) 5.42 (5.49) 4 7.29 (7.47) 4.26 (4.27) 4.25 (4.22) 4.28 (4.29) 4.29 (4.21) 4.22 (4.23) 5 7.26 (7.46) 4.26 (4.27) 4.25 (4.24) 4.23 (4.21) 4.25 (4.21) 4.27 (4.22) 6 7.28 (7.47) 4.27 (4.24) 4.17 (4.23) 4.16 (4.23) 4.24 (4.27) 4.22 (4.20) The PADR is obtained by Algorithm 1 for solving the PADR-based ERM model.
+Numbers listed in brackets represent training costs, and the lowest costs are highlighted in bold and red text.
 2.3 Consistency comparison with existing DR-based methods
@@
 […]
-In comparison, our PADR-based ERM model is nonasymptotically consistent for general SP problems, including nonconvex ones.
-Moreover, if there exists a convex or DC optimal DR, based on the approximation bounds $\mathcal{O}(K^{-2/p})$ of ${\mathcal{H}}_{\mu}^{K}$ from Balazs et al.
-(2015) and Siahkamari et al.
-(2020), our excess risk bound can be improved to $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$ with $K=\mathcal{O}(n^{p/(p+6)})$, which is close to the bound of NV-ERM1 (a special case of our model) when $p$ is large.
-This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$ for SP with a box-constrained set $[-M,M]^{d}$, by employing a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{\mu}$ in Theorem 2.3 could be independent of $\mu$.
-Though obtaining the global optimum $f^{*}_{S,{\mathcal{H}}_{\mu}^{K}}$ of the nonconvex ERM problem (5) is prohibitively challenging, we show in Sections 3 and 4 how to asymptotically compute a stationary solution that could be globally optimal to problem (5) under particular algorithmic settings, upon which our PADR-based ERM framework is asymptotically optimal.
-3 Enhanced SMM Algorithm for Learning PADR
-In this section, we focus on learning PADR by solving problem (5).
-Following the majorization-minimization (MM) framework, we develop a stochastic algorithm that iteratively solves convex subproblems constructed with sampling-based enhanced surrogation functions, referred to as an enhanced SMM algorithm.
-The highlight of our algorithm that distinguishes itself from the deterministic MM-related algorithms in Pang et al.
-(2016) and Cui et al.
-(2018) and stochastic ones in Xu et al.
-(2019), Le Thi et al.
-(2022), and Liu et al.
-(2022) is that it combines the enhanced $\varepsilon$-active surrogation technique with sequential sampling so as to converge to d-stationary solutions with a significant improvement of computational efficiency.
-We first make some remarks on notations to facilitate algorithmic development and convergence analysis.
-With a slight abuse of notation, to emphasize the role of PADR as a function of $\theta$, we represent the PA function as $f(\theta;\xi)=g(\theta;\xi)-h(\theta;\xi)$ with $\xi\in S=\{(x^{s},y^{s})\}_{s=1}^{n}$, where $g(\theta;\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta;\xi)$ and $h(\theta;\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta;\xi)$ are maximum of linear functions.
-With $\widetilde{\xi}_{n}$ representing a random variable uniformly distributed on $S$, the PADR-based ERM problem (5) can be generalized as the following stochastic program with finite support:
-$\min_{\theta\in\Theta}\quad F(\theta)\coloneqq\mathbb{E}\,\left[\varphi(f(\theta;\widetilde{\xi}_{n});\widetilde{\xi}_{n})\right],$ (9)
-where $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set.
-In particular, in problem (5) with $\Theta=[-\mu,\mu]^{q}$ and $q=2dK(p+1)$, the outer function $\varphi(\cdot;\xi)$ with $\xi=(x,y)$ depends solely on $y$, and the inner PA function $f(\cdot;\xi)$ depends solely on $x$.
-3.1 Construction of Surrogation Functions
-Similar to other MM-related algorithms, the cornerstone of the enhanced SMM algorithm is the construction of surrogation functions of $F(\cdot;\xi)\coloneqq\varphi(f(\cdot;\xi);\xi)$.
-For a directionally differentiable function $\phi:\Theta\to\mathbb{R}$, that is, the directional derivative $\phi^{\prime}(\theta;d)\coloneqq\lim_{\tau\to 0}\frac{\phi(\theta+\tau d)-\phi(\theta)}{\tau}$ exists for any $\theta\in\Theta$ and $d\in\mathbb{R}^{q}$, we are interested in constructing a surrogation function $\widehat{\phi}(\cdot;\theta^{\prime})$ at a reference point $\theta^{\prime}\in\Theta$ that may satisfy the following conditions:
+In comparison, our PADR-based ERM model is nonasymptotically consistent in more general cases, including convex SP problems and nonconvex ones.
+Moreover, by Corollary 2.4, if there exists a convex or DC optimal DR, our excess risk bound can be improved to $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$, which is close to the bound of NV-ERM1 (a special case of our model) when $p$ is large.
+This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$, for problem (1) with a box-constrained set $[-M,M]^{d}$ by employing a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{\mu}$ in Theorem 2.3 could be independent of $\mu$.
+Given the statistical consistency merits of the PADR method demonstrated by the analysis above, it is still prohibitively challenging to obtain the globally optimal solution to the nonconvex PADR-based ERM problem (5).
+From the continuous optimization perspective, the PADR-based ERM problem is a composite nonconvex and nondifferentiable optimization problem embedded with a particular piecewise affine structure.
+This leads to the development of a sampling-based surrogate-based algorithm in the next section to efficiently solve problem (5) of large sizes.
+3 Efficient Sampling-based Algorithm for Learning PADR
+In this section, we focus on learning PADR by solving the following parameter estimation problem
+$\min_{\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\forall\iota\in[d]}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi\left(\Big{(}\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x^{s}+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x^{s}+b^{k}_{\iota}\right\}\Big{)}_{\iota=1}^{d},y^{s}\right),$ (9)
+where the objective is the finite sum of composite-PA functions.
+Instead of attempting for global optimality, it is more practical to apply majorization-minimization (MM)-type methods to the contextual optimization problems with the practical scale in seeking “stationary” solutions.
+Among various kinds of stationary concepts, the d-stationary solution is the sharpest type of first-order stationary solutions for general nonconvex optimization.
+For the PADR learning problem (9) with a general nonconvex and nondifferentiable composite structure, our algorithmic development is mainly motivated by the following question:
+Given the potentially large sample size, how to efficiently compute a sharper “stationary” solution to the PADR-based ERM problem (9) so that a global minimizer could be more easily obtained through replications with random initialization?
+Building on this motivation, we develop an MM-type algorithm specializing to the structured PADR-based ERM problem with the theoretical convergence analysis.
+It is further observed with numerous experiments in Section 4 that satisfactory (nearly optimal) solutions can be obtained with a few number of repetitions, which speeds up to 4 times faster than the basic MM method.
+To make our discussion self-contained, we first present the standard MM algorithms in Section 3.1, which serves as a reference for the motivation behind our proposed sampling-based algorithm in Section 3.2.
+We make some remarks on notations to facilitate algorithmic development and convergence analysis.
+To emphasize the role of PADR as a function of $\theta$, with $\xi=(\xi_{x},\xi_{y})\in\Xi$, we denote $g(\theta,\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta,\xi_{x})$ and $h(\theta,\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta,\xi_{x})$ where $\{g_{i}(\theta,\xi_{x})\}$ and $\{h_{i}(\theta,\xi_{x})\}$ are linear functions with respect to $\theta$ and we denote $f(\theta,\xi)\coloneqq g(\theta,\xi)-h(\theta,\xi)$ as the PA function with a slight abuse of notation.
+With $\widetilde{\xi}_{n}$ representing a random variable uniformly distributed on $\Xi$, the PADR-based ERM problem (5) can be generalized as the following stochastic program with finite scenarios:
+$\min_{\theta\in\Theta}\quad F(\theta)\coloneqq\mathbb{E}\,\left[F(\theta,\widetilde{\xi}_{n})\right],$ (10)
+where $F(\theta,\xi)\coloneqq\varphi(f(\theta,{\xi}),{\xi})$ and $\Theta\subseteq\mathbb{R}^{q}$ is a compact convex set.
+In particular, it reduces to problem (5) when $\Theta=[-\mu,\mu]^{q}$ with $q=2dK(p+1)$, the outer function $\varphi(\cdot,\xi)$ with $\xi=(x,y)$ depends solely on $y$, and the inner PA function $f(\cdot,\xi)$ depends solely on $x$.
+We say a function $F:\Theta\subseteq\mathbb{R}^{q}\to\mathbb{R}$ is directionally differentiable if the directional derivative $F^{\prime}(\theta;d)\coloneqq\lim_{\tau\to 0}\frac{F(\theta+\tau d)-F(\theta)}{\tau}$ exists for any $\theta\in\Theta$ and $d\in\mathbb{R}^{q}$.
+Supposing that $\Theta$ is a closed convex set and $F:\Theta\rightarrow\mathbb{R}$ is directionally differentiable, we say that $\theta^{d}\in\Theta$ is a d(irectional)-stationary point of problem (10) if $F^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for any $\theta\in\Theta$.
+3.1 Preliminary analysis: standard MM algorithms
+The core of MM-type methods for solving nonconvex minimization problem is to iteratively solve subproblems with a surrogate function $\widehat{F}(\theta;\theta^{\nu})$ majorizing $F(\theta)$ at an iterate point $\theta^{\nu}\in\Theta$.
+The iterate point is updated with $\eta\geq 0$ as follows,
+$\theta^{\nu+1}\leftarrow\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta^{\nu})+\frac{\eta}{2}\left\lVert\theta-\theta^{\nu}\right\rVert^{2}\right\}.$ (11)
+A surrogate function $\widehat{F}(\theta;\bar{\theta})$ at a reference point $\bar{\theta}\in\Theta$ typically satisfies the following conditions:
   1. (P1)
-    Touching condition: $\widehat{\phi}(\theta^{\prime};\theta^{\prime})=\phi(\theta^{\prime})$.
+    Touching condition: $\widehat{F}(\bar{\theta};\bar{\theta})=F(\bar{\theta})$.
   2. (P2)
-    Majorizing condition: $\widehat{\phi}({\theta};{\theta^{\prime}})\geq\phi(\theta)$ for any $\theta\in\Theta$.
+    Majorizing condition: $\widehat{F}({\theta};{\bar{\theta}})\geq F(\theta)$ for any $\theta\in\Theta$.
   3. (P3)
-    Convexity: $\widehat{\phi}({\cdot};{\theta^{\prime}})$ is convex on $\Theta$.
-  4. (P4)
-    Directional derivative consistency: $\widehat{\phi}(\cdot;\theta^{\prime})$ is directionally differentiable on $\Theta$ and satisfies $\widehat{\phi}(\cdot;\theta^{\prime})^{\prime}(\theta^{\prime};d)=\phi^{\prime}(\theta^{\prime};d)$ for any $d\in\mathbb{R}^{q}$.
-  5. (P5)
-    Quadratic surrogation gap: for some $\kappa>0$, $\phi(\theta)+\displaystyle{\frac{\kappa}{2}}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\geq\widehat{\phi}(\theta;\theta^{\prime})$ for any $\theta\in\Theta$.
-A function satisfying the first two conditions is referred to as an upper surrogation function according to Cui and Pang (2021, Definition 7.1.1).
-Though the enhanced $\varepsilon$-active surrogation functions to be constructed generally do not satisfy (P1), we still refer to ones solely satisfying (P2) as upper surrogation functions and similarly for lower surrogation functions.
-Condition (P3) ensures that surrogation-based subproblems in the algorithm can be solved by convex-programming solvers.
-Condition (P4) connecting the directional derivatives of surrogation and original functions is used to establish stationarity conditions of the limit points produced by surrogation-based algorithms to the original problem.
-Condition (P5) originating from Liu et al.
-(2022) is a technical condition for the nonasymptotic convergence analysis of the enhanced SMM algorithm.
-Note that Condition (P5) implies (P4) provided that $\widehat{\phi}(\cdot;\theta^{\prime})$ is directionally differentiable and satisfies (P2).
-We first present upper and lower surrogation functions of the inner PA function $f(\cdot;\xi)$, which provide the basis of surrogation construction of $F(\cdot;\xi)$ and also play an essential role in the randomized index-mapping selection in the enhanced SMM algorithm.
-For $d=1$ and $\varepsilon\geq 0$, let $\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ denote the set of $\varepsilon$-active index mappings for functions $\{g(\cdot;\xi)\}_{\xi\in S}$ at $\theta^{\prime}\in\Theta$, i.e.,
-$\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})\coloneqq\Big{\{}I_{1}:S\rightarrow[K]\mid\,g_{I_{1}(\xi)}(\theta^{\prime};\xi)\geq g(\theta^{\prime};\xi)-\varepsilon,\,\forall\,\xi\in S\Big{\}},$
-and the set $\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ is similarly defined for $\{h(\cdot;\xi)\}_{\xi\in S}$.
-Each mapping $I_{1}\in\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ assigns an $\varepsilon$-active index for each $\xi\in S$, and thus $\{I_{1}(\xi)\}_{\xi\in S}$ constitutes an $\varepsilon$-active index combination of $\{g(\cdot;\xi)\}_{\xi\in S}$ at $\theta^{\prime}$.
-It is clear to see that $\widehat{f}(\theta;\xi,I_{2})\coloneqq g(\theta;\xi)-h_{I_{2}(\xi)}(\theta;\xi)$ with $I_{2}\in\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ is an upper surrogation function of $f(\cdot;\xi)$ satisfying Conditions (P2) and (P3) with $\varepsilon\geq 0$, and additionally (P1) with $\varepsilon=0$.
-The concave lower surrogation function $\widecheck{f}(\theta;\xi,I_{1})$ with $I_{1}\in\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})$ can be similarly defined.
-When $d>1$, the index mapping sets and surrogates are constructed componentwisely.
-Now, we can construct the upper surrogation function of $F(\cdot;\xi)$ at a reference point $\theta^{\prime}$ based on the surrogation of inner PA functions with the index mapping $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})\coloneqq\mathcal{I}_{g}^{\varepsilon}(\theta^{\prime})\times\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$, which we denote by $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$.
-To facilitate the convergence analysis of the enhanced SMM algorithm, we impose the following assumptions on the family of surrogates $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$:
-  1. [label=(B0), series=B]
-  2. 1.
-    The family of surrogates $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ for any $\xi\in S$ satisfy the following properties:
-    1. (a)
-      $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies Conditions (P2) and (P3) for every $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ with $\varepsilon\geq 0$;
-    2. (b)
-      $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies Condition (P1) for every $I\in\mathcal{I}^{0}(\theta^{\prime})$;
-    3. (c)
-      $\widehat{F}(\cdot;\cdot,\xi,I)$ is continuous on $\Theta\times\Theta$ given a fixed index mapping $I$;
-    4. (d)
-      the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,I)$ satisfies Condition (P4) for any $\varepsilon\geq 0$.
-  3. 2.
-    The minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)$ satisfies Condition (P5) for any $\varepsilon\geq 0$ and $\xi\in S$.
-To illustrate the specific construction of surrogation functions satisfying the above assumptions, we consider the following three cases of the outer function $\varphi$:
+    Convexity: $\widehat{F}({\cdot};{\bar{\theta}})$ is convex on $\Theta$.
+A function satisfying (P1) and (P2) is referred to as an upper surrogate function according to Cui and Pang (2021, Definition 7.1.1).
+Under (P1) and (P2), it yields the sequence of decreasing objective values as follows,
+$F(\theta^{\nu+1})\leq\widehat{F}(\theta^{\nu+1};\theta^{\nu})\leq\widehat{F}(\theta^{\nu};\theta^{\nu})-\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}=F(\theta^{\nu})-\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\leq F(\theta^{\nu}).$
+Condition (P3) ensures that surrogation-based subproblems in the algorithm can be solved by state-of-the-art convex-programming solvers.
+By the basic analysis (see Chapter 7 in Cui and Pang (2021)), the limit point of $\{\theta^{\nu}\}$ produced by the process (11) could be easily shown to be a fixed point of the surrogation map satisfying $\theta^{*}=\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta^{*})+\frac{\eta}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}\right\}$.
+Such a fixed point property could be related to various types of stationarity of the original problem depending on the surrogate functions as illustrated in Cui et al.
+(2018) and Liu et al.
+(2022).
+Regarding the PADR-based ERM problem (10), due to the nonconvex and nondifferentiable piecewise affine functions, a major issue of the standard MM algorithms including Lipp and Boyd (2016) and Le Thi and Pham Dinh (2018) is that the fixed point corresponding to the surrogation map is a (compound) critical solution (see definitions in Pang et al.
+(2016), Liu et al.
+(2022)) which is much weaker than directional stationarity.
+Specifically, based on the composite structure in (10), approaches to construct surrogate functions of $F(\theta,\xi)$ in Cui et al.
+(2018) typically involve linearization of the two max-affine components of the inner piecewise affine function as follows,
+$\widehat{f}(\theta,\xi;I_{2})$ $\coloneqq g(\theta,\xi)-h_{I_{2}}(\theta,\xi)\mbox{ with }I_{2}\in\mathcal{I}_{h}(\bar{\theta},\xi)\coloneqq\mbox{argmax}_{i\in[K]}h_{i}(\bar{\theta},\xi),$ (12)
+$\widecheck{f}(\theta,\xi;I_{1})$ $\coloneqq g_{I_{1}}(\theta,\xi)-h(\theta,\xi)\mbox{ with }I_{1}\in\mathcal{I}_{g}(\bar{\theta},\xi)\coloneqq\mbox{argmax}_{i\in[K]}g_{i}(\bar{\theta},\xi).$ (13)
+They are the upper and lower surrogate functions of $f(\cdot,\xi)$ at $\bar{\theta}\in\Theta$ respectively.
+For $d>1$, the maximum is taken componentwisely.
+We illustrate the construction of the surrogate function when outer function $\varphi$ is linear as an example.
+With $\varphi(z,\xi)=\xi^{\top}z$, we construct the surrogate function of $F(\cdot,\xi)$ as $\widehat{F}(\theta,\xi;\bar{\theta},I)=\max\{\xi,0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+\min\{\xi,0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})$, where $I\coloneqq\big{(}I_{1},I_{2}\big{)}\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$ for any $\xi\in\Xi$.
+It obviously satisfies Conditions (P1) to (P3), and the standard MM algorithm for such a case is exactly a DC algorithm.
+Clearly $f^{\prime}(\cdot,\xi)(\theta;d)=g^{\prime}(\cdot,\xi)(\theta;d)-\max\{\,\nabla h_{i}(\bar{\theta},\xi)^{\top}d:\,i\in\mbox{argmax}_{i\in[K]}h_{i}(\bar{\theta},\xi)\}$, whereas the surrogate function $\widehat{F}(\theta,\xi;\bar{\theta},I)$ employs an arbitrary active index $I\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$.
+Hence, the fixed point of the surrogation map is a critical solution.
+Solutions of criticality are not desirable because in practice solutions may be at the “cliff” of the cost landscape, which is illustrated in Example 3.1 and Figure 1b.
+To obtain a directional stationary solution, Pang et al.
+(2016) considered a class of nondifferentiable DC programs which have a max-smooth structure in the second convex part, and introduced $\varepsilon$-active sets to construct the enhanced surrogate functions for DC functions.
+Specifically, for the piecewise affine function $f(\cdot,\xi)$ given $\xi\in\Xi$, we denote the $\varepsilon$-active set with $\varepsilon>0$ by
+$\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\Big{\{}i\in[K]\mid\,g_{i}(\bar{\theta},\xi)\geq g(\bar{\theta},\xi)-\varepsilon\Big{\}},$ (14)
+and the set $\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$ is similarly defined for $h(\cdot,\xi)$.
+Let $\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$.
+Pang et al.
+(2016) proposes the enhanced MM algorithm in which the subproblem takes the minimum over the family of surrogate functions with indices in $\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$ and provides the theoretical convergence analysis to a d-stationarity point.
+The following example provides an intuitive illustration that the iterate point can escape from a critical point to better local minima and make valid descents via the enhanced technique.
+Example 3.1
+Consider problem (10) with $d=2$, $q=1$, $\Xi=\{\xi_{0}\}$, $\varphi(z,\xi_{0})=z_{1}^{2}+z_{2},f_{1}(\theta,\xi_{0})=\theta$, and $f_{2}(\theta,\xi_{0})=-\max\{-\theta,0\}$.
+Figure 1a illustrates the two candidates of surrogate functions depending on which piece of the inner PA function $f_{2}(\theta,\xi_{0})$ is active at the reference point of interest.
+Figure 1b and Figure 1c illustrate the iterative procedures of the standard MM algorithm and the enhanced MM algorithm respectively, with the initial point $\theta_{0}=1.1$.
+The standard MM algorithm converges to a critical point $\theta_{\infty}=0$.
+For the enhanced MM algorithm with $\varepsilon=0.3$, at the second iterate point $\theta_{2}$, we have $\left\lvert\mathcal{I}^{\varepsilon}(\theta_{2},\xi_{0})\right\rvert=2$ at $\theta_{2}$, which thus leads to the next iterate point passing over 0 and continues to move towards a d-stationary point $\theta^{d}$ that is also globally optimal to the problem.
+[…]
+(a) DC function
+(b) Surrogation with $\varepsilon=0$
+(c) Surrogation with $\varepsilon=0.3$
+Iterative process of the standard MM algorithm and enhanced MM algorithm
+Despite possessing stronger convergence properties, the enhanced MM algorithm applied for large-scale ERM problems could be exceptionally inefficient as it requires iteratively solving a potentially very large number of finite-sum convex subproblems over all the data.
+Sampling-based MM-type algorithms have been recently studied for (partially) smooth stochastic DC programs in Nitanda and Suzuki (2017), Xu et al.
+(2019), and Le Thi et al.
+(2022) and nondifferentiable stochastic composite DC problems in Liu et al.
+(2022) to compute compound critical solutions.
+Based on a natural idea of combining a sequential sampling scheme with the enhanced technique, we propose an enhanced sampling-based MM algorithm with rigorous convergence analysis in the next subsection to solve the stochastic composite DC programs (10) with both computational efficiency and theoretical convergence guarantee.
+3.2 Enhanced sampling-based MM algorithm for learning PADR
+In this subsection, we first introduce the family of enhanced surrogate functions of the composite objective $F$ and provide several function classes of the outer function $\varphi$ as particular instances.
+This provides the basis for the characterization of a so-called composite $(\varepsilon,\rho)$-strong d-stationarity.
+We then propose the enhanced sampling-based MM algorithm (ESMM) which combines the enhanced $\varepsilon$-active technique with sequential sampling and analyze the asymptotic and non-asymptotic convergence properties in terms of (composite strong) d-stationarity under suitable incremental sampling schemes.
+Such a combination enables the algorithm to solve subproblems of small sizes in the early iterations that effectively reduces computational costs.
+The computational improvement of such a technique is further supported by numerical experiments in Section 4.
+3.2.1 Family of enhanced surrogate functions.
+Based on surrogation of $f(\theta,\xi)$ in (12) and (13) and corresponding $\varepsilon$-active index set $\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$ given $\bar{\theta}\in\Theta$, we introduce a family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ for any $\xi\in\Xi$ that satisfy the following conditions:
+  1. (B1)
+    $\widehat{F}(\cdot,\xi;\bar{\theta},I)$ satisfies (P1) for every $I\in\mathcal{I}^{0}(\bar{\theta},\,\xi)$;
+  2. (B2)
+    $\widehat{F}(\cdot,\xi;\bar{\theta},I)$ satisfies (P2) and (P3) for every $I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)$ with $\varepsilon\geq 0$;
+  3. (B3)
+    $\widehat{F}(\cdot,\xi;\,\cdot,I)$ is continuous on $\Theta\times\Theta$ given fixed indices $I=(I_{1},I_{2})\in[K]^{2}$;
+  4. (B4)
+    the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}\widehat{F}(\theta,\xi;\bar{\theta},I)$ satisfies the directional derivative consistency for any $\varepsilon\geq 0$; namely, $\widehat{F}_{\min}^{\varepsilon}(\cdot\,;\bar{\theta})$ is directionally differentiable on $\Theta$ and satisfies $\widehat{F}_{\min}^{\varepsilon}(\cdot\,;\bar{\theta})^{\prime}(\bar{\theta};d)=F^{\prime}(\bar{\theta};d)$ for any $d\in\mathbb{R}^{q}$.
+The directional derivative consistency condition is introduced in Cui and Pang (2021) to relate the directional derivatives of surrogation and original functions.
+We consider the following three cases of the outer function $\varphi$ covering a broad class of cost functions.
   1. Case 1.
-    $\varphi(\cdot;\xi)$ is piecewise affine for any $\xi\in S$.
+    $\varphi(\cdot,\xi)$ is piecewise affine for any $\xi\in\Xi$.
   2. Case 2.
-    $\varphi(\cdot;\xi)$ can be decomposed into monotonic convex functions for any $\xi\in S$.
+    $\varphi(\cdot,\xi)$ can be decomposed into monotonic convex functions for any $\xi\in\Xi$.
   3. Case 3.
-    $\varphi(\cdot;\xi)$ is smooth with Lipschitz gradient for any $\xi\in S$.
-In Appendix 9, we provide the surrogation construction of $F(\cdot;\xi)$ for these cases with verification of Assumptions 1 and 2.
-The three cases cover a broad class of cost functions.
+    $\varphi(\cdot,\xi)$ is smooth with Lipschitz gradient for any $\xi\in\Xi$.
+For the above cases, we can construct the surrogate function of $F(\cdot,\xi)$ satisfying Assumptions (B1) to (B4), by utilizing surrogate functions (12) and (13) of the inner PA function $f(\cdot,\xi)$ with the selected index $I=(I_{1},I_{2})$.
+The specific construction and verification are provided in Online Appendix 7.
 In particular, Case 2 is a multivariate extension of the one considered in Cui et al.
-(2018), which shows that any univariate convex function can be written as the sum of monotonic convex functions.
-The surrogation construction for Case 3 is new and suitable for nonconvex smooth cost functions.
-3.2 Enhanced SMM Algorithm
-We present the enhanced SMM algorithm in Algorithm 1.
+(2018).
+Case 3 is suitable for nonconvex smooth cost functions that are applicable for many practical problems.
+Proposition 3.2
+Let $\Theta$ be a compact set.
+For Case 1 to Case 3, given $\varepsilon\geq 0$ and $\bar{\theta}\in\Theta$, there exists a family of enhanced surrogate functions based on the index set $\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$ of the inner PA function that satisfy Assumptions (B1)-(B4) for any $\xi\in\Xi$.
+3.2.2 Concepts on stationarity.
+For a general nonconvex problem on a convex set with the directional differentiable objective, d-stationarity is identified as the sharpest kind among the first-order stationarity concepts.
+Specialized to the composite DC program (10), according to Cui et al.
+(2018, Proposition 2), d-stationarity is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2.
+Lu et al.
+(2019) proposes the strengthened d-stationarity for structured DC programs which takes into account the $\epsilon$-active index set.
+We extend to the concept of composite $(\varepsilon,\rho)$-strong d-stationarity specialized for the PADR-based ERM problem (9).
+The composite strongly d-stationary point can be interpreted as the fixed point corresponding to the proximal mapping of the expected minimum surrogate function $\mathbb{E}[\,\widehat{F}^{\varepsilon}_{\min}(\theta,\widetilde{\xi}^{n};\bar{\theta})]$ based on the family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ satisfying Assumptions (B1)-(B4).
+Definition 3.3
+Given $\varepsilon\geq 0$ and $\rho>0$, we say that $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (10) if ${\{\theta^{d}\}=\mathop{\arg\min}_{\theta\in\Theta}\left\{\mathbb{E}[\,\widehat{F}^{\varepsilon}_{\min}(\theta,\widetilde{\xi}^{n};{\theta}^{d})\,]+\frac{\rho}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}}$.
+Correspondingly, let $\mathcal{S}^{d}_{\varepsilon,\rho}$ denote the set of composite $(\varepsilon,\rho)$-strongly d-stationary points.
+By Condition (B4), we must have $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}_{d}$ for any $\varepsilon\geq 0$ where $\mathcal{S}_{d}$ denotes the set of d-stationary points.
+Moreover, with $\varepsilon>0$, we have the inclusion property $\mathcal{I}^{0}(x,\xi)\subseteq\mathcal{I}^{0}(\hat{x},\xi)\subseteq\mathcal{I}^{\varepsilon}(x,\xi)$ for $x$ sufficiently close to $\hat{x}$, which is crucial for the convergence analysis of the limit point in terms of (composite strong) d-stationarity.
+With $\mathcal{S}^{*}$ being the set of global minima of (10), the following proposition gives the relation among $\mathcal{S}^{*},\mathcal{S}^{d}_{\varepsilon,\rho}$ and $\mathcal{S}^{d}$.
+The proof is provided in Online Appendix 8.1.
+Proposition 3.4
+$\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$ for any $\varepsilon\geq 0$ and $\rho>0$.
+In particular, $\mathcal{S}^{d}_{0,\rho}=\mathcal{S}^{d}$ for any $\rho>0$.
+Due to the requirement of a unique proximal mapping point in the defining condition, our concept is slightly stronger than $(\alpha,\eta)$-d-stationarity defined in Lu et al.
+(2019) when the problem is reduced to composite DC program.
+Moreover, it is weaker than composite $\varepsilon$-strong d-stationarity introduced in Qi et al.
+(2021b) (see also Cui and Pang (2021, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition which turns out to be essential for the algorithm development and the follow-up convergence analysis.
+Furthermore, in Online Appendix 8.1, we show that $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ under a regularity assumption on $F$.
+It implies that the definition of $\mathcal{S}^{d}_{\varepsilon,\rho}$ serves as a natural connection between d-stationarity and global optimality.
+3.2.3 ESMM algorithm and convergence analysis.
+We propose the ESMM algorithm with a natural idea of combining the randomized $\varepsilon$-active surrogation with sequential sampling presented in Algorithm 1.
+At iteration $\nu$, a subset of samples $\Xi_{\nu}=\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ of size $N_{\nu}$ is generated in step 2.
+Then in step 3 a sampling-based surrogate function $\widehat{F}_{\nu}(\theta;\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})$ is constructed which can be regarded as the upper surrogation function to the sample average approximation function $F_{\nu}(\cdot)\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}F(\cdot,\widetilde{\xi}_{\nu}^{s})$ with the combination of indices $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}=\{\widetilde{I}^{\varepsilon}_{\nu,s}\}_{s=1}^{N_{\nu}}$ randomly selected from the $\varepsilon$-active index sets $\{\mathcal{I}^{\varepsilon}(\theta^{\nu},\widetilde{\xi}_{\nu}^{s})\}_{s=1}^{N_{\nu}}$, respectively.
+The randomized selection combined with the sampling scheme enables the algorithm to solve subproblems of small sizes in the early iterations.
+By the convexity of the surrogate function, in step 4 the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ can be obtained via the state-of-the-art convex programming solvers.
+However, since the indices $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}$ are randomly selected from the $\varepsilon$-active index sets, the touching condition (P1) may not be satisfied when $\varepsilon>0$; namely, the surrogate function value $\widehat{F}_{\nu}(\theta^{\nu};\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})$ might be strictly larger than $F_{\nu}(\theta^{\nu})$.
+Thus the candidate solution $\theta^{\nu+\frac{1}{2}}$ may not lead to a valid update with the desired descent property.
+Hence, in step 5 the candidate solution $\theta^{\nu+\frac{1}{2}}$ is accepted only when sufficient descent is satisfied.
+In the special case of $\varepsilon=0$ for which the touching condition (P1) is satisfied, the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ satisfies the sufficient descent inequality in step 5, and thus we could directly update $\theta^{\nu+1}=\theta^{\nu+\frac{1}{2}}$.
 Input: initial point $\theta^{0}$, iteration length $T$, parameters $\{\eta,\varepsilon\}$, and increasing sequence $\{N_{\nu}\}_{\nu}$;
 1 for $\nu=0$ to $T-1$ do
-2       for $s=1,\dots,N_{\nu}$, draw $\xi_{\nu}^{s}$ independently and uniformly from $S$;
-3       select $I_{\nu}^{\varepsilon}$ uniformly from $\mathcal{I}^{\varepsilon}(\theta^{\nu})$;
-4       compute $\theta^{\nu+\frac{1}{2}}\in\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}_{N_{\nu}}(\theta;\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\left\lVert\theta-\theta^{\nu}\right\rVert^{2}\right\}$;
-5       update the iterate $\theta^{\nu+1}=\begin{cases}\theta^{\nu+\frac{1}{2}},&\text{if}\ \widehat{F}_{N_{\nu}}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{N_{\nu}}(\theta^{\nu}),\\ \theta^{\nu},&\text{otherwise.}\end{cases}$
+2       generate $\Xi_{\nu}\coloneqq\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ independently and uniformly from $\Xi$;
+3       construct the surrogate function $\widehat{F}_{\nu}(\theta;\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}\,\widehat{F}(\theta,\widetilde{\xi}_{\nu}^{s};\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu,s})$ where $\widetilde{\mathbf{I}}_{\nu}^{\varepsilon}\coloneqq\{\widetilde{I}^{\varepsilon}_{\nu,s}\}_{s=1}^{N_{\nu}}$ and $\widetilde{I}^{\varepsilon}_{\nu,s}$ is uniformly generated from $\mathcal{I}^{\varepsilon}(\theta^{\nu},\widetilde{\xi}_{\nu}^{s})$;
+4       compute $\theta^{\nu+\frac{1}{2}}\in\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}_{\nu}(\theta;\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})+\frac{\eta}{2}\left\lVert\theta-\theta^{\nu}\right\rVert^{2}\right\}$;
+5       update the iterate $\theta^{\nu+1}=\begin{cases}\theta^{\nu+\frac{1}{2}},&\text{if}\ \widehat{F}_{\nu}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{\nu}(\theta^{\nu}),\\ \theta^{\nu},&\text{otherwise.}\end{cases}$
 6 end for
 Output: $\widetilde{\theta}^{\,T}$ uniformly sampled from the iterates $\{\theta^{\nu}\}_{\nu=0}^{T-1}$.
-Algorithm 1 Enhanced SMM Algorithm
-At iteration $\nu$, with samples $\{\xi_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ and an $\varepsilon$-active index mapping $I^{\varepsilon}_{\nu}\in\mathcal{I}^{\varepsilon}(\theta^{\nu})$ randomly selected in Steps 2 and 3, we compute in Step 4 the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ with respect to the upper surrogation function $\widehat{F}_{N_{\nu}}(\cdot;\theta^{\nu},I_{\nu}^{\varepsilon})\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}\widehat{F}(\cdot;\theta^{\nu},\xi_{\nu}^{s},I^{\varepsilon}_{\nu})$ of the sample average approximation function $F_{N_{\nu}}(\cdot)\coloneqq\frac{1}{N_{\nu}}\sum_{s=1}^{N_{\nu}}F(\cdot;\xi_{\nu}^{s})$.
-Since the surrogate function value $\widehat{F}_{N_{\nu}}(\theta^{\nu};\theta^{\nu},I_{\nu}^{\varepsilon})$ could be strictly larger than $F_{N_{\nu}}(\theta^{\nu})$, the candidate point $\theta^{\nu+\frac{1}{2}}$ may not lead to a valid update.
-Hence, it is accepted in Step 5 only when sufficient descent is satisfied.
-In the special case of $\varepsilon=0$, the touching condition (P1) is satisfied.
-Thus the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ naturally satisfies the sufficient descent inequality in Step 5, and we directly update $\theta^{\nu+1}=\theta^{\nu+\frac{1}{2}}$.
-It is worth noticing that the randomized index-mapping selection in Step 3 is vital for the ERM problem, without which the number of subproblems to be solved at each iteration, identical with the cardinality of $\mathcal{I}^{\varepsilon}(\theta^{\nu})$, would grow exponentially with the sample size.
-Furthermore, by combining the enhanced $\varepsilon$-active surrogation with incremental sample averaging, the enhanced SMM can achieve satisfactory descent progress in early iterations with relatively small sample sizes, thus improving the computational efficiency; yet such a combination raises the major challenges of convergence analysis to be addressed in the next section.
-4 Convergence Analysis of Enhanced SMM
-In this section, we first introduce preliminaries on stationarity concepts and then analyze the convergence properties of the enhanced SMM algorithm for solving problem (9).
-Since much of the convergence analysis is quite technical, we present the main theoretical results with essential discussions while leaving analytical details in Appendix 10.
-4.1 Preliminaries on Stationarity
-Suppose that the outer function $\varphi$ is Bouligand-differentiable (B-differentiable), i.e., it is locally Lipschitz continuous and directionally differentiable on $\mathbb{R}^{d}$.
-Then the composite function $F$ is also B-differentiable according to Cui and Pang (2021, Proposition 4.1.2).
-We say that $\theta^{d}\in\Theta$ is a d-stationary point of problem (9) if $F^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for any $\theta\in\Theta$.
-For a general nonconvex problem on a convex set, d-stationarity is identified as the sharpest among the first-order stationarity concepts, which gives a necessary condition for local optimality.
-Specialized to the composite DC program (9), it is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2 (Cui et al.
-2018, Proposition 2), and more generally we further define the following strengthened d-stationarity.
-Definition 4.1
-Given $\varepsilon\geq 0$ and $\rho>0$, we say that $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (9) if ${\{\theta^{d}\}=\mathop{\arg\min}_{\theta\in\Theta}\left\{\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}\widehat{F}(\theta;\theta^{d},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}}$, where $\widehat{F}(\theta;\theta^{d},I)\coloneqq\mathbb{E}\,[\widehat{F}(\theta;\theta^{d},\widetilde{\xi}_{n},I)\,]$, and $\{\widehat{F}(\cdot;\theta^{d},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}$ satisfy Assumption 1 for every $\xi\in S$.
-This concept is generalized from the $(\alpha,\eta)$-d-stationarity defined in Lu et al.
-(2019) for structured DC programs.
-Due to the requirement of the unique proximal mapping point in the defining condition, our concept is slightly stronger for a reduced composite DC program with the identity outer function $\varphi$.
-Moreover, it is weaker than the composite $\varepsilon$-strong d-stationarity introduced in Qi et al.
-(2021) (see also Cui and Pang (2021, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition that is necessary for the follow-up convergence analysis.
-Let $\mathcal{S}^{d}$, $\mathcal{S}^{d}_{\varepsilon,\rho}$, and $\mathcal{S}^{*}$ denote the sets of d-stationary points, composite $(\varepsilon,\rho)$-strongly d-stationary points, and global minimizers of (9), respectively.
-We show below that the composite strong d-stationarity is stronger than d-stationarity and necessary for global optimality.
+Algorithm 1 ESMM algorithm
+Next, we establish the convergence analysis of Algorithm 1 in terms of (composite strong) d-stationarity for problem (10).
+Due to the combination of sampling and enhanced technique, we introduce the averaged residual for convergence analysis, which is technically different from the analysis in Pang et al.
+(2016), Cui et al.
+(2018), Lu et al.
+(2019).
+We present the main theoretical results with essential discussions while leaving technical details in Online Appendix 8.
+Throughout the analysis, we make several blanket assumptions: the outer function $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the family of enhanced surrogate functions $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$ and any $\bar{\theta}\in\Theta$.
+With $\Xi=\{\xi^{s}\}_{s\in[n]}=\{(x^{s},y^{s})\}_{s\in[n]}$, let $\mathcal{I}^{\varepsilon}(\bar{\theta})$ denote the set of all $\varepsilon$-active index combinations of $\{(g(\bar{\theta},\xi),h(\bar{\theta},\xi)\}_{\xi\in\Xi}$, i.e.,
+$\mathcal{I}^{\varepsilon}(\bar{\theta})\coloneqq\Big{\{}{\mathbf{I}}^{\varepsilon}=\{(I_{1,s}^{\varepsilon},I_{2,s}^{\varepsilon})\}_{s=1}^{n}\mid I^{\varepsilon}_{1,s}\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi^{s}),I^{\varepsilon}_{2,s}\in\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi^{s}),\,\forall\,s=1,\ldots,n\Big{\}},$
+with $|\mathcal{I}^{\varepsilon}(\bar{\theta})|$ representing the total number of combinations of $\varepsilon$-active index selection for the ERM objective function $\frac{1}{n}\sum_{s=1}^{n}F(\theta,\xi^{s})$.
+For the purpose of analysis, we introduce the surrogate function for such an ERM function, and adjusted proximal mapping point which is motivated by the iterative update in step 5 of Algorithm 1.
+At $\bar{\theta}\in\Theta$ with $\mathbf{I}^{\varepsilon}=\{I_{s}^{\varepsilon}\}_{s\in[n]}\in\mathcal{I}^{\varepsilon}(\bar{\theta})$, let
+$\begin{array}[]{ll}\widehat{F}(\cdot;\bar{\theta},\mathbf{I}^{\varepsilon})&\coloneqq\displaystyle{\frac{1}{n}}\sum_{s=1}^{n}\widehat{F}(\theta,\xi^{s};\bar{\theta},I^{\varepsilon}_{s})\\[14.45377pt] P^{\rho}_{\mathbf{I}^{\varepsilon}}(\bar{\theta})&\coloneqq\begin{cases}\mathrm{prox}_{\widehat{F}/\rho}(\bar{\theta},\mathbf{I}^{\varepsilon})&\text{if}\ \min_{\theta\in\Theta}\,\left\{\widehat{F}(\theta;\bar{\theta},\mathbf{I}^{\varepsilon})+\frac{\rho}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}\leq F(\bar{\theta}),\\ \bar{\theta}&\text{otherwise}.\end{cases}\end{array}$ (15)
+where the proximal map $\mathrm{prox}_{\widehat{F}/\rho}(\bar{\theta},\mathbf{I}^{\varepsilon})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}(\theta;\bar{\theta},\mathbf{I}^{\varepsilon})+\frac{\rho}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}$.
+With the above prerequisite, we introduce the averaged residual
+$\bar{r}^{\varepsilon,\rho}(\bar{\theta})\coloneqq\frac{1}{\left\lvert\mathcal{I}^{\varepsilon}(\bar{\theta})\right\rvert}\sum_{\mathbf{I}^{\varepsilon}\in\mathcal{I}^{\varepsilon}(\bar{\theta})}\left\lVert\bar{\theta}-P^{\rho}_{\mathbf{I}^{\varepsilon}}(\bar{\theta})\right\rVert,$
+which plays a central role in the follow-up convergence analysis.
+Under the sampling size sequence $\{N_{\nu}\}$, we present the convergence result of $\bar{r}^{\varepsilon,\rho}$ in expectation in the proposition below, with the proof provided in Online Appendix 8.5.
 Proposition 3.5
-Under Assumption 1 for the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ for each $\xi\in S$, it holds for problem (9) that $\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$ for any $\varepsilon\geq 0$ and $\rho>0$.
-In particular, $\mathcal{S}^{d}_{0,\rho}=\mathcal{S}^{d}$ for any $\rho>0$.
-The proof is provided in Appendix 10.1.
-Furthermore, we can show that $\mathcal{S}^{*}=\mathcal{S}^{d}_{\varepsilon,\rho}$ for sufficiently large $\varepsilon$ and suitable $\rho$ under Assumption 2 and a regularity condition that d-stationary values of problem (9) are isolated from the optimal value, which naturally hold when $\varphi$ is in Case 1 or Case 2; this is formally stated as Proposition 10.2 in Appendix 10.1.
-4.2 Convergence Analysis of Enhanced SMM Algorithm
-Now we establish the convergence analysis of Algorithm 1 in terms of (composite strong) d-stationarity for problem (9).
-Throughout this section, we make several blanket assumptions that the outer function $\varphi$ is B-differentiable, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ satisfies Assumption 1 for each $\xi\in S$.
-Since the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{d})\coloneqq\underset{I\in\mathcal{I}^{\varepsilon}(\theta^{d})}{\min}\widehat{F}(\theta;\theta^{d},I)$ is only of theoretical interest in defining composite strong d-stationarity, we need an auxiliary residual as a bridge to connect with the sampling-based surrogation function $\widehat{F}_{N_{\nu}}(\cdot;\theta^{\nu},I_{\nu}^{\varepsilon})$ utilized in Algorithm 1 for convergence analysis.
-With $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ and the corresponding surrogate $\widehat{F}(\cdot;\theta^{\prime},I)$ at $\theta^{\prime}\in\Theta$, we denote the proximal map $\mathrm{prox}_{\widehat{F}/\rho}(\theta^{\prime},I)\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}(\theta;\theta^{\prime},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}$ and further define
-$P^{\rho}_{I}(\theta^{\prime})\coloneqq\begin{cases}\mathrm{prox}_{\widehat{F}/\rho}(\theta^{\prime},I)&\text{if}\ \min_{\theta\in\Theta}\,\left\{\widehat{F}(\theta;\theta^{\prime},I)+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}\leq F(\theta^{\prime}),\\ \theta^{\prime}&\text{otherwise}.\end{cases}$ (10)
-The map $P^{\rho}_{I}(\theta^{\prime})$ can be regarded as an iterative update of Algorithm 1 accounting for all $n$ scenarios, with which we define the averaged residual
-$\bar{r}^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\frac{1}{\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert}\sum_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\left\lVert\theta^{\prime}-P^{\rho}_{I}(\theta^{\prime})\right\rVert.$
-The following proposition gives convergence rates of Algorithm 1 in terms of $\bar{r}^{\varepsilon,\rho}$ in expectation with the proof provided in Appendix 10.5.
-Proposition 4.3
-Let $\widetilde{\theta}^{\,T}$ be the output of Algorithm 1 after $T$ iterations with $\varepsilon\geq 0$, $\eta\geq 0$, and $\{N_{\nu}\}_{\nu}$.
-Then there exists $C_{1}>0$ such that for any $\rho>\eta$,
+Let $\widetilde{\theta}^{\,T}$ be the output of Algorithm 1 after $T$ iterations with $\varepsilon\geq 0$, $\eta\geq 0$, and $\{N_{\nu}\}_{\nu=1}^{T}$.
+There exists $C_{1}>0$ such that for any $\rho>\eta$,
 $\big{(}\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]\,\big{)}^{2}\leq\frac{F(\theta^{0})-\min_{\theta\in\Theta}F(\theta)+C_{1}\cdot\sum_{\nu=0}^{T-1}\sqrt{\log N_{\nu}/N_{\nu}}}{\frac{\rho-\eta}{2}\cdot T}.$
@@
 $\big{(}\mathbb{E}\,[\bar{r}^{0,\rho}(\widetilde{\theta}^{\,T})]\,\big{)}^{2}\leq\frac{F(\theta^{0})-\min_{\theta\in\Theta}F(\theta)+C_{2}\cdot\sum_{\nu=0}^{T-1}(1/N_{\nu})}{\frac{\rho-\eta}{2}\cdot T}.$
-By relating $\bar{r}^{\varepsilon,\rho}$ to another residual that measures the proximal mapping gap with respect to $\widehat{F}_{\min}^{\varepsilon}(\cdot;\theta^{\prime})$, we can establish the convergence of Algorithm 1 to composite strong d-stationarity.
-The error-bound analysis for the nonasymptotic convergence to d-stationarity needs the following assumption generalized from Drusvyatskiy and Lewis (2018).
-We denote the closure operator of a set by $\mathrm{cl}(\cdot)$ and $\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)\coloneqq\mathop{\arg\min}_{\theta^{\prime}\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta^{\prime};\theta)+\frac{\rho}{2}\left\lVert\theta^{\prime}-\theta\right\rVert^{2}\right\}$.
-  1. [label=(B0), resume*=B]
-  2. 1.
-    Generalized quadratic growth: given $\varepsilon>0$ and $\rho>0$, there exist positive constants $\delta_{q}$, $\lambda_{q}$, and $c_{q}$ such that for any $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$ and $\max_{\theta^{\prime}\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)}\left\lVert\theta^{\prime}-\theta\right\rVert\leq\lambda_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(t;\theta^{d},\theta)\coloneqq F(\theta^{d}+t(\theta-\theta^{d}))$ is convex on $[0,1]$ and
-    $\bar{F}(t;\theta^{d},\theta)\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{q}}{2}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2},\quad\forall\,t\in[0,1].$
-The condition holds, for example, when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope, which is verified in Appendix 10.4.
-It can be expected to hold in more general cases since it is a local property only requiring the convexity and quadratic growth of $F$ near $\mathcal{S}^{d}$.
-Theorem 4.4 (Convergence of Enhanced SMM)
+By relating $\bar{r}^{\varepsilon,\rho}$ to the proximal mapping gap with respect to $\widehat{F}_{\min}^{\varepsilon}(\cdot;\bar{\theta})$, with $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$, we can establish the convergence of Algorithm 1 in terms of composite strong d-stationarity.
+Theorem 3.6 (Asymptotic convergence of ESMM)
 For problem (10), let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon>0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$.
-Then the following results hold:
-  1. (a)
-    Any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of problem (9) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$.
-  2. (b)
-    Under Assumptions 2 and 1, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$.
-In particular, if $\varphi$ is piecewise affine and $\Theta$ is a polytope, the result holds without the two assumptions.
-Theorem 4.4 shows the asymptotic convergence of the enhanced SMM algorithm to composite strong d-stationarity and nonasymptotic convergence rate of the distance to the d-stationary set.
-We speculate that it is possible, albeit challenging, to achieve nonasymptotic results in terms of composite strong d-stationarity with carefully designed residuals and error-bound analysis, which we will leave for future study.
-When $\varphi$ is piecewise affine and $\Theta$ is a polytope, Assumption 2 holds naturally; furthermore, the nonasymptotic result holds without the need for Assumption 1 since a piecewise affine objective on a polytope has maximum descent rates bounded away from 0 at all points that are not d-stationary.
-Leveraging this special property, we can further establish the nonasymptotic convergence to d-stationarity (also local optimality in this case) of Algorithm 1 with $\varepsilon=0$, which reduces to the standard SMM with randomized index-mapping selection and achieves a faster rate with $\eta>0$.
-Details of the error-bound analysis and convergence proofs can be found in Appendix 10.3.
-Theorem 4.5 (Convergence of SMM)
-For problem (9) with $\varphi$ being piecewise affine and $\Theta$ being a polytope, under the same settings of Algorithm 1 specified in Theorem 4.4 except that $\varepsilon=0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$.
-Moreover, with $\eta>0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\mathcal{O}(T^{-\frac{1}{2}\min\left\{\alpha,1\right\}})$, which is up to a logarithmic factor when $\alpha=1$.
+Then any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ which must exist is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of problem (10) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$
+Theorem 3.6 shows the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity.
+Moreover, under the aforementioned settings where $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ and a regularity assumption on $F$, any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a global minimizer of (10) with probability 1.
+This indicates that the PADR-based ERM model together with the ESMM algorithm is asymptotically optimal to the original problem (2).
+However, evidenced by the numerical implementations, when $\varepsilon$ is large, the candidate solutions produced in step 4 of Algorithm 1 may be rejected by the update rule in step 5 very often, which results in inefficient computation and slow convergence.
+Such a computational issue is resolved by a shrinking-$\varepsilon$ strategy which will be specified in Section 4.
+Particularly, when $\varphi$ is piecewise affine and $\Theta$ is a polytope, we develop error-bound analysis and establish the nonasymptotic convergence to d-stationarity (also local optimality in this case) of Algorithm 1 by leveraging the special property that at all points within a polytope that are not d-stationary, the piecewise affine objective has descent rates bounded away from 0.
+Furthermore, the nonasymptotic convergence in this case is also assured with $\varepsilon=0$, in which the algorithm reduces to the standard SMM with randomized surrogation-index selection.
+Details of the error-bound analysis and convergence proofs can be found in Online Appendix 8.3.
+Theorem 3.7 (Nonasymptotic convergence of ESMM)
+For problem (10) with $\varphi$ being piecewise affine and $\Theta$ being a polytope, under the same settings of Algorithm 1 specified in Theorem 3.6 except that $\varepsilon\geq 0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$.
+Moreover, with $\varepsilon=0$ and $\eta>0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\mathcal{O}(T^{-\frac{1}{2}\min\left\{\alpha,1\right\}})$, which is up to a logarithmic factor when $\alpha=1$.
+Corollary 3.8
+Under the settings specified in Theorem 3.7, to achieve $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]\leq\epsilon$, the optimal iteration complexity of Algorithm 1 is $\widetilde{\mathcal{O}}(\epsilon^{-2})$, which holds with a sampling rate $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ where $\alpha\geq 2$.
+The nonasymptotic convergence to d-stationarity for general cases (including Case 2 and Case 3) can be obtained under additional assumptions on the objective $F$ and its surrogate functions, and we refer interested readers to the relative results with technical details in Online Appendix 8.4.
+We recognize that it is possible, albeit challenging, to achieve nonasymptotic results in terms of composite strong d-stationarity with carefully designed residuals and error-bound analysis, which we will leave for future study.
 Finally, we make several remarks on the convergence results.
@@
 (2022), due to the composite structure of objective functions, the SMM algorithm with $\varepsilon=0$ may yield convergence to compound criticality, which is weaker than d-stationarity.
-Pang et al.
-(2016, Example 4) also illustrates that an MM algorithm with $\varepsilon=0$ minimizing a piecewise quadratic objective could converge to a point isolated from the d-stationary set.
+This is also supported with Example 3.1 in Section 3.1.
 Remark 3.10
-According to Theorems 4.4 and 4.5, the optimal iteration complexity of Algorithm 1 for $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]\leq\epsilon$ is $\mathcal{O}(\epsilon^{-2})$, which is the same as that of the stochastic DC algorithm in Le Thi et al.
+By Corollary 3.8, the iteration complexity of Algorithm 1 is similar to that of the stochastic DC algorithm in Le Thi et al.
 (2022) with incremental sampling for differentiable stochastic DC programs, provided that convex subproblems are solved to the global optimum.
-Notably, the optimal complexity of Algorithm 1 is achieved with the sampling rate $N_{\nu}=\mathcal{O}(\nu^{2})$ that is faster than $\mathcal{O}(\nu)$ in Le Thi et al.
+Notably, the optimal complexity of Algorithm 1 is achieved with the sampling rate that is faster than $\mathcal{O}(\nu)$ in Le Thi et al.
 (2022) due to the nondifferentiable composite structure, and the convergence results of our algorithm hold for a much wider range of nonconvex and nondifferentiable composite problems.
-Moreover, with minor modifications of the convergence analysis, we may obtain the same convergence results for an inexact version of Algorithm 1, in which a $\delta_{\nu}$-suboptimal solution $\theta^{\nu+\frac{1}{2}}$ is computed in Step 4 and only accepted in Step 5 when it satisfies $\widehat{F}_{N_{\nu}}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},I_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{N_{\nu}}(\theta^{\nu})+\delta_{\nu}$ with $\delta_{\nu}=\mathcal{O}(\nu^{-1})$.
+Remark 3.11
+With minor modifications of the convergence analysis, we may obtain the same convergence results for an inexact version of Algorithm 1, in which a $\delta_{\nu}$-suboptimal solution $\theta^{\nu+\frac{1}{2}}$ is computed in Step 4 and only accepted in Step 5 when it satisfies $\widehat{F}_{\nu}(\theta^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}_{\nu}^{\varepsilon})+\frac{\eta}{2}\lVert\theta^{\nu+\frac{1}{2}}-\theta^{\nu}\rVert^{2}\leq F_{\nu}(\theta^{\nu})+\delta_{\nu}$ with $\delta_{\nu}=\mathcal{O}(\nu^{-1})$.
 In such a plausible extension, it is viable to design a two-loop algorithmic scheme addressing the computational complexity in solving convex subproblems.
@@
 Remark 3.12
-Since the output $\widetilde{\theta}^{\,t}$ of Algorithm 1 is uniformly sampled from the iterate sequence $\{\theta^{\nu}\}_{\nu=0}^{t-1}$, the almost sure convergence of $\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,t})$ to 0 in Proposition 4.3 implies that there exists a subsequence of $\{\theta^{\nu}\}_{\nu}$ with $\bar{r}^{\varepsilon,\rho}(\theta^{\nu})$ converging to 0 almost surely.
-With similar convergence analysis, we can obtain the almost sure existence of the limit point of $\{\theta^{\nu}\}_{\nu}$, which satisfies the same asymptotic convergence properties as the limit points of $\{\widetilde{\theta}^{\,t}\}_{t}$.
-Remark 4.9
-By Theorem 4.4 and the equivalence between $\mathcal{S}^{d}_{\varepsilon,\rho}$ and $\mathcal{S}^{*}$ stated in Proposition 10.2, with sufficiently large $\varepsilon$ and the proximal parameter $\eta$ close or equal to 0, any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a global minimizer of (9) with probability 1.
-This implies that the PADR-based ERM model together with the enhanced SMM algorithm under such settings is asymptotically optimal to the original problem (2).
-However, evidenced by the numerical implementations, the candidate solutions produced in Step 4 of Algorithm 1 with large $\varepsilon$ may be rejected in Step 5 very often, which results in inefficient computation and slow convergence.
-Such a computational issue is resolved by a shrinking-$\varepsilon$ strategy in Section 6.
-5 Extension: Constrained SP with Covariate Information
+Since the output $\widetilde{\theta}^{\,t}$ of Algorithm 1 is uniformly sampled from the iterate sequence $\{\theta^{\nu}\}_{\nu=0}^{t-1}$, by the almost sure convergence of $\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,t})$ to 0 in Proposition 3.5, we can easily obtain that there exists a subsequence $\{\theta^{\nu}\}_{\nu\in\kappa}$ with $\lim_{\nu(\in\kappa)\to\infty}\bar{r}^{\varepsilon,\rho}(\theta^{\nu})=0$ almost surely.
+With similar convergence analysis, we can obtain that the limit point of $\{\theta^{\nu}\}_{\nu\in\kappa}$ satisfies the same almost sure asymptotic convergence properties as the limit points of $\{\widetilde{\theta}^{\,t}\}_{t}$ presented in Theorem 3.6.
+4 Numerical Experiments
+In this section, we compare the numerical performance of the PADR-based ERM method with several benchmarks including DR-based methods, PO methods, and multiple prescriptive methods which are listed in Table 4.
+The experiments are conducted on Newsvendor problems with both simulated and real data and a product placement problem on a practical scale.
+For the simulated data, we set the true model of $Y$ as $Y=\overline{Y}(X)+\widetilde{\epsilon}$, where $\overline{Y}(\cdot)$ represents the mean model, and $\widetilde{\epsilon}$ denotes a random variable with a standard normal distribution independent of $X$.
+Throughout all experiments with simulated data, we generate feature data uniformly from $[-1,1]^{p}$, and ensure the mean model satisfies $\min_{x\in[-1,1]^{p}}\overline{Y}(x)\geq 5$ which keeps all samples of $Y$ positive.
+Unless otherwise noted, both training and test datasets contain 1000 samples.
 […]
+Methods implemented in numerical experiments Abbreviation Explanation of method PADR Piecewise-affine decision rule method in this paper LDR Linear decision rule method from Ban and Rudin (2018) RKHS-DR Kernel optimizer prediction method from Bertsimas and Koduri (2021) kNN, kernel, cart, rf Prescriptive methods from Bertsimas and Kallus (2019) based on $k$-nearest neighbiors, kernels, regression trees, and random forests respectively sof Prescriptive method from Kallus and Mao (2023) based on stochastic optimization forests PO-PA, PO-L PO methods with piecewise-affine prediction model and linear prediction model respectively SIMOPT Simulated optimal decision using true conditional distribution
+We refer to the PADR model with two max-affine components consisting of $K_{1}$ and $K_{2}$ pieces respectively as PADR$(K_{1},K_{2})$.
+The hyperparameters $K_{1}$ and $K_{2}$ are searched within the range [0, 4] for the Newsvendor problems in Section 4.1 and Section 4.2, and [0, 6] for the product placement problem in Section 4.3.
+Regarding the ESMM algorithm for learning the PADR, initial points are generated uniformly from $[-50,50]^{d(K_{1}+K_{2})(p+1)}$, and the sampling sequence $\{N_{\nu}\}_{\nu}$ follows $N_{\nu}=40\nu$.The parameter $\varepsilon$ in Algorithm 1 is set to be either constant or shrinking, the latter case of which is referred to as the shrinking-$\varepsilon$ strategy.
+Specifically, the shrinking strategy sets the parameter a constant value $\varepsilon_{0}$ for the first $T_{0}$ iterations and then a reduced value $\varepsilon_{1}(<\varepsilon_{0})$ for the remaining ones.
+The ESMM algorithm produces the iterate point obtaining the minimum ERM cost as the output after $T$ iterations.
+For each instance, the algorithm is executed for $r$ rounds with random initialization, and we report the best performance unless otherwise noted.
+We set $T=10$ and $r=10$ in Sections 4.1 and 4.2, and $T=20$ and $r=5$ in Section 4.3.
+The reported running time of PADR is the total time of all rounds.
+We clarify settings for benchmark methods listed in Table 4.
+We use a Gaussian kernel for kernel methods.
+For RKHS-DR, we adjust the diagonal of the kernel by values no larger than $10^{-3}$ so as to resolve numerical issues without undermining performance.
+For PO methods, we use $\ell_{2}$ loss in the prediction stage and solve the regression problem of PO-PA using the ESMM algorithm.
+We set the uniform scale limit $\mu=50$ in Sections 4.1 and 4.2, and $\mu=100$ in Section 4.3.
+The hyperparameters of all benchmarks methods listed in Table 4 are selected via a standard validation process.
+The performance of LDR, PO-L, and kernel is dominated by PADR, PO-PA, and other prescriptive methods in most cases, and thus their results are omitted from the report if necessary to avoid the distraction.
 […]
-6 Numerical Experiments
-In this section, we evaluate the performance of the PADR-based ERM method through synthetic experiments on the Newsvendor problem and its variations with nonconvex costs or nonconvex constraints.
-We compare our method with PO methods and other two prominent DR-based methods.
-PO methods apply linear and PA regression models (referred to as PO-L and PO-PA respectively) and the squared error loss in the prediction stage; DR-based methods include NV-ERM1 method from Ban and Rudin (2018) that employs LDR and the optimizer prediction method from Bertsimas and Koduri (2021) that employs RKHS-DR.
-We omit comparisons with other data-driven approaches in the integrated scheme such as the objective prediction method from Bertsimas and Koduri (2021) and prescriptive methods from Bertsimas and Kallus (2019) because numerical results in Bertsimas and Koduri (2021) indicate that both two are outperformed by the optimizer prediction method in various cases.
-The Newsvendor problem with covariate information $X=x$ is formulated as follows,
+Hyperparameters of each method Methods Hyperparameters PADR number of pieces (PADR); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM) LDR N/A RKHS-DR kernel bandwidth, regularization parameter, penalty parameter kNN, kernel, cart, rf number of nearest neighbors (kNN); kernel bandwidth, regularization parameter (kernel); depth, minimum samples to split, minimum samples per leaf (cart and rf); sof choice of apx-soln or apx-risk method in Kallus and Mao (2023), depth, minimum samples to split, minimum samples per leaf PO-PA number of pieces (prediction model); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM solving $\ell_{2}$ regression) PO-L N/A SIMOPT N/A
+All methods are implemented in Python 3.9.7 on a MacBook Pro with a macOS Monterey system and an Apple M1 Pro chip.
+All programs involved are built and solved by the optimizer Gurobi, version 9.5.0.
+We provide the code on GitHub: https://github.com/zhangyiy0223/PADR.
+4.1 Newsvendor problem: simulated data
+We consider a Newsvendor problem with covariate information $X=x$ as follows,
 $\min_{z\geq 0}\quad\mathbb{E}\left[c_{b}(Y-z)_{+}+c_{h}(z-Y)_{+}\mid X=x\right],$ (16)
 where $c_{b}$ and $c_{h}$ are the unit back-order and holding costs respectively, and $(\cdot)_{+}$ denotes $\max\{\cdot,0\}$.
-We set the true demand as $Y=\overline{Y}(X)+\epsilon$, where $\overline{Y}(\cdot)$ is the mean model, and $\epsilon$ is a random variable with the standard normal distribution independent of $X$.
-Throughout all experiments, we generate the feature data uniformly from $\mathcal{X}=[-1,1]^{p}$, and we set the mean model satisfying $\min_{x\in\mathcal{X}}\overline{Y}(x)\geq 5$ to keep all corresponding demand data positive.
-Unless otherwise noted, both training and test data sets contain 1000 samples.
-We clarify the settings throughout this section.
-The PADR model with two max-affine components consisting of $K_{1}$ and $K_{2}$ pieces respectively is referred to as PADR$(K_{1},K_{2})$, and we implement multiple settings of $(K_{1},K_{2})$ from $\{0,1,2,3\}^{2}$.
-We set the uniform scale $\mu=50$ for parameters of PADR, PO, and LDR methods.
-In the implementation of the enhanced SMM algorithm, initial points are generated uniformly from $[-\mu,\mu]^{d(K_{1}+K_{2})(p+1)}$, and the sequence $\{N_{\nu}\}_{\nu}$ satisfies $N_{\nu}=\beta_{1}\nu+\beta_{2}$ with $\beta_{1},\beta_{2}>0$.
-Besides Algorithm 1 with constant $\varepsilon$, we also implement a version with a shrinking-$\varepsilon$ strategy, in which $\varepsilon$ is set as $\varepsilon_{0}$ at the first $T_{0}$ iterations and $\varepsilon_{1}(<\varepsilon_{0})$ at the remaining ones.
-We set the iterate point that obtains the minimum ERM cost in $T=10$ iterations as the output of the algorithm.
-For each instance, the algorithm is executed for $10$ rounds with random initialization, and we report the best performance unless otherwise noted.
-The hyperparameters are selected via a validation process with searching ranges listed in Table 6.
+We set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $k=1$ unless otherwise specified:
+$\overline{Y}(X)=k\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (17)
+4.1.1 Basic results.
 […]
-Hyperparameter Selection for PADR-based ERM and Enhanced SMM.
-Hyperparameters Description Searching Ranges $\gamma$ the parameter for the constraints in (5) $[0,1]$ $\lambda$ the penalty parameter for exact penalization $[0,10^{3}]$ $\varepsilon_{0},\varepsilon_{1},T_{0}$ parameters for the shrinking-$\varepsilon$ strategy $[0,10^{4}],[0,10^{4}],\{1,2,\dots,6\}$ $\beta_{1},\beta_{2}$ parameters for the sampling strategy $[5,50],[10,40]$ $\eta$ the proximal parameter $[0,1]$
-We use the Gaussian kernel in RKHS-DR, and the diagonal of the kernel is added by values no larger than $10^{-3}$ so as to resolve numerical issues without undermining its performance.
-The kernel bandwidth, regularization parameter, and penalty parameter are selected via a validation procedure searching from $[10^{-1},10^{3}]$, $[10^{-8},1]$, and $[0,10^{3}]$ respectively.
-We approximate the optimal cost by the empirical cost of the theoretical (in Section 6.1) or simulated (in Sections 6.2 and 6.3) optimal decision, which is treated as the baseline and referred to as SIMOPT.
-For the training time comparisons, we report the total time of the enhanced SMM algorithm for 10 rounds with predetermined hyperparameters and the average time of RKHS-DR over all hyperparameter settings since the running time of RKHS-DR is highly related to the generated kernel and thus quite unstable.
-The training time of PO-L and LDR for each instance is less than 0.5 seconds and thus is omitted.
-All methods are implemented in Python 3.9.7 on a MacBook Pro with a macOS Monterey system and an Apple M1 Pro chip. All programs involved are built and solved by the optimizer Gurobi, version 9.5.0.
-6.1 Unconstrained Newsvendor Problem
-For the Newsvendor problem (11), we set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $k=1$ unless otherwise specified:
-$\overline{Y}(X)=k\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (12)
-a.
-Basic Results.
-In Figure 1a, we report comparison results with the training sample size $n$ increasing from 50 to 1000.
-For a fair comparison, we set the PA function class with $(K_{1},K_{2})=(3,0)$ as the regression model in PO-PA, which contains the true model $\overline{Y}(\cdot)$; the parameters of the PA regression model are estimated by solving the corresponding ERM problem via the enhanced SMM algorithm with the same hyperparameter selection procedure as PADR methods.
-We also implement the quadratic DR and cubic DR as the generalized LDRs, denoted by GLDR-2 and GLDR-3 respectively, the running time of which is less than 0.5 seconds and thus is omitted.
-(a) Basic Results
-(b) Enhanced SMM with Different $\varepsilon$-Strategies
-Unconstrained Newsvendor Problem: Basic Results We also test the performance of PO-L and LDR in this instance, whose costs (around 14 and 9, respectively) are excluded in (a) for readability.
-In (b), solid and dotted lines represent constant-$\varepsilon$ strategies; dashed lines represent shrinking-$\varepsilon$ strategies with $T_{0}=3$ and $\varepsilon_{1}=0$.
-The costs of PADR(3,0) and PADR(3,3) converge to SIMOPT because the optimal DR in this case is a max-affine function with three pieces, and thus PADR(3,0) and PADR(3,3) have no approximation error in principle.
-The RKHS-DR method converges slower than PADR(3,0) and PADR(3,3); moreover, its performance with $n=1000$ is slightly worse than that of PADR(3,0) with $n=400$.
-For PO-PA, given the observation that the prediction model provided by the enhanced SMM algorithm is quite close to the true model $\overline{Y}(\cdot)$, we conjecture that its gap to SIMOPT is mainly due to inconsistency between the prediction loss and Newsvendor cost with imbalanced $c_{b}$ and $c_{h}$.
-This phenomenon is more evident from the comparison between LDR and PO-L, the two methods with misspecified models.
-The training time of RKHS-DR grows roughly exponentially with respect to the sample size, whereas that of PA-related methods is quite stable and significantly shorter when $n=1000$, which can be mainly credited to the sampling strategy used in the enhanced SMM algorithm.
-The numerical results with the parameter ratio $c_{b}:c_{h}$ varying from 1:9 to 9:1 exhibit similar patterns and thus are omitted.
-Figure 1b displays sorted training cost curves of PADR(3,3) provided by the enhanced SMM with constant-$\varepsilon$ and shrinking-$\varepsilon$ strategies over 30 rounds.
-The shrinking-$\varepsilon$ strategy significantly improves the performance of the algorithm compared to the constant-$\varepsilon$ strategy, which typically does not reach satisfactory convergence within 10 iterations.
-This indicates that despite the theoretical convergence of the algorithm to global optimality with sufficiently large $\varepsilon$, it needs to engage with a huge number of index mappings when the sample size is large and thus can hardly make a valid update in each iteration.
-In contrast, the shrinking-$\varepsilon$ strategy with $\varepsilon_{0}=3000$ approximately attains the optimal value in 18 rounds.
-The significant efficiency improvement is because the shrinking-$\varepsilon$ strategy enables flexible adjustments of iterate points at early iterations as a warm start and afterward accelerates the convergence by ensuring valid updates with $\varepsilon=0$.
-b.
-Varying Feature Dimensions.
-We study the effect of the feature dimension $p$ on the performance of PADR methods under sparse and dense demand models $\overline{Y}$ following the settings in Bertsimas and Koduri (2021) and report the results in Figure 1d.
-(c) Sparse Demand Model
-(d) Dense Demand Model
-Unconstrained Newsvendor Problem with Varying Feature Dimensions We also test the performance of PO-L and LDR under dense demand models, whose costs (decreasing from 15 to 5 and 9 to 4, respectively) are excluded in (b) for readability.
-In the sparse case, only the first two components of $X$ contribute to $\overline{Y}$ following the demand model (12); in the dense case, we replace $X_{1}$ in (12) with the average of the first half of the features $\frac{2}{p}\sum_{i=1}^{p/2}X_{i}$ and $X_{2}$ with the average of the other half.
-Figure 1d implies that PADR and PO-PA are generally more robust to redundant features than RKHS-DR in the sparse case, and that PADR performs better than RKHS-DR and PO-PA in the dense case.
-The performance of PADR(3,3) and PADR(3,0) deteriorates as $p$ increases due to the growing estimation error and optimization difficulty.
-The simpler PADR(2,0) performs more stable in both cases, despite the existence of an approximation error.
-c.
-Nonlinearity in Demand Model.
-We study the effect of nonlinearity in the demand model (12) with varying $k$ and report results in Figure 1f.
-It indicates that the performance of PADR(3,0), PADR(3,3), and PO-PA is irrelevant to the nonlinearity level of the demand model, whereas that of PADR(2,0) and RKHS-DR deteriorates as the nonlinearity level increases.
-(e) $p=2\phantom{xxxxxxx}$
-(f) $p=20$
-Unconstrained Newsvendor Problem with Varying Nonlinearity in Demand Model In (b), the demand model is of the dense type.
-We also test the performance of PO-L and LDR, whose costs are excluded for readability.
-When $p=2$, PO-L and LDR increase from 5 to 40 and 3 to 30, respectively; when $p=20$, PO-L and LDR increase from 4 to 16 and 3 to 13, respectively.
-Besides the piecewise affine demand model (12), we also test the performance of the DR-based methods with another nonlinear demand model that includes a sine function to account for seasonality.
-The numerical results are provided in Appendix 12.
-6.2 Constrained Newsvendor Problem
-We next consider a two-product Newsvendor problem with a total inventory capacity $C_{0}$.
-Assume that the uncensored historical demand data is available.
-We denote the unit back-order cost, unit holding cost, and order decision by $c_{bi}$, $c_{hi}$, and $z_{i}$ respectively for product $i=1,2$.
-The capacity constraint is formulated as $z_{1}+z_{2}\leq C_{0}$.
-In this subsection, we set $(c_{b1},c_{h1},c_{b2},c_{h2})=(8,2,2,8)$, $C_{0}=60$, and the mean demand models for the two products as
-$\overline{Y}_{1}(X)=15X_{1}-5X_{2}+30,\quad\overline{Y}_{2}(X)=15X_{1}+5X_{2}+30,$
-under which the probability that the constraint is active at the optimal decision is around 0.5.
-We compare PADR(2,2) with PO-L and the penalized RKHS-DR method from Bertsimas and Koduri (2021, Section 6.2).
-Since DR-based methods may yield infeasible decisions, in Figure 1i we report the feasible frequency of decisions provided by PADR(2,2) and RKHS-DR.
-By projecting the decisions to the convex feasible region when needed, we report the test cost of the feasible decisions in Figure 1i.
-(g) Basic Results
+(d) Basic results
+(e) Running time
+Newsvendor problem: basic results and running time
+Since the optimal DR in this case is a max-affine function with three pieces, PADR$(K_{1},K_{2})$ with $K_{1}\geq 3$ has no approximation error in principle.
+Figure 1d shows that the costs of PADR with $(K_{1},K_{2})=(3,0)$ converge to SIMOPT as the training sample size $n$ increases from 50 to 1000.
+RKHS-DR converges slower than PADR; additionally, its performance with $n=1000$ is worse than that of PADR with $n=400$.
+Prescriptive methods, including kNN, cart, rf, and sof, are dominated by the PADR and RKHS-DR methods.
+For PO-PA, we observe that the prediction model provided by the PO-PA method is quite close to the true model $\overline{Y}(\cdot)$, so its substantial gap to SIMOPT could be mainly due to inconsistency between the prediction loss and Newsvendor cost with imbalanced $c_{b}$ and $c_{h}$.
+Numerical results with the parameter ratio $c_{b}:c_{h}$ varying from 1:9 to 9:1 exhibit similar patterns and thus are omitted.
+Figure 1e shows the training time of each method.
+The training time of RKHS-DR grows roughly exponentially when $n\geq 200$, whereas the time for PA-related methods remains stable and significantly shorter when $n=1000$, which is mainly credited to the enhanced and sampling techniques used in the ESMM algorithm.
+[…]
+(f) $\varepsilon$-Strategies
+(g) Sampling technique
+Performance of ESMM under various settings In (a), solid and dotted lines represent constant-$\varepsilon$ strategies; dashed lines represent shrinking-$\varepsilon$ strategies with $T_{0}=3$ and $\varepsilon_{1}=0$.
+In (b), experiments for $p=20$ are conducted on datasets using the sparse demand model introduced in Section 4.1.2.
+The y-axis is set on a log scale.
+In Figure 1g, we provide results of the ESMM algorithm under different parameter settings that illustrate the effect of the enhanced sampling strategies.
+Figure 1f displays curves of sorted training cost of PADR provided by the ESMM algorithm under the constant-$\varepsilon$ and shrinking-$\varepsilon$ strategies over 30 rounds.
+Under the constant-$\varepsilon$ strategy, the algorithm converges to stationary solutions which can be strengthened to global optimality with $\varepsilon$ being sufficiently large.
+However, Figure 1f shows that the constant-$\varepsilon$ strategy typically yields poor performance over all 30 rounds with 10 iterations.
+We conjecture that this is because when the sample size is large, the algorithm requires engaging with the random selection among a huge number of active indices and thus hardly makes a valid update at each iteration.
+In contrast, the shrinking-$\varepsilon$ strategy enables flexible adjustments of active indices for early iterations with a large positive value of $\varepsilon$ and produces iterate points serving as a warm start; afterward, it accelerates the convergence with a decreased value of $\varepsilon$.
+Figure 1f shows that the shrinking-$\varepsilon$ strategy has significantly better performance which attains the global optimum among roughly half of the total rounds.
+Figure 1g reports the running time of PADR using the ESMM algorithm and the enhanced MM algorithm (EMM) without sampling (including all samples in the subproblem of each iteration).
+We use PADR(3,0) and PADR(6,4) to represent models with low and high complexity, denoted by PADRL and PADRH, respectively.
+The sampling technique significantly saves computational costs without compromising performance.
+The time consumption of ESMM is primarily determined by the sampling rate and remains relatively stable when $n\geq 400$, while the time required for EMM increases exponentially with the training sample size.
+The benefit of the sampling technique becomes more pronounced as the decision rule complexity increases.
+The time required for EMM is approximately 4 times longer than that for ESMM when $n=1000$.
+4.1.2 Varying feature dimensions.
+We study the effect of the feature dimension of $X\in\mathbb{R}^{p}$ on the performance of PADR under sparse and dense demand models of $\overline{Y}$ following the settings in Bertsimas and Koduri (2021).
+Specifically, in the sparse case, only the first two components of $X$ contribute to $\overline{Y}$ following the demand model (17); in the dense case, we replace the two feature variables in (17) by the average of the first half of the features $\frac{2}{p}\sum_{i=1}^{p/2}X_{i}$ and the average of the other half respectively.
+[…]
 (h) Sparse demand model
 (i) Dense demand model
-Newsvendor Problem with a Capacity Constraint
-With the basic demand model, PADR(2,2) converges to SIMOPT and attains 100% feasibility when $n\geq 200$.
-In comparison, RKHS-DR and PO-L display notable gaps to SIMOPT, and the feasible frequency of RKHS-DR is only around 80%.
-Figures 1h and 1i display numerical results with varying feature dimensions.
-In the sparse case, PO-L is markedly better than the other two methods when $p=50$, indicating that regularization for sparsity is needed for learning PADR and RKHS-DR.
-In the dense case, PADR(2,2) performs the best except when $p=50$.
-6.3 Nonconvex Newsvendor Problem
-In this section, we consider two variations of the Newsvendor problem by adding a concave cost $C(z)$ in the objective or as a constraint.
-The cost $C(z)$ added in the objective can be interpreted as the capacity acquisition cost (Atamtürk and Hochbaum 2001); it also can be restricted in a constraint as a resource limitation.
-In the following experiments, we set the concave cost as
-$C(z)=\min\{z,0.6z+0.8,0.4z+15.6\},$
-and then both two variant problems of interest are nonconvex.
-We compare the PADR methods with RKHS-DR and use the deterministic MM algorithm to solve the nonconvex ERM problem for learning RKHS-DR, referred to as RKHS-DR(MM).
-The baseline SIMOPT is computed by solving simulated nonconvex problems via a solver of mix-integer linear programming.
-a.
-Nonconvex Objective.
-We first consider a generalized single-product Newsvendor with $(c_{b},c_{h})=(5,5)$ and $C(z)$ added in the objective of (11).
-The numerical results in Figure 1l indicate that PADR(3,0) performs the best in almost all the settings, which is consistent with the convex case in Section 6.1; moreover, the gap between PADR(3,0) and RKHS-DR(MM) in the sparse case with $p=50$ is much more significant.
-(j) Basic Results
-(k) Sparse Demand Model
-(l) Dense Demand Model
-Newsvendor Problem with a Nonconvex Objective
-b.
-Nonconvex Constraint.
-We next consider the 2-product Newsvendor problem given in Section 6.2 with parameters $(c_{b1},c_{h1},c_{b2},c_{h2})=(7,7,3,3)$ and the replacement of the convex constraint by $C(z_{1})+C(z_{2})\leq C_{0}$.
-We set $C_{0}=50$ to keep the same active probability for the constraint.
-Since the projection of infeasible decisions to the nonconvex feasible region may not yield meaningful solutions, we only report the average test cost of feasible decisions in Figure 1o, with a careful hyperparameter selection to ensure high feasible frequencies.
-(m) Basic Results
-(n) Sparse Demand Model
-(o) Dense Demand Model
-Newsvendor Problem with a Nonconvex Constraint In (b) and (c) with $p=50$, RKHS-DR(MM) takes more than 500 seconds on average, while PADR(2,2) takes less than 200 seconds.
-For the problem with the basic demand model (12), PADR(2,2) performs comparably with RKHS-DR(MM), with a notable gap to SIMOPT.
-Figures 1n and 1o indicate the superiority of PADR(2,2) over the other two methods under both sparse and dense demand models with varying feature dimensions.
-It’s also worth noting that the simpler PADR(1,0) (equivalent to LDR), with poor performance in general, performs comparably well with the others in the high-dimensional case.
-It is more difficult to obtain feasible solutions for cases with sparse models than those with dense models; moreover, with sparse models, PADR(1,0) is more likely to produce feasible solutions than the others.
-Finally, we also conduct experiments on constrained Newsvendor problems in Sections 6.2 and 6.3 with multiple $C_{0}$ values and Newsvendor cost parameters.
-The numerical results exhibit similar patterns and thus are omitted.
+Newsvendor problem with varying feature dimensions For readability, kNN is excluded under the dense demand model, whose costs exceed 8 in most trials.
+Figure 1i displays the results.
+In the sparse case, forest-based prescriptive methods are robust to redundant features, as they only focus on informative features.
+When the dimension is less than 20, PADR method is competitive with comparable performance to sof and much less training time.
+It would be beneficial to employ an effective feature selection technique or regularization in the PADR method to improve the performance in the sparse case with a large feature dimension.
+In the dense case, PADR performs generally well, but its performance slightly deteriorates as $p$ increases due to growing estimation errors and optimization difficulties.
+We also observe that PADR with fewer pieces performs more stably in both cases despite the presence of approximation errors.
+4.1.3 Nonlinearity in the demand model.
+We study the effect of nonlinearity of the demand with respect to the feature variables by varying the coefficient $k$ in the demand model (17) and report results in Figure 1k.
+It indicates that the performance of PADR and PO-PA is irrelevant to the nonlinearity level of the demand model due to the strong approximation ability to the nonlinear structure.
+In contrast, the performance of RKHS-DR and prescriptive methods deteriorates significantly as the nonlinearity level increases.
+These methods require significantly more samples to achieve the similar performance when the level of nonlinearity increases.
+[…]
+(j) $p=2\phantom{xxxxxxx}$
+(k) $p=20$
+Newsvendor problem with varying nonlinearity in demand model In (b), the demand model is of the dense type. For readability, kNN, with costs increasing from 5 to 40, is excluded in (a).
+4.2 Newsvendor problem with nonconvex cost: real-world data
+In this section, we conduct experiments on a real-world dataset, which consists of order records for a specific model of mobile phone from a Chinese electronic product supplier.
+The order record, spanning from September 01, 2021 to November 31, 2023, contains order quantities and time stamps from retailers.
+In order to create a production plan, the supplier needs to determine the packing quantity in each day to fulfill the order requests.
+We first organize the individual order records into daily order quantities.
+The aggregated daily order quantities have been rescaled for confidentiality purposes and also normalized for pretreatment.
+The boxplots of daily orders in Figure 1n show that order quantities are larger on weekdays and during the second quarter of each month, which aligns with retailers’ ordering pattern.
+[…]
+(l) Weekdays
+(m) Month days
+(n) Month quarters
+Real-world dataset: boxplots of orders by time period
+We set observations from the first year as the training set and the remaining data as the test set.
+Here we consider three types of features, namely weekday, month quarter, and order quantities in past days.
+The first two types of features are converted into dummy variables (7 for weekdays and 4 for month quarters) and set as basic features.
+For the third type of features, we incorporate order quantities in past days of different lengths as different instances in conducting the experiments.
+To determine the packing quantity in each day to fulfill the order requests, we formulate a Newsvendor problem with features listed above and report the numerical results in Table 4.2.
+It shows that PADR and forest-based prescriptive methods work well.
+Overall, the PADR method performs slightly better in most instances with features of different length.
+[…]
+Test costs for newsvendor on a real-world dataset Features $p$ DR methods Prescriptive methods PO methods PADR RKHS-DR kNN cart rf sof PO-PA basic features 11 55.95 56.10 56.06 56.00 56.00 56.00 79.02 basic features + order quantities in past 1 day 13 42.67 51.53 50.25 61.72 49.54 48.70 68.18 basic features + order quantities in past 3 days 15 41.65 55.59 48.25 54.64 41.74 42.19 59.73 basic features + order quantities in past 5 days 17 41.40 65.17 49.59 56.74 40.90 41.83 52.39 basic features + order quantities in past 7 days 19 40.93 70.60 49.02 65.62 43.62 43.51 54.52 basic features + order quantities in past 14 days 26 41.03 78.37 46.81 58.39 42.45 42.48 52.10 basic features + order quantities in past 21 days 33 42.43 80.24 46.61 59.45 41.59 41.33 53.44 The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
+We next consider a practical setting where the supplier incurs an additional cost associated with packing quantities in actual production operations, in order to balance the daily inventory cost and the packing cost.
+The cost is approximated by the following nonconvex cost function:
+$C(z)=\min\{z,0.6z+0.8,0.4z+15.6\},$ (18)
+Therefore, we consider a Newsvendor problem with an additional concave cost $C(z)$ in the objective.
+Since the concave cost introduces nonconvexity into the problem.
+we use the deterministic MM algorithm to solve the ERM problem for learning RKHS-based decision rule, referred to as RKHS-DR(MM).
+For prescriptive methods, we solve one-dimensional nonconvex problems by line search.
+We do not implement sof from Kallus and Mao (2023) in this nonconvex Newsvendor problem, as it requires a new development of second-order perturbation analysis of stochastic optimization, which is beyond the scope of this paper.
+[…]
+Test costs for nonconvex newsvendor on a real-world dataset
+Features $p$ DR methods Prescriptive methods PO methods
+PADR RKHS-DR(MM) kNN cart rf PO-PA
+basic features 11 74.28 81.89 84.52 81.56 81.56 90.48
+basic features + order quantities in past 1 day 13 64.75 72.11 77.86 86.58 73.78 77.89
+basic features + order quantities in past 3 days 15 64.10 68.29 75.58 78.31 64.99 73.61
+basic features + order quantities in past 5 days 17 63.20 64.72 76.81 79.54 65.74 71.56
+basic features + order quantities in past 7 days 19 62.01 65.35 75.38 85.95 68.35 69.97
+basic features + order quantities in past 14 days 26 64.86 65.15 73.81 75.52 64.71 70.80
+basic features + order quantities in past 21 days 33 65.65 65.15 73.81 75.52 65.73 71.02
+The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
+[…]
+4.3 Product placement problem
+We consider a product placement problem on a graph $\mathcal{G}=(\mathcal{N},\mathcal{A})$ with the aim to determine the quantities of products $z\in\mathbb{R}^{\left\lvert\mathcal{N}\right\rvert}$ to place at different nodes.
+After placing the products, demands $y\in\mathbb{R}^{\left\lvert\mathcal{N}\right\rvert}$ are observed at nodes $\mathcal{N}$, and then products are shipped across arcs $\mathcal{A}$ to fulfill these demands.
+We denote the node-arc matrix of $\mathcal{G}$ by $W$, the product placement cost by $c\in\mathbb{R}^{\left\lvert\mathcal{N}\right\rvert}$, the shipment cost along edges by $g\in\mathbb{R}^{\left\lvert\mathcal{A}\right\rvert}$, and the penalty for unmet demand by $b\in\mathbb{R}^{\left\lvert\mathcal{N}\right\rvert}$.
+The problem is formulated as follows,
+$\min_{z\geq 0}\quad\mathbb{E}\left[c^{\top}z+V(z,Y)\,\middle|\,X=x\right],$
+where $V(z,Y)=\min_{f\geq 0}\left\{g^{\top}f+b^{\top}(Y+Wf-z)^{+}\right\}$.
+This is a two-stage stochastic linear program, and its ERM problem can be converted into a large linear program by adding the second-stage decision variables for each scenario.
+Following the graph generation settings in Bertsimas and Koduri (2021), we set $\left\lvert\mathcal{N}\right\rvert=5$, $\left\lvert\mathcal{A}\right\rvert=10$, $c=(3,\dots,3)$, $g=(1,\dots,1)$, and $b=(5,\dots,5)$.
+To evaluate PADR’s performance under more complex conditions, we employed the following demand model:
+$\overline{Y}^{\prime}(X)=k\alpha_{1}\sin(\alpha_{2}X_{1})+k\max\{\alpha_{3}X_{2},\alpha_{4}X_{2}\}+\alpha_{5},$ (19)
+where the sine function represents the seasonal dependence of the demand on $X_{1}$, and the max-affine function represents the nonlinear trend with respect to $X_{2}$.
+We set $k=1$ unless otherwise specified.
+We randomly generate parameters $\{\alpha_{i}\}_{i=1}^{5}$ and the node-arc matrix $W$ for each trial, and we report performance of methods over 20 trials for each experimental setting.
+The sof method from Kallus and Mao (2023) is not applied for this two-stage SP, as it would require a new development of second-order perturbation analysis of stochastic optimization, which is beyond the scope of this paper.
+4.3.1 Varying feature dimensions.
+We first study the effect of the feature dimension $p$ on the performance of PADR under sparse and dense demand models similar to the discussion in Section 4.1.2.
+The results are reported in Figure 1p.
+Similar to the results in Section 4.1.2 for the Newsvendor problem, with sparse demand model, PADR performs well in low-dimensional cases but deteriorates as $p$ increases which may necessitate the proper parameter regularization.
+In contrast, forest-based prescriptive methods are robust to redundant features.
+With dense demand model, PADR and PO-PA have the best performance, with the former one slightly outperforming when the feature dimension increases to 50.
+Despite the increased complexity and the challenge in accurately approximating the true model, PADR continues to demonstrate superior performance for instances with dense demand model, high nonlinearity, and sparse demand models with small feature dimensions.
+[…]
+(o) Sparse demand model
+(p) Dense demand model
+Product placement problem with varying feature dimensions
+4.3.2 Nonlinearity in demand model.
+We also study the effect of nonlinearity in the demand model by varying the coefficient $k$.
+Notice that the nonlinear structure of the demand model here is different from the one in Section 4.1 with the sine component function.
+The results for $p=20$ are provided in Figure 1r.
+The increasing gap in Figure 1q indicates that except for PADR and PO-PA, the performance of all the other methods deteriorates as the nonlinearity of the demand model increases.
+Figure 1r indicates that PADR is slightly better than PO-PA and outperforms the other methods, and the relative gap among them remains approximately stable as $k$ increases.
+[…]
+Product placement problem: varying nonlinearity level $k$ (dense model, $p=20$) In (b), with the cost of the cart scaled to 1 and the cost of SIMOPT set to 0, we plot the relative cost of other methods compared to SIMOPT.
 5 Conclusion
-We propose the PADR-based ERM method together with the enhanced SMM algorithm to solve a broad class of SP problems with covariate information.
-For our PADR-based ERM framework, we provide the approximation bound of the PA function class and further establish the nonasymptotic consistency result for unconstrained problems and the asymptotic consistency result for deterministic-constrained ones.
-To solve the ERM problem with a composite DC structure, we develop the enhanced SMM algorithm by combining enhanced $\varepsilon$-active surrogation with sequential sampling.
-We provide the asymptotic convergence of the enhanced SMM algorithm to composite strong d-stationarity, and the nonasymptotic convergence to d-stationarity via the error-bound analysis.
-Numerical experiments indicate the superiority of our method over PO methods and prominent DR-based methods in the Newsvendor problem regarding the test cost, computational time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
+We propose the PADR-based ERM method together with the ESMM algorithm to solve a broad class of SP problems with covariate information.
+For our PADR-based ERM framework, we provide the approximation bound of the PA function class and further establish the nonasymptotic consistency result for unconstrained problems.
+To solve the ERM problem with a composite DC structure, we develop the ESMM algorithm by combining enhanced $\varepsilon$-active surrogation with sequential sampling.
+We provide the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity, and the nonasymptotic convergence to d-stationarity via the error-bound analysis.
+Numerical experiments indicate the advantages of our method over PO methods and prominent DR-based methods for various instances regarding the test cost, computational time, and robustness to nonlinearity of the underlying dependency.
 In conclusion, the PADR is a favorable hypothesis class for DR-based ERM methods, supported by theoretical consistency, algorithmic convergence, and computational efficiency for solving SP with covariate information.
@@
 6.1 Proof of Proposition 2.1
-The idea of the proof is to construct an interpolation PA function to a set of points related to the target function $f_{0}$ and a cover of the set $\mathcal{X}$.
-The structure of the interpolation function is motivated by Siahkamari et al.
-(2020) but with a different parameter, which ensures the Lipschitz continuity of the function with modulus independent of the approximation accuracy.
-Proof 8.1
+Proof 6.1
 […]
 $\left(\frac{\sqrt{p}\bar{X}}{\epsilon}\right)^{p}\leq\left\lvert\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\right\rvert\leq\left(\frac{\sqrt{p}\bar{X}}{\epsilon}+1\right)^{p}\leq\left(\frac{2\sqrt{p}\bar{X}}{\epsilon}\right)^{p}.$
-Given $f_{0}\in\mathcal{F}_{L_{0},M_{0}}$, by McShane (1934, Theorem 1), we can extend $f_{0}$ to an $L_{0}$-Lipschitz function $\bar{f}_{0}:\bar{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying $\left\lVert\bar{f}_{0}\right\rVert_{\infty}\leq M_{0}$ and $\bar{f}_{0}(x)=f_{0}(x)$ for any $x\in\mathcal{X}$.
-Next, we construct a piecewise affine interpolation function to a set of points $\{(\widehat{x},\bar{f}_{0}(\widehat{x}))\mid\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\}$:
+[…]
 $f_{\widehat{\mathcal{C}}}(x)\coloneqq$ $\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}-\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\},$ (20)
@@
 Definition 6.2
-Let $\Phi$ be a class of functions mapping from $\mathcal{U}$ to $\mathbb{R}$ and $S_{N}=\{u^{s}\}_{s=1}^{N}\subseteq\mathcal{U}$ be a sample set of size $N$.
-Then the empirical Rademacher complexity of $\Phi$ with respect to $S_{N}$ is defined as
-^$\widehat{\mathcal{R}}_{S_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
+Let $\Phi$ be a class of functions mapping from $\mathcal{U}$ to $\mathbb{R}$ and $\Xi_{N}=\{u^{s}\}_{s=1}^{N}\subseteq\mathcal{U}$ be a sample set of size $N$.
+Then the empirical Rademacher complexity of $\Phi$ with respect to $\Xi_{N}$ is defined as
+^$\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
 where ${\boldsymbol{\sigma}}=(\sigma_{1},\dots,\sigma_{N})^{\top}$ is a uniform random vector taking values in $\{-1,+1\}^{N}$.
 The Rademacher complexity of $\Phi$ is defined as the expectation of the empirical Rademacher complexity over all samples of size $N$ drawn from the underlying distribution:
-$\mathcal{R}_{N}(\Phi)\coloneqq\mathbb{E}_{S_{N}}[\widehat{\mathcal{R}}_{S_{N}}(\Phi)].$
-It captures the richness of the class $\Phi$ by measuring the degree to which a hypothesis set can fit random noise.
-The following lemma quantifies the Rademacher complexity of a class of parametric functions, which is generalized from Ermoliev and Norkin (2013, Lemma B.2) with minor modifications made to accommodate the developments in this paper.
+$\mathcal{R}_{N}(\Phi)\coloneqq\mathbb{E}_{\Xi_{N}}[\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)].$
+[…]
 Lemma 6.3
@@
 $\mathcal{R}_{N}(\widehat{\Phi})\leq\frac{2\mu L_{\widehat{\Phi}}\sqrt{q}}{N}+M_{\widehat{\Phi}}\sqrt{\frac{q\log N}{N}}.$
-Proof 8.4
-Proof of Lemma 8.3 Given a sample set $S_{N}=\{u^{s}\}_{s=1}^{N}$ of size $N$, following the proof of Ermoliev and Norkin (2013, Lemma B.2), we derive that for any $\epsilon\in(0,2\mu]$,
-^$\widehat{\mathcal{R}}_{S_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
-By setting $\epsilon=2\mu/N$ and taking expectation over $S_{N}$, we obtain the desired upper bound.∎
+Proof 6.4
+[…]
+^$\widehat{\mathcal{R}}_{\Xi_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
+By setting $\epsilon=2\mu/N$ and taking expectation over $\Xi_{N}$, we obtain the desired upper bound.∎
 The following lemma gives the generalization bound based on the Rademacher complexity.
-Lemma 8.5 (Mohri et al. (2018), Theorem 3.3)
+[…]
 Let $\Phi$ be a function class mapping from $\mathcal{U}$ to $[-M,M]$.
-Then for any $\delta>0$, with probability at least $1-\delta$ over the draw of an independently and identically distributed (i.i.d.) sample $S_{N}$ of size $N$, the following holds true for all $\phi\in\Phi$:
+Then for any $\delta>0$, with probability at least $1-\delta$ over the draw of an independently and identically distributed (i.i.d.) sample $\Xi_{N}$ of size $N$, the following holds true for all $\phi\in\Phi$:
 $\mathbb{E}[\phi(\tilde{u})]\leq\frac{1}{N}\sum_{s=1}^{N}\phi(u^{s})+2{\mathcal{R}}_{N}(\Phi)+\sqrt{2}M\sqrt{\frac{\log(1/\delta)}{N}}.$
 6.2.2 Proofs of Proposition 2.2 and Theorem 2.3.
-Proof 8.6
+Proof 6.6
 Proof of Proposition 2.2.
-We first quantify the Rademacher complexity of the composite function class $\Phi_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\left\{\phi:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}\mid\phi(x,y)=\varphi({f(x)};{y}),\,f\in{\mathcal{H}}_{\mu}^{K}\right\}$.
+We first quantify the Rademacher complexity of the composite function class $\Phi_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\left\{\phi:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}\mid\phi(x,y)=\varphi({f(x)},{y}),\,f\in{\mathcal{H}}_{\mu}^{K}\right\}$.
 Notice that any function in $\Phi_{{\mathcal{H}}_{\mu}^{K}}$ is parameterized by some $\theta\in[-\mu,\mu]^{2dK(p+1)}$.
 For any $(x,y)\in\mathcal{X}\times\mathcal{Y}$ and $\theta_{1},\theta_{2}\in[-\mu,\mu]^{2dK(p+1)}$,
-$\left\lvert\varphi(f(x;\theta_{1});y)-\varphi(f(x;\theta_{2});y)\right\rvert\leq L_{\varphi}\left\lVert f(x;\theta_{1})-f(x;\theta_{2})\right\rVert\leq 2L_{\varphi}\sqrt{p\bar{X}^{2}+1}\left\lVert\theta_{1}-\theta_{2}\right\rVert,$
+$\left\lvert\varphi(f(x,\theta_{1}),y)-\varphi(f(x,\theta_{2}),y)\right\rvert\leq L_{\varphi}\left\lVert f(x,\theta_{1})-f(x,\theta_{2})\right\rVert\leq 2L_{\varphi}\sqrt{p\bar{X}^{2}+1}\left\lVert\theta_{1}-\theta_{2}\right\rVert,$
 where the second inequality follows from Assumption (A2). Moreover, for any $f\in{\mathcal{H}}_{\mu}^{K}$,
-$\left\lvert\varphi(f(x);y)\right\rvert\leq\left\lvert\varphi(z_{0};y)\right\rvert+\left\lvert\varphi(f(x);y)-\varphi(z_{0};y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{\mu}.$ (14)
+$\left\lvert\varphi(f(x),y)\right\rvert\leq\left\lvert\varphi(z_{0},y)\right\rvert+\left\lvert\varphi(f(x),y)-\varphi(z_{0},y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{\mu}.$ (21)
 Therefore, by Lemma 6.3, we deduce
@@
 Therefore, by applying Lemma 6.5 again to $\Phi_{{\mathcal{H}}_{\mu}^{K}}^{-}$ and using the union bound, we derive that with probability at least $1-\delta$,
-$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$ (16)
+$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$ (23)
 Finally, by combining (22) and (23), we complete the proof.∎
-Proof 8.7
+Proof 6.7
 Proof of Theorem 2.3.
@@
 $R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})-\min_{f\in\mathcal{F}}R(f)$ $\leq R(f_{app})-R(f^{*})$
-$=\mathbb{E}_{X,Y}\left[\varphi({f_{app}(X)};{Y})-\varphi({f^{*}(X)};{Y})\right]$
+$=\mathbb{E}_{X,Y}\left[\varphi({f_{app}(X)},{Y})-\varphi({f^{*}(X)},{Y})\right]$
 $\leq L_{\varphi}\sup_{x\in\mathcal{X}}\left\lVert f_{app}(x)-f^{*}(x)\right\rVert_{2}$
@@
 where the last inequality follows from Proposition 2.1. For the estimation error, we have
-$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$ $=R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
-$\leq R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{S}(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{S}(f^{*}_{{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
-$\leq 2\left(\max_{f\in{\mathcal{H}}_{\mu}^{K}}\,|R(f)-\widehat{R}_{S}(f)|\right).$
+$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$ $=R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{\Xi}(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
+$\leq R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})+\widehat{R}_{\Xi}(f^{*}_{{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})$
+$\leq 2\left(\max_{f\in{\mathcal{H}}_{\mu}^{K}}\,|R(f)-\widehat{R}_{\Xi}(f)|\right).$
 By Proposition 2.2, with probability at least $1-\delta$, we have
-$R(f^{*}_{S,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
+$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
 By combining these two error bounds, we obtain the desired inequality. ∎
-9 Surrogation Construction of $F(\cdot;\xi)$ for Case 1 to Case 3
-In this section, we first illustrate the construction of upper surrogation functions of $F(\cdot;\xi)\coloneqq\varphi(f(\cdot;\xi);\xi)$ for Case 1 to Case 3, which are directionally differentiable and satisfy the first three properties in Assumption 1.
-Then we show in Claim 1 that $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,I)$ with $\varepsilon\geq 0$ satisfies Condition (P5) for the three cases, which also validates Condition (P4) since $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ by construction is directionally differentiable and satisfies (P2).
-Case 1:
-since $f(\cdot;\xi)$ and $\varphi(\cdot;\xi)$ are both piecewise affine, the composite function $F(\cdot;\xi)=\varphi(f(\cdot;\xi);\xi)$ is also piecewise affine.
-Thus, following the surrogation construction of $f(\cdot;\xi)$ with the difference-of-max-affine representation in Section 3.1, we can construct the upper surrogation function of $F(\cdot;\xi)$, which satisfies the first three properties in Assumption 1.
-Specifically, the first two properties clearly hold, and the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$ is justified since given a fixed index mapping $I$, the surrogate of a PA function is independent of the reference point $\theta^{\prime}$.
-Case 2:
-by assumption, $\varphi(\cdot;\xi)$ can be written as $\varphi(\cdot;\xi)=\varphi^{\uparrow}(\cdot;\xi)+\varphi^{\downarrow}(\cdot;\xi)$, where $\varphi^{\uparrow}(\cdot;\xi)$ and $\varphi^{\downarrow}(\cdot;\xi)$ are nondecreasing and nonincreasing convex functions respectively, that is, for any $\theta_{1},\theta_{2}\in\Theta$ satisfying $\theta_{1}\leq\theta_{2}$ componentwisely, we have $\varphi^{\uparrow}(\theta_{2};\xi)\geq\varphi^{\uparrow}(\theta_{1};\xi)$ and $\varphi^{\downarrow}(\theta_{2};\xi)\leq\varphi^{\downarrow}(\theta_{1};\xi)$.
-With $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$, the surrogation function of $F(\cdot;\xi)$ at $\theta^{\prime}$ is constructed as
-$\widehat{F}(\theta;\theta^{\prime},\xi,I)=\varphi^{\uparrow}(\widehat{f}(\theta;\xi,I_{2});\xi)+\varphi^{\downarrow}(\widecheck{f}(\theta;\xi,I_{1});\xi),$
-where $\widehat{f}(\theta;\xi,I_{2})$ and $\widecheck{f}(\theta;\xi,I_{1})$ are the upper and lower surrogation functions of $f(\theta,\xi)$ respectively.
-By monotonicity of $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$, $\widehat{F}(\theta;\theta^{\prime},\xi,I)\geq\varphi^{\uparrow}(f(\theta;\xi);\xi)+\varphi^{\downarrow}(f(\theta;\xi);\xi)=F(\theta;\xi)$ for any $\theta\in\Theta$, and thus the majorizing condition (P2) holds.
-Moreover, $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ satisfies the convex condition (P3) since a nondecreasing convex function composite with a convex function and a nonincreasing convex function composite with a concave function are both convex.
+7 Surrogation Construction of $F(\cdot,\xi)$ for Case 1 to Case 3
+We provide the construction of surrogation of $F(\cdot,\xi)\coloneqq\varphi(f(\cdot,\xi),\xi)$ for Case 1 to Case 3 in the proof of Proposition 3.2.
+In order to verify Assumption (B4), we prove that the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\cdot,\xi;\bar{\theta})$ at a reference point $\bar{\theta}$ satisfies the following condition for any $\xi\in\Xi$:
+  1. (P4)
+    Quadratic surrogation gap: for some $\kappa>0$, $F(\theta,\xi)+\displaystyle{\frac{\kappa}{2}}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\geq\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ for any $\theta\in\Theta$.
+For directionally differentiable surrogate functions, Condition (P4) is stronger than directional derivative consistency condition because it also limits the growth rate of the surrogate function with respect to the original function, which is essential to develop the nonasymptotic convergence analysis in Section 8.4.
+Proof 7.1
+Proof of Proposition 3.2
+We first construct the surrogate functions for the three cases and verify Assumptions (B1)-(B3), followed by the verification of Assumption (B4).
+Case 1: Suppose that $\varphi(z,\xi)=\varphi_{1}(z,\xi)-\varphi_{2}(z,\xi)$, where $\varphi_{1}(z,\xi)=\max_{i\in[K_{\varphi}]}\varphi_{1i}(z,\xi)$ and $\varphi_{2}(z,\xi)=\max_{i\in[K_{\varphi}]}\varphi_{2i}(z,\xi)$ are both max-affine functions.
+We denote the affine functions by $\{\varphi_{1i}(z,\xi)=a_{1i}^{\top}z+b_{1i}\}_{i\in[K_{\varphi}]}$ and $\{\varphi_{2i}(z,\xi)=a_{2i}^{\top}z+b_{2i}\}_{i\in[K_{\varphi}]}$, where $\xi$ represents the random components of $\{a_{1i},b_{1i},a_{2i},b_{2i}\}_{i\in[K_{\varphi}]}$.
+Then the surrogate function of $F(\cdot,\xi)$ at $\bar{\theta}$ is constructed as follows,
+$\widehat{F}(\theta,\xi;\bar{\theta},I)$ $=\max_{i\in[K_{\varphi}]}\left\{\max\{a_{1i},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})-\max\{-a_{1i},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})+b_{1i}\right\}$
+$\phantom{=}-\left(\max\{a_{2I_{3}},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})-\max\{-a_{2I_{3}},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+b_{2I_{3}}\right),$
+where $I_{1}\in\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$, $I_{2}\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)$, and $I_{3}\in\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$.
+With a alight abuse of notation, we denote $I=(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$.
+Assumptions (B1) and (B2) hold in this case by construction.
+Assumption (B3) is justified since given a fixed surrogation index $I$, the surrogate function is independent of the reference point $\bar{\theta}$.
+Case 2: by assumption, $\varphi(\cdot,\xi)$ can be written as $\varphi(\cdot,\xi)=\varphi^{\uparrow}(\cdot,\xi)+\varphi^{\downarrow}(\cdot,\xi)$, where $\varphi^{\uparrow}(\cdot,\xi)$ and $\varphi^{\downarrow}(\cdot,\xi)$ are nondecreasing and nonincreasing convex functions respectively, that is, for any $\theta_{1},\theta_{2}\in\Theta$ satisfying $\theta_{1}\leq\theta_{2}$ componentwisely, we have $\varphi^{\uparrow}(\theta_{2},\xi)\geq\varphi^{\uparrow}(\theta_{1},\xi)$ and $\varphi^{\downarrow}(\theta_{2},\xi)\leq\varphi^{\downarrow}(\theta_{1},\xi)$.
+With $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$, the surrogate function of $F(\cdot,\xi)$ at $\bar{\theta}$ is constructed as
+$\widehat{F}(\theta,\xi;\bar{\theta},I)=\varphi^{\uparrow}(\widehat{f}(\theta,\xi;I_{2}),\xi)+\varphi^{\downarrow}(\widecheck{f}(\theta,\xi;I_{1}),\xi),$
+where $\widehat{f}(\theta,\xi;I_{2})$ and $\widecheck{f}(\theta,\xi;I_{1})$ are the upper and lower surrogate functions of $f(\theta,\xi)$ respectively.
+By monotonicity of $\varphi^{\uparrow}$ and $\varphi^{\downarrow}$, $\widehat{F}(\theta,\xi;\bar{\theta},I)\geq\varphi^{\uparrow}(f(\theta,\xi),\xi)+\varphi^{\downarrow}(f(\theta,\xi),\xi)=F(\theta,\xi)$ for any $\theta\in\Theta$, and thus the majorizing condition (P2) holds.
+Moreover, $\widehat{F}(\cdot,\xi;\bar{\theta},I)$ satisfies the convex condition (P3) since a nondecreasing convex function composite with a convex function and a nonincreasing convex function composite with a concave function are both convex.
 The touching condition (P1) clearly holds when $\varepsilon=0$.
-With a fixed index mapping $I$, $\widehat{F}(\cdot;\theta^{\prime},\xi,I)$ in this case is also independent of $\theta^{\prime}$, validating the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$.
-Case 3:
-since $S$ is finite, for any $\xi\in S$, $\varphi(\cdot;\xi)$ is Lipschitz smooth with uniform modulus denoted by $L_{\nabla\varphi}$, and $f(\cdot;\xi)$ is Lipschitz continuous with uniform modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$.
-Then the surrogation function of $F(\cdot;\xi)$ at $\theta^{\prime}$ with $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ is constructed by surrogating the inner PA function according to componentwise signs of the gradient $\nabla\varphi(f(\theta^{\prime};\xi);\xi)$ as follows,
-$\widehat{F}(\theta;\theta^{\prime},\xi,I)$ $=\max\{\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widehat{f}(\theta;\xi,I_{2})-\max\{-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widecheck{f}(\theta;\xi,I_{1})$
-$\phantom{=,}-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi)^{\top}f(\theta^{\prime};\xi)+\varphi(f(\theta^{\prime};\xi);\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}.$
+With a fixed index $I$, $\widehat{F}(\cdot,\xi;\bar{\theta},I)$ is also independent of $\bar{\theta}$ in this case, thus validating the bi-variate continuity assumption (B3).
+Case 3: since $\Xi$ is finite, for any $\xi\in\Xi$, $\varphi(\cdot,\xi)$ is Lipschitz smooth with uniform modulus denoted by $L_{\nabla\varphi}$, and $f(\cdot,\xi)$ is Lipschitz continuous with uniform modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$.
+Then the surrogate function of $F(\cdot,\xi)$ at $\bar{\theta}$ with $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$ is constructed by surrogating the inner PA function according to componentwise signs of the gradient $\nabla\varphi(f(\bar{\theta},\xi),\xi)$ as follows,
+$\widehat{F}(\theta,\xi;\bar{\theta},I)$ $=\max\{\nabla{\varphi}(f(\bar{\theta},\xi),\xi),0\}^{\top}\widehat{f}(\theta,\xi;I_{2})-\max\{-\nabla{\varphi}(f(\bar{\theta},\xi),\xi),0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})$
+$\phantom{=,}-\nabla{\varphi}(f(\bar{\theta},\xi),\xi)^{\top}f(\bar{\theta},\xi)+\varphi(f(\bar{\theta},\xi),\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}.$
 The convex condition (P3) holds evidently; the touching condition (P1) also holds when $\varepsilon=0$.
-The surrogate $\widehat{F}(\cdot;\cdot,\xi,I)$ is continuous on $\Theta\times\Theta$ since $\nabla\varphi(\cdot;\xi)$ is continuous.
+The surrogate $\widehat{F}(\cdot,\xi;\cdot,I)$ is continuous on $\Theta\times\Theta$ since $\nabla\varphi(\cdot,\xi)$ is continuous, validating Assumption (B3).
 Based on the quadratic upper bound of a Lipschitz smooth function, the majorizing condition (P2) is verified as follows.
-Since $\varphi(\cdot;\xi)$ is $L_{\nabla\varphi}$-smooth, we have for any $z_{1},z_{2}\in\mathbb{R}^{d}$ that
-$\varphi(z_{1};\xi)\leq\varphi(z_{2};\xi)+\nabla\varphi(z_{2};\xi)^{\top}(z_{1}-z_{2})+\frac{L_{\nabla\varphi}}{2}\left\lVert z_{1}-z_{2}\right\rVert^{2}.$
-By setting $z_{1}=f(\theta;\xi)$ and $z_{2}=f(\theta^{\prime};\xi)$, we obtain that for any $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$,
-$F(\theta;\xi)$ $\leq F(\theta^{\prime};\xi)+\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}(f(\theta;\xi)-f(\theta^{\prime};\xi))+\frac{L_{\nabla\varphi}}{2}\left\lVert f(\theta;\xi)-f(\theta^{\prime};\xi)\right\rVert^{2}$
-$\leq F(\theta^{\prime};\xi)+\max\{\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widehat{f}(\theta;\xi,I_{2})$
-$\phantom{\leq,}-\max\{-\nabla{\varphi}(f(\theta^{\prime};\xi);\xi),0\}^{\top}\widecheck{f}(\theta;\xi,I_{1})-\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}f(\theta^{\prime};\xi)+\frac{L_{\nabla\varphi}L_{f}^{2}}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}$
-$=\widehat{F}(\theta;\theta^{\prime},\xi,I).$
-Claim 1
-Let $\Theta$ be a compact set.
-Given $\varepsilon\geq 0$ and $\theta^{\prime}\in\Theta$, with the surrogation construction for Case 1 to Case 3, $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ satisfies Condition (P5) for any $\xi\in S$.
-Proof 9.1
-Proof of Claim 1. Since $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime},\xi)$ also satisfies the majorizing condition (P2), by the compactness of $\Theta$, it suffices to show that
-$\widehat{\kappa}\coloneqq\limsup_{\theta\rightarrow\theta^{\prime},\theta\neq\theta^{\prime}}\frac{2(\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)-F(\theta;\xi))}{\left\lVert\theta-\theta^{\prime}\right\rVert^{2}}<+\infty.$
-For the piecewise affine function $f(\theta;\xi)\coloneqq g(\theta;\xi)-h(\theta;\xi)$, we have $\mathcal{I}_{h}^{0}(\theta)\subseteq\mathcal{I}_{h}^{\varepsilon}(\theta^{\prime})$ when $\theta\in\Theta$ is sufficiently close to $\theta^{\prime}$.
-It follows from (P2) that for any $\theta\in\Theta$ sufficiently close to $\theta^{\prime}$,
-$f(\theta;\xi)\leq\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2})\leq\min_{I_{2}\in\mathcal{I}_{h}^{0}(\theta)}\widehat{f}(\theta;\xi,I_{2})=f(\theta;\xi).$
-Therefore, by similar arguments for lower surrogation functions of $f(\cdot;\xi)$, we obtain that for any $\theta\in\Theta$ sufficiently close to $\theta^{\prime}$,
-$\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2})=f(\theta;\xi)=\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\widecheck{f}(\theta;\xi,I_{1}).$
-For Case 1, since the objective is piecewise affine, we directly obtain $\widehat{\kappa}=0$. For Case 2, when $\theta$ is sufficiently close to $\theta^{\prime}$, we have
-$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)$ $\coloneqq\min_{(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},\xi,(I_{1},I_{2}))$
-$=\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\varphi^{\uparrow}(\widehat{f}(\theta;\xi,I_{2});\xi)+\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\varphi^{\downarrow}(\widecheck{f}(\theta;\xi,I_{1});\xi)$
-$=\varphi^{\uparrow}\left(\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\prime})}\widehat{f}(\theta;\xi,I_{2});\xi\right)+\varphi^{\downarrow}\left(\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\prime})}\widecheck{f}(\theta;\xi,I_{1});\xi\right)$
-$=F(\theta;\xi),$
-also implying $\widehat{\kappa}=0$. For Case 3, when $\theta$ is sufficiently close to $\theta^{\prime}$, we obtain
-$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)=\nabla\varphi(f(\theta^{\prime};\xi);\xi)^{\top}(f(\theta;\xi)-f(\theta^{\prime};\xi))+\varphi(f(\theta^{\prime};\xi);\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}.$
-Thus, by the second-order Taylor expansion of $\varphi(\cdot;\xi)$ at $f(\theta^{\prime};\xi)$, we obtain
-$\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime},\xi)-F(\theta;\xi)=\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}-\frac{1}{2}v^{\top}\nabla^{2}\varphi(f(\theta^{\prime};\xi);\xi)v+o(\left\lVert v\right\rVert^{2}),$
+Since $\varphi(\cdot,\xi)$ is $L_{\nabla\varphi}$-smooth, we have for any $z_{1},z_{2}\in\mathbb{R}^{d}$ that
+$\varphi(z_{1},\xi)\leq\varphi(z_{2},\xi)+\nabla\varphi(z_{2},\xi)^{\top}(z_{1}-z_{2})+\frac{L_{\nabla\varphi}}{2}\left\lVert z_{1}-z_{2}\right\rVert^{2}.$
+By setting $z_{1}=f(\theta,\xi)$ and $z_{2}=f(\bar{\theta},\xi)$, we obtain that for any $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)$,
+$F(\theta,\xi)$ $\leq F(\bar{\theta},\xi)+\nabla\varphi(f(\bar{\theta},\xi),\xi)^{\top}(f(\theta,\xi)-f(\bar{\theta},\xi))+\frac{L_{\nabla\varphi}}{2}\left\lVert f(\theta,\xi)-f(\bar{\theta},\xi)\right\rVert^{2}$
+$\leq F(\bar{\theta},\xi)+\max\{\nabla{\varphi}(f(\bar{\theta},\xi),\xi),0\}^{\top}\widehat{f}(\theta,\xi;I_{2})$
+$\phantom{\leq,}-\max\{-\nabla{\varphi}(f(\bar{\theta},\xi),\xi),0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})-\nabla\varphi(f(\bar{\theta},\xi),\xi)^{\top}f(\bar{\theta},\xi)+\frac{L_{\nabla\varphi}L_{f}^{2}}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}$
+$=\widehat{F}(\theta,\xi;\bar{\theta},I).$
+Now we prove Assumption (B4) for Case 1 to Case 3.
+We prove that the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\cdot,\xi;\bar{\theta})$ satisfies Condition (P4).
+Assumption (B4) is then verified since $\widehat{F}^{\varepsilon}_{\min}(\cdot,\xi;\bar{\theta})$ by construction is directionally differentiable and satisfies (P2).
+Since $\widehat{F}^{\varepsilon}_{\min}(\cdot,\xi;\bar{\theta})$ also satisfies the majorizing condition (P2), by the compactness of $\Theta$, it suffices to show that
+$\widehat{\kappa}\coloneqq\limsup_{\theta\rightarrow\bar{\theta},\theta\neq\bar{\theta}}\frac{2(\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})-F(\theta,\xi))}{\left\lVert\theta-\bar{\theta}\right\rVert^{2}}<+\infty.$
+For the piecewise affine function $f(\theta,\xi)\coloneqq g(\theta,\xi)-h(\theta,\xi)$, we have $\mathcal{I}_{h}^{0}(\theta,\xi)\subseteq\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$ when $\theta\in\Theta$ is sufficiently close to $\bar{\theta}$.
+It follows from (P2) that for any $\theta\in\Theta$ sufficiently close to $\bar{\theta}$,
+$f(\theta,\xi)\leq\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\widehat{f}(\theta,\xi;I_{2})\leq\min_{I_{2}\in\mathcal{I}_{h}^{0}(\theta,\xi)}\widehat{f}(\theta,\xi;I_{2})=f(\theta,\xi).$
+Therefore, by similar arguments for lower surrogate functions of $f(\cdot,\xi)$, we obtain that for any $\theta\in\Theta$ sufficiently close to $\bar{\theta}$,
+$\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\widehat{f}(\theta,\xi;I_{2})=f(\theta,\xi)=\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)}\widecheck{f}(\theta,\xi;I_{1}).$
+For Case 1, when $\theta$ is sufficiently close to $\bar{\theta}$, we have
+$\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ $\coloneqq\min_{(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}\widehat{F}(\theta,\xi;\bar{\theta},I)$
+$=\underbrace{\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi),I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\max_{i\in[K_{\varphi}]}\left\{\max\{a_{1i},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})-\max\{-a_{1i},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})+b_{1i}\right\}}_{\mathcal{V}_{1}}$
+$\phantom{==}-\underbrace{\max_{(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}\left(\max\{a_{2I_{3}},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})-\max\{-a_{2I_{3}},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+b_{2I_{3}}\right)}_{\mathcal{V}_{2}}$
+$=F(\theta,\xi).$ (24)
+The third equality is validated as follows. For the first component $\mathcal{V}_{1}$, we have
+$\mathcal{V}_{1}$ $\geq\max_{i\in[K_{\varphi}]}\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi),I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\left\{\max\{a_{1i},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})-\max\{-a_{1i},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})+b_{1i}\right\}$
+$=\max_{i\in[K_{\varphi}]}\left\{\max\{a_{1i},0\}^{\top}\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\widehat{f}(\theta,\xi;I_{2})-\max\{-a_{1i},0\}^{\top}\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)}\widecheck{f}(\theta,\xi;I_{1})+b_{1i}\right\}$
+$=\max_{i\in[K_{\varphi}]}\left\{a_{1i}^{\top}f(\theta,\xi)+b_{1i}\right\}\eqqcolon\varphi_{1}(f(\theta,\xi),\xi).$
+Since there exists $\bar{I}_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)$ and $\bar{I}_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)$ such that $\widehat{f}(\theta,\xi;\bar{I}_{2})=f(\theta,\xi)$ and $\widecheck{f}(\theta,\xi;\bar{I}_{1})=f(\theta,\xi)$, we also have
+$\mathcal{V}_{1}$ $\leq\max_{i\in[K_{\varphi}]}\left\{\max\{a_{1i},0\}^{\top}\widehat{f}(\theta,\xi;\bar{I}_{2})-\max\{-a_{1i},0\}^{\top}\widecheck{f}(\theta,\xi;\bar{I}_{1})+b_{1i}\right\}$
+$=\max_{i\in[K_{\varphi}]}\left\{a_{1i}^{\top}f(\theta,\xi)+b_{1i}\right\}\eqqcolon\varphi_{1}(f(\theta,\xi),\xi),$
+indicating $\mathcal{V}_{1}=\varphi_{1}(f(\theta,\xi),\xi)$.
+Moreover, for the second component $\mathcal{V}_{2}$, when $\theta$ is sufficiently close to $\bar{\theta}$, there exists $\bar{I}_{3}\in\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$ such that $\varphi_{2\bar{I}_{3}}(f(\theta,\xi),\xi)=\varphi_{2}(f(\theta,\xi),\xi)$.
+Then, we have
+$\mathcal{V}_{2}$ $\geq\max\{a_{2\bar{I}_{3}},0\}^{\top}\widecheck{f}(\theta,\xi;\bar{I}_{1})-\max\{-a_{2\bar{I}_{3}},0\}^{\top}\widehat{f}(\theta,\xi;\bar{I}_{2})+b_{2\bar{I}_{3}}=\varphi_{2}(f(\theta,\xi),\xi)$
+and
+$\mathcal{V}_{2}$ $\leq\max_{(I_{1},I_{2})\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi),i\in[K_{\varphi}]}\left(\max\{a_{2i},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})-\max\{-a_{2i},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+b_{2i}\right)$
+$\leq\max_{(I_{1},I_{2})\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi),i\in[K_{\varphi}]}\left\{a_{2i}^{\top}f(\theta,\xi)+b_{2i}\right\}=\varphi_{2}(f(\theta,\xi),\xi),$
+indicating $\mathcal{V}_{2}=\varphi_{2}(f(\theta,\xi),\xi)$.
+Therefore, we derive (7.1) and directly obtain $\widehat{\kappa}=0$.
+For Case 2, when $\theta$ is sufficiently close to $\bar{\theta}$, we have
+$\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ $\coloneqq\min_{(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}\widehat{F}(\theta,\xi;\bar{\theta},(I_{1},I_{2}))$
+$=\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\varphi^{\uparrow}(\widehat{f}(\theta,\xi;I_{2}),\xi)+\min_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)}\varphi^{\downarrow}(\widecheck{f}(\theta,\xi;I_{1}),\xi)$
+$=\varphi^{\uparrow}\left(\min_{I_{2}\in\mathcal{I}^{\varepsilon}_{h}(\bar{\theta},\xi)}\widehat{f}(\theta,\xi;I_{2}),\xi\right)+\varphi^{\downarrow}\left(\max_{I_{1}\in\mathcal{I}^{\varepsilon}_{g}(\bar{\theta},\xi)}\widecheck{f}(\theta,\xi;I_{1}),\xi\right)$
+$=F(\theta,\xi),$
+also implying $\widehat{\kappa}=0$. For Case 3, when $\theta$ is sufficiently close to $\bar{\theta}$, we obtain
+$\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})=\nabla\varphi(f(\bar{\theta},\xi),\xi)^{\top}(f(\theta,\xi)-f(\bar{\theta},\xi))+\varphi(f(\bar{\theta},\xi),\xi)+\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}.$
+Thus, by the second-order Taylor expansion of $\varphi(\cdot,\xi)$ at $f(\bar{\theta},\xi)$, we obtain
+$\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})-F(\theta,\xi)=\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}-\frac{1}{2}v^{\top}\nabla^{2}\varphi(f(\bar{\theta},\xi),\xi)v+o(\left\lVert v\right\rVert^{2}),$
 […]
-10 Details of Convergence Analysis in Section 4
+8 Details of Convergence Analysis in Section 3.2.3
 8.1 Proof of Proposition 3.4 and equivalence between $\mathcal{S}^{d}_{\varepsilon,\rho}$ and $\mathcal{S}^{*}$
-Proof 10.1
+Proof 8.1
 Proof of Proposition 3.4.
 Given $\varepsilon\geq 0$ and $\rho>0$, for any $\theta^{*}\in\mathcal{S}^{*}$ it holds that $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{*})+\frac{\rho}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}>F(\theta)\geq F(\theta^{*})=\widehat{F}^{\varepsilon}_{\min}(\theta^{*};\theta^{*})$ for any $\theta\neq\theta^{*}$, showing $\theta^{*}\in\mathcal{S}^{d}_{\varepsilon,\rho}$ and thus $\mathcal{S}^{*}\subseteq\mathcal{S}^{d}_{\varepsilon,\rho}$.
-For any $\theta^{d}\in\mathcal{S}^{d}_{\varepsilon,\rho}$, from Condition (P4) and the first-order optimality condition, it follows that $F^{\prime}(\theta^{d};\theta-\theta^{d})=\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{d})^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for all $\theta\in\Theta$, yielding $\theta^{d}\in\mathcal{S}^{d}$ and thus $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$.
+For any $\theta^{d}\in\mathcal{S}^{d}_{\varepsilon,\rho}$, from Condition (B4) and the first-order optimality condition, it follows that $F^{\prime}(\theta^{d};\theta-\theta^{d})=\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{d})^{\prime}(\theta^{d};\theta-\theta^{d})\geq 0$ for all $\theta\in\Theta$, yielding $\theta^{d}\in\mathcal{S}^{d}$ and thus $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{d}$.
 It remains to show $\mathcal{S}^{d}\subseteq\mathcal{S}^{d}_{0,\rho}$ for any $\rho>0$.
 Assume by contradiction that there exists $\theta^{d}\in\mathcal{S}^{d}\backslash\mathcal{S}^{d}_{0,\widehat{\rho}}$ for some $\widehat{\rho}>0$.
-Then there exist $\widehat{I}\in\mathcal{I}^{0}(\theta^{d})$ and $\widehat{\theta}(\neq\theta^{d})\in\Theta$ such that $\widehat{\theta}\in\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}(\theta;\theta^{d},\widehat{I}\,)+\frac{\widehat{\rho}}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}$.
-By the strong convexity of the function $\theta\mapsto\widehat{F}(\theta;\theta^{d},\widehat{I}\,)+\frac{\widehat{\rho}}{2}\,\left\lVert\theta-\theta^{d}\right\rVert^{2}$, we have $\widehat{F}(\cdot;\theta^{d},\widehat{I})^{\prime}(\theta^{d};\widehat{\theta}-\theta^{d})<0$.
+Then there exist $\{\widehat{I}_{s}\in\mathcal{I}^{0}(\theta^{d},\xi^{s})\mid s=1,\dots,n\}$ and $\widehat{\theta}(\neq\theta^{d})\in\Theta$ such that $\widehat{\theta}\in\mathop{\arg\min}_{\theta\in\Theta}\left\{\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta,\xi;\theta^{d},\widehat{I}_{s}\,)+\frac{\widehat{\rho}}{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}\right\}$.
+By the strong convexity of the function $\theta\mapsto\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta,\xi;\theta^{d},\widehat{I}_{s})+\frac{\widehat{\rho}}{2}\,\left\lVert\theta-\theta^{d}\right\rVert^{2}$, we have $\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\cdot,\xi;\theta^{d},\widehat{I}_{s})^{\prime}(\theta^{d};\widehat{\theta}-\theta^{d})<0$.
 Thus, there exists $t_{0}>0$ such that
-$F(\theta^{d}+t(\widehat{\theta}-\theta^{d}))\leq\widehat{F}(\theta^{d}+t(\widehat{\theta}-\theta^{d});\theta^{d},\widehat{I})<\widehat{F}(\theta^{d};\theta^{d},\widehat{I})=F(\theta^{d}),\quad\forall\,t\in(0,t_{0}],$
+$F(\theta^{d}+t(\widehat{\theta}-\theta^{d}))\leq\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta^{d}+t(\widehat{\theta}-\theta^{d}),\xi^{s};\theta^{d},\widehat{I}_{s})<\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta^{d},\xi^{s};\theta^{d},\widehat{I}_{s})=F(\theta^{d}),\quad\forall\,t\in(0,t_{0}],$
 where the equality holds by the touching condition (P1).
@@
 The following assumption is needed:
-  1. [label=(B0), resume*=B, labelwidth=!, labelindent=0pt]
-  2. 1.
+  1. (EC1)
     There exists $M_{\Delta}>0$ such that $F(\theta^{d})\geq\min_{\theta\in\Theta}F(\theta)+M_{\Delta}$ for any $\theta^{d}\in\mathcal{S}^{d}\backslash\mathcal{S}^{*}$.
-The assumption is naturally satisfied when the outer function of $F$ is in Case 1 or Case 2.
-Specifically, from the proof of Claim 2 afterward, we see that in these two cases, $\Theta$ can be decomposed into finitely many convex regions, on each of which any d-stationary point $\theta^{d}\in\mathcal{S}^{d}$ is a minimizer of $F$ over that region.
-Since the number of such regions is finite, the function $F$ takes finitely many values on $\mathcal{S}^{d}$, verifying Assumption 1.
-Discussions on the finite number of d-stationary values for more general nonconvex problems with piecewise and composite structures can be seen in Cui and Pang (2018).
+[…]
 Proposition 8.2
-For problem (9) with $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$ and Assumption 1 holding for the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ for each $\xi\in S$, the following statements hold:
+Suppose Assumptions (B1)-(B4) hold. For problem (10) with $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$, the following statements hold:
   1. (a)
-    With $\varphi$ in Case 1 or Case 2, $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$, where $\varepsilon_{0}\coloneqq 2\mu\sqrt{q}\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, and $M_{\Delta}$ is the constant defined in 1.
+    With $\varphi$ in Case 1 or Case 2, $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$, where $\varepsilon_{0}\coloneqq 2\mu\sqrt{q}\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, and $M_{\Delta}$ is the constant defined in (EC1).
   2. (b)
-    Suppose that Assumptions 2 and 1 additionally hold with $M_{\Delta}>2\kappa q\mu^{2}$.
+    Suppose that Assumption (EC1) holds with $M_{\Delta}>2\kappa q\mu^{2}$, and that the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ satisfies Condition (P4) for any $\varepsilon\geq 0$ and $\xi\in\Xi$.
 It holds that $\mathcal{S}^{d}_{\varepsilon,\rho}=\mathcal{S}^{*}$ for any $\varepsilon\in[\varepsilon_{0},+\infty)$ and $\rho\in(0,\frac{M_{\Delta}}{2q\mu^{2}}-\kappa]$.
-Proof 10.3
+Proof 8.3
 Proof of Proposition 8.2.
 By Proposition 3.4, it suffices to prove the inclusion $\mathcal{S}^{d}_{\varepsilon,\rho}\subseteq\mathcal{S}^{*}$.
-Since the two max-affine components of $f(\cdot;\xi)$ for any $\xi\in S$ are Lipschitz continuous with modulus $L_{g,h}\coloneqq\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, we see that with $\varepsilon\geq\varepsilon_{0}=2L_{g,h}\sqrt{q}\mu$, the set $\mathcal{I}^{\varepsilon}(\theta^{\prime})$ contains all index mappings, that is, $\mathcal{I}^{\varepsilon}(\theta^{\prime})=\{(I_{1},I_{2})\mid I_{1}:S\rightarrow[K],I_{2}:S\rightarrow[K]\}$ for any $\theta^{\prime}\in\Theta$.
+Since the two max-affine components of $f(\cdot,\xi)$ for any $\xi\in\Xi$ are Lipschitz continuous with modulus $L_{g,h}\coloneqq\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$, we see that with $\varepsilon\geq\varepsilon_{0}=2L_{g,h}\sqrt{q}\mu$, the set $\mathcal{I}^{\varepsilon}(\bar{\theta})\coloneqq\{\{I^{\varepsilon}_{s}\}_{s=1}^{n}\mid I^{\varepsilon}_{s}\in\mathcal{I}(\bar{\theta},\xi^{s}),\,s=1,\dots,n\}$ contains all index combinations from the $\varepsilon$-active index sets $\{\mathcal{I}(\bar{\theta},\xi^{s})\}_{s=1}^{n}$, that is, $\mathcal{I}^{\varepsilon}(\bar{\theta})=[K]^{2n}$ for any $\bar{\theta}\in\Theta$.
 For the first statement, assume by contradiction that there exists $\widehat{\theta}^{d}\in\mathcal{S}^{d}_{\widehat{\varepsilon},\widehat{\rho}}\backslash\mathcal{S}^{*}$ for some $\widehat{\varepsilon}\in[\varepsilon_{0},+\infty)$ and $\widehat{\rho}\in(0,\frac{M_{\Delta}}{2q\mu^{2}}]$.
-Since the set $\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ contains all index mappings for the inner PA functions, by the surrogation construction for Case 1 and Case 2 in Appendix 9, for any $\theta^{*}\in\mathcal{S}^{*}$ there exists $I^{\widehat{\varepsilon}}\in\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ such that $\widehat{F}^{\widehat{\varepsilon}}_{\min}(\theta^{*};\widehat{\theta}^{d})=\widehat{F}(\theta^{*};\widehat{\theta}^{d},I^{\widehat{\varepsilon}})=\min_{\theta\in\Theta}F(\theta)$.
+Since the set $\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ contains all index combinations for the inner PA functions, by the surrogation construction for Case 1 and Case 2 in Online Appendix 7, for any $\theta^{*}\in\mathcal{S}^{*}$ there exists $\{I^{\widehat{\varepsilon}}_{s}\}_{s=1}^{n}\in\mathcal{I}^{\widehat{\varepsilon}}(\widehat{\theta}^{d})$ such that $\widehat{F}^{\widehat{\varepsilon}}_{\min}(\theta^{*};\widehat{\theta}^{d})=\frac{1}{n}\sum_{s=1}^{n}\widehat{F}(\theta^{*},\xi^{s};\widehat{\theta}^{d},I^{\widehat{\varepsilon}}_{s})=\min_{\theta\in\Theta}F(\theta)$.
 It follows that
@@
 8.2 Proof of Proposition 3.5
-The following lemma from Deng and Gao (2021) is rephrased in our context for the proof of Proposition 4.3.
-Lemma 10.4 (Deng and Gao (2021), Theorem 3.2)
-Suppose that the stochastic function $\widehat{F}(\cdot;\theta^{\prime},\xi)$ is convex and Lipschitz continuous on $\Theta$ with modulus $L_{\widehat{F}}$ for any $\xi\in S$ and $\theta^{\prime}\in\Theta$, and that samples $\Xi_{N}\coloneqq\{\xi^{s}\}_{s=1}^{N}$ are i.i.d.
-generated from $S$.
-For $\eta>0$, denote $\theta^{\prime}_{+}\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\frac{1}{N}\sum_{s=1}^{N}\widehat{F}(\theta;\theta^{\prime},\xi^{s})+\frac{\eta}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}$.
+[…]
+[…]
+Suppose that a stochastic function $\phi(\cdot,\xi)$ is convex and Lipschitz continuous on $\Theta$ with modulus $L_{\phi}$ for any $\xi\in\Xi$ and $\bar{\theta}\in\Theta$, and that samples $\Xi_{N}\coloneqq\{\widetilde{\xi}^{s}\}_{s=1}^{N}$ are i.i.d.
+generated from $\Xi$.
+For $\eta>0$, denote $\bar{\theta}_{+}\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\frac{1}{N}\sum_{s=1}^{N}\phi(\theta,\widetilde{\xi}^{s})+\frac{\eta}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}$.
 Then it holds that
-$\left\lvert\mathbb{E}_{\Xi_{N}}\left[\frac{1}{N}\sum_{s=1}^{N}\widehat{F}(\theta^{\prime}_{+};\theta^{\prime},\xi^{s})-\mathbb{E}_{\widetilde{\xi}_{n}}\left[\widehat{F}(\theta^{\prime}_{+};\theta^{\prime},\widetilde{\xi}_{n})\right]\right]\right\rvert\leq\frac{2L_{\widehat{F}}^{2}}{\eta N}.$
-Proof 10.5
+$\left\lvert\mathbb{E}_{\Xi_{N}}\left[\frac{1}{N}\sum_{s=1}^{N}\phi(\bar{\theta}_{+},\widetilde{\xi}^{s})-\mathbb{E}_{\widetilde{\xi}_{n}}\left[\phi(\bar{\theta}_{+},\widetilde{\xi}_{n})\right]\right]\right\rvert\leq\frac{2L_{\phi}^{2}}{\eta N}.$
+Proof 8.5
 Proof of Proposition 3.5.
-Let $\widetilde{I}^{\varepsilon}_{\nu}$ be a random mapping uniformly drawn from $\mathcal{I}^{\varepsilon}(\theta^{\nu})$, $\{\widetilde{\xi}^{s}_{\nu}\}_{s=1}^{N_{\nu}}$ be random vectors independently and uniformly drawn from $S$, $\mathcal{G}_{\nu-1}$ denote the $\sigma$-algebra generated by all past samples $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and mappings $\{\widetilde{I}^{\varepsilon}_{t}\}_{t=0}^{\nu-1}$, and $\mathcal{G}_{\nu-1}^{+}$ denote the $\sigma$-algebra generated by $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and $\{\widetilde{I}^{\varepsilon}_{t}\}_{t=0}^{\nu}$.
-Given the current iterate $\theta^{\nu}$ and the index mapping $I^{\varepsilon}_{\nu}\in\mathcal{I}^{\varepsilon}(\theta^{\nu})$, by the definition of $P_{I^{\varepsilon}_{\nu}}^{\rho}(\cdot)$ in (10), we have
-$\min\left\{\widehat{F}(P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},I^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}+\frac{\rho}{2}\,\big{\lVert}\,{P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{\nu}).$ (17)
+Let $\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}=\{\widetilde{I}^{\varepsilon}_{\xi}\}_{\xi\in\Xi}$ be a random index combination uniformly drawn from $\mathcal{I}^{\varepsilon}(\theta^{\nu})$, $\{\widetilde{\xi}^{s}_{\nu}\}_{s=1}^{N_{\nu}}$ be random vectors independently and uniformly drawn from $\Xi$.
+Let $\mathcal{G}_{\nu-1}$ denote the $\sigma$-algebra generated by all past samples $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and combinations $\{\widetilde{\mathbf{I}}^{\varepsilon}_{t}\}_{t=0}^{\nu-1}$, and $\mathcal{G}_{\nu-1}^{+}$ denote the $\sigma$-algebra generated by $\{\{\widetilde{\xi}_{t}^{s}\}_{s=1}^{N_{t}}\}_{t=0}^{\nu-1}$ and $\{\widetilde{\mathbf{I}}^{\varepsilon}_{t}\}_{t=0}^{\nu}$.
+Given the current iterate $\theta^{\nu}$ and the index combination $\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}\in\mathcal{I}^{\varepsilon}(\theta^{\nu})$, by the definition of $P_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}^{\rho}(\cdot)$ in (15), we have
+$\min\left\{\widehat{F}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}+\frac{\rho}{2}\,\big{\lVert}\,{P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{\nu}).$ (25)
 With the notation $\mathbb{E}_{N_{\nu}}$ representing the expectation with respect to $\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$, by rearranging terms of (25) and taking the conditional expectation on both sides, we obtain that
-$\frac{\rho}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$ $\leq F(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
-$=\underbrace{F(\theta^{\nu})-W(\theta^{\nu})}_{\mathcal{E}_{1}^{\nu}}+\underbrace{W(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]}_{\mathcal{E}_{2}^{\nu}},$ (18)
-where $W(\theta^{\nu})\coloneqq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$.
-From the update rule in Step 5 of Algorithm 1, we have $\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\geq F_{N_{\nu}}(\theta^{\nu+1})$ and $F_{N_{\nu}}(\theta^{\nu})\geq F_{N_{\nu}}(\theta^{\nu+1})$.
+$\frac{\rho}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$ $\leq F(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
+$=\underbrace{F(\theta^{\nu})-W(\theta^{\nu})}_{\mathcal{E}_{1}^{\nu}}+\underbrace{W(\theta^{\nu})-\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]}_{\mathcal{E}_{2}^{\nu}},$ (26)
+where $W(\theta^{\nu})\coloneqq\mathbb{E}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{\nu}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$.
+From the update rule in Step 5 of Algorithm 1, we have $\widehat{F}_{\nu}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\geq F_{\nu}(\theta^{\nu+1})$ and $F_{\nu}(\theta^{\nu})\geq F_{\nu}(\theta^{\nu+1})$.
 It follows that
-$W(\theta^{\nu})$ $\geq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\mid\mathcal{G}_{\nu-1}\right]=\mathbb{E}\left[F_{N\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right].$
+$W(\theta^{\nu})$ $\geq\mathbb{E}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[F_{\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\mid\mathcal{G}_{\nu-1}\right]=\mathbb{E}\left[F_{\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right].$
 With $\Theta$ being a compact set, let $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$. Define
-$\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\{\phi(\cdot;\theta):S\rightarrow\mathbb{R}\mid\phi(\xi;\theta)=\varphi(f(\theta;\xi);\xi),\,\xi\in S,f(\theta;\cdot)\in{\mathcal{H}}_{\mu}^{K},\theta\in\Theta\}.$
-From finiteness of $S$ and continuity of $\varphi(\cdot;\xi)$, it follows that for any $\phi\in\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}$, $\phi(\xi;\cdot)$ is Lipschitz continuous on the compact set $\Theta$ with uniform modulus, and that $\phi(\cdot;\theta)$ is uniformly bounded.
+$\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}\coloneqq\{\phi(\cdot,\theta):\Xi\rightarrow\mathbb{R}\mid\phi(\xi,\theta)=\varphi(f(\theta,\xi),\xi),\,\xi\in\Xi,f(\theta,\cdot)\in{\mathcal{H}}_{\mu}^{K},\theta\in\Theta\}.$
+From finiteness of $\Xi$ and continuity of $\varphi(\cdot,\xi)$, it follows that for any $\phi\in\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}}$, $\phi(\xi,\cdot)$ is Lipschitz continuous on the compact set $\Theta$ with uniform modulus, and that $\phi(\cdot,\theta)$ is uniformly bounded.
 Therefore, for the term $\mathcal{E}_{1}^{\nu}$, there exists a constant $C_{1}>0$ such that
-$\mathcal{E}_{1}^{\nu}$ $\leq F(\theta^{\nu})-\mathbb{E}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
-$=F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[F_{N_{\nu}}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
-$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\sup_{\theta\in\Theta}\left\lvert F(\theta)-F_{N_{\nu}}(\theta)\right\rvert\,\middle|\,\mathcal{G}_{\nu-1}\right]$
+$\mathcal{E}_{1}^{\nu}$ $\leq F(\theta^{\nu})-\mathbb{E}\left[F_{\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
+$=F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[F_{\nu}(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]$
+$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\sup_{\theta\in\Theta}\left\lvert F(\theta)-F_{\nu}(\theta)\right\rvert\,\middle|\,\mathcal{G}_{\nu-1}\right]$
 $\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+2\mathcal{R}_{N_{\nu}}(\widehat{\Phi}_{{\mathcal{H}}_{\mu}^{K}})$
 $\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+C_{1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}},$ (27)
-where the third inequality follows from inequalities (3.8)-(3.13) in Mohri et al. (2018) that are part of the proof of Lemma 8.5, and the last inequality is derived by Lemma 8.3.
-Next, we bound the term $\mathcal{E}_{2}^{\nu}$. Given $I^{\varepsilon}_{\nu}$ and $\{\xi_{\nu}^{s}\}_{s=1}^{N_{\nu}}$, by the optimality of ${\theta}^{\nu+\frac{1}{2}}$, we have
-$\widehat{F}_{N_{\nu}}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\leq\widehat{F}_{N_{\nu}}(P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},I^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{I^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}.$
+[…]
+Next, we bound the term $\mathcal{E}_{2}^{\nu}$.
+Given $\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}$ and $\{\xi_{\nu}^{s}\}_{s=1}^{N_{\nu}}$, by the optimality of ${\theta}^{\nu+\frac{1}{2}}$, we have
+$\widehat{F}_{\nu}({\theta}^{\nu+\frac{1}{2}};\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,{{\theta}^{\nu+\frac{1}{2}}-\theta^{\nu}}\,\big{\rVert}^{2}\leq\widehat{F}_{\nu}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}.$
 It follows that
-$W(\theta^{\nu})$ $\leq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
-$\leq\mathbb{E}_{\widetilde{I}^{\varepsilon}_{\nu}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{N_{\nu}}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
-$=\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{I}^{\varepsilon}_{\nu}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]+\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right],$
-where the equality holds since both $\theta^{\nu}$ and $P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})$ are independent of $\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$. Then we obtain
-$\mathcal{E}_{2}^{\nu}$ $\leq\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$ (20)
+$W(\theta^{\nu})$ $\leq\mathbb{E}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{\nu}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
+$\leq\mathbb{E}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}}\left[\min\left\{\mathbb{E}_{N_{\nu}}\left[\widehat{F}_{\nu}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}^{+}\right],F(\theta^{\nu})\right\}+\frac{\eta}{2}\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
+$=\mathbb{E}\left[\min\left\{\widehat{F}(P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu});\theta^{\nu},\widetilde{\mathbf{I}}^{\varepsilon}_{\theta^{\nu}}),F(\theta^{\nu})\right\}\,\middle|\,\mathcal{G}_{\nu-1}\right]+\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right],$
+where the equality holds since both $\theta^{\nu}$ and $P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})$ are independent of $\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$.
+Then we obtain
+$\mathcal{E}_{2}^{\nu}$ $\leq\frac{\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$ (28)
 Combining bounds (8.5) and (28) of $\mathcal{E}_{1}^{\nu}$ and $\mathcal{E}_{2}^{\nu}$ and rearranging terms, we derive that
-$\frac{\rho-\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+C_{1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
+$\frac{\rho-\eta}{2}\mathbb{E}\left[\,\big{\lVert}\,{P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+C_{1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
 By applying Jensen’s inequality, taking full expectation on both sides, and summing inequalities up over $\nu=0,\dots,T-1$, we have
-$\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\big{(}\,\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\theta^{\nu})]\,\big{)}^{2}\leq\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\mathbb{E}\,\big{\lVert}\,{P^{\rho}_{\widetilde{I}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{0})-\mathbb{E}\left[F(\theta^{T})\right]+C_{1}\cdot\sum_{\nu=0}^{T-1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
+$\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\big{(}\,\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\theta^{\nu})]\,\big{)}^{2}\leq\frac{\rho-\eta}{2}\sum_{\nu=0}^{T-1}\mathbb{E}\,\big{\lVert}\,{P^{\rho}_{\widetilde{\mathbf{I}}^{\varepsilon}_{\nu}}(\theta^{\nu})-\theta^{\nu}}\,\big{\rVert}^{2}\leq F(\theta^{0})-\mathbb{E}\left[F(\theta^{T})\right]+C_{1}\cdot\sum_{\nu=0}^{T-1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}}.$
 Dividing both sides by $T$, applying Jensen’s inequality again, and using the fact that $\mathbb{E}\left[F(\theta^{T})\right]\geq\min_{\theta\in\Theta}F(\theta)$, we finally obtain that
@@
 Now we consider the special case with $\varepsilon=0$ and $\eta>0$. In such a case, ${\theta}^{\nu+\frac{1}{2}}=\theta^{\nu+1}$, and the intermediate term $W(\theta^{\nu})$ degenerates into
-$W(\theta^{\nu})=\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$
+$W(\theta^{\nu})=\mathbb{E}\left[\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right].$
 The term $\mathcal{E}_{2}^{\nu}$ is still bounded by (28).
@@
 Specifically, there exists $C_{2}>0$ such that
-$\mathcal{E}_{1}^{\nu}$ $=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
-$\leq F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
-$=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
-$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})-\widehat{F}_{N_{\nu}}(\theta^{\nu+1};\theta^{\nu},\widetilde{I}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
+$\mathcal{E}_{1}^{\nu}$ $=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\nu})+\frac{\eta}{2}\left\lVert\theta^{\nu+1}-\theta^{\nu}\right\rVert^{2}\,\middle|\,\mathcal{G}_{\nu-1}\right]$
+$\leq F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
+$=F(\theta^{\nu})-\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\theta_{\nu}})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\theta_{\nu}})\,\middle|\,\mathcal{G}_{\nu-1}\right]-\mathbb{E}\left[\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\theta_{\nu}})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
+$\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\mathbb{E}\left[\widehat{F}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\theta_{\nu}})-\widehat{F}_{\nu}(\theta^{\nu+1};\theta^{\nu},\widetilde{\mathbf{I}}^{0}_{\nu})\,\middle|\,\mathcal{G}_{\nu-1}\right]$
 $\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\,\middle|\,\mathcal{G}_{\nu-1}\right]+\frac{C_{2}}{N_{\nu}},$
-where the last inequality holds by Lemma 10.4, and the Lipschitz continuity of surrogates follows from the bi-variate continuity of $\widehat{F}(\cdot;\cdot,\xi,I)$, compactness of $\Theta$, and finiteness of $S$.
+where the last inequality holds by Lemma 8.4, and the Lipschitz continuity of surrogates follows from the bi-variate continuity of $\widehat{F}(\cdot,\xi;\cdot,I)$, compactness of $\Theta$, and finiteness of $\Xi$.
 Then the second inequality in the proposition can be proved with arguments similar to those above.∎
-10.3 Proofs of Theorems 4.4 and 4.5
-In this section, we prove the convergence results in Theorems 4.4 and 4.5.
-As in Section 4.2, we suppose throughout this section that $\varphi$ is B-differentiable, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot;\theta^{\prime},\xi,I)\}_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}$ of $F(\cdot;\xi)$ satisfies Assumption 1 for each $\xi\in S$.
+8.3 Proofs of Theorems 3.6 and 3.7
+In this section, we prove the convergence results in Theorems 3.6 and 3.7.
+As in Section 3.2.3, we suppose throughout this section that $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}$ of $F(\cdot,\xi)$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$.
 The convergence analysis requires the following two residual functions.
-The first residual of a reference point $\theta^{\prime}\in\Theta$ with $\varepsilon\geq 0$ and $\rho>0$ measures the proximal mapping gap with respect to the surrogation function $\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime})\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})}\widehat{F}(\theta;\theta^{\prime},I)$, which is defined as
-$r^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\max_{\theta\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta^{\prime})}\left\lVert\theta-\theta^{\prime}\right\rVert,\text{\ where\ }\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta^{\prime})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta;\theta^{\prime})+\frac{\rho}{2}\left\lVert\theta-\theta^{\prime}\right\rVert^{2}\right\}.$ (21)
+The first residual of a reference point $\bar{\theta}\in\Theta$ with $\varepsilon\geq 0$ and $\rho>0$ measures the proximal mapping gap with respect to the surrogate function $\widehat{F}^{\varepsilon}_{\min}(\theta;\bar{\theta})\coloneqq\min_{\mathbf{I}^{\varepsilon}\in\mathcal{I}^{\varepsilon}(\bar{\theta})}\widehat{F}(\theta;\bar{\theta},\mathbf{I}^{\varepsilon})$, which is defined as
+$r^{\varepsilon,\rho}(\bar{\theta})\coloneqq\max_{\theta\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\bar{\theta})}\left\lVert\theta-\bar{\theta}\right\rVert,\text{\ where\ }\mathcal{P}_{\min}^{\varepsilon,\rho}(\bar{\theta})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta;\bar{\theta})+\frac{\rho}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}.$ (29)
 Clearly, $\theta^{d}\in\Theta$ is a composite $(\varepsilon,\rho)$-strongly d-stationary point of problem (10) if and only if $r^{\varepsilon,\rho}(\theta^{d})=0$.
@@
 This residual is utilized to derive the global error bound in terms of $r^{\varepsilon,\rho}$ in the second statement of Proposition 8.6.
-In Figure 10.3, we summarize a roadmap for the convergence analysis of the enhanced SMM algorithm.
-A Roadmap for Convergence Analysis of Enhanced SMM.
+In Figure 8.3, we summarize a roadmap for the convergence analysis of the ESMM algorithm.
+[…]
+A roadmap for convergence analysis of ESMM
 We first provide in the following proposition the limiting property of the residual $r^{\varepsilon,\rho}$ and the global error bound of distance to $\mathcal{S}^{d}$ in terms of $r^{\varepsilon,\rho}$.
@@
   2. (b)
-    Under Assumptions 2 and 1, for $\varepsilon>0$ and $\rho>0$, there exists a constant $\eta_{r}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}r^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$.
-In particular, if $\varphi$ is piecewise affine and $\Theta$ is a polytope, the error bound holds for any $\varepsilon\geq 0$ and $\rho>0$ without the two assumptions.
-Before proving this proposition, we show that Theorem 4.4 could be derived straightforwardly by combining Propositions 4.3 and 10.6, similarly for Theorem 4.5.
-Proof 10.7
+    For $\varepsilon\geq 0$ and $\rho>0$, if $\varphi$ is piecewise affine and $\Theta$ is a polytope, there exists a constant $\eta_{r}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}r^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$.
+Before proving this proposition, we show that Theorem 3.6 could be derived straightforwardly by combining Propositions 3.5 and 8.6, similarly for Theorem 3.7.
+Proof 8.7
 Proof of Theorem 3.6.
-According to the definition in (21), for any $\theta\in\mathcal{P}^{\varepsilon,\rho}_{\min}(\theta^{\prime})$, there exists $I\in\mathcal{I}^{\varepsilon}(\theta^{\prime})$ such that $\theta=P^{\rho}_{I}(\theta^{\prime})$.
-Due to the finite-sum structure of the objective $F$, there exists a uniform constant $C_{\mathcal{I}}>0$ such that the cardinality $\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert\leq C_{\mathcal{I}}$ for any $\theta^{\prime}\in\Theta$, and thus ${r}^{\varepsilon,\rho}(\theta^{\prime})\leq\left\lvert\mathcal{I}^{\varepsilon}(\theta^{\prime})\right\rvert\bar{r}^{\varepsilon,\rho}(\theta^{\prime})\leq C_{\mathcal{I}}\bar{r}^{\varepsilon,\rho}(\theta^{\prime})$, which implies that $\mathbb{E}\,[{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ converges to 0 at least at the same rate as $\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ in terms of $\{N_{\nu}\}_{\nu}$ and $T$.
-Therefore, the statements of the theorem follow from Propositions 4.3 and 10.6.∎
-We next present the proof of Proposition 10.6.
-The overall scheme of the error-bound analysis follows the error-bound theory developed in Liu et al.
-(2022) for the post-convergence analysis of the SMM algorithm.
-To justify the global error bound in the second statement of Proposition 10.6, we first give the error bounds of distance to $\mathcal{S}^{d}$ in terms of the residual $r_{d}$ in Proposition 10.8 and then relate $r_{d}$ with $r^{\varepsilon,\rho}$ in Proposition 10.10.
+According to the definition in (29), for any $\theta\in\mathcal{P}^{\varepsilon,\rho}_{\min}(\bar{\theta})$, there exists $\mathbf{I}^{\varepsilon}\in\mathcal{I}^{\varepsilon}(\bar{\theta})$ such that $\theta=P^{\rho}_{\mathbf{I}^{\varepsilon}}(\bar{\theta})$.
+Due to the finite-sum structure of the objective $F$, there exists a uniform constant $C_{\mathcal{I}}>0$ such that the cardinality $\left\lvert\mathcal{I}^{\varepsilon}(\bar{\theta})\right\rvert\leq C_{\mathcal{I}}$ for any $\bar{\theta}\in\Theta$, and thus ${r}^{\varepsilon,\rho}(\bar{\theta})\leq\left\lvert\mathcal{I}^{\varepsilon}(\bar{\theta})\right\rvert\bar{r}^{\varepsilon,\rho}(\bar{\theta})\leq C_{\mathcal{I}}\bar{r}^{\varepsilon,\rho}(\bar{\theta})$, which implies that $\mathbb{E}\,[{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ converges to 0 at least at the same rate as $\mathbb{E}\,[\bar{r}^{\varepsilon,\rho}(\widetilde{\theta}^{\,T})]$ in terms of $\{N_{\nu}\}_{\nu}$ and $T$.
+Therefore, the statements of the theorem follow from Propositions 3.5 and 8.6.∎
+[…]
 Proposition 8.8
-For problem (9), the following statements hold:
-  1. (a)
-    If $\varphi$ is piecewise affine and $\Theta$ is a polyhedral set, then there exists $\eta_{d}>0$ such that
+For problem (10), if $\varphi$ is piecewise affine and $\Theta$ is a polyhedral set, then there exists $\eta_{d}>0$ such that
 $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{d}r_{d}(\theta),\quad\forall\,\theta\in\Theta.$
-  2. (b)
-    Under Assumption 1 with $\varepsilon>0$, $\rho>0$, and the corresponding constants $\delta_{q}$, $\lambda_{q}$, and $c_{q}$,
-    $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\frac{2}{c_{q}}r_{d}(\theta),\quad\text{$\forall\,\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$ and $r^{\varepsilon,\rho}(\theta)\leq\lambda_{q}$.}$
-Proof 10.9
+Proof 8.9
 Proof of Proposition 8.8.
-For the first statement, since the objective $F$ is a piecewise affine function on a polyhedral set $\Theta$, $\Theta$ can be decomposed into finitely many polyhedral sets such that $F$ is affine on each of them, showing that $r_{d}(\cdot)$ takes finitely many values on $\Theta$.
+Since the objective $F$ is a piecewise affine function on a polyhedral set $\Theta$, $\Theta$ can be decomposed into finitely many polyhedral sets such that $F$ is affine on each of them, showing that $r_{d}(\cdot)$ takes finitely many values on $\Theta$.
 Notice that $\theta\in\mathcal{S}^{d}$ if and only if $r_{d}(\theta)=\max_{d\in\mathcal{T}(\theta;\Theta),\left\lVert d\right\rVert\leq 1}\{-F^{\prime}(\theta;d)\}=0$.
@@
 By setting $\eta_{d}\coloneqq\frac{2}{l_{0}}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, we deduce $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq 2\max_{\theta\in\Theta}\left\lVert\theta\right\rVert\leq\eta_{d}r_{d}(\theta)$.
-For the second statement, the inequality holds trivially if $\theta\in\mathrm{cl}(\mathcal{S}^{d})$.
-By Assumption 1, for any $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$ satisfying $r^{\varepsilon,\rho}(\theta)\leq\lambda_{q}$ and $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(\cdot;\theta^{d},\theta)$ is convex and $\bar{F}(t;\theta^{d},\theta)\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{q}}{2}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2}$ for all $t\in[0,1].$ According to Drusvyatskiy and Lewis (2018, Theorem 3.3), it follows that
-$\left\lVert\theta-\theta^{d}\right\rVert\leq\frac{2}{c_{q}}\mathrm{dist}(0,\partial\bar{F}(1;\theta^{d},\theta))\leq\frac{2}{c_{q}}\left[-\bar{F}(\cdot;\theta^{d},\theta)^{\prime}(1;-1)\right]=\frac{2}{c_{q}}\left[-F^{\prime}\left(\theta;\frac{\theta^{d}-\theta}{\left\lVert\theta^{d}-\theta\right\rVert}\right)\right]\leq\frac{2}{c_{q}}r_{d}(\theta).$
-Thus, we obtain $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\left\lVert\theta-\theta^{d}\right\rVert\leq\frac{2}{c_{q}}r_{d}(\theta)$, which completes the proof.∎
+∎
 Proposition 8.10
-Under Assumption 2, for $\varepsilon\geq 0$ and $\rho>0$, it holds that for any $\theta\in\Theta$, there exists $\bar{\theta}\in\Theta$ such that $\left\lVert\bar{\theta}-\theta\right\rVert\leq 2\,r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\bar{\theta})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}(\theta)$.
-A similar result to Proposition 10.10 is provided in Liu et al.
-(2022, Lemma 4) under the assumption that the relevant surrogation function satisfies Condition (P5).
-With the surrogation function $\widehat{F}^{\varepsilon}_{\min}(\cdot;\theta^{\prime})$ associated with $r^{\varepsilon,\rho}(\theta^{\prime})$ satisfying (P5), we can prove Proposition 10.10 following the lines of Liu et al.
-(2022, Lemma 4), and thus the details are omitted here.
-Proof 10.11
+Suppose that the minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ satisfies the quadratic surrogation gap condition (P4) for any $\varepsilon\geq 0$ and $\xi\in\Xi$.
+For problem (10) with $\varepsilon\geq 0$ and $\rho>0$, it holds that for any $\theta\in\Theta$, there exists $\theta^{\prime}\in\Theta$ such that $\left\lVert\theta^{\prime}-\theta\right\rVert\leq 2\,r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\theta^{\prime})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}(\theta)$.
+[…]
+Proof 8.11
 Proof of Proposition 8.6.
@@
 Given $\varepsilon^{\prime}\in[0,\varepsilon)$, it follows from the Lipschitz continuity of $f$ that $\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\subseteq\mathcal{I}^{\varepsilon}(\theta_{t})$ for sufficiently large $t\in\mathcal{K}$.
-With $\widehat{P}^{\rho}_{I}(\theta_{t})\coloneqq\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta_{t},I)+\frac{\rho}{2}\left\lVert\theta-\theta_{t}\right\rVert^{2}\right\}$, by definition of the residual $r^{\varepsilon,\rho}$ there exists $I^{*}_{t}\in\underset{I\in\mathcal{I}^{\varepsilon}(\theta_{t})}{\arg\min}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon,\rho}(\theta_{t})=\,\big{\lVert}\,{\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}}\,\big{\rVert}$.
+With $\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t})\coloneqq\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta_{t},\mathbf{I})+\frac{\rho}{2}\left\lVert\theta-\theta_{t}\right\rVert^{2}\right\}$, by definition of the residual $r^{\varepsilon,\rho}$ there exists $\mathbf{I}^{*}_{t}\in\underset{\mathbf{I}\in\mathcal{I}^{\varepsilon}(\theta_{t})}{\arg\min}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t});\theta_{t},\mathbf{I})+\frac{{\rho}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t})-\theta_{t}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon,\rho}(\theta_{t})=\,\big{\lVert}\,{\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})-\theta_{t}}\,\big{\rVert}$.
 Then for sufficiently large $t\in\mathcal{K}$, we have that
-$\widehat{F}^{*}(\theta_{t})$ $\coloneqq\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
-$\geq\min_{I\in\mathcal{I}^{\varepsilon}(\theta_{t})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{t});\theta_{t},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
-$=\widehat{F}(\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t});\theta_{t},I^{*}_{t})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}$
-$\geq F(\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t}))+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}.$ (22)
-Since $\lim_{t(\in\mathcal{K})\rightarrow\infty}r^{{\varepsilon},{\rho}}(\theta_{t})=0$, we deduce $\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{P}^{\rho}_{I^{*}_{t}}(\theta_{t})=\theta_{\infty}$.
-Additionally, for a fixed mapping $I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})$, the surrogation function $\widehat{F}(\cdot;\cdot,I)$ is continuous on $\Theta\times\Theta$ by assumption, and thus by Shapiro et al.
-(2021, Theorem 7.23), $\widehat{P}^{\rho}_{I}(\cdot)$ is continuous at $\theta_{\infty}$.
-It follows from (10.11) that
-$F(\theta_{\infty})\leq\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{F}^{*}(\theta_{t})=\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{I}(\theta_{\infty});\theta_{\infty},I)+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{I}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\},$
-which leads to $F(\theta_{\infty})<\displaystyle{\min_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\backslash\mathcal{I}^{0}(\theta_{\infty})}}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{I}(\theta_{\infty});\theta_{\infty},I)+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,\widehat{P}^{\rho^{\prime}}_{I}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\}$ for $\rho^{\prime}>\rho$, yielding $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$.
+$\widehat{F}^{*}(\theta_{t})$ $\coloneqq\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t});\theta_{t},\mathbf{I})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
+$\geq\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon}(\theta_{t})}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t});\theta_{t},\mathbf{I})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}\right\}$
+$=\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t});\theta_{t},\mathbf{I}^{*}_{t})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}$
+$\geq F(\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t}))+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})-\theta_{t}\,\big{\rVert}^{2}.$ (30)
+[…]
+$F(\theta_{\infty})\leq\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{F}^{*}(\theta_{t})=\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{\infty});\theta_{\infty},\mathbf{I})+\frac{{\rho}}{2}\,\big{\lVert}\,\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\},$
+which leads to $F(\theta_{\infty})<\displaystyle{\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})\backslash\mathcal{I}^{0}(\theta_{\infty})}}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\mathbf{I}}(\theta_{\infty});\theta_{\infty},\mathbf{I})+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,\widehat{P}^{\rho^{\prime}}_{\mathbf{I}}(\theta_{\infty})-\theta_{\infty}\,\big{\rVert}^{2}\right\}$ for $\rho^{\prime}>\rho$, yielding $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$.
 We next prove the second statement.
-For the special case of $\varphi$ being piecewise affine, Assumption 2 holds naturally.
-Given $\varepsilon\geq 0$ and $\rho>0$, by Proposition 10.8 there exists $\eta_{d}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{d}r_{d}(\theta)$ for any $\theta\in\Theta$; by Proposition 10.10, for any $\theta\in\Theta$, there exists $\bar{\theta}\in\Theta$ such that $\left\lVert\bar{\theta}-\theta\right\rVert\leq 2r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\bar{\theta})\leq(2\rho+\frac{5\kappa}{2})r^{\varepsilon,\rho}(\theta)$.
+If $\varphi$ is piecewise affine, by Proposition 3.2 in Appendix 7, the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ satisfies Condition (P4) for any $\varepsilon\geq 0$ and $\xi\in\Xi$.
+Given $\varepsilon\geq 0$ and $\rho>0$, by Proposition 8.8 there exists $\eta_{d}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{d}r_{d}(\theta)$ for any $\theta\in\Theta$; by Proposition 8.10, for any $\theta\in\Theta$, there exists $\theta^{\prime}\in\Theta$ such that $\left\lVert\theta^{\prime}-\theta\right\rVert\leq 2r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\theta^{\prime})\leq(2\rho+\frac{5\kappa}{2})r^{\varepsilon,\rho}(\theta)$.
 Setting $\eta_{r}=2+\eta_{d}\left(2\rho+\frac{5\kappa}{2}\right)$, we have that
-$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\eta_{d}r_{d}(\bar{\theta})+2r^{\varepsilon,\rho}(\theta)\leq\eta_{r}r^{\varepsilon,\rho}(\theta).$
-To prove the desired statement for general B-differentiable $\varphi$ under Assumptions 2 and 1, we first show that given $\varepsilon>\varepsilon^{\prime}\geq 0$ and $\rho^{\prime}>\rho>0$, for any $\lambda^{\prime}_{d}>0$ there exists $\lambda_{d}>0$ such that
+$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\theta^{\prime},\mathcal{S}^{d})+\left\lVert\theta-\theta^{\prime}\right\rVert\leq\eta_{d}r_{d}(\theta^{\prime})+2r^{\varepsilon,\rho}(\theta)\leq\eta_{r}r^{\varepsilon,\rho}(\theta),$
+which completes the proof.∎
+8.4 Nonasymptotic convergence result for general cases
+In this section, we extend the nonasymptotic convergence result for the general case where the outer function $\varphi$ is directionally differentiable and locally Lipschitz continuous, and the surrogate functions of $F$ satisfy Assumptions (B1)-(B4).
+The error-bound analysis in this case requires the following two assumptions.
+We denote the closure operator of a set by $\mathrm{cl}(\cdot)$ and $\mathcal{P}_{\min}^{\varepsilon,\rho}(\bar{\theta})\coloneqq\mathop{\arg\min}_{\theta\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta;\bar{\theta})+\frac{\rho}{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}\right\}$.
+  1. (B5)
+    The minimum surrogate $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})$ satisfies the quadratic surrogation gap condition (P4) for any $\varepsilon\geq 0$ and $\xi\in\Xi$;
+  2. (B6)
+    generalized quadratic growth: given $\varepsilon>0$ and $\rho>0$, there exist positive constants $\delta_{q}$, $\lambda_{q}$, and $c_{q}$ such that for any $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$ and $\max_{\theta^{\prime}\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)}\left\lVert\theta^{\prime}-\theta\right\rVert\leq\lambda_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(t;\theta^{d},\theta)\coloneqq F(\theta^{d}+t(\theta-\theta^{d}))$ is convex on $[0,1]$ and
+    $\bar{F}(t;\theta^{d},\theta)\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{q}}{2}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2},\quad\forall\,t\in[0,1].$
+Assumption (B5) has appeared in Proposition 8.10, and we state it here officially.
+Condition (P4) originating from Liu et al.
+(2022) is a technical condition for the nonasymptotic convergence analysis of the ESMM algorithm.
+Note that Condition (P4) is verified for Case 1-Case 3 in Online Appendix 7.
+Assumption (B6) is generalized from Drusvyatskiy and Lewis (2018).
+The condition holds, for example, when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope, which is verified in Online Appendix 8.5.
+It can be expected to hold in more general cases since it is a local property only requiring the convexity and quadratic growth of $F$ near $\mathcal{S}^{d}$.
+With those assumptions, the nonasymptotic convergence result of ESMM for general cases is stated as follows.
+Theorem 8.12 (Nonasymptotic convergence of ESMM for general cases)
+Under Assumptions (B1)-(B6), for problem (10) under the same settings of Algorithm 1 specified in Theorem 3.6, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{\,T},\mathcal{S}^{d})]=\widetilde{\mathcal{O}}\Big{(}T^{-\frac{1}{2}\min\left\{\frac{\alpha}{2},1\right\}}\Big{)}$.
+Following the error-bound analysis developed in Section 8.3, the convergence result can be established with the following proposition.
+Proposition 8.13
+Under Assumptions (B1)-(B6), for $\varepsilon>0$ and $\rho>0$, there exists a constant $\eta_{r}>0$ such that $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}r^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$.
+Similar to the proof of the second statement of Proposition 8.6, to prove the above error bound, we require the following proposition, which is a generalization of Proposition 8.8.
+Proposition 8.14
+Under Assumptions (B1)-(B6), for problem (10)with $\varepsilon>0$, $\rho>0$, and the corresponding constants $\delta_{q}$, $\lambda_{q}$, and $c_{q}$ in Assumption (B6),
+$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\frac{2}{c_{q}}r_{d}(\theta),\quad\text{$\forall\,\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\delta_{q}$ and $r^{\varepsilon,\rho}(\theta)\leq\lambda_{q}$.}$
+Proof 8.15
+[…]
+$\left\lVert\theta-\theta^{d}\right\rVert\leq\frac{2}{c_{q}}\mathrm{dist}(0,\partial\bar{F}(1;\theta^{d},\theta))\leq\frac{2}{c_{q}}\left[-\bar{F}(\cdot;\theta^{d},\theta)^{\prime}(1;-1)\right]=\frac{2}{c_{q}}\left[-F^{\prime}\left(\theta;\frac{\theta^{d}-\theta}{\left\lVert\theta^{d}-\theta\right\rVert}\right)\right]\leq\frac{2}{c_{q}}r_{d}(\theta).$
+Thus, we obtain $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\left\lVert\theta-\theta^{d}\right\rVert\leq\frac{2}{c_{q}}r_{d}(\theta)$, which completes the proof.∎
+Proof 8.16
+Proof of Proposition 8.13. We first show that given $\varepsilon>\varepsilon^{\prime}\geq 0$ and $\rho^{\prime}>\rho>0$, for any $\lambda^{\prime}_{d}>0$ there exists $\lambda_{d}>0$ such that
 $\Big{\{}\theta^{\prime}\in\Theta:r^{\varepsilon,\rho}(\theta)\leq\lambda_{d},\ \lVert\theta^{\prime}-\theta\rVert\leq 2\lambda_{d},\ \theta\in\Theta\Big{\}}\subseteq\left\{\theta\in\Theta:r^{\varepsilon^{\prime},\rho^{\prime}}(\theta)\leq\lambda_{d}^{\prime}\right\}.$ (31)
-By contradiction, suppose that there exists $\bar{\lambda}^{\prime}_{d}>0$ such that for any $\lambda>0$, there exist points $\theta,\theta^{\prime}\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq\lambda$, $\left\lVert\theta-\theta^{\prime}\right\rVert\leq 2\lambda$, and $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime})>\bar{\lambda}^{\prime}_{d}$.
-Then there exist two sequences $\{\theta_{k}\}_{k}$ and $\{\theta^{\prime}_{k}\}_{k}$ in $\Theta$ with $\inf_{k\in\mathbb{N}}r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime}_{k})\geq\bar{\lambda}^{\prime}_{d}$ such that both $r^{\varepsilon,\rho}(\theta_{k})$ and $\left\lVert\theta_{k}-\theta^{\prime}_{k}\right\rVert$ converge to 0.
-Since $\Theta$ is compact, by passing to a subsequence if necessary, we can assume that both $\{\theta_{k}\}_{k}$ and $\{\theta^{\prime}_{k}\}_{k}$ converge to $\theta_{\infty}\in\Theta$.
+By contradiction, suppose that there exists $\bar{\lambda}^{\prime}_{d}>0$ such that for any $\lambda>0$, there exist points $\theta,\bar{\theta}\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq\lambda$, $\left\lVert\theta-\bar{\theta}\right\rVert\leq 2\lambda$, and $r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta})>\bar{\lambda}^{\prime}_{d}$.
+Then there exist two sequences $\{\theta_{k}\}_{k}$ and $\{\bar{\theta}_{k}\}_{k}$ in $\Theta$ with $\inf_{k\in\mathbb{N}}r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta}_{k})\geq\bar{\lambda}^{\prime}_{d}$ such that both $r^{\varepsilon,\rho}(\theta_{k})$ and $\left\lVert\theta_{k}-\bar{\theta}_{k}\right\rVert$ converge to 0.
+Since $\Theta$ is compact, by passing to a subsequence if necessary, we can assume that both $\{\theta_{k}\}_{k}$ and $\{\bar{\theta}_{k}\}_{k}$ converge to $\theta_{\infty}\in\Theta$.
 By the first statement of the proposition, we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})=0$.
-Let $I_{k}\in{\arg\min}_{I\in\mathcal{I}^{\varepsilon^{\prime}}(\theta^{\prime}_{k})}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{I}(\theta^{\prime}_{k});\theta^{\prime}_{k},I)+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{I}(\theta^{\prime}_{k})-\theta^{\prime}_{k}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime}_{k})=\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{I_{k}}(\theta^{\prime}_{k})-\theta^{\prime}_{k}}\,\big{\rVert}>\bar{\lambda}^{\prime}_{d}$.
-Since the total number of index mappings is finite, by passing to a subsequence if necessary, we can assume that $I_{k}\equiv\widehat{I}$ for any $k\in\mathbb{N}$.
-By continuity of $\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\cdot)$ at $\theta_{\infty}$, we derive $\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}\geq\bar{\lambda}^{\prime}_{d}$ and $\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty});\theta_{\infty},\widehat{I})+\frac{\rho^{\prime}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{I}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}^{2}\leq F(\theta_{\infty})$, implying $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})>0$, a contradiction.
-Given $\varepsilon^{\prime}\in(0,\varepsilon)$ and $\rho^{\prime}\in(\rho,+\infty)$, let $\delta_{q}^{\prime},\lambda_{q}^{\prime}$, and $c_{q}^{\prime}$ be the corresponding constants in Assumption 1.
+Let $\mathbf{I}_{k}\in{\arg\min}_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\bar{\theta}_{k})}\left\{\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\mathbf{I}}(\bar{\theta}_{k});\bar{\theta}_{k},\mathbf{I})+\frac{{\rho^{\prime}}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\mathbf{I}}(\bar{\theta}_{k})-\bar{\theta}_{k}}\,\big{\rVert}^{2}\right\}$ such that $r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta}_{k})=\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\mathbf{I}_{k}}(\bar{\theta}_{k})-\bar{\theta}_{k}}\,\big{\rVert}>\bar{\lambda}^{\prime}_{d}$.
+Since the total number of index combination is finite, by passing to a subsequence if necessary, we can assume that $\mathbf{I}_{k}\equiv\widehat{\mathbf{I}}$ for any $k\in\mathbb{N}$.
+By continuity of $\widehat{P}^{\rho^{\prime}}_{\widehat{\mathbf{I}}}(\cdot)$ at $\theta_{\infty}$, we derive $\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{\mathbf{I}}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}\geq\bar{\lambda}^{\prime}_{d}$ and $\widehat{F}(\widehat{P}^{\rho^{\prime}}_{\widehat{\mathbf{I}}}(\theta_{\infty});\theta_{\infty},\widehat{\mathbf{I}})+\frac{\rho^{\prime}}{2}\,\big{\lVert}\,{\widehat{P}^{\rho^{\prime}}_{\widehat{\mathbf{I}}}(\theta_{\infty})-\theta_{\infty}}\,\big{\rVert}^{2}\leq F(\theta_{\infty})$, implying $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta_{\infty})>0$, a contradiction.
+Given $\varepsilon^{\prime}\in(0,\varepsilon)$ and $\rho^{\prime}\in(\rho,+\infty)$, let $\delta_{q}^{\prime},\lambda_{q}^{\prime}$, and $c_{q}^{\prime}$ be the corresponding constants in Assumption (B6).
 There exists $\lambda_{q}>0$ such that (31) holds for $\lambda_{q}^{\prime}$.
-For $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\frac{\delta_{q}^{\prime}}{3}$ and $r^{\varepsilon,\rho}(\theta)\leq\lambda_{0}\coloneqq\min\left\{\lambda_{q},\frac{\delta_{q}^{\prime}}{3}\right\}$, by Proposition 10.10 let $\bar{\theta}$ be the point satisfying $\left\lVert\bar{\theta}-{\theta}\right\rVert\leq 2\,r^{\varepsilon,\rho}({\theta})\leq 2\lambda_{q}$ and $r_{d}(\bar{\theta})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}({\theta})$.
-Then we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\bar{\theta})\leq\lambda_{q}^{\prime}$ from (23) and
-$\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})\leq\mathrm{dist}(\theta,\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\frac{\delta_{q}^{\prime}}{3}+2r^{\varepsilon,\rho}(\theta)\leq\delta_{q}^{\prime}.$
-It follows from Proposition 10.8 that $\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})\leq\frac{2}{c_{q}^{\prime}}r_{d}(\bar{\theta})$. We further obtain that
-$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\bar{\theta},\mathcal{S}^{d})+\left\lVert\theta-\bar{\theta}\right\rVert\leq\frac{2}{c_{q}^{\prime}}r_{d}(\bar{\theta})+2r^{\varepsilon,\rho}(\theta)\leq\left(\frac{4\rho+5\kappa}{c_{q}^{\prime}}+2\right)r^{\varepsilon,\rho}(\theta).$
+For $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\frac{\delta_{q}^{\prime}}{3}$ and $r^{\varepsilon,\rho}(\theta)\leq\lambda_{0}\coloneqq\min\left\{\lambda_{q},\frac{\delta_{q}^{\prime}}{3}\right\}$, by Proposition 8.10 let $\theta^{\prime}$ be the point satisfying $\left\lVert\theta^{\prime}-{\theta}\right\rVert\leq 2\,r^{\varepsilon,\rho}({\theta})\leq 2\lambda_{q}$ and $r_{d}(\theta^{\prime})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}({\theta})$.
+Then we have $r^{\varepsilon^{\prime},\rho^{\prime}}(\theta^{\prime})\leq\lambda_{q}^{\prime}$ from (31) and
+$\mathrm{dist}(\theta^{\prime},\mathcal{S}^{d})\leq\mathrm{dist}(\theta,\mathcal{S}^{d})+\left\lVert\theta-\theta^{\prime}\right\rVert\leq\frac{\delta_{q}^{\prime}}{3}+2r^{\varepsilon,\rho}(\theta)\leq\delta_{q}^{\prime}.$
+It follows from Proposition 8.8 that $\mathrm{dist}(\theta^{\prime},\mathcal{S}^{d})\leq\frac{2}{c_{q}^{\prime}}r_{d}(\theta^{\prime})$. We further obtain that
+$\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\mathrm{dist}(\theta^{\prime},\mathcal{S}^{d})+\left\lVert\theta-\theta^{\prime}\right\rVert\leq\frac{2}{c_{q}^{\prime}}r_{d}(\theta^{\prime})+2r^{\varepsilon,\rho}(\theta)\leq\left(\frac{4\rho+5\kappa}{c_{q}^{\prime}}+2\right)r^{\varepsilon,\rho}(\theta).$
 Next, for $\theta\in\Theta$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{d})>\frac{\delta_{q}^{\prime}}{3}$, by the first statement of the proposition, there exists $\lambda_{1}>0$ such that $r^{\varepsilon,\rho}(\theta)\geq\lambda_{1}$.
@@
 Therefore, we obtain $\mathrm{dist}(\theta,\mathcal{S}^{d})\leq\eta_{r}{r}^{\varepsilon,\rho}(\theta)$ for any $\theta\in\Theta$ with $\eta_{r}=\max\left\{\frac{4\rho+5\kappa}{c_{q}^{\prime}}+2,\frac{2\sqrt{q}\mu}{\lambda_{1}},\frac{2\sqrt{q}\mu}{\min\{\lambda_{q},\delta_{q}^{\prime}/3\}}\right\}$, which completes the proof.∎
-10.4 Verification of Assumption 1
-We verify that Assumption 1 holds when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope by proving a stronger result in Claim 2 with the following lemma.
-Lemma 10.12 (Cui et al. (2018), Proposition 2 (ii))
+8.5 Verification of Assumption (B6)
+We verify that Assumption (B6) holds when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope by proving a stronger result in Claim 1 with the following lemma.
+[…]
 Let $\Theta\subseteq\mathbb{R}^{q}$ be a closed convex set.
@@
 Recall the notations $\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)\coloneqq\mathop{\arg\min}_{\theta^{\prime}\in\Theta}\left\{\widehat{F}^{\varepsilon}_{\min}(\theta^{\prime};\theta)+\frac{\rho}{2}\left\lVert\theta^{\prime}-\theta\right\rVert^{2}\right\}$ and $\bar{F}(t;\theta_{1},\theta_{2})\coloneqq F(\theta_{1}+t(\theta_{2}-\theta_{1}))$ for $\theta_{1},\theta_{2}\in\Theta$.
-In the analysis below, let $\partial(S)$ denote the boundary of $S$, $\mathrm{int}(S)$ denote the interior of $S$, and $\mathrm{rint}(S)$ denote the relative interior of $S$.
+In the analysis below, let $\partial(S)$ denote the boundary of $\Xi$, $\mathrm{int}(S)$ denote the interior of $\Xi$, and $\mathrm{rint}(S)$ denote the relative interior of $\Xi$.
 Claim 1
-Let $\Theta\subseteq\mathbb{R}^{q}$ be a polytope and $\varphi(\cdot;\xi)$ be a strongly convex function with modulus $c_{\varphi}$ for any $\xi\in S$.
+Let $\Theta\subseteq\mathbb{R}^{q}$ be a polytope and $\varphi(\cdot,\xi)$ be a strongly convex function with modulus $c_{\varphi}$ for any $\xi\in\Xi$.
 Given $\varepsilon>0$ and $\rho>0$, there exist $\lambda_{q}>0$ and $c_{q}>0$ such that for any $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\coloneqq\underset{\theta^{\prime}\in\mathcal{P}_{\min}^{\varepsilon,\rho}(\theta)}{\max}\,\left\lVert\theta-\theta^{\prime}\right\rVert\leq\lambda_{q}$, there exists $\theta^{d}\in\mathrm{cl}(\mathcal{S}^{d})$ such that $\bar{F}(\cdot;\theta^{d},\theta)$ is convex on $[0,1]$ and
 $\bar{F}(t;\theta^{d},\theta)\geq\bar{F}(0;\theta^{d},\theta)+\frac{c_{q}}{2}t^{2}\left\lVert\theta-\theta^{d}\right\rVert^{2},\quad\forall\,t\in[0,1].$
-Proof 10.13
+Proof 8.18
 Proof of Claim 1.
@@
 Given $\mathcal{C}\in\mathfrak{C}$, $f$ restricted to $\mathcal{C}$ can be represented as $f(\theta)=A\theta$ for some $A\in\mathbb{R}^{d\times q}$.
-Since the strongly convex function $\varphi$ has a unique minimizer on $\{A\theta\in\mathbb{R}^{d}\mid\theta\in\mathcal{C}\}$, the set of minimizers of $F$ on $\mathcal{C}$ denoted by $\mathcal{S}^{*}_{\mathcal{C}}$ can be represented as $\mathcal{S}^{*}_{\mathcal{C}}=[\{\bar{\theta}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}$ for any $\bar{\theta}\in\mathcal{S}^{*}_{\mathcal{C}}$, where $\mathcal{N}_{A}$ denotes the null space of $A$.
+Since the strongly convex function $\varphi$ has a unique minimizer on $\{A\theta\in\mathbb{R}^{d}\mid\theta\in\mathcal{C}\}$, the set of minimizers of $F$ on $\mathcal{C}$ denoted by $\mathcal{S}^{*}_{\mathcal{C}}$ can be represented as $\mathcal{S}^{*}_{\mathcal{C}}=[\{\theta^{\prime}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}$ for any $\theta^{\prime}\in\mathcal{S}^{*}_{\mathcal{C}}$, where $\mathcal{N}_{A}$ denotes the null space of $A$.
 We first show that $\mathcal{S}^{*}_{\mathcal{C}}=\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$ if $\mathcal{S}^{d}\bigcap\mathcal{C}\neq\emptyset$.
@@
 Since $\mathcal{S}^{d}\bigcap\mathrm{int}(\mathcal{C})=\emptyset$, by arguments above there exists a polytope $\mathcal{C}^{\prime}\in\mathfrak{C}$ such that $\mathcal{S}^{*}_{\mathcal{C}}\subseteq\mathcal{C}^{\prime}$ and $\min_{\theta\in\mathcal{S}^{*}_{\mathcal{C}}}\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}^{\prime}})>0$.
-Let $I_{\mathcal{C}^{\prime}}$ denote the index mapping of $f$ corresponding to $\mathcal{C}^{\prime}$.
-Due to the Lipschitz continuity of $F$ on $\Theta$, for any $\varepsilon>0$ there exist $\delta>0$ and $\Delta>0$ such that any $\theta\in\mathcal{C}$ with $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})\leq\delta$ satisfies $I_{\mathcal{C}^{\prime}}\in\mathcal{I}^{\varepsilon}(\theta)$ and $F(\theta)>\min_{\theta^{\prime}\in\mathcal{C}^{\prime}}F(\theta^{\prime})+\Delta$; consequently, for any given $\rho>0$, there exists $\lambda_{1}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{1}$.
+Let $\mathbf{I}_{\mathcal{C}^{\prime}}$ denote the index combination of $f$ corresponding to $\mathcal{C}^{\prime}$.
+Due to the Lipschitz continuity of $F$ on $\Theta$, for any $\varepsilon>0$ there exist $\delta>0$ and $\Delta>0$ such that any $\theta\in\mathcal{C}$ with $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})\leq\delta$ satisfies $\mathbf{I}_{\mathcal{C}^{\prime}}\in\mathcal{I}^{\varepsilon}(\theta)$ and $F(\theta)>\min_{\theta^{\prime}\in\mathcal{C}^{\prime}}F(\theta^{\prime})+\Delta$; consequently, for any given $\rho>0$, there exists $\lambda_{1}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{1}$.
 On the other hand, for any point $\theta\in\mathcal{C}$ satisfying $\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})>\delta$, by the first statement of Proposition 8.6, there exists $\lambda_{2}>0$ such that $r^{\varepsilon,\rho}(\theta)>\lambda_{2}$.
@@
 It yields that there exists $\lambda_{q}>0$ such that for any $\theta\in\Theta$ satisfying $r^{\varepsilon,\rho}(\theta)\leq{\lambda}_{q}$, we have $\mathcal{S}^{d}\bigcap\mathcal{C}\neq\emptyset$ and $\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}=\mathcal{S}^{*}_{\mathcal{C}}$ for any $\mathcal{C}\in\mathfrak{C}_{\theta}\coloneqq\{\mathcal{C}^{\prime}\in\mathfrak{C}\mid\theta\in\mathcal{C}^{\prime}\}$.
-Now, consider a point $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$ satisfying $r^{\varepsilon,\rho}(\theta)\leq{\lambda}_{q}$.
-Given a polytope $\mathcal{C}\in\mathfrak{C}_{\theta}$ denoted by $\mathcal{C}\coloneqq\{\theta\in\mathbb{R}^{q}\mid B\theta\leq b\}$, since $\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}=\mathcal{S}^{*}_{\mathcal{C}}$, we also have $\theta\in\mathcal{C}\backslash\mathcal{S}^{*}_{\mathcal{C}}$.
-With $f(\theta)\coloneqq A\theta$ over $\mathcal{C}$ for some matrix $A$ and $a^{*}\coloneqq A\theta^{*}$ for $\theta^{*}\in\mathcal{S}^{*}_{\mathcal{C}}$, we obtain $\mathcal{S}^{*}_{\mathcal{C}}=\{\theta\in\mathbb{R}^{q}\mid B\theta\leq b,A\theta\leq a^{*},-A\theta\leq-a^{*}\}$.
-By compactness of $\mathcal{C}$ and Hoffman’s lemma (Hoffman 1952), there exists $c>0$ depending on $\mathcal{C}$ such that for any $\theta\in\mathcal{C}$, there exists $\theta^{d}\in\mathcal{S}^{*}_{\mathcal{C}}=\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$ satisfying
+[…]
 $\left\lVert\theta-\theta^{d}\right\rVert=\mathrm{dist}(\theta,\mathcal{S}^{*}_{\mathcal{C}})\leq c\,\big{(}\left\lVert(B\theta-b)_{+}\right\rVert+\left\lVert A\theta-a^{*}\right\rVert\big{)}=c\left\lVert A(\theta-\theta^{d})\right\rVert.$
@@
 By the finiteness of $\mathfrak{C}$, we thus prove the desired claim.∎
-11 Analysis of PADR-based ERM for Constrained SP
+9 Constrained SP with Covariate Information
+In this section, we study problem (1) with deterministic constraints $\psi_{j}(z)\leq 0$ for $j\in[J]$.
+We first provide the formulation of the PADR-based ERM method.
+Then, we establish the asymptotic consistency of the constrained PADR-based ERM model in Section 9.1 and develop ESMM algorithm with exact penalization in Section 9.2.
+We present the numerical results on the Newsvendor problem with convex and nonconvex constraints in Section 9.3.
+[…]
+  1. (C1)
+    Finite scenarios: when the support set $\mathcal{X}$ of the feature vector $x$ contains a finite number of known scenarios, feasibility can be guaranteed by imposing constraints on the decision rule under all scenarios.
+  2. (C2)
+    Box constraints: when deterministic constraints lead to a box-constrained set $\{z\in\mathbb{R}^{d}:\left\lVert z\right\rVert_{\infty}\leq M\}$, a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ can be used, where ${\mathcal{H}}_{\mu}^{K}$ represents the PA hypothesis class, and the max/min are element-wise operators.
+This ensures feasibility of the ERM solution.
+Following the excess risk decomposition in Section 2.2, it is straightforward to see that the PADR-based ERM model for the three special cases (C1) and (C2) is nonasymptotically consistent with the original SP problem (9).
+[…]
+We next show that the PADR-based ERM model (9) is asymptotically consistent with (9).
 9.1 Asymptotic consistency of constrained PADR-based ERM
@@
 […]
-  1. (${}^{\prime}$)
-    There exists a continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (11.1).
-By Assumptions (A2) and (${}^{\prime}$), $f^{*}_{\gamma_{0}}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for some $L_{z}^{\prime}>0$ and $M_{z}^{\prime}>0$.
-We denote by ~${\widetilde{\mathcal{H}}}^{K}_{\mu,\gamma}\coloneqq{\mathcal{H}}_{\mu}^{K}\bigcap\left\{f\in\mathcal{F}\mid\psi_{j}(f(x^{s}))+\gamma\leq 0,\,s\in[n],\,j\in[J]\right\}$ and $f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ the feasible PADR class and solution to the constrained ERM problem (5) respectively.
-Then asymptotic consistency of the constrained PADR-based ERM model is presented below.
-Theorem 11.1 (Asymptotic Consistency of Constrained PADR-based ERM)
-Suppose that Assumptions (A1)-(A3) and (A5) hold, and that Assumption (${}^{\prime}$) holds for some $\gamma_{0}>\Gamma\coloneqq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}$.
-With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $S$,
-$R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})\leq$ $R(f^{*}_{\gamma_{0}})+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}^{\prime}\bar{X}K^{-1/p}$
+  1. (A4)
+    Given $\gamma_{0}>0$, there exists a Lipschitz continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (9.1).
+By Assumptions (A2) and (A4), $f^{*}_{\gamma_{0}}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for some $L_{z}^{\prime}>0$ and $M_{z}^{\prime}>0$.
+We first provide the excess risk bound of the restricted ERM model (9) to the restricted SP problem (9.1).
+Proposition 9.1 (Excess risk bound of the restricted constrained ERM)
+Suppose that Assumptions (A1)-(A3) and (A5) hold, and that Assumption (A4) holds for some $\gamma_{0}>\Gamma\coloneqq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}$.
+With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
+$R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})\leq$ $R(f^{*}_{\gamma_{0}})+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}^{\prime}\bar{X}K^{-1/p}$
 $+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
-Furthermore, with $\mu=\mathcal{O}(K^{1/p})$, $\gamma=\mathcal{O}(K^{-1/p})$, $\gamma_{0}=\mathcal{O}(K^{-1/p})$, and $K=\mathcal{O}(n^{p/(p+4)})$, suppose that there exists $\bar{\gamma}_{0}>0$ such that for any $\gamma\in[0,\bar{\gamma}_{0}]$, Assumption (${}^{\prime}$) holds and $f^{*}_{\gamma}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$.
-Then $R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})$ converges to the optimal cost of (5) in probability, as $n$ goes to infinity.
-Proof 11.2
+Proof 9.2
 […]
-$R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$ $=R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
-$\leq R(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{S}(f^{*}_{S,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{S}(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
-$\leq\underbrace{2\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{S}(f)\right\rvert}_{\mathcal{E}_{1}}+\underbrace{R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})}_{\mathcal{E}_{2}},$
+$R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$ $=R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{\Xi}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
+$\leq R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{\Xi}(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
+$\leq\underbrace{2\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert}_{\mathcal{E}_{1}}+\underbrace{R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})}_{\mathcal{E}_{2}},$
 where the last inequality holds since ~$\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}\subseteq{\mathcal{H}}_{\mu}^{K}$.
-The first term $\mathcal{E}_{1}$ can be directly bounded by Proposition 2.2, that is, given $\delta>0$, with probability at least $1-\delta$ over the draw of $S$,
+The first term $\mathcal{E}_{1}$ can be directly bounded by Proposition 2.2, that is, given $\delta>0$, with probability at least $1-\delta$ over the draw of $\Xi$,
 $\mathcal{E}_{1}\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
@@
 $\mathcal{E}_{22}\leq L_{\varphi}\sqrt{d}\left\lVert f_{app,\gamma_{0}}-f^{*}_{\gamma_{0}}\right\rVert_{\infty}\leq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}^{\prime}\bar{X}K^{-1/p}.$
-By combining bounds of $\mathcal{E}_{1}$, $\mathcal{E}_{21}$, and $\mathcal{E}_{22}$, we derive the desired inequality in the theorem, and asymptotic consistency follows instantly.∎
-11.2 Convergence of Enhanced SMM to B-stationarity
-According to (Cui and Pang 2021), for a constrained problem with a B-differentiable objective, a B-stationary point is defined as follows.
+By combining bounds of $\mathcal{E}_{1}$, $\mathcal{E}_{21}$, and $\mathcal{E}_{22}$, we derive the desired bound.∎
+With the excess risk bound in Proposition 9.1, and the probabilistic feasibility guarantee according to Luedtke and Ahmed (2008, Theorem 10), we obtain the asymptotic consistency of restricted PADR-based ERM model to the original constrained SP (9).
+Theorem 9.3 (Asymptotic consistency of constrained PADR-based ERM)
+Suppose that Assumptions (A1)-(A3) and (A5) hold, and that $f^{*}_{\gamma}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for any $\gamma\in[0,\bar{\gamma}]$ with some $\bar{\gamma}>0$.
+Then for any $n>0$, there exist $\delta_{n}>0$ and $\epsilon_{n}>0$ such that $f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ satisfies $\mathbb{P}_{X}\big{\{}\psi_{j}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}(X))\leq 0,\,j\in[J]\big{\}}\geq 1-\epsilon_{n}$ with probability $1-\delta_{n}$, where $\epsilon_{n}$ and $\delta_{n}$ converge to 0 as $n$ goes to infinity.
+Moreover, with $K=\mathcal{O}(n^{p/(p+4)})$, $\mu=\mathcal{O}(K^{1/p})$, and $\gamma=\mathcal{O}(K^{-1/p})$, $R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})$ converges to the optimal cost of (9) in probability, as $n$ goes to infinity.
+9.2 Convergence of ESMM to B-stationarity
+[…]
 Definition 9.4
 We say that $\theta^{b}$ is a B-stationary point of the objective $F(\cdot)$ on a closed set $\bar{\Theta}\subseteq\Theta$ if $\theta^{b}\in\bar{\Theta}$ and $F^{\prime}(\theta^{b};v)\geq 0$ for any $v\in\mathcal{T}(\theta^{b};\bar{\Theta})$.
+where $\mathcal{T}(\theta^{b};\bar{\Theta})$ is the tangent cone of $\bar{\Theta}$ at $\theta^{b}$.
 The B-stationarity coincides with d-stationarity if $\bar{\Theta}$ is convex.
-Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (5) are B-differentiable.
-With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (5) by $\mathcal{S}^{b}$.
-With Assumptions 1-1 imposed in the context of the penalized problem (5), we can apply the enhanced SMM algorithm to solve (5) and obtain the same convergence results in terms of d-stationarity as in Theorems 4.4 and 4.5.
-To further obtain the convergence to B-stationarity, we intend to connect d-stationarity of (5) with B-stationarity of (5) by applying the following exact penalty result from Cui and Pang (2021).
-Note that by the finiteness of $S$ and compactness of $\Theta$, for any $\xi\in S$, $f(\cdot;\xi)$ is Lipschitz continuous with modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$ and uniformly bounded with $\left\lVert f(\cdot;\xi)\right\rVert_{\infty}\leq M_{f}\coloneqq L_{f}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, and $\varphi(\cdot;\xi)$ is Lipschitz continuous on $[-M_{f},M_{f}]^{d}$ with uniform modulus denoted by $L^{\prime}_{\varphi}$.
-Lemma 11.4 (Cui and Pang (2021), Proposition 9.2.2)
+Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (9.2) are B-differentiable.
+With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (9.2) by $\mathcal{S}^{b}$.
+[…]
+[…]
 Let $\Theta$ be a compact convex set. For $\gamma>0$, if there exists $l_{c}>0$ such that
 $\sup_{\theta\in\Theta\setminus\bar{\Theta}}\left[\min_{v\in\mathcal{T}(\theta;\Theta);\left\lVert v\right\rVert\leq 1}G_{\gamma}^{\prime}(\theta;v)\right]\leq-l_{c},$ (32)
-then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (5) is feasible to (5) and thus is a B-stationary point of (5).
-Condition (24) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (5) are bounded away from 0.
-Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (5) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (5).
+then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (9.2) is feasible to (9.2) and thus is a B-stationary point of (9.2).
+Condition (32) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (9.2) are bounded away from 0.
+Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (9.2) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (9.2).
 Nonetheless, it is not easy to verify the technical condition (32) when $G_{\gamma}$ is in a composite structure.
@@
 Then condition (32) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$.
-Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (5) is a B-stationary point of problem (5).
-Proof 11.6
+Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (9.2) is a B-stationary point of problem (9.2).
+Proof 9.7
 Proof of Proposition 9.6.
 By Lemma 9.5, it suffices to show that condition (32) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$.
-For any $\xi\in S$ and $j\in[J]$, denote $G_{\gamma}(\theta;\xi,j)\coloneqq\max\left\{\psi_{j}(f(\theta;\xi))+\gamma,0\right\}$.
-For a point $\bar{\theta}\in\Theta$ infeasible to the constraint $G_{\gamma}(\theta)\coloneqq\mathbb{E}\left[\sum_{j=1}^{J}G_{\gamma}(\theta;\widetilde{\xi}_{n},j)\right]\leq 0$, there exist $\bar{j}\in[J]$ and $\bar{\xi}\in S$ such that $\psi_{\bar{j}}(f(\bar{\theta};\bar{\xi}))+\gamma>0$.
-For any $\xi\in S$, we denote $\bar{z}=f(\bar{\theta};\xi)$ and $\bar{v}=-\frac{\bar{\theta}}{\left\lVert\bar{\theta}\right\rVert}$ for simplicity.
-Since $f(\cdot;\xi)$ is piecewise linear, there exists $A\in\mathbb{R}^{d\times q}$ dependent on $\xi$ and $\bar{\theta}$ such that $f(\bar{\theta};\xi)=A\bar{\theta}$ and $f(\cdot;\xi)^{\prime}(\bar{\theta};\bar{v})=A\bar{v}$.
+For any $\xi\in\Xi$ and $j\in[J]$, denote $G_{\gamma}(\theta,\xi;j)\coloneqq\max\left\{\psi_{j}(f(\theta,\xi))+\gamma,0\right\}$.
+For a point $\bar{\theta}\in\Theta$ infeasible to the constraint $G_{\gamma}(\theta)\coloneqq\mathbb{E}\left[\sum_{j=1}^{J}G_{\gamma}(\theta,\widetilde{\xi}_{n};j)\right]\leq 0$, there exist $\bar{j}\in[J]$ and $\bar{\xi}\in\Xi$ such that $\psi_{\bar{j}}(f(\bar{\theta},\bar{\xi}))+\gamma>0$.
+For any $\xi\in\Xi$, we denote $\bar{z}=f(\bar{\theta},\xi)$ and $\bar{v}=-\frac{\bar{\theta}}{\left\lVert\bar{\theta}\right\rVert}$ for simplicity.
+Since $f(\cdot,\xi)$ is piecewise linear, there exists $A\in\mathbb{R}^{d\times q}$ dependent on $\xi$ and $\bar{\theta}$ such that $f(\bar{\theta},\xi)=A\bar{\theta}$ and $f(\cdot,\xi)^{\prime}(\bar{\theta};\bar{v})=A\bar{v}$.
 Let
 $\mathcal{A}_{<}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma<0\},\ \mathcal{A}_{>}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma>0\},\text{ and }\mathcal{A}_{=}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma=0\}.$
-For $j\in\mathcal{A}_{<}$, it is clear that $G_{\gamma}(\cdot;\xi,j)^{\prime}(\bar{\theta};\bar{v})=0$. By Assumption (B1), we have that for $j\in\mathcal{A}_{>}$,
-$G_{\gamma}(\cdot;\xi,j)^{\prime}(\bar{\theta};\bar{v})=\psi_{j}^{\prime}(f(\bar{\theta};\xi);f(\cdot;\xi)^{\prime}(\bar{\theta};\bar{v}))=\psi_{j}^{\prime}\left(\bar{z};-\frac{\bar{z}}{\left\lVert\bar{\theta}\right\rVert}\right)\leq-\frac{c_{\psi}l_{\psi}}{\left\lVert\bar{\theta}\right\rVert}\leq-\frac{c_{\psi}l_{\psi}}{\sqrt{q}\mu},$
+For $j\in\mathcal{A}_{<}$, it is clear that $G_{\gamma}(\cdot,\xi;j)^{\prime}(\bar{\theta};\bar{v})=0$. By Assumption (B7), we have that for $j\in\mathcal{A}_{>}$,
+$G_{\gamma}(\cdot,\xi;j)^{\prime}(\bar{\theta};\bar{v})=\psi_{j}^{\prime}(f(\bar{\theta},\xi);f(\cdot,\xi)^{\prime}(\bar{\theta};\bar{v}))=\psi_{j}^{\prime}\left(\bar{z};-\frac{\bar{z}}{\left\lVert\bar{\theta}\right\rVert}\right)\leq-\frac{c_{\psi}l_{\psi}}{\left\lVert\bar{\theta}\right\rVert}\leq-\frac{c_{\psi}l_{\psi}}{\sqrt{q}\mu},$
 and that for $j\in\mathcal{A}_{=}$,
@@
 Since $\bar{j}\in A_{>}$, we obtain that
-$\min_{v\in\mathcal{T}(\bar{\theta};\Theta);\left\lVert v\right\rVert\leq 1}G_{\gamma}^{\prime}(\bar{\theta};v)\leq G_{\gamma}^{\prime}(\bar{\theta};\bar{v})\leq\frac{1}{n}G_{\gamma}(\cdot;\bar{\xi},\bar{j})^{\prime}(\bar{\theta};\bar{v})\leq-\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu},$
+$\min_{v\in\mathcal{T}(\bar{\theta};\Theta);\left\lVert v\right\rVert\leq 1}G_{\gamma}^{\prime}(\bar{\theta};v)\leq G_{\gamma}^{\prime}(\bar{\theta};\bar{v})\leq\frac{1}{n}G_{\gamma}(\cdot,\bar{\xi};\bar{j})^{\prime}(\bar{\theta};\bar{v})\leq-\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu},$
 which completes the proof.∎
-Finally, from Proposition 11.5 and convergence results in Section 4, we summarize the convergence properties of the enhanced SMM with exact penalization in the theorem below, where Assumptions 1 and 2 are adjusted to surrogate families of $F(\cdot;\xi)$ and $\psi_{j}(f(\cdot;\xi))$ for any $\xi\in S$ and $j\in[J]$, and Assumption 1 is adjusted to $V(\cdot;\lambda)$.
-The proof is similar to those of Theorems 4.4 and 4.5 and thus is omitted here.
-Theorem 11.7 (Convergence of Enhanced SMM with exact penalization)
-Under Assumption 1 and the same settings of Proposition 11.5, for problem (5) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$.
+Finally, from Proposition 9.6 and convergence results in Sections 3.2.3 and 8.4, we summarize the convergence properties of the ESMM with exact penalization in the theorem below, where Assumptions (B1) to (B5) are adjusted to surrogate families of $F(\cdot,\xi)$ and $\psi_{j}(f(\cdot,\xi))$ for any $\xi\in\Xi$ and $j\in[J]$, and Assumption (B6) is adjusted to $V(\cdot;\lambda)$.
+The proof is similar to those of Theorems 3.6-8.12, and thus is omitted here.
+Theorem 9.8 (Convergence of ESMM with exact penalization)
+Under Assumptions (B1) to (B4) and the same settings of Proposition 9.6, for problem (9.2) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$.
 The following statements hold:
   1. (a)
-    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (5) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (5).
+    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (9.2) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (9.2).
   2. (b)
-    With $\varepsilon>0$, under Assumptions 2 and 1, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\widetilde{\mathcal{O}}(T^{-\frac{1}{2}\min\{\frac{\alpha}{2},1\}})$.
+    With $\varepsilon>0$, under Assumptions (B5) and (B6), $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\widetilde{\mathcal{O}}(T^{-\frac{1}{2}\min\{\frac{\alpha}{2},1\}})$.
 In particular, if $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ are piecewise affine and $\Theta$ is a polytope, the same convergence rate holds without the two assumptions.
@@
     With $\varepsilon=0$, functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ being piecewise affine, and $\Theta$ being a polytope, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\widetilde{\mathcal{O}}(T^{-\frac{1}{2}\min\{\frac{\alpha}{2},1\}})$; moreover, with $\eta>0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\mathcal{O}(T^{-\frac{1}{2}\min\{\alpha,1\}})$, which is up to a logarithmic factor when $\alpha=1$.
-12 Additional Experimental Results
-For the unconstrained Newsvendor problem in Section 6.1, we set $(c_{b},c_{h})=(5,5)$ and consider another demand model
-$\overline{Y}^{\prime}(X)=4\sin(\pi X_{1})+\max\{16X_{2},-20X_{2}\}+10,$
-where the sine function can be interpreted as the seasonal dependence of the demand on $X_{1}$, and the max-affine function represents another nonlinear trend with respect to $X_{2}$.
-In practice, we may distinguish seasonal features from the others and build the improved DR represented as the sum of multiple DRs that are constructed for separate groups of features.
-Thus, we employ the sum of two PADR(2,2) models and the sum of two RKHS-DRs as the improved DRs in this instance and denote them by PADR(2,2)+ and RKHS-DR+, respectively.
-a.
-Basic Results.
-In Figure 1p, we report results with the training sample size $n$ increasing from 50 to 1000.
-All PADR and RKHS-DR methods except PADR(2,2) converge close to SIMOPT.
-Figure 1q shows the results of running PADR(4,4) and PADR(2,2)+ for 30 rounds each, where we report the mean cost of the top 15 with standard deviation for each trial.
-It can be seen that although PADR(4,4) has a stronger approximation ability, PADR(2,2)+, designed with prior information, obtains lower and stabler costs.
-(p) Comparison with Benchmarks
-(q) PADR(4,4) vs. PADR(2,2)+
-Unconstrained Newsvendor Problem with $\overline{Y}^{\prime}$: Basic Results
-b.
-Varying Feature Dimensions.
-We set sparse and dense models in the same way as in Section 6.1 and report the comparison results with RKHS-DR and RKHS-DR+ in Figure 1s.
-PADR(2,2)+ outperforms all the others in both cases.
-It is worth mentioning that PADR(2,2) is close to PADR(2,2)+ in high-dimensional cases, showing that the training difficulty gradually dominates compared to the value of prior information as the dimension grows.
-But RKHS-DR+ consistently performs better than RKHS-DR, and it performs better than PADR(2,2) when $p$ is small.
-As for the running time, PADR(4,4) and PADR(2,2)+ grow linearly with $p$ and are longer than RKHS-DR and RKHS-DR+ only when $p\geq 30$.
-(r) Sparse Demand Model
-(s) Dense Demand Model
-Unconstrained Newsvendor Problem with $\overline{Y}^{\prime}$ and Varying Feature Dimensions
+For the ESMM algorithm with exact penalization, the limit point of generated sequence is a composite strong d-stationarity to (9.2) with probability 1 according to Theorem 9.8, and thus is a global minimizer of (9.2) when $\varepsilon$ is sufficiently large and $\eta$ is close to 0, by the equivalence between composite strong d-stationarity and global optimality in Proposition 8.2 with appropriate assumptions when needed.
+Furthermore, such a global minimizer of the penalized problem (9.2) should be also a global minimizer of (9.2) since it is feasible according to Proposition 9.6.
+Under such settings, the asymptotic consistency of the PADR framework for the constrained case in Theorem 9.3 turns into asymptotic optimality.
+9.3 Experiments on constrained newsvendor problem
+We compare PADR and other benchmarks on two-product constrained Newsvendor problems, assuming that the uncensored historical demand data is available.
+We denote the unit back-order cost, unit holding cost, and order decision by $c_{bi}$, $c_{hi}$, and $z_{i}$, respectively, for product $i=1,2$.
+The demand models for the two products are set as follows:
+$\overline{Y}_{1}(X)=15X_{1}-5X_{2}+30,\quad\overline{Y}_{2}(X)=15X_{1}+5X_{2}+30.$
+In this subsection, we compare PADR with prescriptive methods and the penalized RKHS-DR method from Bertsimas and Koduri (2021, Section 6.2).
+9.3.1 Newsvendor problem with a linear constraint.
+We first consider an inventory capacity constraint for the two products, which is formulated as $z_{1}+z_{2}\leq C_{0}$.
+In this subsection, we set $(c_{b1},c_{h1},c_{b2},c_{h2})=(8,2,2,8)$, $C_{0}=60$, such that the probability that the constraint is active at the optimal decision is around 0.5.
+Since DR-based methods may yield infeasible decisions, in Figure 1u we report the test cost of the order decisions, projected onto the convex feasible region when necessary, along with the feasible frequency of the original decisions provided by PADR and RKHS-DR.
+[…]
+(s) Basic results
+(t) Sparse demand model
+(u) Dense demand model
+Newsvendor problem with a linear constraint
+With the basic demand model, PADR converges to SIMOPT and attains 100% feasibility when $n\geq 200$.
+In comparison, RKHS-DR and prescriptive methods display notable gaps to SIMOPT, with the feasible frequency of RKHS-DR being only around 80%.
+Figures 1t and 1u show the numerical results with varying feature dimensions.
+Similar to the unconstrained cases, PADR outperforms the other methods in the sparse case when $p<20$ and in the dense case when $p<50$.
+Notably, with penalization, PADR achieve 100% feasibility when $p<20$.
+Although PADR can not guarantee 100% feasibility of the decision compared to prescriptive methods, it still provides satisfactory performance via projection.
+9.3.2 Newsvendor problem with a nonconvex constraint.
+We next consider a capacity constraint $C(z_{1})+C(z_{2})\leq C_{0}$ with the concave capacity cost function (18) for the two products.
+We set $(c_{b1},c_{h1},c_{b2},c_{h2})=(7,7,3,3)$ and $C_{0}=50$ with the same probability for the constraint being active as the experiment in Section 9.3.1.
+Since the projection of infeasible decisions to the nonconvex feasible region may be far away to the optimal solutions, we only report the average test cost of feasible decisions in Figure 1x, with a careful hyperparameter selection to ensure high feasible frequencies.
+We use the standard MM algorithm to solve the penalized ERM problem of RKHS-DR, referred to here as RKHS-DR(MM).
+For prescriptive methods and SIMOPT, we solve the two-dimensional nonconvex-constrained problems using grid search.
+The sof method in Kallus and Mao (2023)is not suitable for this nonconvex-constrained problem, and it may require the study of the second-order perturbation analysis for stochastic optimization, which is beyond the scope of this paper.
+For the problem with the basic demand model, both PADR and RKHS-DR attain 100% feasibility on the test set.
+PADR converges to SIMOPT with significantly fewer training samples than other benchmarks.
+Figures 1w and 1x show the results across varying feature dimensions.
+PADR outperforms the other methods in both sparse and dense cases when $p\leq 30$, achieving 90 % $\sim$ 100% feasibility in the sparse case when $p\leq 20$ and in the dense case when $p\leq 30$.
+Under the dense demand model, RKHS-DR performs well in low-dimensional cases, but its performance deteriorates as $p$ increases.
+Prescriptive methods—kNN (in the dense case), cart, and rf—maintain a stable yet noticeable gap to SIMOPT, which is possibly due to their limitations to approximate conditional distributions with the sample set that is not sufficiently large.
+The relatively small gap of PADR to SIMOPT shows PADR’s capability to learn the underlying decision rule structure in problems with nonconvex constraints while maintaining a high probability of feasibility when $p\leq 30$.
+[…]
+(v) Basic results
+(w) Sparse demand model
+(x) Dense demand model
+Newsvendor problem with a nonconvex constraint
+[…]
+[…]

### v5

@@
 Zhang, Liu, and Zhao
-[…]
 […]
@@
 To solve the nonconvex and nondifferentiable ERM problem, we develop an enhanced stochastic majorization-minimization (ESMM) algorithm and establish the convergence to (composite strong) directional stationarity along with convergence rate analysis.
-We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational traceability.
+We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational tractability.
 Our numerical study demonstrates the effectiveness of PADR-based ERM methods with comparisons to state-of-the-art approaches under various settings, with significantly lower costs, less computation time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
@@
 With the fast-growing real-world applications and availability of large datasets, the field of stochastic programming (SP) has flourished with many theoretical and algorithmic advances (Birge and Louveaux 2011, Shapiro et al.
-2021), in which the empirical probability distribution based on historical scenarios plays a central role in approximating the true distribution.
+2021a), in which the empirical probability distribution based on historical scenarios plays a central role in approximating the true distribution.
 However, it is common in practice that some covariate information (also referred to as contextual information or features) that may affect the random variable could be observed before the decision is made.
-In such situations, the empirical probability distribution and classical SP methodologies ignoring the covariate information leads to the poor performance.
+In such situations, the empirical probability distribution and classical SP methodologies ignoring the covariate information lead to the poor performance.
 Given the covariate information $X=x\in\mathbb{R}^{p}$, an unconstrained SP is formulated as follows,
@@
 By the interchangeability principle (Shapiro et al.
-2021, Theorem 7.80), if a mapping $f^{*}\in\mathcal{F}$ is optimal to problem (2), then the decision $f^{*}(x)$ is optimal to (1) for almost every $x\in\mathcal{X}$.
+2021a, Theorem 7.80), if a mapping $f^{*}\in\mathcal{F}$ is optimal to problem (2), then the decision $f^{*}(x)$ is optimal to (1) for almost every $x\in\mathcal{X}$.
 In general, problem (2) is intractable due to the infinite-dimensional function space and the unknown joint distribution of $X$ and $Y$.
@@
 $f(x,\theta)=\max_{k\in[K]}\left\{(\alpha^{k})^{\top}x+a^{k}\right\}-\max_{k\in[K]}\left\{(\beta^{k})^{\top}x+b^{k}\right\},$ (4)
-where $\theta\coloneqq\left(\alpha^{k},a^{k},\beta^{k},b^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)}$ and $[K]\coloneqq\{1,\dots,K\}$.
-The PADR of such a form is beneficial because it requires minimal parameter tuning in multi-dimensional cases without the need to specify breakpoints for piecewise linear parts in each dimension, and it can asymptotically approximate any continuous functions according to Royset (2020).
+where $\theta\coloneqq\left(\alpha^{k},a^{k},\beta^{k},b^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)}$ and $K]\coloneqq\{1,\dots,K\}$.
+The PADR of such a form is beneficial because it requires minimal parameter tuning in multi-dimensional cases without the need to specify breakpoints for piecewise linear parts in each dimension, and it can asymptotically approximate any continuous function according to Royset (2020).
 Throughout the theoretical analysis in this paper, we assume that the numbers of pieces in the two max-affine components of (4) are equal to simplify the exposition, which could be easily extended to the case with different numbers of pieces.
@@
 In Appendix 9, we illustrate that the obtained decision rule at a new scenario is feasible with high probability and is asymptotically consistent with the original SP problem, and we also discuss several special cases where the feasibility of the decision rule can be guaranteed.
-In the rest of this section, we give a brief literature review and present the distinguished features of the proposed PADR method over other relative work.
-We will also summarize major contributions of the present paper.
+In the rest of this section, we give a brief literature review and present the distinguished features of the proposed PADR method over other related work.
+We will also summarize the major contributions of the present paper.
 1.2 Literature review
@@
 (2023) point out that, when the model class is well-specified and there is sufficient data, methods that directly estimate the conditional distribution outperform those that consider the downstream optimization problem; in a misspecified scenario, the reverse holds.
-However, as claimed by Bertsimas and Koduri (2021), the prescriptive methods are highly affected by the data with features close to the current observation and thus suffers from the curse of dimensionality.
+However, as claimed by Bertsimas and Koduri (2021), the prescriptive methods are highly affected by the data with features close to the current observation and thus suffer from the curse of dimensionality.
 The DR-based ERM framework presented in (3), which is another one-step approach in the integrated scheme, is relieved from the aforementioned issue due to its functional approximation in nature.
-For constrained SP with general nonlinear constraints, the current study of decision rule methods could only guarantee feasibility with high probability, and the predict-then-optimize framework may be more suitable for solving the contextual problems when the feasibility is essential to decision-makers.
+For constrained SP with general nonlinear constraints, the current study of decision rule methods could only guarantee feasibility with high probability, and the predict-then-optimize framework may be more suitable for solving the contextual problems when feasibility is essential to decision-makers.
+We refer interested readers to Qi and Shen (2022) and Sadana et al.
+(2025) for comprehensive reviews on recent advances in both two-stage and integrated approaches in contextual SP.
 Several works on DR-based ERM methods are most relevant to this paper.
 Ban and Rudin (2018) employ LDR for the Newsvendor problem with nonasymptotic optimality guarantees under the assumption of the linear demand model.
-Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems.
-Compared to the above existing DR methods, given the asymptotic approximation of PA functions to any continuous functions, the PADR method is applicable to more general decision-making problems with nonconvex cost function or nonlinear dependency with asymptotic optimality.
 Besides the classical LDR, several generalized LDRs are proposed for standard SP problems which lift random vectors into a higher-dimensional probability space to describe nonlinear relations (Chen et al.
@@
 2015).
-However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, they usually lack theoretical guarantees on approximation accuracy except for the posterior estimation provided in Kuhn et al.
+However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, they usually lack theoretical guarantees on approximation accuracy except for the posterior estimation in Kuhn et al.
 (2011).
+Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems.
+With the rise of deep neural networks (NN), several studies apply NN-based decision rules (NN-DR) in practical applications, such as inventory management (Zhang et al.
+2020, Oroojlooyjadid et al.
+2020, Qi et al.
+2023) and portfolio optimization (Zhang et al.
+2020).
+NN-DRs are powerful in approximating the underlying relationship between decisions and covariate information; however, they may face numerical challenges for complex decision-making problems and also lack theoretical consistency guarantees.
+Compared to the above existing DR methods, given the asymptotic approximation of PA functions to any continuous function, we show that the PADR method is applicable to a class of more general decision-making problems with nonconvex cost function with both numerical efficiency and theoretical consistency.
 Regarding solving the PADR-based ERM problem with potentially large sample size $n$, it is prohibitively challenging to obtain a global minimizer to such a coupled nonconvex and nondifferentiable optimization problem.
 As provided by Bertsimas and Georghiou (2015, Section 2.4), the PADR-based ERM problem can be formulated as a mixed-integer nonlinear program using Big-M method with additional $2ndK$ binary variables to describe the PADR of interest.
-For a PADR-based ERM problem of the modest size with data size $n=1000$, the obtained mixed-integer nonlinear program could have $10^{4}\sim 10^{5}$ binary variables and the state-of-the-art solvers are hard to solve within a few hours to achieve an approximately optimal solution.
+For a PADR-based ERM problem of the modest size with data size $n=1000$, the obtained mixed-integer nonlinear program could have $10^{4}\sim 10^{5}$ binary variables and state-of-the-art solvers find it hard to solve within a few hours to achieve an approximately optimal solution.
 Instead of seeking global optimality, majorization-minimization (MM) methods provide a unified framework that is practically efficient and theoretically convergent for solving the PADR-based ERM problem with the finite-sum composite objective.
@@
 According to variational analysis by Rockafellar and Wets (2009), there are many kinds of stationary concepts which differ in the ability to characterize necessary optimality conditions for local/global optimality.
-As pointed in Cui and Pang (2021, Chapter 7), MM algorithms may converge to different types of stationary solutions depending on the properties of surrogate and objective functions.
-Algorithms with convergence to a stronger stationarity is favorable not only theoretically but also computationally because the obtained solution from numerical implementation with random initialization has the greater potential to be a global optimum.
+As pointed out in Cui and Pang (2021a, Chapter 7), MM algorithms may converge to different types of stationary solutions depending on the properties of surrogate and objective functions.
+Algorithms with convergence to a stronger stationarity are favorable not only theoretically but also computationally because the obtained solution from numerical implementation with random initialization has the greater potential to be a global optimum.
 Difference-of-convex (DC) algorithms, which is a special MM method, have been extensively studied over the last several decades (Le Thi and Pham Dinh 2018) and can compute a d(irectional)-stationary point for DC programs with a differentiable concave component.
@@
 (2016) and Cui et al.
-(2018) propose an enhanced surrogation technique that iteratively solves convex subproblems that are constructed based on $\varepsilon$-active pieces of the concave component.
+(2018a) propose an enhanced surrogation technique that iteratively solves convex subproblems that are constructed based on $\varepsilon$-active pieces of the concave component.
 However, the enhanced MM algorithm for large-scale ERM problems requires iteratively solving finite-sum convex subproblems over all the data, which raises the fundamental issue of computational efficiency.
@@
 (2022) addressing smooth or partially smooth objectives, and in Liu et al.
-(2022) addressing nonconvex and nonsmooth compound SP to compute critical solutions.
-There still lacks a comprehensive study on whether the $\varepsilon$-active set technique is amenable for sampling-based methods to enhance the convergence properties for nondifferentiable stochastic DC programs and PADR-based ERM problems.
+(2022a) addressing nonconvex and nonsmooth compound SP to compute critical solutions.
+There is still a lack of a comprehensive study on whether the $\varepsilon$-active set technique is amenable for sampling-based methods to enhance the convergence properties for nondifferentiable stochastic DC programs and PADR-based ERM problems.
 1.3 Contributions
-We summarize our major contributions as follows.
+Our work contributes to the literature on contextual SP along three major dimensions:
 1.
+Modeling aspect.
 We propose a PADR-based ERM method for a general class of SP problems with covariate information, where the cost function could be nonconvex and nondifferentiable.
-We quantify the approximation accuracy of the PA function class to any Lipschitz continuous function, which to the best of our knowledge is the first quantification result for such a general case.
-Without any parametric assumption on the optimal decision rule or the latent dependency, we establish the nonasymptotic consistency guarantee with the excess risk bound of our proposed framework, which illustrates the role of piece number in balancing the approximation and estimation errors.
+PADR provides a compact structure to approximate non-parametric decision rules.
+Though PA hypothesis class is not new in classical SP, we provide the first quantification of the approximation error and the excess risk bound of the data-driven PADR model for general nonconvex contextual SP problems.
+The excess risk bound illustrates the role of the piece numbers $K_{1}$ and $K_{2}$ in balancing the approximation and estimation errors, which is further evidenced by numerical experiments.
+In contrast, previous DR methods for contextual SP provide asymptotic guarantees only limited to convex or linear settings.
+The PADR method thereby provides a rigorous framework to solve nonconvex contextual SP problems with theoretical guarantees.
 2.
-For efficiently solving the PADR-based ERM problem, we develop an enhanced sampling-based MM (ESMM) algorithm with the natural combination of sequential sampling and the enhanced technique over the $\varepsilon$-active set.
-With sequential sampling, the ESMM algorithm effectively reduces computational costs by solving subproblems of small sizes in the early iterations.
-We show that under the proper control of increasing sample rate, ESMM achieves the asymptotic convergence in terms of composite strong d-stationarity, which is strictly stronger than d-stationarity and thus has greater potential to be a global minimizer.
-Numerical results show that the ESMM algorithm could speed up to 4 times faster than the basic MM method.
-We also would like to highlight several distinguishing features of the proposed PADR method for contextual SP: (a) from the modeling aspect, the PADR method is suitable for contextual SP problems with nonlinear or nonconvex cost functions.
-All other relevant works on contextual SP with asymptotic optimality are restricted to linear programs or convex programs.
-For instance, the RKHS-based DR method by Bertsimas and Koduri (2021) requires convexity of the cost function for asymptotic optimality.
-The PADR method opens up the application to nonconvex contextual SP problems; (b) from the theoretical aspect, PADR provides a compact structure to asymptotically approximate non-parametric decision rule with minimal hyperparameter tuning.
-Though PA hypothesis class is not new in classical SP, we provide the first quantification of the approximation error and the excess risk bound of the data-driven PADR model for general nonconvex SP problems.
-Except for the LDR method, all other methods for contextual SP do not provide the excess risk bound, including the RKHS-based DR in Bertsimas and Koduri (2021) and Notz and Pibernik (2021); (c) regarding to the computational aspect, the proposed ESMM algorithm is developed for efficiently solving the PADR-based ERM problem of the large scale.
-Under the suitable sampling scheme, we show that the ESMM method significantly improves computational efficiency and enhances the convergence property compared to the existing MM methods in the literature; (d) moreover, the PADR method possesses several prominent numerical advantages.
-Numerous experiments indicate that it requires significantly fewer data points for satisfactory performance, thus is particularly suitable for cases with insufficient data sets.
-Additionally, it yields superior performance in terms of the test costs with both synthetic and real data sets when the feature dimension is less than 20.
-The nonlinearity level of the demand model has almost no effect on the performance of the PADR method, while other methods suffer significantly from the high level of nonlinearity.
+Computational aspect.
+By exploiting the compact max-affine structure of PADR, we develop the ESMM algorithm to solve the highly nonconvex and nonsmooth PADR-based ERM problem, with the natural combination of sequential sampling and the enhanced technique over the $\varepsilon$-active set within the majorization-minimization framework.
+We analyze the asymptotic convergence of ESMM in terms of composite strong d-stationarity which is strictly stronger than d-stationarity, the sharpest type of first-order stationary solutions for general nonconvex optimization.
+Furthermore, with the design of a suitable sampling scheme, the ESMM method significantly improves computational efficiency compared to existing MM approaches.
+Numerical results show that the ESMM algorithm for PADR scales well to large problem sizes (e.g., up to 80 pieces in PADR) and could be up to 4 times faster than the basic MM method.
+3.
+Empirical aspect.
+Numerical experiments with comparisons to a variety of prominent methods including NN-DR and RKHS-DR (see Table 4) show that PADR requires fewer data points to achieve satisfactory performance, and thus it is suitable for cases with insufficient data sets.
+Additionally, it yields superior performance in terms of the test costs with both synthetic and real data sets in various instances including the Newsvendor problem and the product placement problem with feature dimension less than 20.
+Finally, PADR is more robust to the level of nonlinearity in demand models and consistently obtains satisfactory performance with highly nonlinear demand models, compared to all other benchmarks.
 The rest of this paper is organized as follows.
@@
 (2020) respectively.
-PA functions which are of “nonparametric” nature are known to asymptotically approximate Lipschitz continuous functions (Royset 2020); nevertheless, the approximation error of PA functions with finite pieces remains unsolved.
+PA functions that are of “nonparametric” nature are known to asymptotically approximate Lipschitz continuous functions (Royset 2020); nevertheless, the approximation error of PA functions with finite pieces remains unsolved.
 For the class of Lipschitz continuous functions that allow a more general setting for SP problems of interest, we provide the approximation error bound of the PA function class, which to the best of our knowledge is the first quantification result for such a general case.
 With $L>0$ and $M>0$, let
-[…]
+$\mathcal{F}_{L,M}\coloneqq\left\{f:\mathcal{X}\rightarrow\mathbb{R}^{d}\mid\left\lVert f\right\rVert_{\infty}\leq M,\,\text{$f$ is $L$-Lipschitz continuous}\right\},$
 where $\left\lVert f\right\rVert_{\infty}\coloneqq\sup_{x\in\mathcal{X}}\max_{\iota\in[d]}|f_{\iota}(x)|$. We define the PADR class ${\mathcal{H}}_{\mu}^{K}$ as
@@
 The boundedness of $\mathcal{X}$ in Assumption (A2) is common in practical applications.
+Assumption (A4) may appear to be strict, however, in Online Appendix 10, we show that it applies to a broad class of problems.
+We also provide examples of problems that violate this assumption to further clarify its scope of applicability.
 Under Assumption (A2), the optimal DR $f^{*}$ in Assumption (A4) is bounded on $\mathcal{X}$, and thus $f^{*}\in\mathcal{F}_{L_{z},M_{z}}$ for some $L_{z}>0$ and $M_{z}>0$.
@@
 $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)}_{\text{estimation error}}+\underbrace{\min_{f\in{\mathcal{H}}_{\mu}^{K}}R(f)-\inf_{f\in\mathcal{F}}R(f)}_{\text{approximation error}}.$ (7)
-[…]
+The approximation error, measuring the gap between the minimal expected costs over the PADR class and the class of all measurable functions, can thus be bounded based on the approximation error bound of PADR class provided in Proposition 2.1.
+We provide the uniform generalization bound based on the Rademacher complexity theory in Proposition 2.2 which is used to bound the estimation error.
+Combining the approximation and estimation error bounds, we thus obtain the excess risk bound of PADR-based ERM in Theorem 2.3.
+The proof of the following two results are provided in Online Appendix 6.2.
+Note that in the present paper the notation $\left\lVert\cdot\right\rVert$ without a subscript refers to the Euclidean norm.
 Proposition 2.2 (Uniform generalization bound of PADR-based ERM)
-Under Assumptions (A1)-(A3), with $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
-$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq\frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
+Under Assumptions (A1)-(A3), with $M_{p,d}(\mu)\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
+$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq$ $\ \frac{8\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+2M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}$
+$\ +\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
 Theorem 2.3 (Excess risk bound of PADR-based ERM)
-Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, with probability at least $1-\delta$ over the draw of $\Xi$, we have
+Under Assumptions (A1)-(A4), with $\mu\geq\max\left\{\frac{1}{2}L_{z}K^{1/p},\frac{1}{4}pL_{z}\bar{X}K^{1/p}+\frac{1}{2}M_{z}\right\}$ and $M_{p,d}(\mu)\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, with probability at least $1-\delta$ over the draw of $\Xi$, we have
 $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})\,\leq\,$ $\inf_{f\in\mathcal{F}}R(f)+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}\bar{X}K^{-1/p}$
-$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
+$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
 The above theorem implies that with $\mu=\mathcal{O}(K^{1/p})$, the complexity of the excess risk bound with respect to the sample size $n$ and parameter $K$ is
 $R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-\inf_{f\in\mathcal{F}}R(f)=\underbrace{\widetilde{\mathcal{O}}\left(K^{\frac{p+2}{2p}}n^{-\frac{1}{2}}\right)}_{\begin{subarray}{c}\text{estimation}\\ \text{error bound}\end{subarray}}+\underbrace{\mathcal{O}\left(K^{-\frac{1}{p}}\right)}_{\begin{subarray}{c}\text{approximation}\\ \text{error bound}\end{subarray}},$ (8)
-where $\widetilde{\mathcal{O}}(\cdot)$ suppresses logarithmic factors.
-It can be seen that the sample size $n$ only affects the estimation error, whereas $K$, the number of max-affine pieces for PA functions in ${\mathcal{H}}_{\mu}^{K}$, plays the central role of balancing the two error terms.
-As $K$ grows, the PA function class becomes larger, so the approximation error decreases due to the stronger approximation ability of ${\mathcal{H}}_{\mu}^{K}$, and the estimation error increases due to the more pronounced overfitting issue.
-With an appropriate choice of $K$ that balances these two errors, we can derive the optimal excess risk bound.
-Furthermore, with the existence of a convex or DC optimal DR, this bound can be improved based on the approximation bounds $\mathcal{O}(K^{-2/p})$ of ${\mathcal{H}}_{\mu}^{K}$ from Balazs et al.
-(2015) and Siahkamari et al.
-(2020).
-The extensions are summarized in the following corollary.
+[…]
 Corollary 2.4
@@
 Noticing that $\inf_{f\in\mathcal{F}}R(f)$ is irrelevant to the PA hypothesis class, the performance of excess risk could be revealed through the pattern of validation costs in the table below.
-For $K_{1}$ and $K_{2}$ being small, the validation cost decreases along with increasing values of $K_{1}$ and $K_{2}$ which may be credited to the decreasing approximation error.
-For $K_{1}$ and $K_{2}$ being larger than 4, the validation cost slightly increases, suggesting that as the PADR model becomes more complex, the increasing estimation error gradually dominates the decreasing approximation error.
+The training cost decreases consistently as $K_{1}$ and $K_{2}$ increase, indicating the stronger fitting capacity of PADR on the training data.
+For small values of $K_{1}$ and $K_{2}$, the validation cost also decreases significantly, which can be attributed to the reduction in approximation error.
+However, this decreasing trend does not hold as $K_{1}$ and $K_{2}$ become large; instead, the validation cost begins to plateau or even slightly increase.
+For instance, when $K_{2}$ is fixed at 40, the validation cost first decreases to 4.13 (at $K_{1}=5$) and then increases to 4.24 (at $K_{1}=20$).
+This suggests that as the PADR model becomes more complex, the increasing estimation error gradually dominates the decreasing approximation error.
 Hence, the best validation costs shall be achieved with moderate values of $K_{1}$ and $K_{2}$, which are usually considerably smaller than the ones implied by the theoretical results.
 […]
-Validation and training costs for basic Newsvendor with demand model (19) $K_{1}\backslash K_{2}$ 1 2 3 4 5 6 1 24.06 (24.09) 23.62 (23.57) 23.60 (23.63) 23.64 (23.55) 23.72 (23.66) 23.64 (23.62) 2 8.68 (8.73) 7.01 (7.09) 7.01 (7.09) 7.02 (7.08) 7.00 (7.10) 7.02 (7.09) 3 7.82 (7.95) 5.41 (5.53) 5.45 (5.51) 5.44 (5.51) 5.43 (5.49) 5.42 (5.49) 4 7.29 (7.47) 4.26 (4.27) 4.25 (4.22) 4.28 (4.29) 4.29 (4.21) 4.22 (4.23) 5 7.26 (7.46) 4.26 (4.27) 4.25 (4.24) 4.23 (4.21) 4.25 (4.21) 4.27 (4.22) 6 7.28 (7.47) 4.27 (4.24) 4.17 (4.23) 4.16 (4.23) 4.24 (4.27) 4.22 (4.20) The PADR is obtained by Algorithm 1 for solving the PADR-based ERM model.
-Numbers listed in brackets represent training costs, and the lowest costs are highlighted in bold and red text.
+Validation and training costs for basic Newsvendor with demand model (19) K1 $\backslash$ K2 1 2 3 4 5 10 20 40 1 24.06 (24.09) 23.62 (23.57) 23.61 (23.61) 23.64 (23.55) 23.56 (23.62) 23.55 (23.64) 23.54 (23.62) 23.56 (23.64) 2 8.68 (8.73) 7.01 (7.09) 7.07 (7.00) 7.02 (7.08) 7.08 (6.99) 7.05 (7.02) 7.05 (7.00) 7.05 (7.01) 3 7.95 (7.79) 5.60 (5.40) 5.47 (5.39) 5.40 (5.35) 5.42 (5.34) 5.36 (5.33) 5.41 (5.37) 5.39 (5.35) 4 7.29 (7.47) 4.26 (4.27) 4.25 (4.14) 4.28 (4.29) 4.14 (4.17) 4.13 (4.25) 4.15 (4.15) 4.14 (4.21) 5 7.41 (7.23) 4.21 (4.17) 4.34 (4.14) 4.32 (4.15) 4.17 (4.13) 4.34 (4.16) 4.14 (4.15) 4.11 (4.13) 10 7.41 (7.25) 4.15 (4.18) 4.15 (4.11) 4.10 (4.13) 4.14 (4.14) 4.06 (4.24) 4.04 (4.25) 4.07 (4.14) 20 7.41 (7.24) 4.16 (4.21) 4.15 (4.14) 4.10 (4.14) 4.32 (4.11) 4.04 (4.14) 4.00 (4.15) 4.03 (4.24) 40 7.40 (7.23) 4.15 (4.14) 4.12 (4.15) 4.08 (4.15) 4.11 (4.13) 4.01 (4.24) 3.99 (4.26) 4.02 (4.15) The PADR is obtained by Algorithm 1 for solving the PADR-based ERM model.
+Numbers listed in brackets represent training costs, and the lowest costs are highlighted in bold and blue text.
 2.3 Consistency comparison with existing DR-based methods
+We compare the consistency result of the PADR-based ERM model with two prominent types of DR-based ERM methods in Table 2.3.
+NV-ERM1 and NV-ERM2 represent the two LDR methods from Ban and Rudin (2018), the latter of which additionally applies regularization in the objective.
+$\text{RKHS-DR}_{\text{NP}}$ and $\text{RKHS-DR}_{\text{BK}}$ represent RKHS-DR methods from Notz and Pibernik (2021) and Bertsimas and Koduri (2021) respectively, with the former imposing regularization via constraints and the latter applying regularization in the objective.
 […]
-[…]
-[…]
-[…]
+Comparison of error bounds for multiple DR-based ERM models.
+Problem type Method Approximation error Estimation error Excess risk Convex NV-ERM1 0 (Linear asm.) $\widetilde{\mathcal{O}}(n^{-\frac{1}{2+p/2}})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{2+p/2}})$ NV-ERM2 0 (Linear asm.) $\geq\mathcal{O}(n^{-\frac{1}{2+p/4}})$ / $\text{RKHS-DR}_{\text{NP}}$ / $\mathcal{O}(n^{-\frac{1}{2}})$ / $\text{RKHS-DR}_{\text{BK}}$ 0 $\geq\mathcal{O}(n^{-\frac{1}{2}})$ / Nonconvex PADR $\mathcal{O}(K^{-\frac{1}{p}})$ $\widetilde{\mathcal{O}}(K^{\frac{p+2}{2p}}n^{-\frac{1}{2}})$ $\widetilde{\mathcal{O}}(n^{-\frac{1}{4+p}})$ The complexity with $\geq$” means that the error is only bounded partly.
+The notation “/” means that the error bound is not provided.
+Out of the four benchmarks considered in the literature, only Ban and Rudin (2018) provide an excess risk bound for NV-ERM1, leveraging the special structure of the Newsvendor problem with a linear true demand model.
+The estimation error bounds for NV-ERM2 and $\text{RKHS-DR}_{\text{BK}}$ are not provided due to the unknown regularization bias.
+Since an RKHS with a universal kernel is dense in the space of continuous functions, there is no approximation error for $\text{RKHS-DR}_{\text{BK}}$ , while such an error of $\text{RKHS-DR}_{\text{NP}}$ with a norm constraint is not analyzed in Notz and Pibernik (2021).
+It should be noted that the theoretical consistency results for all benchmarks in the aforementioned references, except $\text{RKHS-DR}_{\text{NP}}$ , are restricted to convex SP problems.
 In comparison, our PADR-based ERM model is nonasymptotically consistent in more general cases, including convex SP problems and nonconvex ones.
 Moreover, by Corollary 2.4, if there exists a convex or DC optimal DR, our excess risk bound can be improved to $\widetilde{\mathcal{O}}(n^{-1/(3+p/2)})$, which is close to the bound of NV-ERM1 (a special case of our model) when $p$ is large.
-This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$, for problem (1) with a box-constrained set $[-M,M]^{d}$ by employing a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{\mu}$ in Theorem 2.3 could be independent of $\mu$.
+This gap will vanish, that is, the bound can be further refined to $\widetilde{\mathcal{O}}(n^{-1/(2+p/2)})$, for problem (1) with a box-constrained set $M,M]^{d}$ by employing a truncated PADR class ${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ in (3), under which $M_{p,d}(\mu)$ in Theorem 2.3 could be independent of $\mu$.
 Given the statistical consistency merits of the PADR method demonstrated by the analysis above, it is still prohibitively challenging to obtain the globally optimal solution to the nonconvex PADR-based ERM problem (5).
@@
 where the objective is the finite sum of composite-PA functions.
-Instead of attempting for global optimality, it is more practical to apply majorization-minimization (MM)-type methods to the contextual optimization problems with the practical scale in seeking “stationary” solutions.
+Instead of attempting to achieve global optimality, it is more practical to apply majorization-minimization (MM)-type methods to the contextual optimization problems with the practical scale in seeking “stationary” solutions.
 Among various kinds of stationary concepts, the d-stationary solution is the sharpest type of first-order stationary solutions for general nonconvex optimization.
@@
 Building on this motivation, we develop an MM-type algorithm specializing to the structured PADR-based ERM problem with the theoretical convergence analysis.
-It is further observed with numerous experiments in Section 4 that satisfactory (nearly optimal) solutions can be obtained with a few number of repetitions, which speeds up to 4 times faster than the basic MM method.
+It is further observed with numerous experiments in Section 4 that satisfactory (nearly optimal) solutions can be obtained with a few number of repetitions, which is up to 4 times faster than the basic MM method.
 To make our discussion self-contained, we first present the standard MM algorithms in Section 3.1, which serves as a reference for the motivation behind our proposed sampling-based algorithm in Section 3.2.
 We make some remarks on notations to facilitate algorithmic development and convergence analysis.
-To emphasize the role of PADR as a function of $\theta$, with $\xi=(\xi_{x},\xi_{y})\in\Xi$, we denote $g(\theta,\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta,\xi_{x})$ and $h(\theta,\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta,\xi_{x})$ where $\{g_{i}(\theta,\xi_{x})\}$ and $\{h_{i}(\theta,\xi_{x})\}$ are linear functions with respect to $\theta$ and we denote $f(\theta,\xi)\coloneqq g(\theta,\xi)-h(\theta,\xi)$ as the PA function with a slight abuse of notation.
+To emphasize the role of PADR as a function of $\theta$, with $\xi=(x,y)\in\Xi$, we denote $g(\theta,\xi)\coloneqq\max_{i\in[K]}g_{i}(\theta,x)$ and $h(\theta,\xi)\coloneqq\max_{i\in[K]}h_{i}(\theta,x)$ where $\{g_{i}(\theta,x)\}$ and $\{h_{i}(\theta,x)\}$ are linear functions with respect to $\theta$, and we denote $f(\theta,\xi)\coloneqq g(\theta,\xi)-h(\theta,\xi)$ as the PA function with a slight abuse of notation.
 With $\widetilde{\xi}_{n}$ representing a random variable uniformly distributed on $\Xi$, the PADR-based ERM problem (5) can be generalized as the following stochastic program with finite scenarios:
@@
     Convexity: $\widehat{F}({\cdot};{\bar{\theta}})$ is convex on $\Theta$.
-A function satisfying (P1) and (P2) is referred to as an upper surrogate function according to Cui and Pang (2021, Definition 7.1.1).
+A function satisfying (P1) and (P2) is referred to as an upper surrogate function according to Cui and Pang (2021a, Definition 7.1.1).
 Under (P1) and (P2), it yields the sequence of decreasing objective values as follows,
@@
 Condition (P3) ensures that surrogation-based subproblems in the algorithm can be solved by state-of-the-art convex-programming solvers.
-By the basic analysis (see Chapter 7 in Cui and Pang (2021)), the limit point of $\{\theta^{\nu}\}$ produced by the process (11) could be easily shown to be a fixed point of the surrogation map satisfying $\theta^{*}=\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta^{*})+\frac{\eta}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}\right\}$.
+By the basic analysis (see Chapter 7 in Cui and Pang (2021a)), the limit point of $\{\theta^{\nu}\}$ produced by the process (11) could be easily shown to be a fixed point of the surrogation map satisfying $\theta^{*}=\underset{\theta\in\Theta}{\mathop{\arg\min}}\left\{\widehat{F}(\theta;\theta^{*})+\frac{\eta}{2}\left\lVert\theta-\theta^{*}\right\rVert^{2}\right\}$.
 Such a fixed point property could be related to various types of stationarity of the original problem depending on the surrogate functions as illustrated in Cui et al.
-(2018) and Liu et al.
-(2022).
+(2018a) and Liu et al.
+(2022a).
 Regarding the PADR-based ERM problem (10), due to the nonconvex and nondifferentiable piecewise affine functions, a major issue of the standard MM algorithms including Lipp and Boyd (2016) and Le Thi and Pham Dinh (2018) is that the fixed point corresponding to the surrogation map is a (compound) critical solution (see definitions in Pang et al.
 (2016), Liu et al.
-(2022)) which is much weaker than directional stationarity.
+(2022a)) which is much weaker than directional stationarity.
 Specifically, based on the composite structure in (10), approaches to construct surrogate functions of $F(\theta,\xi)$ in Cui et al.
-(2018) typically involve linearization of the two max-affine components of the inner piecewise affine function as follows,
+(2018a) typically involve linearization of the two max-affine components of the inner piecewise affine function as follows,
 $\widehat{f}(\theta,\xi;I_{2})$ $\coloneqq g(\theta,\xi)-h_{I_{2}}(\theta,\xi)\mbox{ with }I_{2}\in\mathcal{I}_{h}(\bar{\theta},\xi)\coloneqq\mbox{argmax}_{i\in[K]}h_{i}(\bar{\theta},\xi),$ (12)
 $\widecheck{f}(\theta,\xi;I_{1})$ $\coloneqq g_{I_{1}}(\theta,\xi)-h(\theta,\xi)\mbox{ with }I_{1}\in\mathcal{I}_{g}(\bar{\theta},\xi)\coloneqq\mbox{argmax}_{i\in[K]}g_{i}(\bar{\theta},\xi).$ (13)
-They are the upper and lower surrogate functions of $f(\cdot,\xi)$ at $\bar{\theta}\in\Theta$ respectively.
-For $d>1$, the maximum is taken componentwisely.
-We illustrate the construction of the surrogate function when outer function $\varphi$ is linear as an example.
-With $\varphi(z,\xi)=\xi^{\top}z$, we construct the surrogate function of $F(\cdot,\xi)$ as $\widehat{F}(\theta,\xi;\bar{\theta},I)=\max\{\xi,0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+\min\{\xi,0\}^{\top}\widecheck{f}(\theta,\xi;I_{1})$, where $I\coloneqq\big{(}I_{1},I_{2}\big{)}\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$ for any $\xi\in\Xi$.
-It obviously satisfies Conditions (P1) to (P3), and the standard MM algorithm for such a case is exactly a DC algorithm.
+Utilizing the above upper and lower surrogate functions of $f(\cdot,\xi)$ at $\bar{\theta}\in\Theta$, the following example illustrates the construction of the surrogate function when outer function $\varphi(\cdot,\xi)$ is linear.
+Example 3.1
+For $\varphi(z,\xi)=c_{\xi}^{\top}z$, where $c_{\xi}\in\mathbb{R}^{d}$ is a parameter dependent on $\xi$, the composite objective function is $F(\theta,\xi)=c_{\xi}^{\top}f(\theta,\xi)$, and its surrogate function is constructed as
+$\widehat{F}(\theta,\xi;\bar{\theta},I)=\max\{c_{\xi},0\}^{\top}\widehat{f}(\theta,\xi;I_{2})+\min\{c_{\xi},0\}^{\top}\widecheck{f}(\theta,\xi;I_{1}),$
+where $I\coloneqq\big(I_{1},I_{2}\big)\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$. For $d>1$, the maximum is taken componentwisely.
+Surrogation in Example 3.1 obviously satisfies Conditions (P1) to (P3), and the standard MM algorithm for such a case is exactly a DC algorithm.
 Clearly $f^{\prime}(\cdot,\xi)(\theta;d)=g^{\prime}(\cdot,\xi)(\theta;d)-\max\{\,\nabla h_{i}(\bar{\theta},\xi)^{\top}d:\,i\in\mbox{argmax}_{i\in[K]}h_{i}(\bar{\theta},\xi)\}$, whereas the surrogate function $\widehat{F}(\theta,\xi;\bar{\theta},I)$ employs an arbitrary active index $I\in\mathcal{I}_{g}(\bar{\theta},\xi)\times\mathcal{I}_{h}(\bar{\theta},\xi)$.
@@
 Consider problem (10) with $d=2$, $q=1$, $\Xi=\{\xi_{0}\}$, $\varphi(z,\xi_{0})=z_{1}^{2}+z_{2},f_{1}(\theta,\xi_{0})=\theta$, and $f_{2}(\theta,\xi_{0})=-\max\{-\theta,0\}$.
+In this case, the composite objective function is $F(\theta)=\theta^{2}+\max\{-\theta,0\}$.
 Figure 1a illustrates the two candidates of surrogate functions depending on which piece of the inner PA function $f_{2}(\theta,\xi_{0})$ is active at the reference point of interest.
@@
 (2022) and nondifferentiable stochastic composite DC problems in Liu et al.
-(2022) to compute compound critical solutions.
+(2022a) to compute compound critical solutions.
 Based on a natural idea of combining a sequential sampling scheme with the enhanced technique, we propose an enhanced sampling-based MM algorithm with rigorous convergence analysis in the next subsection to solve the stochastic composite DC programs (10) with both computational efficiency and theoretical convergence guarantee.
@@
     the minimum surrogate function $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})\coloneqq\min_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\,\xi)}\widehat{F}(\theta,\xi;\bar{\theta},I)$ satisfies the directional derivative consistency for any $\varepsilon\geq 0$; namely, $\widehat{F}_{\min}^{\varepsilon}(\cdot\,;\bar{\theta})$ is directionally differentiable on $\Theta$ and satisfies $\widehat{F}_{\min}^{\varepsilon}(\cdot\,;\bar{\theta})^{\prime}(\bar{\theta};d)=F^{\prime}(\bar{\theta};d)$ for any $d\in\mathbb{R}^{q}$.
-The directional derivative consistency condition is introduced in Cui and Pang (2021) to relate the directional derivatives of surrogation and original functions.
+The directional derivative consistency condition is introduced in Cui and Pang (2021a) to relate the directional derivatives of surrogation and original functions.
 We consider the following three cases of the outer function $\varphi$ covering a broad class of cost functions.
@@
 In particular, Case 2 is a multivariate extension of the one considered in Cui et al.
-(2018).
+(2018a).
 Case 3 is suitable for nonconvex smooth cost functions that are applicable for many practical problems.
@@
 Specialized to the composite DC program (10), according to Cui et al.
-(2018, Proposition 2), d-stationarity is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2.
+(2018a, Proposition 2), d-stationarity is equivalent to local optimality if $\varphi$ is in Case 1 or Case 2.
 Lu et al.
@@
 Moreover, it is weaker than composite $\varepsilon$-strong d-stationarity introduced in Qi et al.
-(2021b) (see also Cui and Pang (2021, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition which turns out to be essential for the algorithm development and the follow-up convergence analysis.
+(2021b) (see also Cui and Pang (2021a, Definition 6.1.3)) due to the presence of an additional proximal term in the defining condition which turns out to be essential for the algorithm development and the follow-up convergence analysis.
 Furthermore, in Online Appendix 8.1, we show that $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ under a regularity assumption on $F$.
@@
 In the special case of $\varepsilon=0$ for which the touching condition (P1) is satisfied, the proximal mapping point $\theta^{\nu+\frac{1}{2}}$ satisfies the sufficient descent inequality in step 5, and thus we could directly update $\theta^{\nu+1}=\theta^{\nu+\frac{1}{2}}$.
+We remark that under such an update scheme, since the size of $\varepsilon$-active index set grows exponentially along with the number of pieces of PADR, the iterate point would be more easily rejected for PADR with larger pieces.
 Input: initial point $\theta^{0}$, iteration length $T$, parameters $\{\eta,\varepsilon\}$, and increasing sequence $\{N_{\nu}\}_{\nu}$;
-1 for $\nu=0$ to $T-1$ do
+1 for $\nu=0$ to T1T-1 do
 2   generate $\Xi_{\nu}\coloneqq\{\widetilde{\xi}_{\nu}^{s}\}_{s=1}^{N_{\nu}}$ independently and uniformly from $\Xi$;
@@
 (2016), Cui et al.
-(2018), Lu et al.
+(2018a), Lu et al.
 (2019).
@@
 Then any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ which must exist is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of problem (10) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$
-Theorem 3.6 shows the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity.
-Moreover, under the aforementioned settings where $\mathcal{S}^{d}_{\varepsilon,\rho}$ is equivalent to $\mathcal{S}^{*}$ with sufficiently large $\varepsilon$ and a regularity assumption on $F$, any limit point of $\{\widetilde{\theta}^{\,t}\}_{t}$ is a global minimizer of (10) with probability 1.
+Theorem 3.7 shows the asymptotic convergence of the ESMM algorithm to composite strong d-stationarity that is stronger than d-stationarity by Proposition 3.5.
+Moreover, Proposition 8.2 in Online Appendix 8.1 shows that $\mathcal{S}^{d}_{\varepsilon,\rho}$ coincides with $\mathcal{S}^{*}$ under a sufficiently large $\varepsilon$ and a regularity assumption on $F$.
+Consequently, any limit point of ${\widetilde{\theta}^{,t}}_{t}$ generated by the algorithm is a global minimizer of (10) with probability one.
 This indicates that the PADR-based ERM model together with the ESMM algorithm is asymptotically optimal to the original problem (2).
@@
 As noted in Liu et al.
-(2022), due to the composite structure of objective functions, the SMM algorithm with $\varepsilon=0$ may yield convergence to compound criticality, which is weaker than d-stationarity.
-This is also supported with Example 3.1 in Section 3.1.
+(2022a), due to the composite structure of objective functions, the SMM algorithm with $\varepsilon=0$ may yield convergence to compound criticality, which is weaker than d-stationarity.
+This is also supported by Example 3.2 in Section 3.1.
 Remark 3.11
@@
 For the simulated data, we set the true model of $Y$ as $Y=\overline{Y}(X)+\widetilde{\epsilon}$, where $\overline{Y}(\cdot)$ represents the mean model, and $\widetilde{\epsilon}$ denotes a random variable with a standard normal distribution independent of $X$.
-Throughout all experiments with simulated data, we generate feature data uniformly from $[-1,1]^{p}$, and ensure the mean model satisfies $\min_{x\in[-1,1]^{p}}\overline{Y}(x)\geq 5$ which keeps all samples of $Y$ positive.
+Throughout all experiments with simulated data, we generate feature data uniformly from $1,1]^{p}$, and ensure the mean model satisfies $\min_{x\in[-1,1]^{p}}\overline{Y}(x)\geq 5$ which keeps all samples of $Y$ positive.
 Unless otherwise noted, both training and test datasets contain 1000 samples.
 […]
-Methods implemented in numerical experiments Abbreviation Explanation of method PADR Piecewise-affine decision rule method in this paper LDR Linear decision rule method from Ban and Rudin (2018) RKHS-DR Kernel optimizer prediction method from Bertsimas and Koduri (2021) kNN, kernel, cart, rf Prescriptive methods from Bertsimas and Kallus (2019) based on $k$-nearest neighbiors, kernels, regression trees, and random forests respectively sof Prescriptive method from Kallus and Mao (2023) based on stochastic optimization forests PO-PA, PO-L PO methods with piecewise-affine prediction model and linear prediction model respectively SIMOPT Simulated optimal decision using true conditional distribution
-We refer to the PADR model with two max-affine components consisting of $K_{1}$ and $K_{2}$ pieces respectively as PADR$(K_{1},K_{2})$.
-The hyperparameters $K_{1}$ and $K_{2}$ are searched within the range [0, 4] for the Newsvendor problems in Section 4.1 and Section 4.2, and [0, 6] for the product placement problem in Section 4.3.
-Regarding the ESMM algorithm for learning the PADR, initial points are generated uniformly from $[-50,50]^{d(K_{1}+K_{2})(p+1)}$, and the sampling sequence $\{N_{\nu}\}_{\nu}$ follows $N_{\nu}=40\nu$.The parameter $\varepsilon$ in Algorithm 1 is set to be either constant or shrinking, the latter case of which is referred to as the shrinking-$\varepsilon$ strategy.
-Specifically, the shrinking strategy sets the parameter a constant value $\varepsilon_{0}$ for the first $T_{0}$ iterations and then a reduced value $\varepsilon_{1}(<\varepsilon_{0})$ for the remaining ones.
-The ESMM algorithm produces the iterate point obtaining the minimum ERM cost as the output after $T$ iterations.
-For each instance, the algorithm is executed for $r$ rounds with random initialization, and we report the best performance unless otherwise noted.
-We set $T=10$ and $r=10$ in Sections 4.1 and 4.2, and $T=20$ and $r=5$ in Section 4.3.
-The reported running time of PADR is the total time of all rounds.
+Methods implemented in numerical experiments Abbreviation Explanation of method PADR Piecewise-affine decision rule method in this paper RKHS-DR Kernel optimizer prediction method from Bertsimas and Koduri (2021) NN-DR Neural network-based decision rule method with two or three fully connected hidden layers LDR Linear decision rule method from Ban and Rudin (2018) kNN, KERNEL, CART, RF Prescriptive methods from Bertsimas and Kallus (2019) based on $k$-nearest neighbiors, kernels, regression trees, and random forests respectively SOF Prescriptive method from Kallus and Mao (2023) based on stochastic optimization forests PO-PA, PO-L PO methods with piecewise-affine prediction model and linear prediction model respectively SIMOPT Simulated optimal decision using true conditional distribution
+[…]
 We clarify settings for benchmark methods listed in Table 4.
+For NN-DR, we use ReLU function ($\max\{\cdot,0\}$) as activation, and the number of nodes in each hidden layer is set to half the size of the preceding layer.
+In each experiment, the model is trained for 100 epochs over 3 independent runs using the Adam optimizer (Kingma and Ba 2014) to minimize the empirical cost with $\ell_{1}$-regularization.
 We use a Gaussian kernel for kernel methods.
@@
 […]
-Hyperparameters of each method Methods Hyperparameters PADR number of pieces (PADR); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM) LDR N/A RKHS-DR kernel bandwidth, regularization parameter, penalty parameter kNN, kernel, cart, rf number of nearest neighbors (kNN); kernel bandwidth, regularization parameter (kernel); depth, minimum samples to split, minimum samples per leaf (cart and rf); sof choice of apx-soln or apx-risk method in Kallus and Mao (2023), depth, minimum samples to split, minimum samples per leaf PO-PA number of pieces (prediction model); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM solving $\ell_{2}$ regression) PO-L N/A SIMOPT N/A
+Hyperparameters of each method Methods Hyperparameters PADR number of pieces (PADR); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM) RKHS-DR kernel bandwidth, regularization parameter, penalty parameter NN-DR first-hidden-layer size, regularization parameter, learning rate, weight decay, batch size LDR N/A kNN, KERNEL, CART, RF number of nearest neighbors (kNN); kernel bandwidth, regularization parameter (KERNEL); depth, minimum samples to split, minimum samples per leaf (CART and RF); SOF choice of apx-soln or apx-risk method in Kallus and Mao (2023), depth, minimum samples to split, minimum samples per leaf PO-PA number of pieces (prediction model); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM solving $\ell_{2}$ regression) PO-L N/A SIMOPT N/A
 All methods are implemented in Python 3.9.7 on a MacBook Pro with a macOS Monterey system and an Apple M1 Pro chip.
+NN-DR is implemented in Pytorch version 2.2.2.
 All programs involved are built and solved by the optimizer Gurobi, version 9.5.0.
@@
 where $c_{b}$ and $c_{h}$ are the unit back-order and holding costs respectively, and $(\cdot)_{+}$ denotes $\max\{\cdot,0\}$.
-We set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $k=1$ unless otherwise specified:
-$\overline{Y}(X)=k\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (17)
+We set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $\omega=1$ unless otherwise specified:
+$\overline{Y}(X)=\omega\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (17)
 4.1.1 Basic results.
@@
 (e) Running time
-Newsvendor problem: basic results and running time
-Since the optimal DR in this case is a max-affine function with three pieces, PADR$(K_{1},K_{2})$ with $K_{1}\geq 3$ has no approximation error in principle.
-Figure 1d shows that the costs of PADR with $(K_{1},K_{2})=(3,0)$ converge to SIMOPT as the training sample size $n$ increases from 50 to 1000.
-RKHS-DR converges slower than PADR; additionally, its performance with $n=1000$ is worse than that of PADR with $n=400$.
-Prescriptive methods, including kNN, cart, rf, and sof, are dominated by the PADR and RKHS-DR methods.
-For PO-PA, we observe that the prediction model provided by the PO-PA method is quite close to the true model $\overline{Y}(\cdot)$, so its substantial gap to SIMOPT could be mainly due to inconsistency between the prediction loss and Newsvendor cost with imbalanced $c_{b}$ and $c_{h}$.
-Numerical results with the parameter ratio $c_{b}:c_{h}$ varying from 1:9 to 9:1 exhibit similar patterns and thus are omitted.
-Figure 1e shows the training time of each method.
-The training time of RKHS-DR grows roughly exponentially when $n\geq 200$, whereas the time for PA-related methods remains stable and significantly shorter when $n=1000$, which is mainly credited to the enhanced and sampling techniques used in the ESMM algorithm.
+Newsvendor problem: basic results and running time PO-L, KERNEL, and LDR are excluded in (a) for readability, whose costs are around 14, 11, and 9 respectively.
+[…]
 […]
@@
 (g) Sampling technique
+(h) Time across $K_{1}$ and $K_{2}$
 Performance of ESMM under various settings In (a), solid and dotted lines represent constant-$\varepsilon$ strategies; dashed lines represent shrinking-$\varepsilon$ strategies with $T_{0}=3$ and $\varepsilon_{1}=0$.
@@
 The y-axis is set on a log scale.
-In Figure 1g, we provide results of the ESMM algorithm under different parameter settings that illustrate the effect of the enhanced sampling strategies.
+In (c), each point represents the running time of PADR with $K_{1}$ and $K_{2}$ varying from 0 to 40 for the Newsvendor problem with $p=2$ and $n=1000$.
+In Figure 1h, we provide results of the ESMM algorithm under different parameter settings that illustrate the effect of the enhanced sampling strategies.
 Figure 1f displays curves of sorted training cost of PADR provided by the ESMM algorithm under the constant-$\varepsilon$ and shrinking-$\varepsilon$ strategies over 30 rounds.
@@
 The time required for EMM is approximately 4 times longer than that for ESMM when $n=1000$.
+Figure 1h illustrates the running time of PADR for multiple settings of $(K_{1},K_{2})$ with each value ranging from 1 to 40.
+The running time for 10 rounds remains within 100 seconds when $\max\{K_{1},K_{2}\}\leq 40$.
+An interesting observation is that, with the same total number of pieces (i.e., total number of parameters), the time consumption of PADR is relatively higher when its two components have more balanced piece numbers, and considerably lower when the decision rule consists of only one component (i.e., $K_{1}=0$ or $K_{2}=0$).
+This can be explained by the complex structure of the surrogate function for the PADR with balanced piece numbers.
 4.1.2 Varying feature dimensions.
@@
 […]
-(h) Sparse demand model
-(i) Dense demand model
+(i) Sparse demand model
+(j) Dense demand model
 Newsvendor problem with varying feature dimensions For readability, kNN is excluded under the dense demand model, whose costs exceed 8 in most trials.
-Figure 1i displays the results.
+[…]
+(k) Sparse demand model
+(l) Dense demand model
+Running time in Newsvendor problem with varying feature dimensions
+Figure 1j and Figure 1l report the test cost and running time, respectively.
 In the sparse case, forest-based prescriptive methods are robust to redundant features, as they only focus on informative features.
-When the dimension is less than 20, PADR method is competitive with comparable performance to sof and much less training time.
-It would be beneficial to employ an effective feature selection technique or regularization in the PADR method to improve the performance in the sparse case with a large feature dimension.
+It would be beneficial to employ an effective feature selection technique in the PADR method to improve the performance in the sparse case with a large feature dimension.
 In the dense case, PADR performs generally well, but its performance slightly deteriorates as $p$ increases due to growing estimation errors and optimization difficulties.
 We also observe that PADR with fewer pieces performs more stably in both cases despite the presence of approximation errors.
+When $p<20$, PADR consistently achieves slightly better performance than benchmarks, while requiring substantially less training time.
+For instance, at $p=8$, PADR attains a 10% lower test cost compared to NN-DR and SOF, with a 90% less runtime than these methods.
+This highlights the efficiency of PADR in learning decision rules in settings with moderate feature dimensions.
 4.1.3 Nonlinearity in the demand model.
-We study the effect of nonlinearity of the demand with respect to the feature variables by varying the coefficient $k$ in the demand model (17) and report results in Figure 1k.
+We study the effect of nonlinearity of the demand with respect to the feature variables by varying the coefficient $\omega$ in the demand model (17) and report results in Figure 1o.
 It indicates that the performance of PADR and PO-PA is irrelevant to the nonlinearity level of the demand model due to the strong approximation ability to the nonlinear structure.
 In contrast, the performance of RKHS-DR and prescriptive methods deteriorates significantly as the nonlinearity level increases.
-These methods require significantly more samples to achieve the similar performance when the level of nonlinearity increases.
+These methods require significantly more samples to achieve similar performance when the level of nonlinearity increases.
+NN-DR shows the robustness towards non-linearity for the demand model with $p=20$; however, additional experiments in Figure 1o reporting relative cost gaps to SIMOPT for PADR, RKHS-DR, and NN-DR with multiple dimension settings, show that the performance of RKHS-DR deteriorates as $p$ increases, while NN-DR becomes less sensitive to nonlinearity as $p$ increases.
+We speculate that when $p$ is small, the instability of gradient-based optimization becomes more pronounced, and this effect is further amplified under highly nonlinear demand functions.
+To further test the effect of nonlinearity, we consider demand models with quadratic, cubic, and sine terms (in addition to a shared linear component).
+Results in Online Appendix 12 show consistent patterns across methods, validating PADR’s robustness to model misspecification under varying nonlinear structures.
 […]
-(j) $p=2\phantom{xxxxxxx}$
-(k) $p=20$
+(m) $p=2\phantom{xxxxxxx}$
+(n) $p=20$
+(o) Relative Cost Gap ($\omega=3$)
 Newsvendor problem with varying nonlinearity in demand model In (b), the demand model is of the dense type. For readability, kNN, with costs increasing from 5 to 40, is excluded in (a).
@@
 In this section, we conduct experiments on a real-world dataset, which consists of order records for a specific model of mobile phone from a Chinese electronic product supplier.
-The order record, spanning from September 01, 2021 to November 31, 2023, contains order quantities and time stamps from retailers.
+The order record, spanning from September 01, 2021 to November 30, 2023, contains order quantities and time stamps from retailers.
 In order to create a production plan, the supplier needs to determine the packing quantity in each day to fulfill the order requests.
@@
 The aggregated daily order quantities have been rescaled for confidentiality purposes and also normalized for pretreatment.
-The boxplots of daily orders in Figure 1n show that order quantities are larger on weekdays and during the second quarter of each month, which aligns with retailers’ ordering pattern.
+The boxplots of daily orders in Figure 1r show that order quantities are larger on weekdays and during the second quarter of each month, which aligns with retailers’ ordering pattern.
 […]
-(l) Weekdays
-(m) Month days
-(n) Month quarters
+(p) Weekdays
+(q) Month days
+(r) Month quarters
 Real-world dataset: boxplots of orders by time period
@@
 For the third type of features, we incorporate order quantities in past days of different lengths as different instances in conducting the experiments.
+Note that incorporating retailer’s past order quantities into the feature vector introduces temporal dependence, violating the i.i.d.
+assumption in Assumption (A1).
+This issue can be addressed using generalization bounds for non-stationary $\beta$-mixing processes, as established by Kuznetsov and Mohri (2017).
+Since only a fixed-length history is used in the feature, the resulting data sequence is $\beta$-mixing with finite memory, i.e., there exists a lag $a_{0}$ such that $\beta(a)=0$ for all $a>a_{0}$.
+Under such settings, Theorem 1 and Corollary 1 from Kuznetsov and Mohri (2017) ensure that the generalization bound still holds, and thus the excess risk analysis remains valid.
 To determine the packing quantity in each day to fulfill the order requests, we formulate a Newsvendor problem with features listed above and report the numerical results in Table 4.2.
-It shows that PADR and forest-based prescriptive methods work well.
-Overall, the PADR method performs slightly better in most instances with features of different length.
+The results show that PADR and NN-DR exhibit competitive performance.
+PADR outperforms NN-DR when $p\leq 15$, whereas NN-DR demonstrates greater robustness as the feature dimension increases.
 […]
-Test costs for newsvendor on a real-world dataset Features $p$ DR methods Prescriptive methods PO methods PADR RKHS-DR kNN cart rf sof PO-PA basic features 11 55.95 56.10 56.06 56.00 56.00 56.00 79.02 basic features + order quantities in past 1 day 13 42.67 51.53 50.25 61.72 49.54 48.70 68.18 basic features + order quantities in past 3 days 15 41.65 55.59 48.25 54.64 41.74 42.19 59.73 basic features + order quantities in past 5 days 17 41.40 65.17 49.59 56.74 40.90 41.83 52.39 basic features + order quantities in past 7 days 19 40.93 70.60 49.02 65.62 43.62 43.51 54.52 basic features + order quantities in past 14 days 26 41.03 78.37 46.81 58.39 42.45 42.48 52.10 basic features + order quantities in past 21 days 33 42.43 80.24 46.61 59.45 41.59 41.33 53.44 The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
-We next consider a practical setting where the supplier incurs an additional cost associated with packing quantities in actual production operations, in order to balance the daily inventory cost and the packing cost.
-The cost is approximated by the following nonconvex cost function:
-$C(z)=\min\{z,0.6z+0.8,0.4z+15.6\},$ (18)
-Therefore, we consider a Newsvendor problem with an additional concave cost $C(z)$ in the objective.
-Since the concave cost introduces nonconvexity into the problem.
+Test costs for Newsvendor on a real-world dataset Added features $p$ DR methods Prescriptive methods PO methods (all include basic features) PADR RKHS-DR NN-DR kNN CART RF SOF PO-PA None 11 55.95 56.10 56.74 56.06 56.00 56.00 56.00 79.02 + order quantities in past 1 day 13 42.67 51.53 45.13 50.25 61.72 49.54 48.70 68.18 + order quantities in past 3 days 15 41.65 55.59 42.46 48.25 54.64 41.74 42.19 59.73 + order quantities in past 5 days 17 41.40 65.17 41.78 49.59 56.74 40.90 41.83 52.39 + order quantities in past 7 days 19 40.93 70.60 40.27 49.02 65.62 43.62 43.51 54.52 + order quantities in past 14 days 26 41.03 78.37 41.04 46.81 58.39 42.45 42.48 52.10 + order quantities in past 21 days 33 42.43 80.24 40.12 46.61 59.45 41.59 41.33 53.44 The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
+We next consider a practical setting in which the supplier incurs an additional cost due to packaging requirements in actual production operations, aiming to balance daily inventory and packing costs.
+The packing cost follows an incremental discount scheme (Tersine and Barman 1991), where the unit price decreases for quantities exceeding certain thresholds.
+To align with the rescaled order quantities, we specify the following piecewise affine concave cost function:
+$C(z)=\min\{z,0.6z+8,0.4z+15.6\}.$ (18)
+Accordingly, we consider a Newsvendor problem with an additional term $C(z)$ in the objective function.
+The inclusion of this concave term introduces nonconvexity into the optimization problem.
 we use the deterministic MM algorithm to solve the ERM problem for learning RKHS-based decision rule, referred to as RKHS-DR(MM).
 For prescriptive methods, we solve one-dimensional nonconvex problems by line search.
-We do not implement sof from Kallus and Mao (2023) in this nonconvex Newsvendor problem, as it requires a new development of second-order perturbation analysis of stochastic optimization, which is beyond the scope of this paper.
+As SOF from Kallus and Mao (2023) requires a new development of second-order perturbation analysis of stochastic optimization to accommodate the nonconvex term, we instead replace $C(z)$ by its affine piece $0.6z+8$ when computing weights for training samples.
+We refer to this linearized version as SOF(L).
 […]
 Test costs for nonconvex Newsvendor on a real-world dataset
-Features $p$ DR methods Prescriptive methods PO methods
-PADR RKHS-DR(MM) kNN cart rf PO-PA
-basic features 11 74.28 81.89 84.52 81.56 81.56 90.48
-basic features + order quantities in past 1 day 13 64.75 72.11 77.86 86.58 73.78 77.89
-basic features + order quantities in past 3 days 15 64.10 68.29 75.58 78.31 64.99 73.61
-basic features + order quantities in past 5 days 17 63.20 64.72 76.81 79.54 65.74 71.56
-basic features + order quantities in past 7 days 19 62.01 65.35 75.38 85.95 68.35 69.97
-basic features + order quantities in past 14 days 26 64.86 65.15 73.81 75.52 64.71 70.80
-basic features + order quantities in past 21 days 33 65.65 65.15 73.81 75.52 65.73 71.02
+Added features $p$ DR methods Prescriptive methods PO methods
+(all include basic features) PADR RKHS-DR NN-DR kNN CART RF SOF(L) PO-PA
+None 11 74.28 81.89 80.67 84.52 81.56 81.56 82.50 90.48
++ order quantities in past 1 day 13 64.75 72.11 68.19 77.86 86.58 73.78 74.45 77.89
++ order quantities in past 3 days 15 64.10 68.29 66.62 75.58 78.31 64.99 66.54 73.61
++ order quantities in past 5 days 17 63.20 64.72 64.07 76.81 79.54 65.74 67.66 71.56
++ order quantities in past 7 days 19 62.01 65.35 62.25 75.38 85.95 68.35 69.23 69.97
++ order quantities in past 14 days 26 64.86 65.15 63.45 73.81 75.52 64.71 65.58 70.80
++ order quantities in past 21 days 33 65.65 65.15 63.33 73.81 75.52 65.73 66.01 71.02
 The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
-[…]
+Table 4.2 demonstrates that PADR consistently outperforms other methods when $p\leq 17$, with performance gains exceeding 5% for $p\leq 13$, and achieves the lowest test cost at $p=19$.
+These findings highlight PADR’s effectiveness in addressing nonconvex contextual SP problems.
+On the other hand, NN-DR maintains superior performance when $p\geq 26$.
+Forest-based methods, however, perform relatively poorly in this nonconvex setting.
+The performance gap between SOF(L) and RF becomes more pronounced than in the convex case, since linearizing the concave cost introduces considerable model misspecification.
+Moreover, the persistently weak performance of PO-PA across both convex and nonconvex scenarios indicates that its sequential prediction and decision-making process may not be suitable for this real-world dataset.
 4.3 Product placement problem
@@
 To evaluate PADR’s performance under more complex conditions, we employed the following demand model:
-$\overline{Y}^{\prime}(X)=k\alpha_{1}\sin(\alpha_{2}X_{1})+k\max\{\alpha_{3}X_{2},\alpha_{4}X_{2}\}+\alpha_{5},$ (19)
+$\overline{Y}^{\prime}(X)=\omega\alpha_{1}\sin(\alpha_{2}X_{1})+\omega\max\{\alpha_{3}X_{2},\alpha_{4}X_{2}\}+\alpha_{5},$ (19)
 where the sine function represents the seasonal dependence of the demand on $X_{1}$, and the max-affine function represents the nonlinear trend with respect to $X_{2}$.
-We set $k=1$ unless otherwise specified.
-We randomly generate parameters $\{\alpha_{i}\}_{i=1}^{5}$ and the node-arc matrix $W$ for each trial, and we report performance of methods over 20 trials for each experimental setting.
-The sof method from Kallus and Mao (2023) is not applied for this two-stage SP, as it would require a new development of second-order perturbation analysis of stochastic optimization, which is beyond the scope of this paper.
+We set $\omega=1$ unless otherwise specified.
+We randomly generate parameters $\{\alpha_{i}\}_{i=1}^{5}$ and the node-arc matrix $W$ for each trial and report performance of methods over 10 trials for each experimental setting.
+For NN-DR, incorporating the second-stage problem into the loss function within the standard training framework is nontrivial.
+To address this difficulty, we treat the second-stage decision variables as additional model parameters and jointly optimize them along with the NN-DR parameters.
+It is worth noting that this implementation is tailored specifically to the problem at hand, and it underscores the limitation of NN–based approaches for multi-stage stochastic programs.
+The SOF method of Kallus and Mao (2023) is not included in this experiment, as applying it to a two-stage stochastic program would require substantial extensions involving second-order perturbation analysis, which are beyond the scope of this paper.
 4.3.1 Varying feature dimensions.
 We first study the effect of the feature dimension $p$ on the performance of PADR under sparse and dense demand models similar to the discussion in Section 4.1.2.
-The results are reported in Figure 1p.
+The results are reported in Figure 1t.
 Similar to the results in Section 4.1.2 for the Newsvendor problem, with sparse demand model, PADR performs well in low-dimensional cases but deteriorates as $p$ increases which may necessitate the proper parameter regularization.
+NN-DR performs poorly in the sparse case, also suffering from overfitting.
 In contrast, forest-based prescriptive methods are robust to redundant features.
-With dense demand model, PADR and PO-PA have the best performance, with the former one slightly outperforming when the feature dimension increases to 50.
+With dense demand model, PADR and NN-DR have the best performance, with the former one slightly outperforming when $p$ is small.
+We also report the running time results in Figure 1v.
+PADR requires substantially more computational time in this problem, as the underlying optimization becomes more complex and involves multiple decision variables.
+In the current implementation, the decision rule is constructed separately for each dimension.
+Given so, we believe the computational efficiency of PADR could be further improved through vectorized multi-dimensional decision rule implementation and engineering optimizations in the ESMM surrogate construction.
 Despite the increased complexity and the challenge in accurately approximating the true model, PADR continues to demonstrate superior performance for instances with dense demand model, high nonlinearity, and sparse demand models with small feature dimensions.
 […]
-(o) Sparse demand model
-(p) Dense demand model
+(s) Sparse demand model
+(t) Dense demand model
 Product placement problem with varying feature dimensions
+[…]
+(u) Sparse demand model
+(v) Dense demand model
+Running time in product placement problem with varying feature dimensions
 4.3.2 Nonlinearity in demand model.
-We also study the effect of nonlinearity in the demand model by varying the coefficient $k$.
+We also study the effect of nonlinearity in the demand model by varying the coefficient $\omega$.
 Notice that the nonlinear structure of the demand model here is different from the one in Section 4.1 with the sine component function.
-The results for $p=20$ are provided in Figure 1r.
-The increasing gap in Figure 1q indicates that except for PADR and PO-PA, the performance of all the other methods deteriorates as the nonlinearity of the demand model increases.
-Figure 1r indicates that PADR is slightly better than PO-PA and outperforms the other methods, and the relative gap among them remains approximately stable as $k$ increases.
+The results for $p=20$ are provided in Figure 1x.
+The increasing gap in Figure 1w indicates that except for PADR, NN-DR and PO-PA, the performance of all the other methods deteriorates as the nonlinearity of the demand model increases.
+Figure 1x indicates that PADR and NN-DR are slightly better than PO-PA and outperforms the other methods, and the relative gap among them remains approximately stable as $\omega$ increases.
 […]
-Product placement problem: varying nonlinearity level $k$ (dense model, $p=20$) In (b), with the cost of the cart scaled to 1 and the cost of SIMOPT set to 0, we plot the relative cost of other methods compared to SIMOPT.
+Product placement problem: varying nonlinearity level $\omega$ (dense model, $p=20$) In (b), with the cost of the CART scaled to 1 and the cost of SIMOPT set to 0, we plot the relative cost of other methods compared to SIMOPT.
 5 Conclusion
@@
 In conclusion, the PADR is a favorable hypothesis class for DR-based ERM methods, supported by theoretical consistency, algorithmic convergence, and computational efficiency for solving SP with covariate information.
-Extensions of the PADR-based ERM framework for more complicated SP problems are deserved to be explored in the future, such as problems with chance constraints or endogenous random variables, which necessitate a further investigation of model construction, consistency analysis, and data-driven algorithmic development for problems with more complex composite structures.
+Extensions of the PADR-based ERM framework for more complicated SP problems deserve to be explored in the future, such as problems with chance constraints or endogenous random variables, which necessitate a further investigation of model construction, consistency analysis, and data-driven algorithmic development for problems with more complex composite structures.
+[…]
 Supplementary Material for
@@
 Proof 6.1
+Proof of Proposition 2.1.
+It suffices to justify the bound when $d=1$.
+For $\epsilon>0$ satisfying $(\sqrt{p}\bar{X}/\epsilon)\geq 1$, we first construct a $(2\epsilon/\sqrt{p})$-grid of the cube $\bar{\mathcal{X}}\coloneqq[-\bar{X},\bar{X}]^{p}$, which is also an $\epsilon\text{-cover}$ of $\bar{\mathcal{X}}$ with $\ell_{2}$-norm, denoted by $\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$.
+Let $d_{\min}\coloneqq\underset{\widehat{x}_{1},\widehat{x}_{2}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}}),\,\widehat{x}_{1}\neq\widehat{x}_{2}}{\min}\left\lVert\widehat{x}_{1}-\widehat{x}_{2}\right\rVert=\frac{2\epsilon}{\sqrt{p}}$.
+The cardinality of $\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$ satisfies
+$\left(\frac{\sqrt{p}\bar{X}}{\epsilon}\right)^{p}\leq\left\lvert\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\right\rvert\leq\left(\frac{\sqrt{p}\bar{X}}{\epsilon}+1\right)^{p}\leq\left(\frac{2\sqrt{p}\bar{X}}{\epsilon}\right)^{p}.$
+Given $f_{0}\in\mathcal{F}_{L_{0},M_{0}}$, by McShane (1934, Theorem 1), we can extend $f_{0}$ to an $L_{0}$-Lipschitz function $\bar{f}_{0}:\bar{\mathcal{X}}\rightarrow\mathbb{R}$ satisfying $\left\lVert\bar{f}_{0}\right\rVert_{\infty}\leq M_{0}$ and $\bar{f}_{0}(x)=f_{0}(x)$ for any $x\in\mathcal{X}$.
+Next, we construct a piecewise affine interpolation function to a set of points $\{(\widehat{x},\bar{f}_{0}(\widehat{x}))\mid\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\}$:
 […]
-$\left(\frac{\sqrt{p}\bar{X}}{\epsilon}\right)^{p}\leq\left\lvert\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\right\rvert\leq\left(\frac{\sqrt{p}\bar{X}}{\epsilon}+1\right)^{p}\leq\left(\frac{2\sqrt{p}\bar{X}}{\epsilon}\right)^{p}.$
-[…]
-$f_{\widehat{\mathcal{C}}}(x)\coloneqq$ $\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}-\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\},$ (20)
 where $C\coloneqq L_{0}/d_{\min}=L_{0}\sqrt{p}/(2\epsilon)$.
@@
 Specifically, for any $\widehat{y}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$,
-$\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{C\widehat{x}^{\top}(\widehat{y}-\widehat{x})+\frac{1}{2}C\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$ $=\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}C\left\lVert\widehat{x}-\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
-$\leq\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}L_{0}\left\lVert\widehat{x}-\widehat{y}\right\rVert+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
-$\leq\max_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}\left\{\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}\left\lvert\bar{f}_{0}(\widehat{x})-\bar{f}_{0}(\widehat{y})\right\rvert+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}$
-$\leq\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{y}),$
-where the first inequality holds since $C=L_{0}/d_{\min}\geq L_{0}/\left\lVert\widehat{x}-\widehat{y}\right\rVert$.
-Moreover, this upper bound is attained by taking $\widehat{x}=\widehat{y}$ in the max operator.
-Similarly, the value of the second max-affine component of $f_{\widehat{\mathcal{C}}}$ at $\widehat{y}$ is $\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{y})$.
-Hence, we obtain $f_{\widehat{\mathcal{C}}}(\widehat{y})=\bar{f}_{0}(\widehat{y})$, verifying the interpolation.
-With $K\coloneqq\left\lvert\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})\right\rvert$, for any $\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$, we obtain $\left\lVert C\widehat{x}\right\rVert_{\infty}\leq\frac{L_{0}\sqrt{p}}{2\epsilon}\bar{X}\leq\frac{L_{0}}{2}K^{1/p}$ and
+[…]
+[…]
+[…]
+[…]
+[…]
 $\left\lvert-C\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}C\left\lVert\widehat{x}\right\rVert^{2}\pm\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\rvert\leq\frac{1}{2}Cp\left\lVert\widehat{x}\right\rVert_{\infty}^{2}+\frac{1}{2}M_{0}\leq\frac{1}{4}pL_{0}\bar{X}K^{1/p}+\frac{1}{2}M_{0}.$
@@
 For any $x\in\bar{\mathcal{X}}$, let
-$\begin{array}[]{ll}&\mathcal{A}_{1}(x)\coloneqq\displaystyle{\mathop{\arg\max}_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\},\\ &\mathcal{A}_{2}(x)\coloneqq\displaystyle{\mathop{\arg\max}_{\widehat{x}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})}}\left\{C\widehat{x}^{\top}(x-\widehat{x})+\frac{C}{2}\left\lVert\widehat{x}\right\rVert^{2}-\frac{1}{2}\bar{f}_{0}(\widehat{x})\right\}.\end{array}$
+[…]
 It suffices to show that $\left\lVert C(\widehat{x}_{1}^{*}-\widehat{x}_{2}^{*})\right\rVert\leq(\sqrt{p}+2)L_{0}$ for any $\widehat{x}_{1}^{*}\in\mathcal{A}_{1}(x)$ and $\widehat{x}_{2}^{*}\in\mathcal{A}_{2}(x)$.
@@
 For any $\widehat{y}\in\widehat{\mathcal{C}}_{2}(\epsilon;\bar{\mathcal{X}})$ satisfying $\left\lVert x-\widehat{y}\right\rVert>\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$, we have that
-$C\widehat{y}^{\top}(x-\widehat{y})+\frac{1}{2}C\left\lVert\widehat{y}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{y})$ $=\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}-C\left\lVert x-\widehat{y}\right\rVert^{2}+\bar{f}_{0}(\widehat{y})\right]$
+[…]
 $=\frac{1}{2}\left[\bar{f}_{0}(\widehat{y})-\bar{f}_{0}(x)-C\left\lVert x-\widehat{y}\right\rVert^{2}\right]+\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}+\bar{f}_{0}(x)\right]$
@@
 $\leq\frac{1}{2}\left[\bar{f}_{0}(\widehat{x})-\bar{f}_{0}(x)-C\left\lVert x-\widehat{x}\right\rVert^{2}\right]+\frac{1}{2}\left[C\left\lVert x\right\rVert^{2}+\bar{f}_{0}(x)\right]$
-$=C\widehat{x}^{\top}(x-\widehat{x})+\frac{1}{2}C\left\lVert\widehat{x}\right\rVert^{2}+\frac{1}{2}\bar{f}_{0}(\widehat{x}),$
-where the second inequality is derived by the strict monotonicity of the quadratic function $t\mapsto L_{0}t-\frac{\sqrt{p}L_{0}}{2\epsilon}t^{2}$ on $[\epsilon/\sqrt{p},+\infty)$.
+[…]
+where the second inequality is derived by the strict monotonicity of the quadratic function $t\mapsto L_{0}t-\frac{\sqrt{p}L_{0}}{2\epsilon}t^{2}$ on $\epsilon/\sqrt{p},+\infty)$.
 The above result shows $\left\lVert\widehat{x}_{1}^{*}-x\right\rVert\leq\frac{\sqrt{p}+2}{\sqrt{p}}\epsilon$ for all $\widehat{x}_{1}^{*}\in\mathcal{A}_{1}(x)$.
@@
 Then the empirical Rademacher complexity of $\Phi$ with respect to $\Xi_{N}$ is defined as
-^$\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
+$\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)\coloneqq\mathbb{E}_{\boldsymbol{\sigma}}\left[\sup_{\phi\in\Phi}\frac{1}{N}\sum_{s=1}^{N}\sigma_{s}\phi(u^{s})\right],$
 where ${\boldsymbol{\sigma}}=(\sigma_{1},\dots,\sigma_{N})^{\top}$ is a uniform random vector taking values in $\{-1,+1\}^{N}$.
@@
 $\mathcal{R}_{N}(\Phi)\coloneqq\mathbb{E}_{\Xi_{N}}[\widehat{\mathcal{R}}_{\Xi_{N}}(\Phi)].$
+It captures the richness of the class $\Phi$ by measuring the degree to which a hypothesis set can fit random noise.
+The following lemma quantifies the Rademacher complexity of a class of parametric functions, which is generalized from Ermoliev and Norkin (2013, Lemma B.2) with minor modifications made to accommodate the developments in this paper.
+Lemma 6.3
 […]
-Lemma 6.3
-Let $\widehat{\Phi}\coloneqq\{\phi(\cdot;\theta):\mathcal{U}\rightarrow\mathbb{R}\mid\theta\in\Theta\}$ be a class of functions parameterized by vectors in a compact set $\Theta\subseteq\mathbb{R}^{q}$ with $\max_{\theta\in\Theta}\left\lVert\theta\right\rVert_{\infty}=\mu$.
-Suppose that there exist $M_{\widehat{\Phi}}>0$ and $L_{\widehat{\Phi}}>0$ such that $\max_{\theta\in\Theta}\left\lVert\phi(\cdot;\theta)\right\rVert_{\infty}\leq M_{\widehat{\Phi}}$ and $\left\lVert\phi(\cdot;\theta_{1})-\phi(\cdot;\theta_{2})\right\rVert_{\infty}\leq L_{\widehat{\Phi}}\left\lVert\theta_{1}-\theta_{2}\right\rVert$ for any $\theta_{1},\theta_{2}\in\Theta$.
-Then
 $\mathcal{R}_{N}(\widehat{\Phi})\leq\frac{2\mu L_{\widehat{\Phi}}\sqrt{q}}{N}+M_{\widehat{\Phi}}\sqrt{\frac{q\log N}{N}}.$
 Proof 6.4
-[…]
-^$\widehat{\mathcal{R}}_{\Xi_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
+Proof of Lemma 6.3 Given a sample set $\Xi_{N}=\{u^{s}\}_{s=1}^{N}$ of size $N$, following the proof of Ermoliev and Norkin (2013, Lemma B.2), we derive that for any $\epsilon\in(0,2\mu]$,
+$\widehat{\mathcal{R}}_{\Xi_{N}}(\widehat{\Phi})\leq L_{\widehat{\Phi}}\sqrt{q}\epsilon+M_{\widehat{\Phi}}\sqrt{\frac{2q\log(2\mu/\epsilon)}{N}}.$
 By setting $\epsilon=2\mu/N$ and taking expectation over $\Xi_{N}$, we obtain the desired upper bound.∎
 The following lemma gives the generalization bound based on the Rademacher complexity.
-[…]
-Let $\Phi$ be a function class mapping from $\mathcal{U}$ to $[-M,M]$.
+Lemma 6.5 (Mohri et al. (2018), Theorem 3.3)
+Let $\Phi$ be a function class mapping from $\mathcal{U}$ to [M,M][-M,M].
 Then for any $\delta>0$, with probability at least $1-\delta$ over the draw of an independently and identically distributed (i.i.d.) sample $\Xi_{N}$ of size $N$, the following holds true for all $\phi\in\Phi$:
@@
 where the second inequality follows from Assumption (A2). Moreover, for any $f\in{\mathcal{H}}_{\mu}^{K}$,
-$\left\lvert\varphi(f(x),y)\right\rvert\leq\left\lvert\varphi(z_{0},y)\right\rvert+\left\lvert\varphi(f(x),y)-\varphi(z_{0},y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{\mu}.$ (21)
+$\left\lvert\varphi(f(x),y)\right\rvert\leq\left\lvert\varphi(z_{0},y)\right\rvert+\left\lvert\varphi(f(x),y)-\varphi(z_{0},y)\right\rvert\leq C_{\varphi}+L_{\varphi}\left\lVert f(x)-z_{0}\right\rVert\leq M_{p,d}(\mu).$ (21)
 Therefore, by Lemma 6.3, we deduce
-$\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{4\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}.$ (22)
-Now we prove the uniform generalization bound. By Lemma 6.5 and the inequality (21), with probability at least $21-\delta/$ 2,
-$\max_{\phi\in\Phi_{{\mathcal{H}}_{\mu}^{K}}}\left(\mathbb{E}_{X,Y}\phi(X,Y)-\frac{1}{n}\sum_{s=1}^{n}\phi(x^{s},y^{s})\right)\leq 2{\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
+$\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{4\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}.$ (22)
+Now we prove the uniform generalization bound. By Lemma 6.5 and the inequality (21), with probability at least $1-\delta/2$,
+$\max_{\phi\in\Phi_{{\mathcal{H}}_{\mu}^{K}}}\left(\mathbb{E}_{X,Y}\phi(X,Y)-\frac{1}{n}\sum_{s=1}^{n}\phi(x^{s},y^{s})\right)\leq 2{\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
 For $\Phi_{{\mathcal{H}}_{\mu}^{K}}^{-}\coloneqq\{-\phi:\phi\in\Phi_{{\mathcal{H}}_{\mu}^{K}}\}$, we have ${\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}}^{-})={\mathcal{R}}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})$.
 Therefore, by applying Lemma 6.5 again to $\Phi_{{\mathcal{H}}_{\mu}^{K}}^{-}$ and using the union bound, we derive that with probability at least $1-\delta$,
-$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$ (23)
+$\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert\leq 2\mathcal{R}_{n}(\Phi_{{\mathcal{H}}_{\mu}^{K}})+\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$ (23)
 Finally, by combining (22) and (23), we complete the proof.∎
@@
 By Proposition 2.2, with probability at least $1-\delta$, we have
-$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
+$R(f^{*}_{\Xi,{\mathcal{H}}_{\mu}^{K}})-R(f^{*}_{{\mathcal{H}}_{\mu}^{K}})\leq$ $\ \frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}$
+$\ +2\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
 By combining these two error bounds, we obtain the desired inequality. ∎
@@
 where $I_{1}\in\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)$, $I_{2}\in\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)$, and $I_{3}\in\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$.
-With a alight abuse of notation, we denote $I=(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$.
+With a slight abuse of notation, we denote $I=(I_{1},I_{2},I_{3})\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)\coloneqq\mathcal{I}_{g}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{h}^{\varepsilon}(\bar{\theta},\xi)\times\mathcal{I}_{\varphi_{2}}^{\varepsilon}(f(\bar{\theta},\xi),\xi)$.
 Assumptions (B1) and (B2) hold in this case by construction.
@@
 $\widehat{F}^{\varepsilon}_{\min}(\theta,\xi;\bar{\theta})-F(\theta,\xi)=\frac{1}{2}L_{\nabla\varphi}L_{f}^{2}\left\lVert\theta-\bar{\theta}\right\rVert^{2}-\frac{1}{2}v^{\top}\nabla^{2}\varphi(f(\bar{\theta},\xi),\xi)v+o(\left\lVert v\right\rVert^{2}),$
-[…]
+where $\left\lVert v\right\rVert=\left\lVert f(\theta,\xi)-f(\bar{\theta},\xi)\right\rVert\leq L_{f}\left\lVert\theta-\bar{\theta}\right\rVert$ by the Lipschitz continuity of $f$.
+By dividing the both sides by $\left\lVert\theta-\bar{\theta}\right\rVert^{2}$ and taking $\limsup$ as $\theta\rightarrow\bar{\theta}$ and $\theta\neq\bar{\theta}$, we obtain that $\widehat{\kappa}$ is finite.∎
 8 Details of Convergence Analysis in Section 3.2.3
@@
     There exists $M_{\Delta}>0$ such that $F(\theta^{d})\geq\min_{\theta\in\Theta}F(\theta)+M_{\Delta}$ for any $\theta^{d}\in\mathcal{S}^{d}\backslash\mathcal{S}^{*}$.
-[…]
+The assumption is naturally satisfied when the outer function $\varphi$ of $F$ is in Case 1 or Case 2.
+Specifically, from the proof of Claim 1 afterward, we see that in these two cases, $\Theta$ can be decomposed into finitely many convex regions, on each of which any d-stationary point $\theta^{d}\in\mathcal{S}^{d}$ is a minimizer of $F$ over that region.
+Since the number of such regions is finite, the function $F$ takes finitely many values on $\mathcal{S}^{d}$, verifying Assumption (EC1).
+Discussions on the finite number of d-stationary values for more general nonconvex problems with piecewise and composite structures can be seen in Cui and Pang (2018).
 Proposition 8.2
@@
 8.2 Proof of Proposition 3.6
-[…]
-[…]
+The following lemma from Deng and Gao (2021) is rephrased in our context for the proof of Proposition 3.6.
+Lemma 8.4 (Deng and Gao (2021), Theorem 3.2)
 Suppose that a stochastic function $\phi(\cdot,\xi)$ is convex and Lipschitz continuous on $\Theta$ with modulus $L_{\phi}$ for any $\xi\in\Xi$ and $\bar{\theta}\in\Theta$, and that samples $\Xi_{N}\coloneqq\{\widetilde{\xi}^{s}\}_{s=1}^{N}$ are i.i.d.
@@
 $\leq F(\theta^{\nu})-\mathbb{E}\left[F(\theta^{\nu+1})\mid\mathcal{G}_{\nu-1}\right]+C_{1}\sqrt{\frac{\log N_{\nu}}{N_{\nu}}},$ (27)
-[…]
+where the third inequality follows from inequalities (3.8)-(3.13) in Mohri et al. (2018) that are part of the proof of Lemma 6.5, and the last inequality is derived by Lemma 6.3.
 Next, we bound the term $\mathcal{E}_{2}^{\nu}$.
@@
 Then the second inequality in the proposition can be proved with arguments similar to those above.∎
-8.3 Proofs of Theorems 3.6 and 3.7
-In this section, we prove the convergence results in Theorems 3.6 and 3.7.
+8.3 Proofs of Theorems 3.7 and 3.8
+In this section, we prove the convergence results in Theorems 3.7 and 3.8.
 As in Section 3.2.3, we suppose throughout this section that $\varphi$ is directionally differentiable and locally Lipschitz continuous, $\Theta$ is a compact convex set, and the surrogate family $\{\widehat{F}(\cdot,\xi;\bar{\theta},I)\}_{I\in\mathcal{I}^{\varepsilon}(\bar{\theta},\xi)}$ of $F(\cdot,\xi)$ satisfies Assumptions (B1)-(B4) for each $\xi\in\Xi$.
@@
 Therefore, the statements of the theorem follow from Propositions 3.6 and 8.6.∎
-[…]
+We next present the proof of Proposition 8.6.
+The overall scheme of the error-bound analysis follows the error-bound theory developed in Liu et al.
+(2022b) for the post-convergence analysis of the SMM algorithm.
+To justify the global error bound in the second statement of Proposition 8.6, we first give the error bounds of distance to $\mathcal{S}^{d}$ in terms of the residual $r_{d}$ in Proposition 8.8 and then relate $r_{d}$ with $r^{\varepsilon,\rho}$ in Proposition 8.10.
 Proposition 8.8
@@
 For problem (10) with $\varepsilon\geq 0$ and $\rho>0$, it holds that for any $\theta\in\Theta$, there exists $\theta^{\prime}\in\Theta$ such that $\left\lVert\theta^{\prime}-\theta\right\rVert\leq 2\,r^{\varepsilon,\rho}(\theta)$ and $r_{d}(\theta^{\prime})\leq\left(2\rho+\frac{5}{2}\kappa\right)\,r^{\varepsilon,\rho}(\theta)$.
-[…]
+A similar result to Proposition 8.10 is provided in Liu et al.
+(2022b, Lemma 4) under the assumption that the relevant surrogate function satisfies Condition (P4).
+With the surrogate function $\widehat{F}^{\varepsilon}_{\min}(\cdot;\bar{\theta})$ associated with $r^{\varepsilon,\rho}(\bar{\theta})$ satisfying (P4), we can prove Proposition 8.10 following the lines of Liu et al.
+(2022b, Lemma 4), and thus the details are omitted here.
 Proof 8.11
@@
 $\geq F(\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t}))+\frac{{\rho}}{2}\,\big\lVert\,\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})-\theta_{t}\,\big\rVert^{2}.$ (30)
-[…]
+Since $\lim_{t(\in\mathcal{K})\rightarrow\infty}r^{{\varepsilon},{\rho}}(\theta_{t})=0$, we deduce $\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{P}^{\rho}_{\mathbf{I}^{*}_{t}}(\theta_{t})=\theta_{\infty}$.
+Additionally, for a fixed index combination $\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})$, the surrogate function $\widehat{F}(\cdot;\cdot,\mathbf{I})$ is continuous on $\Theta\times\Theta$ by assumption, and thus by Shapiro et al.
+(2021b, Theorem 7.23), $\widehat{P}^{\rho}_{\mathbf{I}}(\cdot)$ is continuous at $\theta_{\infty}$.
+It follows from (8.11) that
 $F(\theta_{\infty})\leq\lim_{t(\in\mathcal{K})\rightarrow\infty}\widehat{F}^{*}(\theta_{t})=\min_{\mathbf{I}\in\mathcal{I}^{\varepsilon^{\prime}}(\theta_{\infty})}\left\{\widehat{F}(\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{\infty});\theta_{\infty},\mathbf{I})+\frac{{\rho}}{2}\,\big\lVert\,\widehat{P}^{\rho}_{\mathbf{I}}(\theta_{\infty})-\theta_{\infty}\,\big\rVert^{2}\right\},$
@@
 Condition (P4) originating from Liu et al.
-(2022) is a technical condition for the nonasymptotic convergence analysis of the ESMM algorithm.
+(2022a) is a technical condition for the nonasymptotic convergence analysis of the ESMM algorithm.
 Note that Condition (P4) is verified for Case 1-Case 3 in Online Appendix 7.
-Assumption (B6) is generalized from Drusvyatskiy and Lewis (2018).
+Assumption (B6) is generalized from Drusvyatskiy and Lewis (2018a).
 The condition holds, for example, when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope, which is verified in Online Appendix 8.5.
@@
 We verify that Assumption (B6) holds when the outer function $\varphi$ is strongly convex and $\Theta$ is a polytope by proving a stronger result in Claim 1 with the following lemma.
-[…]
+Lemma 8.17 (Cui et al. (2018b), Proposition 2 (ii))
 Let $\Theta\subseteq\mathbb{R}^{q}$ be a closed convex set.
@@
 Proof of Claim 1.
-For $\theta\in\mathrm{cl}(\mathcal{S}^{d})$, the claim holds clearly with $\theta^{d}=\theta$.
-For $\theta\in\Theta\backslash\mathrm{cl}(\mathcal{S}^{d})$, it suffices to consider the case with $n=1$, for which we denote $F(\theta)=\varphi(f(\theta))$, where $f:\mathbb{R}^{q}\rightarrow\mathbb{R}^{d}$ is piecewise linear and $\varphi$ is strongly convex.
-The polytope $\Theta$ can be partitioned into a finite number of $q$-polytopes, denoted by $\mathfrak{C}\coloneqq\{\mathcal{C}_{i}\}_{i=1}^{N}$, such that $f$ is linear on each of them and the intersection of any two polytopes in $\mathfrak{C}$ is their common face.
-Given $\mathcal{C}\in\mathfrak{C}$, $f$ restricted to $\mathcal{C}$ can be represented as $f(\theta)=A\theta$ for some $A\in\mathbb{R}^{d\times q}$.
-Since the strongly convex function $\varphi$ has a unique minimizer on $\{A\theta\in\mathbb{R}^{d}\mid\theta\in\mathcal{C}\}$, the set of minimizers of $F$ on $\mathcal{C}$ denoted by $\mathcal{S}^{*}_{\mathcal{C}}$ can be represented as $\mathcal{S}^{*}_{\mathcal{C}}=[\{\theta^{\prime}\}+\mathcal{N}_{A}]\bigcap\mathcal{C}$ for any $\theta^{\prime}\in\mathcal{S}^{*}_{\mathcal{C}}$, where $\mathcal{N}_{A}$ denotes the null space of $A$.
+[…]
 We first show that $\mathcal{S}^{*}_{\mathcal{C}}=\mathrm{cl}(\mathcal{S}^{d})\bigcap\mathcal{C}$ if $\mathcal{S}^{d}\bigcap\mathcal{C}\neq\emptyset$.
@@
 We present the numerical results on the Newsvendor problem with convex and nonconvex constraints in Section 9.3.
+The optimization problem for learning the optimal decision rule can be formulated as
+$\underset{\displaystyle\begin{subarray}{c}f\in\mathcal{F}\end{subarray}}{\mathrm{inf}}\quad\mathbb{E}_{X,Y}\left[\varphi(f(X),Y)\right]\hfil\hfil\hfil\hfil$ (32)
 […]
+Constraints bring up the feasibility issue for data-driven decision rule methods.
+We show that in the following three special cases, probabilistic constraints can be equivalently converted into deterministic ones or embedded into the structure of PADR, with which we can construct the PADR-based ERM problem while retaining the feasibility of the obtained PADR:
   1. (C1)
@@
   2. (C2)
-    Box constraints: when deterministic constraints lead to a box-constrained set $\{z\in\mathbb{R}^{d}:\left\lVert z\right\rVert_{\infty}\leq M\}$, a truncated PADR class ¯${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ can be used, where ${\mathcal{H}}_{\mu}^{K}$ represents the PA hypothesis class, and the max/min are element-wise operators.
+    Box constraints: when deterministic constraints lead to a box-constrained set $\{z\in\mathbb{R}^{d}:\left\lVert z\right\rVert_{\infty}\leq M\}$, a truncated PADR class ${\bar{\mathcal{H}}}_{\mu}^{K}\coloneqq\{\bar{f}\mid\bar{f}=\max\{\min\{f,M\},-M\},\,f\in{\mathcal{H}}_{\mu}^{K}\}$ can be used, where ${\mathcal{H}}_{\mu}^{K}$ represents the PA hypothesis class, and the max/min are element-wise operators.
 This ensures feasibility of the ERM solution.
 Following the excess risk decomposition in Section 2.2, it is straightforward to see that the PADR-based ERM model for the three special cases (C1) and (C2) is nonasymptotically consistent with the original SP problem (32).
+For Case (C1), problems with finitely many scenarios can be addressed using the exact penalization formulation (37) developed in Section 9.2, and the feasibility of the solution is guaranteed by Lemma 9.7.
+For the case with general constraints, it is difficult to convert them into deterministic ones, so we impose constraints only on finite scenarios in $\Xi$ and we formulate a restricted PADR-based ERM problem with an additional constant $\gamma>0$ as below,
+$\underset{\displaystyle\begin{subarray}{c}f\in{\mathcal{H}}_{\mu}^{K}\end{subarray}}{\mathrm{min}}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi({f(x^{s})},{y^{s}})\hfil\hfil\hfil\hfil$ (33)
 […]
+Let ${\widetilde{\mathcal{H}}}^{K}_{\mu,\gamma}\coloneqq{\mathcal{H}}_{\mu}^{K}\bigcap\left\{f\in\mathcal{F}\mid\psi_{j}(f(x^{s}))+\gamma\leq 0,\,s\in[n],\,j\in[J]\right\}$ , and let $f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ denote an optimal PADR to the constrained ERM problem (33).
+The idea of imposing the restricted constraints on finite scenarios follows the scenario approximation method in Calafiore and Campi (2005).
+This ensures that the obtained PADR $f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ at a new scenario is feasible to the original constraint set with high probability according to Luedtke and Ahmed (2008, Theorem 10).
+The result is provided in the following proposition.
+Proposition 9.1
+Suppose that the support set $\mathcal{X}\subseteq\mathbb{R}^{p}$ of $X$ is a compact set with $\max_{x\in\mathcal{X}}\left\lVert x\right\rVert_{\infty}=\bar{X}$, and that $\psi_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L_{\psi}$-Lipschitz continuous for every $j\in[J]$.
+Then, with probability at least $1-\delta$ over the draw of $\Xi$, we have:
+$\mathbb{P}_{X}\left[\max_{j\in[J]}\left\{\psi_{j}\left(f^{*}_{\Xi,\widetilde{\mathcal{H}}_{\mu,\gamma}^{K}}(X)\right)\right\}>0\right]<$ $\ \sqrt{\frac{4dK(p+1)}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil\right)}$
+$+\sqrt{\frac{\log(2n)}{n}}+\sqrt{\frac{2\log(1/\delta)}{n}}.$
+Proof 9.2
+Proof of Proposition 9.1 Denote $G(\theta,x)\coloneqq\max_{j\in[J]}\psi_{j}(f(\theta,x))$, where $\theta\in[-\mu,\mu]^{2dK(p+1)}$ is the parameter of PADR $f\in{\mathcal{H}}_{\mu}^{K}$.
+Since each $\psi_{j}$ for $j\in[J]$ is $L_{\psi}$-Lipschitz continuous, we have that for any $x\in\mathcal{X}$ and $\theta_{1},\theta_{2}\in[-\mu,\mu]^{2dK(p+1)}$,
+$\left\lvert G(\theta_{1},x)-G(\theta_{2},x)\right\rvert$ $\leq L_{\psi}\left\lVert f(\theta_{1},x)-f(\theta_{2},x)\right\rVert$
+$\leq L_{\psi}\sqrt{d}\left\lVert f(\theta_{1},x)-f(\theta_{2},x)\right\rVert_{\infty}$
+$\leq 2L_{\psi}\sqrt{d(p\bar{X}^{2}+1)(p+1)}\cdot\left\lVert\theta_{1}-\theta_{2}\right\rVert_{\infty}.$
+Define the sets
+$\Theta^{\Xi}_{\gamma}\coloneqq\left\{\theta\in[-\mu,\mu]^{2dK(p+1)}\mid G(\theta,x^{s})\leq 0,\ s=1,\dots,n\right\},$
+$\Theta_{\epsilon}\coloneqq\left\{\theta\in[-\mu,\mu]^{2dK(p+1)}\mid\mathbb{P}_{X}(G(\theta,X)\leq 0)\geq 1-\epsilon\right\}.$
+By Luedtke and Ahmed (2008, Theorem 10), it follows that
+$\mathbb{P}_{\Xi}\left(\Theta^{\Xi}_{\gamma}\subseteq\Theta_{\epsilon}\right)\geq 1-\left\lceil\frac{2}{\epsilon}\right\rceil\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil^{2dK(p+1)}e^{-n\epsilon^{2}/2}.$
+With $\epsilon\in(0,1]$, let
+$M\coloneqq\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil^{2dK(p+1)},\ \delta\coloneqq\left\lceil\frac{2}{\epsilon}\right\rceil Me^{-n\epsilon^{2}/2}\geq\frac{2}{\epsilon}Me^{-n\epsilon^{2}/2}.$
+Without loss of generality, we can assume that $L_{\psi}\geq\gamma$, which leads to $M>2$ and $\log(\delta/(2M))<\log(1/4)<-1$. Therefore, we have that
+$\frac{\delta}{2M}\geq\frac{1}{\epsilon}e^{-\frac{n}{2}\epsilon^{2}}>e^{(\log(\frac{\delta}{2M})+1)\frac{n}{2}\epsilon^{2}}\cdot e^{-\frac{n}{2}\epsilon^{2}}=e^{\log(\frac{\delta}{2M})\frac{n}{2}\epsilon^{2}},$
+implying $\epsilon>\sqrt{2/n}$.
+Furthermore, denote $\delta^{\prime}\coloneqq Me^{-n\epsilon^{2}/2}$, it follows that with probability at least $1-\delta$ over $\Xi$, the solution $f^{*}_{\Xi,\widetilde{\mathcal{H}}_{\mu,\gamma}^{K}}$ satisfies:
+$\mathbb{P}_{X}\left[\max_{j\in[J]}\left\{\psi_{j}\left(f^{*}_{\Xi,\widetilde{\mathcal{H}}_{\mu,\gamma}^{K}}(X)\right)\right\}>0\right]$
+$\leq\sqrt{\frac{2}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil^{2dK(p+1)}/\delta^{\prime}\right)}$
+$\leq\sqrt{\frac{4dK(p+1)}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil\right)}+\sqrt{\frac{2\log(1/\delta^{\prime})}{n}}$
+$\leq\sqrt{\frac{4dK(p+1)}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil\right)}+\sqrt{\frac{2\log(1/\delta)}{n}}+\sqrt{\frac{2\log(2/\epsilon)}{n}}$
+$<\sqrt{\frac{4dK(p+1)}{n}\log\left(\left\lceil\frac{8L_{\psi}}{\gamma}\sqrt{2K(p\bar{X}^{2}+1)}\cdot d(p+1)\mu\right\rceil\right)}+\sqrt{\frac{2\log(1/\delta)}{n}}+\sqrt{\frac{\log(2n)}{n}},$
+which completes the proof. ∎
+It can be observed that the out-of-sample feasibility of the PADR-based ERM solution exhibits the same order of complexity $\widetilde{\mathcal{O}}\left(\sqrt{\frac{4dK(p+1)}{n}}\right)$ in $n$ and $K$ as the generalization bound in the unconstrained case.
+Similar to the excess risk result for unconstrained problems, with a larger piece number $K$, model misspecification reduces but the likelihood of out-of-sample infeasibility simultaneously increases.
 We next show that the PADR-based ERM model (33) is asymptotically consistent with (32).
@@
     For every $j\in[J]$, $\psi_{j}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L_{\psi}$-Lipschitz continuous.
+By strengthening the constraints by a constant parameter $\gamma_{0}>0$, we obtain a restricted SP problem as follows:
+$\underset{\displaystyle\begin{subarray}{c}f\in\mathcal{F}\end{subarray}}{\mathrm{inf}}\quad R(f)\coloneqq\mathbb{E}_{X,Y}\left[\varphi(f(X),Y)\right]\hfil\hfil\hfil\hfil$ (34)
 […]
+for which we consider the following assumption, as an adaptation of Assumption (A4) for constrained problems:
   1. (A4)
-    Given $\gamma_{0}>0$, there exists a Lipschitz continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (9.1).
+    Given $\gamma_{0}>0$, there exists a Lipschitz continuous function $f^{*}_{\gamma_{0}}$ that is optimal to problem (34).
 By Assumptions (A2) and (A4), $f^{*}_{\gamma_{0}}\in\mathcal{F}_{L_{z}^{\prime},M_{z}^{\prime}}$ for some $L_{z}^{\prime}>0$ and $M_{z}^{\prime}>0$.
-We first provide the excess risk bound of the restricted ERM model (9) to the restricted SP problem (9.1).
+We first provide the excess risk bound of the restricted ERM model (33) to the restricted SP problem (34).
 Proposition 9.3 (Excess risk bound of the restricted constrained ERM)
 Suppose that Assumptions (A1)-(A3) and (A5) hold, and that Assumption (A4) holds for some $\gamma_{0}>\Gamma\coloneqq 2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\psi}L_{z}^{\prime}\bar{X}K^{-1/p}$.
-With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{\mu}\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
+With $\gamma\in(0,\gamma_{0}-\Gamma)$, $\mu\geq\max\{\frac{1}{2}L_{z}^{\prime}K^{1/p},\frac{1}{4}pL_{z}^{\prime}\bar{X}K^{1/p}+\frac{1}{2}M_{z}^{\prime}\}$, and $M_{p,d}(\mu)\coloneqq C_{\varphi}+L_{\varphi}\left(2\sqrt{d}\mu(p\bar{X}+1)+\left\lVert z_{0}\right\rVert\right)$, the following holds with probability at least $1-\delta$ over the draw of $\Xi$,
 $R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})\leq$ $R(f^{*}_{\gamma_{0}})+2d^{1/2}(p^{1/2}+3)p^{1/2}L_{\varphi}L_{z}^{\prime}\bar{X}K^{-1/p}$
-$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
-Proof 9.2
+$+\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
+Proof 9.4
+Proof of Proposition 9.3. Denote by $f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}}$ the optimal solution to the following problem
+$\underset{\displaystyle\begin{subarray}{c}f\in{\mathcal{H}}_{\mu}^{K}\end{subarray}}{\mathrm{min}}\quad\mathbb{E}_{X,Y}\left[\varphi(f(X),Y)\right]\hfil\hfil\hfil\hfil$ (35)
 […]
+Similar to analysis for unconstrained problems, we deduce the following decomposition:
 $R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$ $=R(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-\widehat{R}_{\Xi}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+\widehat{R}_{\Xi}(f^{*}_{\Xi,\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})+R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})$
@@
 $\leq\underbrace{2\max_{f\in{\mathcal{H}}_{\mu}^{K}}\left\lvert R(f)-\widehat{R}_{\Xi}(f)\right\rvert}_{\mathcal{E}_{1}}+\underbrace{R(f^{*}_{\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}})-R(f^{*}_{\gamma_{0}})}_{\mathcal{E}_{2}},$
-where the last inequality holds since ~$\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}\subseteq{\mathcal{H}}_{\mu}^{K}$.
+where the last inequality holds since $\widetilde{\mathcal{H}}^{K}_{\mu,\gamma}\subseteq{\mathcal{H}}_{\mu}^{K}$.
 The first term $\mathcal{E}_{1}$ can be directly bounded by Proposition 2.2, that is, given $\delta>0$, with probability at least $1-\delta$ over the draw of $\Xi$,
-$\mathcal{E}_{1}\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{\mu}\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{\mu}\sqrt{\frac{\log(2/\delta)}{n}}.$
+$\mathcal{E}_{1}\leq\frac{16\mu L_{\varphi}\sqrt{2dK(p+1)(p\bar{X}^{2}+1)}}{n}+4M_{p,d}(\mu)\sqrt{\frac{2dK(p+1)\log n}{n}}+2\sqrt{2}M_{p,d}(\mu)\sqrt{\frac{\log(2/\delta)}{n}}.$
 For the second term $\mathcal{E}_{2}$, we further decompose it by
@@
 By combining bounds of $\mathcal{E}_{1}$, $\mathcal{E}_{21}$, and $\mathcal{E}_{22}$, we derive the desired bound.∎
-With the excess risk bound in Proposition 9.1, and the probabilistic feasibility guarantee according to Luedtke and Ahmed (2008, Theorem 10), we obtain the asymptotic consistency of restricted PADR-based ERM model to the original constrained SP (9).
+With the excess risk bound in Proposition 9.3, and the probabilistic feasibility guarantee according to Proposition 9.1, we obtain the asymptotic consistency of restricted PADR-based ERM model to the original constrained SP (32).
 Theorem 9.5 (Asymptotic consistency of constrained PADR-based ERM)
@@
 9.2 Convergence of ESMM to B-stationarity
+The constrained ERM problem (33)can be reformulated as follows with $\widetilde{\xi}_{n}$ representing a random variable uniformly distributed on $\Xi$:
 […]
+[…]
+Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ are Bouligand-differentiable (B-differentiable), i.e., they are locally Lipschitz continuous and directionally differentiable on $\mathbb{R}^{d}$, and that $\Theta$ is a compact convex set.
+Since $G_{\gamma}$ is nonconvex, we are interested in computing Bouligand-stationary (B-stationary) points of (36), the defining condition of which is an extension of d-stationarity to nonconvex-constrained optimization problems.
+With $\mathcal{I}>0$, we construct the following penalized problem
+$\underset{\displaystyle\begin{subarray}{c}\theta\in\Theta\end{subarray}}{\mathrm{min}}\quad V(\theta;\lambda)\coloneqq F(\theta)+\lambda G_{\gamma}(\theta).\hfil\hfil\hfil\hfil$ (37)
+According to Cui and Pang (2021a, Proposition 9.2.2), problem (33) with some regularity conditions has the exact penalty property such that with a finite penalty parameter $\lambda$, any d-stationary point of (37) is a B-stationary point of (36).
+Since $\max\{\cdot,0\}$ is nondecreasing and convex, it is clear that the upper surrogate functions of $G_{\gamma}(\cdot)$ can be constructed by plugging surrogates of $\{\psi_{j}(f(\cdot,\xi))\}_{j\in[J]}$ into $\max\{\cdot,0\}$, maintaining the corresponding surrogation properties.
+Therefore, with the exact penalty property and Assumptions (B1)-(B6) imposed in the context of problem (37), the ESMM algorithm is ready to solve the penalized problem (37) with convergence to B-stationary points of (36).
+According to (Cui and Pang 2021b), for a constrained problem with a B-differentiable objective, a B-stationary point is defined as follows.
 Definition 9.6
@@
 The B-stationarity coincides with d-stationarity if $\bar{\Theta}$ is convex.
-Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (9.2) are B-differentiable.
-With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (9.2) by $\mathcal{S}^{b}$.
-[…]
-[…]
+Suppose that functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ in problem (36) are B-differentiable.
+With $\bar{\Theta}=\left\{\theta\in\Theta\mid G_{\gamma}(\theta)\leq 0\right\}$, we denote the set of B-stationary points of (36) by $\mathcal{S}^{b}$.
+With Assumptions (B1)-(B6) imposed in the context of the penalized problem (37), we can apply the ESMM algorithm to solve (37) and obtain the same convergence results in terms of d-stationarity as in Theorems 3.7-8.12.
+To further obtain the convergence to B-stationarity, we intend to connect d-stationarity of (37) with B-stationarity of (36) by applying the following exact penalty result from Cui and Pang (2021b).
+Note that by the finiteness of $\Xi$ and compactness of $\Theta$, for any $\xi\in\Xi$, $f(\cdot,\xi)$ is Lipschitz continuous with modulus $L_{f}\coloneqq 2\sqrt{\max_{s\in[n]}\left\lVert x^{s}\right\rVert^{2}+1}$ and uniformly bounded with $\left\lVert f(\cdot,\xi)\right\rVert_{\infty}\leq M_{f}\coloneqq L_{f}\max_{\theta\in\Theta}\left\lVert\theta\right\rVert$, and $\varphi(\cdot,\xi)$ is Lipschitz continuous on $M_{f},M_{f}]^{d}$ with uniform modulus denoted by $L^{\prime}_{\varphi}$.
+Lemma 9.7 (Cui and Pang (2021b), Proposition 9.2.2)
 Let $\Theta$ be a compact convex set. For $\gamma>0$, if there exists $l_{c}>0$ such that
 $\sup_{\theta\in\Theta\setminus\bar{\Theta}}\left[\min_{v\in\mathcal{T}(\theta;\Theta);\left\lVert v\right\rVert\leq 1}G_{\gamma}^{\prime}(\theta;v)\right]\leq-l_{c},$ (38)
-then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (9.2) is feasible to (9.2) and thus is a B-stationary point of (9.2).
-Condition (32) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (9.2) are bounded away from 0.
-Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (9.2) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (9.2).
+then with $\lambda>L^{\prime}_{\varphi}L_{f}/l_{c}$, any d-stationary point of (37) is feasible to (36) and thus is a B-stationary point of (36).
+Condition (38) requires that the maximum descent rates of $G_{\gamma}$ at points infeasible to (36) are bounded away from 0.
+Thus by the Lipschitz continuity of $F$, at any infeasible point, we can always find a descent direction in the penalized problem (37) with a sufficiently large (but still finite) penalty parameter $\lambda$, ensuring that the point can not be a d-stationary point of (37).
 Nonetheless, it is not easy to verify the technical condition (38) when $G_{\gamma}$ is in a composite structure.
@@
 Then condition (38) holds with $l_{c}=\frac{c_{\psi}l_{\psi}}{n\sqrt{q}\mu}$.
-Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (9.2) is a B-stationary point of problem (9.2).
-Proof 9.7
+Moreover, with $\lambda>\bar{\lambda}\coloneqq\frac{n\sqrt{q}\mu L^{\prime}_{\varphi}L_{f}}{c_{\psi}l_{\psi}}$, any d-stationary point of (37) is a B-stationary point of problem (36).
+Proof 9.9
 Proof of Proposition 9.8.
@@
 Let
-$\mathcal{A}_{<}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma<0\},\ \mathcal{A}_{>}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma>0\},\text{ and }\mathcal{A}_{=}\coloneqq\{j\in[J]:\psi_{j}(\bar{z})+\gamma=0\}.$
+[…]
 For $j\in\mathcal{A}_{<}$, it is clear that $G_{\gamma}(\cdot,\xi;j)^{\prime}(\bar{\theta};\bar{v})=0$. By Assumption (B7), we have that for $j\in\mathcal{A}_{>}$,
@@
 and that for $j\in\mathcal{A}_{=}$,
-[…]
+$G_{\gamma}(\cdot,\xi;j)^{\prime}(\bar{\theta};\bar{v})=\max\big\{\psi_{j}^{\prime}(f(\bar{\theta},\xi);f(\cdot,\xi)^{\prime}(\bar{\theta};\bar{v})),\,0\big\}=\max\left\{\psi_{j}^{\prime}\left(\bar{z};-\frac{\bar{z}}{\left\lVert\bar{\theta}\right\rVert}\right),0\right\}=0.$
 Since $\bar{j}\in A_{>}$, we obtain that
@@
 Theorem 9.10 (Convergence of ESMM with exact penalization)
-Under Assumptions (B1) to (B4) and the same settings of Proposition 9.6, for problem (9.2) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$.
+Under Assumptions (B1) to (B4) and the same settings of Proposition 9.8, for problem (37) with $\lambda\geq\bar{\lambda}$, let $\{\widetilde{\theta}^{\,t}\}_{t}$ be a sequence of points produced by Algorithm 1 with $\varepsilon\geq 0$, $\eta\geq 0$, and $N_{\nu}=\mathcal{O}(\nu^{\alpha})$ for some $\alpha>0$.
 The following statements hold:
   1. (a)
-    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (9.2) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (9.2).
+    With $\varepsilon>0$, any limit point of $\{\widetilde{\theta}^{t}\}_{t}$ is a composite $(\varepsilon^{\prime},\rho)$-strongly d-stationary point of (37) with probability 1 for any $\varepsilon^{\prime}\in[0,\varepsilon)$ and $\rho\in(\eta,+\infty)$, and thus a B-stationary point of (36).
   2. (b)
@@
     With $\varepsilon=0$, functions $\varphi$ and $\{\psi_{j}\}_{j\in[J]}$ being piecewise affine, and $\Theta$ being a polytope, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\widetilde{\mathcal{O}}(T^{-\frac{1}{2}\min\{\frac{\alpha}{2},1\}})$; moreover, with $\eta>0$, $\mathbb{E}\,[\mathrm{dist}(\widetilde{\theta}^{T},\mathcal{S}^{b})]=\mathcal{O}(T^{-\frac{1}{2}\min\{\alpha,1\}})$, which is up to a logarithmic factor when $\alpha=1$.
-For the ESMM algorithm with exact penalization, the limit point of generated sequence is a composite strong d-stationarity to (9.2) with probability 1 according to Theorem 9.8, and thus is a global minimizer of (9.2) when $\varepsilon$ is sufficiently large and $\eta$ is close to 0, by the equivalence between composite strong d-stationarity and global optimality in Proposition 8.2 with appropriate assumptions when needed.
-Furthermore, such a global minimizer of the penalized problem (9.2) should be also a global minimizer of (9.2) since it is feasible according to Proposition 9.6.
+For the ESMM algorithm with exact penalization, the limit point of generated sequence is a composite strong d-stationarity to (37) with probability 1 according to Theorem 9.10, and thus is a global minimizer of (37) when $\varepsilon$ is sufficiently large and $\eta$ is close to 0, by the equivalence between composite strong d-stationarity and global optimality in Proposition 8.2 with appropriate assumptions when needed.
+Furthermore, such a global minimizer of the penalized problem (37) should be also a global minimizer of (36) since it is feasible according to Proposition 9.8.
 Under such settings, the asymptotic consistency of the PADR framework for the constrained case in Theorem 9.5 turns into asymptotic optimality.
@@
 $\overline{Y}_{1}(X)=15X_{1}-5X_{2}+30,\quad\overline{Y}_{2}(X)=15X_{1}+5X_{2}+30.$
-In this subsection, we compare PADR with prescriptive methods and the penalized RKHS-DR method from Bertsimas and Koduri (2021, Section 6.2).
+In this subsection, we implement the prescriptive methods, which ensure feasibility by solving the individual constrained problem for each sample.
+Additionally, we implement NN-DR equipped with $\ell_{1}$-penalization and RKHS-DR methods following the penalized formulation proposed by Bertsimas and Koduri (2021, Section 6.2).
 9.3.1 Newsvendor problem with a linear constraint.
@@
 In this subsection, we set $(c_{b1},c_{h1},c_{b2},c_{h2})=(8,2,2,8)$, $C_{0}=60$, such that the probability that the constraint is active at the optimal decision is around 0.5.
-Since DR-based methods may yield infeasible decisions, in Figure 1u we report the test cost of the order decisions, projected onto the convex feasible region when necessary, along with the feasible frequency of the original decisions provided by PADR and RKHS-DR.
+Since DR-based methods (i.e., PADR, NN-DR, and RKHS-DR) may yield infeasible decisions, in Figure 1aa we report the test cost of the order decisions, projected onto the convex feasible region when necessary, along with the feasible frequency of their original decisions.
 […]
-(s) Basic results
-(t) Sparse demand model
-(u) Dense demand model
+(y) Basic results
+(z) Sparse demand model
+($a$) Dense demand model
 Newsvendor problem with a linear constraint
-With the basic demand model, PADR converges to SIMOPT and attains 100% feasibility when $n\geq 200$.
-In comparison, RKHS-DR and prescriptive methods display notable gaps to SIMOPT, with the feasible frequency of RKHS-DR being only around 80%.
-Figures 1t and 1u show the numerical results with varying feature dimensions.
-Similar to the unconstrained cases, PADR outperforms the other methods in the sparse case when $p<20$ and in the dense case when $p<50$.
-Notably, with penalization, PADR achieve 100% feasibility when $p<20$.
-Although PADR can not guarantee 100% feasibility of the decision compared to prescriptive methods, it still provides satisfactory performance via projection.
+Figure 1y shows that with the basic demand model, PADR approaches SIMOPT and achieves 100% feasibility when $n\geq 200$.
+In contrast, other benchmarks exhibit noticeable performance gaps relative to SIMOPT.
+Particularly, RKHS-DR achieves roughly 80% feasibility and NN-DR exhibits unstable feasibility performance.
+Figures 1z and 1aa present the results under varying feature dimensions.
+Consistent with the unconstrained setting, PADR outperforms other methods in the sparse model when $p<20$ and in the dense model when $p<50$.
+Notably, by the utilization of $\ell_{1}$ penalization, PADR achieves 100% feasibility when $p<20$.
+Moreover, in the dense demand model, PADR exhibits the clear edge over prescriptive methods by achieving better empirical performance on test costs, while it fails to be feasible at 10% - 20% probability.
 9.3.2 Newsvendor problem with a nonconvex constraint.
@@
 We set $(c_{b1},c_{h1},c_{b2},c_{h2})=(7,7,3,3)$ and $C_{0}=50$ with the same probability for the constraint being active as the experiment in Section 9.3.1.
-Since the projection of infeasible decisions to the nonconvex feasible region may be far away to the optimal solutions, we only report the average test cost of feasible decisions in Figure 1x, with a careful hyperparameter selection to ensure high feasible frequencies.
+Since the projection of infeasible decisions to the nonconvex feasible region may be far away to the optimal solutions, we only report the average test cost of feasible decisions in Figure 1ad, with a careful hyperparameter selection to ensure high feasible frequencies.
 We use the standard MM algorithm to solve the penalized ERM problem of RKHS-DR, referred to here as RKHS-DR(MM).
@@
 The SOF method in Kallus and Mao (2023)is not suitable for this nonconvex-constrained problem, and it may require the study of the second-order perturbation analysis for stochastic optimization, which is beyond the scope of this paper.
+For the other prescriptive methods, we solve two-dimensional nonconvex constrained problems by grid search.
 For the problem with the basic demand model, both PADR and RKHS-DR attain 100% feasibility on the test set.
 PADR converges to SIMOPT with significantly fewer training samples than other benchmarks.
-Figures 1w and 1x show the results across varying feature dimensions.
+Figures 1ac and 1ad show the results across varying feature dimensions.
 PADR outperforms the other methods in both sparse and dense cases when $p\leq 30$, achieving 90 % $\sim$ 100% feasibility in the sparse case when $p\leq 20$ and in the dense case when $p\leq 30$.
@@
 The relatively small gap of PADR to SIMOPT shows PADR’s capability to learn the underlying decision rule structure in problems with nonconvex constraints while maintaining a high probability of feasibility when $p\leq 30$.
+NN-DR obtains relatively lower test costs than PADR and RKHS-DR; however, its feasibility rate is substantially lower than those of the other two methods.
+This suggests that the decision rule yielded by NN-DR may face difficulties in effectively addressing nonconvex-constrained optimization problems.
 […]
-(v) Basic results
-(w) Sparse demand model
-(x) Dense demand model
+(ab) Basic results
+(ac) Sparse demand model
+(ad) Dense demand model
 Newsvendor problem with a nonconvex constraint
+Thanks to the Associate Editor who suggested a hybrid idea during the review process to address the feasibility issue of DR-based methods, we implement a hybrid strategy by combining the decision rule with an auxiliary method that ensures feasibility (such as RF).
+Specifically, for the Newsvendor problem with a nonconvex constraint, we combine PADR and NN-DR with the prescriptive method RF, referred to as PADR(hybrid) and NN-DR(hybrid).
+Each method applies the corresponding decision rule first; if the output is infeasible, the RF solution is used as a fallback.
+We compare the performance of these hybrid methods with pure RF and CART methods, as shown in Figure 1af.
+Both PADR(hybrid) and NN-DR(hybrid) consistently outperform RF, with the exception when $p=50$.
+Under the sparse demand model, NN-DR(hybrid) outperforms PADR(hybrid) and shows greater improvements over RF.
+Under the dense demand model, the performance of both hybrid methods is close to SIMOPT, with PADR(hybrid) slightly better.
+Overall, these preliminary results demonstrate the promise of hybrid approaches in improving both feasibility and performance.
+This is especially beneficial for contextual SP problems with complex, nonconvex constraints, and we believe this avenue deserves further investigation.
 […]
+(ae) Sparse demand model
+(af) Dense demand model
+Newsvendor problem with a nonconvex constraint: hybrid approaches
+10 Examples Satisfying or Violating Assumption (A4)
+In this section, we present several examples in which the SP problems either satisfy or violate Assumption (A4).
+We conclude with a brief discussion on a potential analytical framework to establish sufficient conditions for the Lipschitz continuity of the optimal decision rule.
+10.1 Problems satisfying Assumption (A4)
+A canonical class of problems ensuring Assumption (A4) is quantile regression, such as the Newsvendor problem, which in fact served as the motivation for the Lipschitz continuity assumption in our paper.
+Additionally, as extensively discussed in Shapiro et al.
+(2021a, Chapter 6), risk-averse optimization also yields quantile-based solutions.
+For example, Average Value-at-Risk (AV@R) attains its minimum at the Value-at-Risk (V@R), which by definition is a quantile of the underlying distribution (see Section 6.2.4).
+Section 6.4.2 in the above reference further shows that in the risk-averse Newsvendor problem, where the objective is a composition of a risk measure and the inventory cost, if the risk measure is real-valued and coherent (e.g., AV@R or the mean-upper-semideviation of order $p$ in Section 6.3.2), the optimal order quantity remains a quantile of a distribution that depends on the risk measure and the inventory parameters.
+In these situations, determining $z^{*}$ reduces to estimating (or “regressing”) the conditional quantile function based on $X$.
+Thus, the Lipschitz continuity of $z^{*}$ hinges on that of the conditional quantile function.
+Under the additive model $Y(x)=\bar{Y}(x)+\epsilon$, where $\epsilon$ is an independent and zero-mean random variable, and $X$ is compact, the continuity of the mean function $\bar{Y}(x)$ implies the desired Lipschitz continuity.
+Another class of problems satisfying Assumption (A4) is a two-stage stochastic optimization problem:
+$\min_{z\in\mathcal{Z}}\quad g(z)+\mathbb{E}\left[Q(Az,Y)\mid X=x\right],$
+where
+$Q(u,Y)=\min\{q^{\top}v:Wv=Y-u,v\geq 0\}.$
+Here, $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is convex, $\mathcal{Z}$ is a closed convex set, and $\mathbb{P}_{Y|X=x}$ is a Borel probability measure with finite first moment.
+Under assumptions of complete recourse and dual feasibility for the second-stage problem, Römisch and Schultz (1993, Theorem 2.2) show that if the function $\mathbb{E}\left[Q(\cdot,Y)\mid X=x\right]$ is strongly convex for all $x\in\mathcal{X}$, then the optimal solution set is a singleton, and there exist constants $L>0$ and $\delta>0$ such that
+$\left\lVert Af^{*}(x_{1})-Af^{*}(x_{2})\right\rVert\leq L\mathcal{W}_{1}(\mathbb{P}_{Y\mid x_{1}},\mathbb{P}_{Y\mid x_{2}})^{1/2},$
+whenever $\mathcal{W}_{1}(\mathbb{P}_{Y\mid x_{1}},\mathbb{P}_{Y\mid x_{2}})\leq\delta$, where $\mathcal{W}_{1}$ denotes the $L_{1}$-Wasserstein distance.
+Consequently, if $A$ is invertible, $\mathcal{X}$ is compact, and $Y(x)=\bar{Y}(x)+\epsilon$ with continuity of $\bar{Y}(\cdot)$, the Lipschitz continuity of the optimal policy follows.
+A subsequent work by Römisch and Schultz (1996) further investigates the Lipschitz stability of the optimal solution sets with respect to a pseudo-metric over the space of probability measures, which may facilitate the establishment of Lipschitz continuous solution mappings in problems with non-unique solutions.
+10.2 Problems not satisfying Assumption (A4)
+In this part, we present examples where Assumption (A4) does not hold.
+Consider the additive random variable model $Y(x)=\bar{Y}(x)+\epsilon$ with a discontinuous mean function $\bar{Y}(x)$ on $\mathcal{X}$.
+Then the resulting optimal decision rule will be discontinuous.
+A notable instance arises in regression problems with the mean squared error loss as the cost function, where the optimal decision rule coincides with $\bar{Y}(x)$ under normally distributed noise $\epsilon$.
+If $\bar{Y}(x)$ is discontinuous or discrete, e.g., due to constraints that depend on $x$, the corresponding optimal decision rule will also be discontinuous.
+For many problems, the set of optimal decisions is inherently disconnected across scenarios.
+One case is when the feasible region $\mathcal{Z}$ itself is disconnected, so that the optimal solutions conditional on different features may lie in different connected components.
+Another case is when the problem structure yields discrete optimal solutions, such as when $\mathcal{Z}$ is a polytope and the optimal decision always lies at an extreme point.
+In these situations, small perturbations of the input can cause abrupt jumps in the optimal decision rule, thereby not satisfying Assumption (A4).
+Addressing such cases calls for the development of Heaviside function classes (Fang et al.
+2025) including decision trees to capture the discontinuous components of the decision rule.
+Advancing such methods constitutes a promising and important direction for future research.
+To conclude this section, we briefly outline a potential direction for analyzing sufficient conditions for the Lipschitz continuity of the optimal decision rule.
+A principled approach is to first characterize conditions on the function $\Phi(z,x)\coloneqq\mathbb{E}_{Y}[\varphi(z,Y)\mid X=x]$ (defined over $\mathcal{Z}\times\mathcal{X}$) that ensure Lipschitz continuity of the solution mapping $z^{*}(x)\in\arg\min_{z\in\mathcal{Z}}\Phi(z,x)$, using implicit function theorems for variational inequalities.
+In particular, with the function $\Phi(z,x)$ being continuously differentiable in $z$ and the generalized equation $0\in\nabla_{z}\Phi(z,x)+\mathcal{N}_{\mathcal{Z}}(z)$ being sufficient for optimality, Dontchev and Rockafellar (2009) provide conditions on the mapping $x\mapsto\nabla_{z}\Phi(\bar{z},x)$ under which the mapping $z^{*}(x)$ is locally Lipschitz continuous (e.g., Dontchev and Rockafellar (2009, Theorem 2B.5)).
+It thus reduces to identify structural properties of $\varphi(z,Y)$ and conditional distributional properties of $Y$ such that those regularity conditions hold, such as growth properties of $\Phi$.
+While we recognize that verifying these conditions is nontrivial in general, this framework offers a promising direction for theoretically analyzing the regularity of optimal policies in contextual SP, which deserves further exploration.
+11 Surrogation Construction for Problems in Section 4
+In this section we provide the specific surrogation construction of $F(\cdot,\xi^{s})$ with $\xi^{s}\in\Xi$ for problems in Section 4, including the Newsvendor problem, the Newsvendor problem with additional concave cost, and the product placement problem.
+For the Newsvendor problem, the cost function $\varphi(z,\xi^{s})=c_{b}\max\{y^{s}-z,0\}+c_{h}\max\{z-y^{s},0\}$ is the sum of two monotonic convex functions.
+By utilizing surrogation construction for Case 2, the surrogate function of $F(\theta,\xi^{s})=\varphi(f(\theta,\xi^{s}),\xi^{s})$ at $\theta^{\nu}$ is constructed as follows,
+$\widehat{F}(\theta,\xi^{s};\theta^{\nu},I)=c_{b}\max\left\{\widehat{f}(\theta,\xi^{s};I_{2})-y^{s},0\right\}+c_{h}\max\left\{y^{s}-\widecheck{f}(\theta,\xi^{s};I_{1}),0\right\},$
+where $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\nu},\xi^{s})$ and $\widehat{f}(\theta,\xi^{s};I_{2})$ and $\widecheck{f}(\theta,\xi^{s};I_{1})$ are the upper and lower surrogate functions of $f(\theta,\xi^{s})$ respectively.
+For the Newsvendor problem with the additional concave piecewise-affine cost function, we denote the concave cost function by the general form $C(z)\coloneqq\min_{i\in[K_{C}]}\{a_{i}z+b_{i}\}$.
+In this case, the cost function is
+$\varphi(z,\xi^{s})$ $=c_{b}\max\{y^{s}-z,0\}+c_{h}\max\{z-y^{s},0\}+\min_{i\in[K_{C}]}\{a_{i}z+b_{i}\}$
+$=c_{b}\max\{y^{s}-z,0\}+c_{h}\max\{z-y^{s},0\}-\max_{i\in[K_{C}]}\{-a_{i}z-b_{i}\},$
+which satisfies Case 1. Following the construction for Case 1, the surrogate function of $F(\theta,\xi^{s})=\varphi(f(\theta,\xi^{s}),\xi^{s})$ at $\theta^{\nu}$ is constructed as below,
+$\widehat{F}(\theta,\xi^{s};\theta^{\nu},I)=$ $c_{b}\max\left\{\widehat{f}(\theta,\xi^{s};I_{2})-y^{s},0\right\}+c_{h}\max\left\{y^{s}-\widecheck{f}(\theta,\xi^{s};I_{1}),0\right\}$
+$-\left(\max\{(-a_{I_{3}}),0\}^{\top}\widecheck{f}(\theta,\xi^{s};I_{1})+\min\{(-a_{I_{3}}),0\}^{\top}\widehat{f}(\theta,\xi^{s};I_{2})-b_{I_{3}}\right),$
+where $I_{1}\in\mathcal{I}^{\varepsilon}_{h}(\theta^{\nu},\xi^{s})$, $I_{2}\in\mathcal{I}^{\varepsilon}_{g}(\theta^{\nu},\xi^{s})$, and $I_{3}\in\mathcal{I}^{\varepsilon}_{C}(f(\theta^{\nu},\xi^{s}),\xi^{s})$.
+Consistent with the surrogation for Case 1, the surrogation index $I=(I_{1},I_{2},I_{3})$ in this case includes the surrogation index for the outer function $C(\cdot)$ at the point $f(\theta^{\nu},\xi^{s})$.
+Finally, for the product placement problem, we first convert the two-stage problem into a single-stage convex problem by extending the auxiliary variable $f$:
+$\min_{\begin{subarray}{c}z\geq 0\\ v^{s}\geq 0,\ s\in[n]\end{subarray}}\quad\frac{1}{n}\sum_{s=1}^{n}\left\{c^{\top}z+g^{\top}v^{s}+b^{\top}\max\{y^{s}+Wv^{s}-z,0\}\right\}.$
+Since $c\geq 0$ and $b\geq 0$, for the component $\varphi(z,v,\xi^{s})=c^{\top}z+g^{\top}v+b^{\top}\max\{y^{s}+Wv-z,0\}$, we construct the surrogate function of $F(\theta,v,\xi^{s})\coloneqq\varphi(f(\theta,\xi^{s}),v,\xi^{s})$ at $(\theta^{\nu})$ as
+$\widehat{F}(\theta,v,\xi^{s};\theta^{\nu},I)=c^{\top}\widehat{f}(\theta,\xi^{s},I_{2})+g^{\top}v+b^{\top}\max\left\{y^{s}+Wv-\widecheck{f}(\theta,\xi^{s};I_{1}),0\right\},$
+where $I=(I_{1},I_{2})\in\mathcal{I}^{\varepsilon}(\theta^{\nu},\xi^{s})$.
+12 Additional Experimental Results
+In this section, we further conduct experiments on the Newsvendor problem using alternative demand models that share a fixed linear component but with different nonlinear structures, including quadratic, cubic, piecewise-affine and sine functions.
+These experiments are designed to assess the robustness of PADR and benchmark methods across varying nonlinear models.
+By incorporating an explicit linear term in all models, we aim to examine whether the nonlinear effects still manifest when combined with a linear component.
+Specifically, we consider the following demand models:
+$\bar{Y}_{\text{ec1}}(X)=3(X_{1}+X_{2})+\omega\cdot\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+20,$ (39)
+$\bar{Y}_{\text{ec2}}(X)=3(X_{1}+X_{2})+\omega\cdot 5(X_{1}-X_{2})^{2}+10,$ (40)
+$\bar{Y}_{\text{ec3}}(X)=3(X_{1}+X_{2})+\omega\cdot[5(X_{1}-X_{2})^{3}-15(X_{1}-X_{2})]+50,$ (41)
+$\bar{Y}_{\text{ec4}}(X)=3(X_{1}+X_{2})+\omega\cdot[4\sin(\pi X_{1})+\max\{16X_{2},-20X_{2}\}]+30.$ (42)
+We report the results under the above models in Figures 1ah-1an.
 […]
+(ag) $p=2\phantom{xxxxxxx}$
+(ah) $p=20$
+Newsvendor problem with varying nonlinearity in demand model (39)
+[…]
+(ai) $p=2\phantom{xxxxxxx}$
+(aj) $p=20$
+Newsvendor problem with varying nonlinearity in demand model (40)
+[…]
+(ak) $p=2\phantom{xxxxxxx}$
+(al) $p=20$
+Newsvendor problem with varying nonlinearity in demand model (41)
+[…]
+(am) $p=2\phantom{xxxxxxx}$
+(an) $p=20$
+Newsvendor problem with varying nonlinearity in demand model (42)
+In general, the comparison results remain similar to those observed in Section 4.1.3.
+Despite the inclusion of an explicit linear component in each demand model, the nonlinear terms continue to significantly affect the performance of benchmark methods.
+Across all settings, PADR consistently ranks among the top-performing methods.
+The observed degradation in NN-DR’s performance in low-dimensional cases persists.
+In contrast, RKHS-DR is competitive at $p=2$, but its performance deteriorates as the feature dimension increases, likely due to the higher sample complexity required for kernel-based methods to generalize in high-dimensional spaces.
+Furthermore, RKHS-DR performs noticeably worse under demand models with nonsmooth components (i.e., models (39) and (42)) than under smooth ones (i.e., models (40) and (41)), suggesting that kernel-based smooth approximations struggle to capture nonsmooth structures without sufficient data.
+These results confirm the robustness of PADR to various nonlinear demand models with superior performance, compared to all other benchmarks, including NN-DR and RKHS-DR.
+[…]
+The second author gratefully acknowledge the support by National Natural Science Foundation of China under grant number 72201151, and 72188101.